### v1

# Semilinear spaces of matrices with transitivity, rank or spectral constraints

Abstract
Let $\mathbb{D}$ be a division ring with finite degree $d$ over a central subfield $\mathbb{F}$. Under mild cardinality assumptions on $\mathbb{F}$, we determine the greatest possible dimension for an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ in which no matrix has a nonzero fixed vector, and we elucidate the structure of the spaces that have the greatest dimension. The classification involves the associative composition algebras over $\mathbb{F}$ and requires the study of intransitive spaces of operators between finite-dimensional vector spaces.
These results are applied to solve various problems of the same flavor, including the determination of the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which all the matrices have rank at least $r$ (for some fixed $r$), as well as the structure of the spaces that have the greatest possible dimension if $n=p=r$; and finally the structure of large $\mathbb{F}$-linear subspaces of $\mathbb{F}$-diagonalisable matrices is elucidated.
AMS MSC: 15A30, 16K20
Keywords: affine space, rank, matrices over division rings, intransitivity, trivial spectrum spaces, simultaneous reduction, associative composition algebras
List of notation
[…]
1 Introduction
1.1 The problem
Throughout, we let $\mathbb{D}$ be a (unital and associative) division ring, $\mathbb{F}$ be a subfield of the center of $\mathbb{D}$, and we assume that $\mathbb{D}$ has finite degree $d\geq 1$ over $\mathbb{F}$. All the vector spaces over $\mathbb{D}$ we manipulate will be finite-dimensional and right vector spaces, with the notable exception of dual spaces which are naturally left vector spaces. This way, we can safely practice matrix computations in the usual sense, with vectors represented by columns.
The space $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ of all endomorphisms of the $\mathbb{D}$-vector space $V$ has no natural structure of $\mathbb{D}$-vector space (because we are avoiding any commutativity assumption), but it can nevertheless be seen as a vector space over $\mathbb{F}$. An element $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have $\mathbb{F}$-trivial spectrum when it has no nonzero (right) eigenvalue in $\mathbb{F}$ (we adopt a similar definition for matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$). A subset of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have $\mathbb{F}$-trivial spectrum whenever all its elements have $\mathbb{F}$-trivial spectrum. For an $\mathbb{F}$-linear subspace $S$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, to have $\mathbb{F}$-trivial spectrum is equivalent to the condition that no element of $S$ has $1$ as eigenvalue, i.e., no element of $S$ has a nonzero fixed vector.
In matrix terms, an example of large $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with $\mathbb{F}$-trivial spectrum is the set of all upper-triangular matrices
$\begin{bmatrix}x_{1}&&*\\ &\ddots&\\ (0)&&x_{n}\end{bmatrix}$
in which the $i$-th diagonal entry $x_{i}$ is required to belong to a given $\mathbb{F}$-linear hyperplane $H_{i}$ of $\mathbb{D}$ such that $\mathbb{F}\oplus H_{i}=\mathbb{D}$. In case $\mathbb{D}=\mathbb{F}$, all the hyperplanes $H_{i}$ are zero and we have simply described the strictly upper-triangular matrices over $\mathbb{D}$. Note that the dimension over $\mathbb{F}$ of the above space is
$\alpha(n,d):=n(d-1)+d\,\frac{n(n-1)}{2},$
and it will be regularly useful to observe that
$\forall m,\;\forall n,\quad\alpha(m+n,d)=\alpha(m,d)+\alpha(n,d)+mnd.$
Trivial spectrum subspaces have been studied in the last decade [15, 18], but only in the special case $\mathbb{D}=\mathbb{F}$. One motivation for considering such spaces is that they are a natural variation of linear subspaces of nilpotent operators: it is unavoidable to cite the celebrated Gerstenhaber theorem [7], which states that the greatest possible dimension for a linear subspace of nilpotent $n$-by-$n$ matrices over a field is $\frac{n(n-1)}{2}$, with equality attained only for the conjugates of the subspace of all strictly upper-triangular matrices. See [17] for a generalization to arbitrary division rings, and [9, 26] for alternative proofs.
There is a more profound justification for studying trivial spectrum subspaces: they are intimately connected with $\mathbb{F}$-affine subspaces of invertible elements of the ring $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Let indeed $\mathcal{S}$ be an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, with translation vector space denoted by $\overrightarrow{S}$ (hence $\overrightarrow{S}$ is a $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$). Assume that $\mathcal{S}$ contains $\operatorname{\operatorname{id}}_{V}$. Then $\mathcal{S}\subseteq\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$ if and only if $\operatorname{\operatorname{id}}_{V}-u$ is invertible for all $u\in\overrightarrow{S}$, i.e., no element of $\overrightarrow{S}$ has a nonzero fixed vector. As noted in the above, this is equivalent to the condition that $\overrightarrow{S}$ has $\mathbb{F}$-trivial spectrum. It follows that the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is exactly the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that is included in $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$. And by a method due to Meshulam [10], this can also be used to obtain the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which all the elements have their rank greater than or equal to some fixed $r\in\mathopen{[\![}1,\min(n,p)\mathclose{]\!]}$.
In [23], we have also discovered a surprising connection between $\mathbb{F}$-trivial spectrum subspaces (when $\mathbb{D}=\mathbb{F}$) and the apparently opposite notion of an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, meaning an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ in which every element is diagonalisable over $\mathbb{F}$ (i.e., every element is represented in some basis by a diagonal matrix with all diagonal entries in $\mathbb{F}$, or alternatively is annihilated by a polynomial of $\mathbb{F}[t]$ that is split with simple roots over $\mathbb{F}$). In particular, it was shown that the classification of the trivial spectrum subspaces over $\mathbb{R}$ (there, $\mathbb{D}=\mathbb{F}=\mathbb{R}$) with the greatest possible dimension yields the one of the $\mathbb{R}$-diagonalisable linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ with maximal dimension, which simply turn out to be the conjugates of the subspace $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ of all $n$-by-$n$ symmetric matrices. As a special case of the results on $\mathbb{F}$-trivial spectrum subspaces, we will obtain both the Hermitian and quaternionic cases, and in particular we will prove that every $\mathbb{R}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ has dimension over $\mathbb{R}$ at most $n^{2}$, with equality attained only for the conjugates of the subspace $\mathcal{H}_{n}(\mathbb{C})$ of all $n$-by-$n$ (complex) Hermitian matrices.
1.2 Reduction to the irreducible case
An $\mathbb{F}$-trivial spectrum (linear) subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is called optimal when it has the greatest possible dimension among such spaces. A subset $\mathcal{X}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is called irreducible when $V$ is nonzero and $\mathcal{X}$ has no nontrivial invariant $\mathbb{D}$-linear subspace, and is called reducible otherwise. The study of optimal $\mathbb{F}$-trivial spectrum subspaces can entirely be reduced to the one of irreducible optimal subspaces. The reduction process is not specific to $\mathbb{F}$-trivial spectrum subspaces of endomorphisms, and is actually shared by many properties on matrices: an account of this is given in the appendix of [25], and here we will simply retrieve the results. Key to this is the observation that the $\mathbb{F}$-trivial spectrum property satisfies the following two conditions:
  • It is invariant under similarity, i.e., if $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ has this property then so does $\varphi\circ u\circ\varphi^{-1}$ for every isomorphism $\varphi:V\overset{\simeq}{\longrightarrow}V^{\prime}$ of $\mathbb{D}$-vector spaces;
  • It satisfies the strong inheritance condition, meaning that for every $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ and every nontrivial $\mathbb{D}$-linear subspace $W$ of $V$ that is invariant under $u$, the endomorphism $u$ has $\mathbb{F}$-trivial spectrum if and only if the endomorphisms of $W$ and $V/W$ induced by $u$ have $\mathbb{F}$-trivial spectrum (this follows from the observation that $u$ can be seen as an endomorphism of the $\mathbb{F}$-vector space $V$, which is finite-dimensional).
By lemma A.1 in [25], we deduce the following result:
Proposition 1.1.
Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then the $\mathcal{S}$-invariant subspaces of $V$ are totally ordered under inclusion.
A flag of ($\mathbb{D}$-linear subspaces of) $V$ is an increasing list $V_{0}\subset\cdots\subset V_{p}$ of $\mathbb{D}$-linear subspaces with $V_{0}=\{0\}$ and $V_{p}=V$. Beware that we do not require that $\dim_{\mathbb{D}}V_{i}=i$ for all $i\in\mathopen{[\![}0,p\mathclose{]\!]}$ (in which case we would say that this flag is complete). The previous result can then be restated as saying that when $\mathcal{S}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, the $\mathcal{S}$-invariant subspaces form a flag.
Next, let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, and consider the $\mathcal{S}$-invariant subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$. Given respective subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{1}/V_{0}),\dots,\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{p}/V_{p-1})$, one denotes by $\mathcal{X}_{1}\vee\cdots\vee\mathcal{X}_{p}$ the set of all $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that leave all the $V_{i}$’s invariant and induce, for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an element of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that belongs to $\mathcal{X}_{i}$. Then, as a consequence of theorem A.2 in [25], we obtain:
Theorem 1.2.
Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum (linear) subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, with invariant subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$. Denote, for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, by $\mathcal{S}_{i}$ the set of all endomorphisms of $V_{i}/V_{i-1}$ induced by the elements of $\mathcal{S}$. Then
$\mathcal{S}=\mathcal{S}_{1}\vee\cdots\vee\mathcal{S}_{p},$
and for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the set $\mathcal{S}_{i}$ is an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$.
Let us now give a matrix interpretation of the above. Two subsets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are called similar, and we write $\mathcal{X}\simeq\mathcal{X}^{\prime}$, when there exists $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that $\mathcal{X}^{\prime}=P\mathcal{X}P^{-1}$.
Let $n_{1},\dots,n_{p}$ be positive integers. For respective subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$ of $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, we denote by $\mathcal{X}_{1}\vee\cdots\vee\mathcal{X}_{p}$ the set of all matrices of the form
$M=\begin{bmatrix}M_{1}&&[?]\\ &\ddots&\\ [0]&&M_{p}\end{bmatrix}\quad\text{where $M_{1}\in\mathcal{X}_{1},\dots,M_{p}\in\mathcal{X}_{p}$.}$
We call it the joint of the sets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$.
Let $\mathcal{M}$ be an optimal $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then there exists a unique partition $n=n_{1}+\cdots+n_{p}$ into positive integers, and, for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an irreducible optimal $\mathbb{F}$-linear subspace $\mathcal{M}_{i}$ of $\operatorname{\operatorname{M}}_{n_{i}}(\mathbb{D})$ such that
$\mathcal{M}\simeq\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}.$
Moreover, each $\mathcal{M}_{i}$ is uniquely determined by $\mathcal{M}$ up to similarity.
As a conclusion, we find that understanding the optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ essentially amounts to understanding the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{k}(\mathbb{D})$ for all integers $k\in\mathopen{[\![}1,n\mathclose{]\!]}$. There is a small difficulty however: given a partition $n=n_{1}+\cdots+n_{p}$ and respective irreducible optimal $\mathbb{F}$-trivial spectrum subspaces $\mathcal{M}_{1},\dots,\mathcal{M}_{p}$ of $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, it is possible in theory that $\mathcal{M}_{1}\vee\dots\vee\mathcal{M}_{p}$, which is an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is not an optimal one. We call this the composition problem. Fortunately, in the situations we will consider, it will turn out that the joint of irreducible optimal $\mathbb{F}$-trivial spectrum subspaces is always optimal, and this will stem from the observation that
$\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}=\alpha(n,d).$
1.3 Basic considerations on trivial spectrum subspaces, and a review of known results
Now, we recall the main known results on trivial spectrum subspaces in the case $\mathbb{D}=\mathbb{F}$.
Theorem 1.3 (See [12, 15]).
Let $V$ be a vector space over $\mathbb{F}$ with finite dimension $n$. Then the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}(V)$ is $\dbinom{n}{2}=\alpha(n,1)$.
In particular, the composition problem stated in the previous section has a positive answer when $\mathbb{D}=\mathbb{F}$.
As far as the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces are concerned, their structure is described by the main theorem of [18] when $\mathbb{D}=\mathbb{F}$, which excludes the case $|\mathbb{F}|=2$ (in that case there are enough counterexamples to discourage anyone to try formulating a plausible conjecture [22]). Before we restate this theorem, some additional definitions and notation are required. Beware that throughout we manipulate bilinear forms that are possibly nonsymmetric.
Definition 1.1.
Let $b$ be a bilinear form on an $\mathbb{F}$-vector space $V$. We say that $b$ is nonisotropic whenever
$\forall x\in V,\;b(x,x)\neq 0.$
An endomorphism $u\in\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ is called $b$-alternating (on the right) whenever
$\forall x\in V,\;b(x,u(x))=0,$
i.e., the bilinear form $(x,y)\in V^{2}\mapsto b(x,u(y))$ is alternating.
We denote by $\mathcal{A}_{b}$ the set of all $b$-alternating $\mathbb{F}$-endomorphisms of $V$ (it is a linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$). If $V$ has an additional structure of $\mathbb{D}$-vector space (which extends its structure of $\mathbb{F}$-vector space), we set
$\mathcal{A}_{b,\mathbb{D}}:=\mathcal{A}_{b}\cap\operatorname{\operatorname{End}}_{\mathbb{D}}(V).$
If $V$ has dimension $n$ over $\mathbb{D}$ and $b$ is nondegenerate, then $\mathcal{A}_{b}$ is naturally isomorphic to the space of all alternating bilinear forms on $V$, and hence has dimension $\frac{nd(nd-1)}{2}$ over $\mathbb{F}$. Moreover, if $b$ is nonisotropic then the linear subspace $\mathcal{A}_{b}$ has $\mathbb{F}$-trivial spectrum, as for all $x\in V$ and $u\in\mathcal{A}_{b}$ the equality $u(x)=x$ implies $b(x,x)=b(x,u(x))=0$ and hence $x=0$. Now, we can state the main result of [18], reformulated in geometric terms.
Theorem 1.4 (See theorems 3 and 4 in [18]).
Let $V$ be a nonzero (finite-dimensional) $\mathbb{F}$-vector space, with $|\mathbb{F}|>2$. Then:
  1. (a)
    The irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are the spaces of the form $\mathcal{A}_{b}$, where $b$ is a nonisotropic bilinear form on $V$.
  2. (b)
    Given two nonisotropic bilinear forms $b$ and $b^{\prime}$ on $V$, the equality $\mathcal{A}_{b}=\mathcal{A}_{b^{\prime}}$ holds if and only if $b^{\prime}$ is a scalar multiple of $b$, and $\mathcal{A}_{b}$ is similar [^3] Two subsets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are called similar when there exists a vector space automorphism $\varphi$ of $V$ such that $\mathcal{X}^{\prime}=\{\varphi\circ u\circ\varphi^{-1}\mid u\in\mathcal{X}\}$. to $\mathcal{A}_{b^{\prime}}$ if and only if $b^{\prime}$ is congruent to $b$ (i.e., equivalent to a scalar multiple of $b$).
Denote by $\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ the space of all alternating matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ (a matrix $M$ is alternating if and only if it represents an alternating bilinear form in a basis, or in other words $M^{T}=-M$ and all the diagonal entries of $M$ are equal). In matrix terms, point (a) of Theorem 1.4 means that the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are the spaces of the form $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ for the nonisotropic matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$. And the second part of point (b) means that two such spaces $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ and $Q^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ are similar if and only if $P$ is congruent to a scalar multiple of $Q$ (i.e., $P=\lambda\,RQR^{T}$ for some $\lambda\in\mathbb{F}^{\times}$ and some $R\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$). Moreover, for the more general case of (potentially reducible) optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$, we have the following corollary:
Theorem 1.5.
Let $V$ be an $\mathbb{F}$-vector space, where $|\mathbb{F}|>2$, and $\mathcal{S}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$. Then the $\mathcal{S}$-invariant linear subspaces form a flag of subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and there exists a list $(b_{1},\dots,b_{p})$ of nonisotropic bilinear forms, with $b_{i}$ defined on $(V_{i}/V_{i-1})^{2}$, such that $\mathcal{S}$ is the space of all $u\in\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ that leave each subspace $V_{i}$ invariant and induce a $b_{i}$-alternating endomorphism $u_{i}$ of $V_{i}/V_{i-1}$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$. Moreover each form $b_{i}$ is uniquely determined up to multiplication by a nonzero scalar.
In matrix terms this yields that for every optimal trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$, there is a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of nonisotropic matrices such that $\mathcal{M}$ is conjugated to the space of all matrices of the form
$\begin{bmatrix}P_{1}^{-1}A_{1}&[?]&&[?]\\ [0]&P_{2}^{-1}A_{2}&\ddots&\\ \vdots&\ddots&\ddots&[?]\ \\ [0]&\cdots&[0]&P_{p}^{-1}A_{p}\end{bmatrix}$
where $A_{1}\in\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}),\dots,A_{p}\in\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F})$ are variable alternating matrices, and the question marks represent arbitrary blocks. Moreover, the list $(n_{1},\dots,n_{p})$ is uniquely determined by $\mathcal{M}$, and each $P_{i}$ is uniquely determined by $\mathcal{M}$ up to congruence and multiplication with a nonzero scalar. Conversely, every space of the above form is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
Hence the structure of the optimal trivial spectrum subspaces is intimately connected with the quadratic structure of $\mathbb{F}$. If $\mathbb{F}$ is quadratically closed, irreducible optimal subspaces only exist in dimension $1$, and we obtain the same result as in Gerstenhaber’s theorem (simultaneous triangularization). If $\mathbb{F}$ is finite then irreducible optimal subspaces only exist in dimensions $1$ and $2$, and there is a very limited number of similarity classes of such spaces. If $\mathbb{F}$ is the field of reals, there are irreducible optimal subspaces in all dimensions, and their classification up to conjugation is rather easy (it can be reduced to the determination of the orbits, under orthogonal conjugation, of the skewsymmetric matrices).
1.4 Main new results on trivial spectrum subspaces
We will now state a generalization of the previous results, however with a cardinality assumption on $\mathbb{F}$.
Theorem 1.6.
Let $\mathbb{D}$ be a division ring with degree $d$ over a field $\mathbb{F}$, and $V$ be a (right) vector space over $\mathbb{D}$ with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq dn$. Then the greatest possible dimension for an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or, equivalently, of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is
$\alpha(n,d)=d\,\dbinom{n}{2}+n(d-1).$
An immediate consequence is that when $|\mathbb{F}|\geq dn$ the composition problem stated in the end of Section 1.2 has a positive answer in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$: Indeed in that case $|\mathbb{F}|\geq dk$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$, and for every partition $n=n_{1}+\cdots+n_{p}$ the number $\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}$ equals $\alpha(n,d)$.
Under the same assumption as in Theorem 1.6, the classification of the irreducible optimal spaces is partly given by the following theorem:
Theorem 1.7 (Partial classification theorem for irreducible optimal $\mathbb{F}$-trivial spectrum subspaces).
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$.
  1. (i)
    If $n=1$ then the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ are the $\mathbb{F}$-linear hyperplanes of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that do not contain $\operatorname{\operatorname{id}}_{V}$.
  2. (ii)
    If $n\geq 2$ then every irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ equals $\mathcal{A}_{b,\mathbb{D}}$ for some nonisotropic $\mathbb{F}$-bilinear form $b$ on $V$.
Note that point (i) is essentially obvious.
Of course, the above two theorems look on the surface as plain generalizations of the case $\mathbb{D}=\mathbb{F}$, and in some sense they are. But there is a subtlety: although $\mathcal{A}_{b,\mathbb{D}}$ is always an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, in general it is not an optimal one! Hence it is necessary to understand when optimality happens. Some additional terminology is necessary for this.
Definition 1.2.
We say that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ whenever one of the following conditions holds:
  1. (i)
    $d=1$;
  2. (ii)
    $d=2$ and $\mathbb{D}$ is a separable extension of $\mathbb{F}$;
  3. (iii)
    $d=4$ and $\mathbb{D}$ is a quaternion algebra over $\mathbb{F}$;
  4. (iv)
    $d=2^{n}$ for some $n\geq 1$, $\operatorname{\operatorname{char}}(\mathbb{F})=2$ and $\mathbb{D}$ is a hyper-radicial [^4] We prefer this terminology over the one of purely inseparable extensions of exponent $1$, since it is more compact and quite natural. extension of $\mathbb{F}$, i.e., $\mathbb{D}$ is a field and $\forall x\in\mathbb{D},\;x^{2}\in\mathbb{F}$.
Cases (i) to (iii) are called the separable types.
Definition 1.3.
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$. Denote by $\sigma$ either the identity of $\mathbb{F}$ in cases (i) and (iv), the non-identity automorphism in case (ii), and the quaternionic conjugation in case (iii); we say that $\sigma$ is the standard involution on $\mathbb{D}$.
In cases (i) and (ii), we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the standard trace of $\mathbb{D}$ over $\mathbb{F}$ (see Section 2.1 for details), and in case (iii) we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the reduced trace $x\mapsto x+\sigma(x)$ (the reduced trace of $x$ is either $2x$ if $x$ belongs to $\mathbb{F}1_{\mathbb{D}}$, otherwise it equals the opposite of the coefficient on $t$ of the minimal polynomial of $x$ over $\mathbb{F}$).
We say that $(\sigma,e)$ is an associated pair for $\mathbb{D}$ over $\mathbb{F}$ when $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, $\sigma$ is its standard involution, and $e$ is either $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ when $\mathbb{D}$ is of separable type, or an arbitrary nonzero $\mathbb{F}$-linear form on $\mathbb{D}$ otherwise.
Finally, if $\sigma$ is an antiautomorphism of the $\mathbb{F}$-algebra $\mathbb{D}$, we define a sesquilinear form on $V$ as a mapping $B:V^{2}\rightarrow\mathbb{D}$ that is right-linear and left-$\sigma$-quasilinear, meaning that $B(\alpha x+y,-)=\sigma(\alpha)B(x,-)+B(y,-)$ and $B(-,\alpha x+y)=\alpha B(-,x)+B(-,y)$ for all $\alpha\in\mathbb{D}$ and $x,y$ in $V$. Two such forms $B$ and $B^{\prime}$ are called equivalent if there exists an automorphism $\varphi$ of the $\mathbb{D}$-vector space $V$ such that $B^{\prime}(\varphi(x),\varphi(y))=B(x,y)$ for all $(x,y)\in V^{2}$.
Theorem 1.8 (Full classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$. For $\mathcal{A}_{b,\mathbb{D}}$ to be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, it is necessary and sufficient that all the following conditions hold:
  1. (i)
    $\mathbb{D}$ is of quadratic type over $\mathbb{F}$;
  2. (ii)
    Given an associated pair $(\sigma,e)$ for $\mathbb{D}$ over $\mathbb{F}$, there is a $\sigma$-sesquilinear form $B$ on the $\mathbb{D}$-vector space $V$ such that
    $\forall(x,y)\in V^{2},\;b(x,y)=e(B(x,y)).$
In that case, $B$ is uniquely determined by $b$.
Hence, if $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ with associated pair $(\sigma,e)$, and we take any $\sigma$-sesquilinear form $B$ on $V$, the space $\mathcal{A}_{b,\mathbb{D}}$ is an irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$ provided that $B$ is $e$-nonisotropic , meaning that $\forall x\in V,\;e(B(x,x))\neq 0$ (for this very result the provision $|\mathbb{F}|\geq nd$ is not needed, see the proof in Section 5.6).
As a consequence of Theorems 1.7 and 1.8, and still under the cardinality assumption from Theorem 1.6, we can refine our understanding of the irreducible optimal trivial spectrum subspaces:
Corollary 1.9.
Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. The space $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ contains an irreducible optimal $\mathbb{F}$-trivial spectrum subspace only if $n=1$.
In case $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, the following result slightly reinforces Theorem 1.7 by taking the $1$-dimensional spaces into account.
Theorem 1.10.
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Then every irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ equals $\mathcal{A}_{b,\mathbb{D}}$ for some nonisotropic $\mathbb{F}$-bilinear form $b$ on $V$. Moreover, for two such forms $b$ and $b^{\prime}$, the equality $\mathcal{A}_{b,\mathbb{D}}=\mathcal{A}_{b^{\prime},\mathbb{D}}$ is satisfied if and only if $b^{\prime}$ is a scalar multiple of $b$.
The next result deals with the conjugacy class of $\mathcal{A}_{b,\mathbb{D}}$ inside the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
Theorem 1.11.
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Let $B$ and $B^{\prime}$ be $e$-nonisotropic $\sigma$-sesquilinear forms on $V$, and set $b:(x,y)\mapsto e(B(x,y))$ and $b^{\prime}:(x,y)\mapsto e(B^{\prime}(x,y))$. Then $\mathcal{A}_{b,\mathbb{D}}$ is conjugated to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ in the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ if and only if $B^{\prime}$ is equivalent to $\alpha B$ as a $\sigma$-sesquilinear form for some $\alpha\in\mathbb{F}^{\times}$.
It is a good idea now to express the above results in matrix terms. So, let us consider the space $V=\mathbb{D}^{n}$, with $n\geq 1$, and let us identify $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with the matrix space $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, so that the matrix $A$ corresponds to $X\in\mathbb{D}^{n}\mapsto AX\in\mathbb{D}^{n}$. Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. For $A\in\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$, write $A^{\star}=(A^{\sigma})^{T}$ as the $\sigma$-starconjugate of $A$ (so that $(AB)^{\star}=B^{\star}A^{\star}$ whenever $AB$ makes sense). The $\sigma$-sesquilinear forms on $\mathbb{D}^{n}$ are the mappings of the form
$B_{P}:(X,Y)\mapsto X^{\star}PY\quad\text{with $P\in\operatorname{\operatorname{M}}_{n}(\mathbb{D})$},$
and such a form is nondegenerate if and only if $P$ is invertible. Of course, such a form is $e$-nonisotropic only if it is nondegenerate. Now, letting $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ and setting $b:(X,Y)\in(\mathbb{D}^{n})^{2}\mapsto e(X^{\star}PY)$, the alternating space $\mathcal{A}_{b,\mathbb{D}}$ is the set of all $A\in\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ such that $PA$ satisfies the following equation with unknown $M$:
$\forall X\in\mathbb{D}^{n},\;e(X^{\star}MX)=0.$ (1)
Proposition 1.12.
The solution set of (1) is the $\mathbb{F}$-linear subspace
$\mathcal{S}\mathcal{H}_{n}(\mathbb{D}):=\{M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D}):M^{\star}=-M\quad\text{and}\quad\forall i\in\mathopen{[\![}1,n\mathclose{]\!]},\;e(m_{i,i})=0\}$
of all skew-Hermitian matrices whose diagonal entries belong to $\operatorname{\operatorname{Ker}}e$.
Proof.
Let $M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ satisfy (1). The identity $\forall i\in\mathopen{[\![}1,n\mathclose{]\!]},\;e(m_{i,i})=0$ is then obtained by applying (1) to the vectors of the standard basis of $\mathbb{D}^{n}$. Next, let $(X,Y)\in(\mathbb{D}^{n})^{2}$. We find by polarizing that $e(X^{\star}MY)+e(Y^{\star}MX)=0$, and hence
$e(X^{\star}(M+M^{\star})Y)=e(X^{\star}MY)+e((Y^{\star}MX)^{\star})=e(X^{\star}MY)+e(Y^{\star}MX)=0.$
Letting $a\in\mathbb{D}$ and applying this to $Ya$ instead of $Y$ leads to $X^{\star}(M+M^{\star})Y=0$. Then, varying $X$ and $Y$ yields $M+M^{\star}=0$.
Conversely, let $M\in\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$. By the same computation as in the above, we see that the identity $M^{\star}=-M$ leads to $X\in(\mathbb{D}^{n})\mapsto e(X^{\star}MX)$ being additive (its polar form vanishes). As a consequence, for all $X=\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$, we find
$e(X^{\star}MX)=\sum_{i=1}^{n}e(x_{i}^{\star}m_{i,i}x_{i})=\sum_{i=1}^{n}e(\underbrace{x_{i}x_{i}^{\star}}_{\in\mathbb{F}}m_{i,i})=\sum_{i=1}^{n}x_{i}x_{i}^{\star}e(m_{i,i})=0.$
Next, we can give a clearer description of $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ by discussing the precise kind of pair $(\mathbb{D},\mathbb{F})$ under consideration.
  • If $d=1$ then $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})=\operatorname{\operatorname{A}}_{n}(\mathbb{F})$, and the constraint $\forall i\in\mathopen{[\![}1,n\mathclose{]\!]},\;m_{i,i}=0$ is critical to deal with fields with characteristic $2$.
  • If $d\geq 2$ and $\mathbb{D}$ is separable over $\mathbb{F}$ then $e=\operatorname{\operatorname{id}}+\sigma$ and hence we simply have $\mathcal{S}\mathcal{H}_{n}(\mathbb{D}):=\{M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D}):M^{\star}=-M\}$ in that case.
  • If $d\geq 2$ and $\mathbb{D}$ is inseparable over $\mathbb{F}$ then $\sigma=\operatorname{\operatorname{id}}$ and $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is the space of all symmetric matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ whose diagonal entries belong to $\operatorname{\operatorname{Ker}}e$ (in that case the constraint $\forall i\in\mathopen{[\![}1,n\mathclose{]\!]},\;e(m_{i,i})=0$ is unavoidable).
Hence, according to the type of $\mathbb{D}$ over $\mathbb{F}$, we can partially translate the previous results in matrix terms:
Corollary 1.13.
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Assume that $|\mathbb{F}|\geq nd$. Then:
  1. (a)
    The irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces of the form $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ where $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ is $e$-nonisotropic.
  2. (b)
    Given $e$-nonisotropic matrices $P_{1}$ and $P_{2}$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$, the spaces $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ and $P_{2}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ are conjugated under an element of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ if and only if $P_{2}$ is star-congruent to $\alpha P_{1}$ for some $\alpha\in\mathbb{F}^{\times}$ (two matrices $M$ and $M^{\prime}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are star-congruent if and only if there exists $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that $M^{\prime}=QMQ^{\star}$).
At this point, the reader might be puzzled by the hyper-radicial case and doubt the existence of such kinds of optimal trivial spectrum spaces. Yet, they actually exist in arbitrary dimensions. Take for instance, for $n\geq 2$ and $p\geq 1$, a list $(X_{1},\dots,X_{n+p})$ of algebraically independent indeterminates over $\mathbb{F}_{2}$, and consider the field $\mathbb{D}=\mathbb{F}_{2}(X_{1},\dots,X_{n+p})$ and the subfields
$\mathbb{F}:=\mathbb{F}_{2}(X_{1},\dots,X_{n},(X_{n+1})^{2},\dots,(X_{n+p})^{2})$
and
$\mathbb{K}:=\mathbb{F}_{2}((X_{1})^{2},\dots,(X_{n})^{2},(X_{n+1})^{2},\dots,(X_{n+p})^{2}).$
Classically $\mathbb{D}:\mathbb{F}]=2^{p}$, $\mathbb{F}:\mathbb{K}]=2^{n}$ and $\mathbb{K}$ is the set of all squares of elements of $\mathbb{D}$; moreover $X_{1},\dots,X_{n}$ are linearly independent over $\mathbb{K}$. Now, take an arbitrary $\mathbb{F}$-linear form $e$ on $\mathbb{D}$ that maps $X_{1},\dots,X_{n}$ to themselves, and consider the diagonal matrix $P:=\operatorname{\operatorname{Diag}}(X_{1},\dots,X_{n})$. For all $Y=\begin{bmatrix}y_{1}&\cdots&y_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$ we have $e(Y^{T}PY)=\sum_{i=1}^{p}(y_{i})^{2}X_{i}$, which vanishes if and only if $y_{1}=\cdots=y_{n}=0$. Hence $P$ is $e$-nonisotropic. Therefore $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and both $n$ and $p$ can be taken as large as we wish.
Let us conclude by combining Theorem 1.2 with the previous results, so as to give a full description of the optimal $\mathbb{F}$-trivial spectrum spaces. We distinguish between two cases, whether $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ or not.
Theorem 1.14.
Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then the $\mathcal{S}$-invariant $\mathbb{D}$-linear subspaces form a complete flag $(V_{0},\dots,V_{n})$ of $V$, and $\mathcal{S}=H_{1}\vee\cdots\vee H_{n}$, where for each $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, $H_{i}$ is an $\mathbb{F}$-linear hyperplane of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that does not contain the identity of $V_{i}/V_{i-1}$.
Conversely, given a complete flag $(V_{0},\dots,V_{n})$ of $V$, and for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$ an $\mathbb{F}$-linear hyperplane $H_{i}$ of $V_{i}/V_{i-1}$ that does not contain the identity of $V_{i}/V_{i-1}$, the space $H_{1}\vee\cdots\vee H_{n}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
Here is a restatement in matrix terms:
Theorem 1.15.
Let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$, and assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$.
The optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces that are conjugated to a space of the form
$H_{1}\vee\cdots\vee H_{n}=\left\{\begin{bmatrix}\alpha_{1}&&[*]\\ &\ddots&\\ [0]&&\alpha_{n}\end{bmatrix}\mid(\alpha_{1},\dots,\alpha_{n})\in H_{1}\times\cdots\times H_{n}\right\}$
where $H_{1},\dots,H_{n}$ are $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$. Two such spaces $H_{1}\vee\cdots\vee H_{n}$ and $H^{\prime}_{1}\vee\cdots\vee H^{\prime}_{n}$ are conjugated in the algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ if and only if, for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, the hyperplanes $H_{i}$ and $H^{\prime}_{i}$ are conjugated in $\mathbb{D}$.
Theorem 1.16 (Full classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Let $V$ be a $\mathbb{D}$-vector space with finite dimension $n>0$, where $|\mathbb{F}|\geq nd$, and $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
Then the $\mathcal{S}$-invariant $\mathbb{D}$-linear subspaces constitute a (potentially incomplete) flag $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there exists an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$ such that, for $b_{i}:=e\circ B_{i}$,
$\mathcal{S}=\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}.$
Moreover, each form $B_{i}$ is uniquely determined by $\mathcal{S}$ up to multiplication by an element of $\mathbb{F}^{\times}$, and the equivalence class of $B_{i}$ is uniquely determined by the conjugacy class of $\mathcal{S}$ in $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ up to multiplication by an element of $\mathbb{F}^{\times}$.
Conversely, given a (potentially incomplete) flag $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$, with associated $\mathbb{F}$-bilinear form $b_{i}:=e\circ B_{i}$, the space $\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
We finish with a restatement in matrix terms:
Theorem 1.17 (Classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$.
Let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$. For every optimal $\mathbb{F}$-trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, there exists a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{D})$ of $e$-nonisotropic matrices such that
$\mathcal{M}\simeq P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}).$
The integers $n_{1},\dots,n_{p}$ are uniquely determined by $n$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the matrix $P_{i}$ is uniquely determined by $\mathcal{M}$ up to star-congruence and multiplication by an element of $\mathbb{F}^{\times}$.
Conversely, for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of $e$-nonisotropic matrices, the space $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D})$, as well as all its conjugates in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
1.5 The strategy, and intransitive operator spaces
At this point, we have stated the main results of the present article, which are the ones on trivial spectrum spaces. They are far from the only results here, and we will expand on the other results in a moment.
But before we come to the other results, let us explain the main strategy for proving those on trivial spectrum spaces. So far, we have not succeeded in generalizing the techniques of [15, 18] to the case $d>1$. This has led us to reconsider an alternative proof of Theorem 1.4, limited to fields with more than n1n-1 elements, which we discovered ten years ago [19]. Critical to the proof is the observation that the $\mathbb{F}$-trivial spectrum property implies a lack of transitivity from the said spaces of operators, meaning that if $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum then $x\not\in\mathcal{S}x$ for every nonzero vector $x\in V\smallsetminus\{0\}$, and in particular $\mathcal{S}x\subsetneq V$. The key insight is then to introduce the dual operator space $\widehat{\mathcal{S}}$ of $\mathcal{S}$, which consists of all the evaluation mappings
$\widehat{x}:u\in\mathcal{S}\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
Then $\widehat{\mathcal{S}}$ is a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$, and every element of $\widehat{\mathcal{S}}$ is nonsurjective and hence has its rank less than $\dim_{\mathbb{F}}V$. The key in [19] was to rely upon the theory of primitive spaces of bounded rank matrices, studied first by Atkinson and Lloyd [1, 2] and rediscovered a few years later by Eisenbud and Harris [4]. Thanks to Atkinson’s classification of primitive matrix spaces that satisfy an extremal condition between the numbers of rows and columns (theorem C of [2]), and more precisely to our own generalization to semi-primitive spaces (see section 5 of [21] and a restatement as theorem 3.1 in [19]), it was then possible to find an easy derivation of Theorem 1.4 in case $|\mathbb{F}|\geq n$ (an unavoidable assumption if one wants to use Atkinson’s theorem).
Yet, for the case $d>1$ a major issue is that the space $\widehat{\mathcal{S}}$ is far from satisfying Atkinson’s extremal condition between the respective dimensions of the source and target spaces, even if $\mathcal{S}$ is optimal; hence nothing can be obtained from the theory of primitive spaces of bounded rank matrices. Nevertheless, a version of Atkinson’s theorem (theorem 5.3 of [21]) can be entirely reformulated in the viewpoint of spaces of operators that lack transitivity, as we shall now see.
A few basic definitions and notation will help:
Notation 1.4.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $V^{\prime}$ be a $\mathbb{D}$-linear subspace of $V$. Given a subset $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, we denote by $\mathcal{S}^{V^{\prime}}:=\mathcal{S}\cap\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$ the subset consisting of the elements of $\mathcal{S}$ that map into $V^{\prime}$.
Definition 1.5.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is:
  • intransitive whenever $\forall x\in U,\;\mathcal{S}x\neq V$ ;
  • deeply intransitive whenever it is intransitive and, for every nonzero linear subspace $V^{\prime}$ of $V$, the set $\mathcal{S}^{V^{\prime}}$ is intransitive as an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$.
In any case, the transitive rank $\operatorname{\operatorname{trk}}(\mathcal{S})$ of $\mathcal{S}$ is defined as the greatest possible dimension $\dim_{\mathbb{F}}(\mathcal{S}x)$ when $x$ ranges over $U$.
Note that $\mathcal{S}$ is intransitive only if $V\neq\{0\}$, and in the definition of deep intransitivity we can replace the condition that $\mathcal{S}$ be intransitive with the one that $V\neq\{0\}$.
It is obvious that if $\mathcal{S}$ is deeply intransitive then so is $\mathcal{S}^{V^{\prime}}$ for every nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$. Constructing large deeply intransitive spaces can be achieved by using nonzero bilinear forms, as we shall now explain.
Definition 1.6.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. An $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$ is called an alternator of $\mathcal{S}$ whenever
$\forall x\in U,\;\forall f\in\mathcal{S},\;b(x,f(x))=0.$
The set of all such forms is denoted by $\operatorname{\operatorname{Alt}}(\mathcal{S})$ (it is obviously a linear subspace of the set of all $\mathbb{F}$-bilinear forms on $U\times V$).
For an $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$, recall that the left radical of $b$ is the $\mathbb{F}$-linear subspace
$\operatorname{\operatorname{Lrad}}(b):=\{x\in U:\;b(x,-)=0\}\subseteq U,$
and the right radical of $b$ is the $\mathbb{F}$-linear subspace
$\operatorname{\operatorname{Rrad}}(b):=\{y\in V:\;b(-,y)=0\}\subseteq V.$
We say that $b$ is right-nondegenerate whenever $\operatorname{\operatorname{Rrad}}(b)=\{0\}$. Now, given $b$, we define $\mathcal{A}_{b}$ as the set of all $\mathbb{F}$-linear mappings $f:U\rightarrow V$ such that $\forall x\in U,\;b(x,f(x))=0$, and
$\mathcal{A}_{b,\mathbb{D}}:=\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)\cap\mathcal{A}_{b}$
as the subset of all such mappings that are $\mathbb{D}$-linear.
We now make two key observations, which we gather in the next proposition.
Proposition 1.18.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
  1. (a)
    If $\mathcal{S}$ has a nonzero alternator then $\mathcal{S}$ is intransitive.
  2. (b)
    If $\mathcal{S}$ has a right-nondegenerate alternator and $V\neq\{0\}$ then $\mathcal{S}$ is deeply intransitive.
Proof.
Assume that $\mathcal{S}$ has a nonzero alternator $b$. It follows by polarizing that $\forall(x,x^{\prime})\in U^{2},\;\forall f\in\mathcal{S},\;b(x,f(x^{\prime}))+b(x^{\prime},f(x))=0$, from which we observe that all the elements of $\mathcal{S}$ map $\operatorname{\operatorname{Lrad}}(b)$ into $\operatorname{\operatorname{Rrad}}(b)$. The latter is not the full space $V$ because $b$ is nonzero. Now, let $x\in U$. If $x\in\operatorname{\operatorname{Lrad}}(b)$ then $\mathcal{S}x\subseteq\operatorname{\operatorname{Rrad}}(b)\subset V$. Otherwise $\mathcal{S}x$ is included in $\operatorname{\operatorname{Ker}}b(x,-)$, an $\mathbb{F}$-linear hyperplane of $V$. Hence $\mathcal{S}$ is intransitive.
Assume now that $\mathcal{S}$ has a right-nondegenerate alternator $b$ and $V\neq\{0\}$. Let $V^{\prime}$ be a nonzero $\mathbb{D}$-linear subspace of $V$. The restriction $b^{\prime}$ of $b$ to $U\times V^{\prime}$ is nonzero because $b$ is right-nondegenerate, and obviously $b^{\prime}$ is an alternator of $\mathcal{S}^{V^{\prime}}$. Hence $\mathcal{S}^{V^{\prime}}$ is intransitive, by point (a). ∎
Hence, whenever we take a right-nondegenerate bilinear form $b:U\times V\rightarrow\mathbb{F}$ and $V\neq\{0\}$, every $\mathbb{F}$-linear subspace of $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive. Atkinson’s theorem can be essentially summed up in saying that all large enough operator spaces are of this kind. Let us now be more precise:
Theorem 1.19 (Atkinson’s theorem for deeply intransitive operator spaces).
Let $U$ and $V$ be finite-dimensional $\mathbb{F}$-vector spaces and $\mathcal{S}$ be a deeply intransitive linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$. Set $n:=\dim V$ and assume that $|\mathbb{F}|\geq n$. Then:
  1. (a)
    $\dim\mathcal{S}\leq\frac{n(n-1)}{2}$, and even $\dim\mathcal{S}\leq\frac{(n-1)(n-2)}{2}$ if $\operatorname{\operatorname{trk}}(\mathcal{S})<n-1$.
  2. (b)
    If $\dim\mathcal{S}\geq\frac{n(n-1)}{2}-(n-3)$ then $\mathcal{S}$ has a right-nondegenerate alternator $b$, and its alternator space $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$.
As an extension of point (b), we note, with the same polarization argument as in the proof of Proposition 1.18, that all the operators in $\mathcal{S}$ then vanish on the left radical $\operatorname{\operatorname{Lrad}}(b)$, and hence $\mathcal{S}$ can naturally be identified with a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U/\operatorname{\operatorname{Lrad}}(b),V)$, for which the form induced by $b$ on $(U/\operatorname{\operatorname{Lrad}}(b))\times V$ is a nondegenerate alternator. This fundamentally reduces the situation to the one where $\dim U=\dim V$, in which case we could go further and simply consider spaces of endomorphisms of $V$.
We will now make several comments on Atkinson’s theorem. First of all, we must state that Atkinson’s formulation of the theorem has several additional assumptions that turn out to be useless. Atkinson frames his result in the language of matrix spaces with bounded rank, and makes an assumption of primitivity. Primitivity consists of as many as four conditions that roughly mean that the rank constraint cannot be easily localized. In contrast, here we replace all four assumptions in the definition of primitivity by a single assumption, which is deep intransitivity (plain intransitivity is of course a far too weak hypothesis, allowing, e.g., $\mathcal{S}$ to be the space of all operators from $U$ to a given linear hyperplane of $V$). The deep intransitivity assumption is a dual counterpart of what Atkinson and Lloyd [1] called the row condition for matrix spaces with bounded rank. The row condition for such a matrix space $\mathcal{M}\subseteq\operatorname{\operatorname{M}}_{n,p}(\mathbb{F})$ means that, whenever we have invertible matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ and $Q\in\operatorname{\operatorname{GL}}_{p}(\mathbb{F})$, and integers $r\in\mathopen{[\![}1,n\mathclose{]\!]}$ and $s\in\mathopen{[\![}1,p\mathclose{]\!]}$ such that every matrix $M$ of $\mathcal{M}$ looks like
$M=P\begin{bmatrix}[?]_{r\times s}&K(M)\\ [?]_{(n-r)\times s}&[0]_{(n-r)\times(p-s)}\end{bmatrix}Q,$
all the matrices in the extracted matrix space $K(\mathcal{M})$ have their rank less than $r$ (the number of rows).
Now, if we take an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, consider its dual operator space $\widehat{\mathcal{S}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$, and represent the latter by a matrix space $\mathcal{M}$ in bases of $\mathcal{S}$ and $V$, the row condition on $\mathcal{M}$ is equivalent to the fact that whenever we take a nonzero linear subspace $\mathcal{S}^{\prime}$ of $\mathcal{S}$ and we know that all the operators in $\mathcal{S}^{\prime}$ map into some nonzero $\mathbb{F}$-linear subspace $V^{\prime}$ of $V$, then the rank of $s^{\prime}\in\mathcal{S}^{\prime}\mapsto s^{\prime}(x)\in V^{\prime}$ is less than the dimension of $V^{\prime}$ for all $x\in U$, which is exactly saying that $\mathcal{S}^{\prime}$ is intransitive as a linear subspace of mappings from $U$ to $V^{\prime}$. Hence, $\mathcal{M}$ satisfies the row condition if and only if $\mathcal{S}$ is deeply intransitive.
Among the four conditions that define primitivity, two are worth mentioning:
Definition 1.7.
An operator space $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ between $\mathbb{F}$-vector spaces $U$ and $V$ is called:
  • source-reduced when $\underset{f\in\mathcal{S}}{\bigcap}\operatorname{\operatorname{Ker}}f=\{0\}$, i.e., no nonzero vector is annihilated by all the operators in $\mathcal{S}$;
  • target-reduced when $\underset{f\in\mathcal{S}}{\sum}\operatorname{\operatorname{Im}}f=V$, i.e., no proper linear subspace of $V$ includes the range of all the operators in $\mathcal{S}$.
Finally, $\mathcal{S}$ is called reduced when it is both source-reduced and target-reduced.
Atkinson’s theorem is frequently formulated by adding the provision that the operator space $\mathcal{S}$ under consideration is reduced. However, this assumption is not necessary for the formulation in terms of intransitive spaces. Indeed, on the one hand the dual operator space $\widehat{\mathcal{S}}$ of an operator space $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}(U,V)$ is automatically source-reduced (as only the zero operator can be annihilated by all the evaluation mappings). On the other hand, if $\mathcal{S}$ is not target-reduced we can take the sum $V^{\prime}$ of all the ranges of the operators in $\mathcal{S}$, then observe that $\widehat{\mathcal{S}}$ can be viewed as a reduced subspace of $\operatorname{\operatorname{Hom}}(\mathcal{S},V^{\prime})$; applying point (a) in Atkinson’s theorem then readily shows that $\dim\mathcal{S}\leq\frac{\dim V^{\prime}(\dim V^{\prime}-1)}{2}\cdot$
Atkinson’s theorem deals with vector spaces over $\mathbb{F}$, but here we need to extend the situation to $\mathbb{F}$-linear subspaces of $\mathbb{D}$-linear mappings. In that context, it makes no sense to try to extend Atkinson’s formulation, as by representing the operators of $\widehat{\mathcal{S}}$ by matrices with entries in $\mathbb{F}$, we forget that the operators in $\mathcal{S}$ are $\mathbb{D}$-linear, and not only $\mathbb{F}$-linear. The correct way to generalize the results seems to be as a theorem on intransitive operator spaces. We now state the generalization:
Theorem 1.20 (Main theorem on deeply intransitive operator spaces).
Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces and $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Set $n:=\dim_{\mathbb{D}}V$ and assume that $|\mathbb{F}|\geq nd$. Then:
  1. (a)
    $\dim\mathcal{S}\leq\alpha(n,d)$.
  2. (b)
    $\dim\mathcal{S}\leq\alpha(n,d)-n+\max(2-d,0)$ if $\operatorname{\operatorname{trk}}(\mathcal{S})<nd-1$.
  3. (c)
    If $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ then $\mathcal{S}$ has a right-nondegenerate alternator $b$, and its alternator space is $1$-dimensional.
Remark 1.1.
Due to point (a), the inequality $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ of point (c) requires that $n\geq 2$.
The next theorem further elucidates the structure of the alternators:
Theorem 1.21.
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Set $n:=\dim_{\mathbb{D}}V$, assume that $|\mathbb{F}|\geq nd$, that $\mathcal{S}$ has a right-nondegenerate alternator $b$ and that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
Then $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and for any associated pair $(\sigma,e)$ the form $b$ equals $(x,y)\mapsto e(B(x,y))$ for a unique right-nondegenerate $\sigma$-sesquilinear form $B$ on $U\times V$, and of course $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$.
Here is a converse statement:
Proposition 1.22.
Assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, let $(\sigma,e)$ be an associated pair, and let $B:U\times V\rightarrow\mathbb{D}$ be a right-nondegenerate $\sigma$-sesquilinear form. Set $b:(x,y)\in U\times V\mapsto e(B(x,y))$. Then $\mathcal{A}_{b,\mathbb{D}}$ has dimension $\alpha(n,d)$ (over $\mathbb{F}$), and it is deeply intransitive.
Having these results, it will be fairly easy to derive the ones on optimal irreducible trivial spectrum subspaces, by adapting the method of [19]. The main difficulty is that although a trivial spectrum subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (with $V$ nonzero) is obviously intransitive (since $\mathcal{S}x$ can never contain $x$ when $x$ is a nonzero vector of $V$), it is however not deeply intransitive in general.
A classical way of obtaining the deep intransitivity is through the following property, which is essentially the analogue of Atkinson and Lloyd’s condition (C) in the definition of a primitive space:
Definition 1.8.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is weakly primitively intransitive whenever there is no $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi\mathcal{S})\leq\operatorname{\operatorname{tr}}(\mathcal{S})-d$.
In fact, for most applications we will prefer the following stronger property, which is easier to work with:
Definition 1.9.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is primitively intransitive whenever it is intransitive and there is no nonzero linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ is intransitive.
The following implication is easy to see:
Proposition 1.23.
Every primitively intransitive operator space is weakly primitively intransitive.
Proof.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces, $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that there exists a $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi\mathcal{S})\leq\operatorname{\operatorname{tr}}(\mathcal{S})-d$. If $\mathcal{S}$ is intransitive, then we deduce that $\operatorname{\operatorname{trk}}(\pi\mathcal{S})<\dim_{\mathbb{F}}(V/V^{\prime})$ and hence $\pi\mathcal{S}$ is intransitive as well: this shows that $\mathcal{S}$ is not primitively intransitive. ∎
A key insight of Atkinson and Lloyd can be rephrased in saying that, provided that the field $\mathbb{F}$ has enough elements, every weakly primitively intransitive space is deeply intransitive. The primitively intransitivity condition has the shortcoming that it is generally not inherited by $\mathcal{S}^{V^{\prime}}$ when $V^{\prime}$ is a linear subspace of $V$, in contrast with the deep intransitivity condition. However, it is the most convenient condition when applying the previous theorem. Indeed, we shall prove – under the usual cardinality conditions on $\mathbb{F}$ – that every irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is primitively intransitive, whereas we are not aware of any direct proof that it must be deeply intransitive.
1.6 Structure of the article
As hinted at in the previous paragraph, our focus is now shifting from the trivial spectrum subspaces, which are the prime motivation for this article, to the deeply intransitive operator spaces, which can be viewed as a more general topic. We expect that studying the later will yield many applications in the future, just like Atkinson’s theorem has found many unexpected applications in the recent past (see e.g. [19, 25]). In an early version of the present article, we wrote a proof of the result on trivial spectrum subspaces alone. Surprising as it may seem, we found later that the result on deeply intransitive operator spaces is easier to prove. The icing on the cake is that it also holds more potential for future applications.
Let us now discuss the structure of the remainder of the article. We will start with a small general section that regroups the basic facts we need on division algebras, including the classification of multiplicative quadratic forms on such algebras. Then we will immediately give a proof of the dimension theorem for trivial spectrum subspaces (Theorem 1.6), along with two key technical lemmas. We could of course have waited for the proofs of the theorems on intransitive operator spaces, but the proof is so short that the reader will certainly enjoy reading it immediately.
The next section (Section 4) introduces the key technical tools to study the intransitive operator spaces: there we make no claims of originality, as we will use the generic matrix machinery that allowed Atkinson to prove his theorem, a machinery which has already demonstrated its remarkable efficiency for such problems. The main results on intransitive operator spaces are obtained in Section 5: there, the technical details for a general division ring $\mathbb{D}$ are substantially more difficult than in Atkinson’s article.
In Section 6 we derive all the main results on optimal trivial spectrum subspaces from the ones on intransitive operator spaces.
The remaining sections are devoted to applications of trivial spectrum subspaces. In Section 7, we obtain the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ with rank at least $r$ (Section 7.1), and in Section 7.2 we characterize the spaces with the greatest possible dimension when $n=p=r$ (i.e., affine subspaces of nonsingular matrices). The characterization for a general $r$ is very difficult (see [16] for the case $d=1$, which takes no less than twenty-three pages of proof), so we believe the reader will be relieved to learn that this issue is postponed to a subsequent article.
The last section is devoted to the applications of $\mathbb{F}$-trivial spectrum subspaces to subspaces of $\mathbb{F}$-diagonalisable matrices. There, we successfully adapt the method featured in [23], and we also give a brand new result for real vector spaces of complex diagonalisable matrices. This new application involves the main theorem on deeply intransitive operator spaces in a case where the space under consideration has its dimension slightly lower than the critical one. We finish with an adaptation to spaces of semi-simple endomorphisms and with considerations on spaces of endomorphisms in which only the zero endomorphism is nilpotent.
2 On $\mathbb{D}$-vector spaces seen as $\mathbb{F}$-vector spaces
2.1 Traces
Throughout, we will need to consider $\mathbb{F}$-linear forms on $\mathbb{D}$. There is a canonical choice of such a linear form: the trace of $\mathbb{D}$ over $\mathbb{F}$, denote by $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ (with capital T), maps every $a\in\mathbb{D}$ to the trace of the $\mathbb{F}$-endomorphism $x\mapsto ax$ of $\mathbb{D}$. Consequently, it is an $\mathbb{F}$-linear form that enjoys the property that $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ab)=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ba)$ for all $a,b$ in $\mathbb{D}$. Let $a\in\mathbb{D}$. Denoting by $\pi=t^{n}-\alpha t^{n-1}+\cdots$ the minimal polynomial of $a$ in the $\mathbb{F}$-algebra $\mathbb{D}$, it is easily seen that $n$ divides $d$ and that $x\mapsto ax$ is the direct sum of $\frac{d}{n}$ cyclic endomorphisms of $\mathbb{F}$-vector spaces with minimal polynomial $\pi$, which yields $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)=\frac{d}{n}\,\alpha.$ Applying the same line of reasoning to $x\mapsto xa$ yields that $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)$ is also the trace of the $\mathbb{F}$-linear endomorphism $x\mapsto xa$. Moreover $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(\varphi(a))=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)$ for every automorphism $\varphi$ of the $\mathbb{F}$-algebra $\mathbb{D}$.
Alas, in some cases $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ might be useless. This is the case for example if $\mathbb{D}$ is a totally radicial extension of $\mathbb{F}$: in that case $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ is identically zero. The same holds if $\mathbb{D}$ is a quaternion algebra and $\operatorname{\operatorname{char}}(\mathbb{F})=2$: in that case the minimal polynomial of every element of $\mathbb{D}$ has degree $1$ or $2$, and hence the trace of such an element is either $2$ times or $4$ times the trace of this polynomial, so it is zero in any case. If however $\operatorname{\operatorname{char}}(\mathbb{F})=0$ then $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}$ is nonzero and we can safely use $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ as a nonzero linear form on $\mathbb{D}$.
2.2 The socle of an $\mathbb{F}$-linear subspace
Definition 2.1.
Let $U$ be a $\mathbb{D}$-vector space, and $U_{0}$ be an $\mathbb{F}$-linear subspace of it. Then there is a greatest $\mathbb{D}$-linear subspace of $U$ that is included in $U_{0}$: we call it the socle of $U_{0}$.
It suffices to take the sum of all $\mathbb{D}$-linear subspaces of $U$ that are included in $U$: it is obviously included in $U$, and it is a $\mathbb{D}$-linear subspace.
2.3 $\mathbb{D}$-linear forms vs $\mathbb{F}$-linear forms
Now, we fix a nonzero linear form $e$ on the $\mathbb{F}$-vector space $\mathbb{D}$. Consider an arbitrary finite-dimensional $\mathbb{D}$-vector space $V$. We have a homomorphism of $\mathbb{F}$-vector spaces
$\Phi_{e}:\begin{cases}\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V,\mathbb{D})&\longrightarrow\operatorname{\operatorname{Hom}}_{\mathbb{F}}(V,\mathbb{F})\\ f&\longmapsto e\circ f,\end{cases}$
and we will see that it is actually an isomorphism. First, let us see that $\Phi_{e}$ is injective: for all $f\in\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V,\mathbb{D})\smallsetminus\{0\}$, the mapping $f$ is surjective and hence cannot map into $\operatorname{\operatorname{Ker}}e$, therefore $e\circ f\neq 0$. We conclude that $\Phi_{e}$ is an isomorphism because both its source and target spaces have dimension $d\dim_{\mathbb{D}}V$ over $\mathbb{F}$.
Here is a first consequence which will be frequently used:
Lemma 2.1.
Let $V$ be a finite-dimensional $\mathbb{D}$-vector space, and $H$ be an $\mathbb{F}$-linear hyperplane of it. Then the socle of $H$ is a $\mathbb{D}$-linear hyperplane of $V$.
Proof.
Choose a nonzero $\mathbb{F}$-linear form $\varphi$ with kernel $H$. Then $\varphi=e\circ f$ for a unique nonzero $\mathbb{D}$-linear form $f$ on $V$. Putting $G:=\operatorname{\operatorname{Ker}}f$, we see that $G$ is a $\mathbb{D}$-linear hyperplane and $G\subseteq H$. The socle $S$ of $H$ is a strict subspace of $V$ (since $H\neq V$), so $\dim_{\mathbb{D}}S\leq\dim_{\mathbb{D}}V-1$, and since $G\subseteq S$ we deduce that $G=S$. ∎
A second consequence lies in the representation of $\mathbb{F}$-bilinear forms, to be used in the study of alternators. Let $b:U\times V\rightarrow\mathbb{F}$ be an $\mathbb{F}$-bilinear form, where $U$ and $V$ are finite-dimensional $\mathbb{D}$-vector spaces. We can then consider the associated $\mathbb{F}$-linear mapping
$x\in U\mapsto b(x,-)\in\operatorname{\operatorname{Hom}}_{\mathbb{F}}(V,\mathbb{F}),$
and then compose it with the inverse of $\Phi_{e}$ to obtain an $\mathbb{F}$-linear mapping
$\widetilde{b}:U\longrightarrow\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V,\mathbb{D}),$
yielding an $\mathbb{F}$-bilinear and right-$\mathbb{D}$-linear mapping
$b^{\mathbb{D},e}:(x,y)\in U\times V\mapsto\widetilde{b}(x)[y],$
so that
$\forall(x,y)\in U\times V,\;b(x,y)=e(b^{\mathbb{D},e}(x,y)).$
Moreover $b^{\mathbb{D},e}$ is uniquely determined by $b$, and it depends $\mathbb{F}$-linearly on $b$.
Here is a simple result in the special case where $b^{\mathbb{D},e}$ is sesquilinear for some involution $\sigma$ of $\mathbb{D}$ over $\mathbb{F}$.
Lemma 2.2.
Assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and consider an associated pair $(\sigma,e)$. Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector space, $B:U\times V\rightarrow\mathbb{D}$ be a $\sigma$-sesquilinear form, and set $b:=e\circ B$. Then the left radical (respectively, the right radical) of $b$ coincides with the one of $B$, and in particular it is a $\mathbb{D}$-linear subspace of $U$ (respectively, of $V$).
Proof.
Let $y\in\operatorname{\operatorname{Rrad}}(b)$. Then
$\forall x\in U,\;\forall a\in\mathbb{D},\;e(aB(x,y))=e(B(x\sigma(a),y))=b(x\sigma(a),y)=0$
and hence B(,y)=0B(-,y)=0. Hence $\operatorname{\operatorname{Rrad}}(b)$ is included in the right radical of $B$; the converse inclusion is obvious. Likewise one proves that the left radical of $B$ is the left radical of $b$. ∎
2.4 Multiplicative quadratic forms on division algebras
Our study will require the following result. It can be extracted from the general theory of composition algebras, but since we are in the very specific context of a division ring $\mathbb{D}$, the proof can be considerably shortened. We include a proof below for non-specialists, but advise the reader to skip it at first reading.
Theorem 2.3.
Let $q$ be a nonisotropic quadratic form on the $\mathbb{F}$-vector space $\mathbb{D}$ such that $\forall(x,y)\in\mathbb{D}^{2},\;q(xy)=q(x)q(y)$. Then $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, and $q:x\mapsto x\sigma(x)$ for the standard involution $\sigma$ of $\mathbb{D}$ over $\mathbb{F}$.
Proof.
Throughout we consider the polar form $b:(x,y)\in\mathbb{D}^{2}\mapsto q(x+y)-q(x)-q(y)$ of $q$, and note that the multiplicativity of $q$ yields
$\forall(a,x,y)\in\mathbb{D}^{3},\;b(ax,ay)=q(a)\,b(x,y)=b(xa,ya).$ (2)
Assume first that $q$ is degenerate. Then $\operatorname{\operatorname{char}}(\mathbb{F})=2$. We start by proving that $q$ is totally degenerate (i.e., additive). Let us consider the radical $R$ of $b$. We have $R\neq\{0\}$ by assumption. Let us choose $x\in R\smallsetminus\{0\}$. For all $a\in\mathbb{D}\smallsetminus\{0\}$ we have
$\forall y\in\mathbb{D},\;b(ax,ay)=q(a)\,b(x,y)=0,$
and since $\{ay\mid y\in\mathbb{D}\}=\mathbb{D}$ we deduce that $ax\in R$. Hence $R=\mathbb{D}$, meaning that $q$ is additive.
Next, for all $x,y$ in $\mathbb{D}$ we find
q(xyyx)=q(xy)q(yx)=q(x)q(y)q(y)q(x)=0,q(xy-yx)=q(xy)-q(yx)=q(x)q(y)-q(y)q(x)=0,
and hence xyyx=0xy-yx=0 because $q$ is nonisotropic. Hence $\mathbb{D}$ is commutative.
Let $x\in\mathbb{D}$. Then
$q(x^{2}-q(x))=q(x^{2})-q(q(x).1_{\mathbb{D}})=q(x)^{2}-q(x)^{2}q(1_{\mathbb{D}})=0,$
and hence $x^{2}=q(x)\in\mathbb{F}$ because $q$ is nonisotropic. Therefore $\mathbb{D}$ is a hyper-radicial extension of $\mathbb{F}$, and $q(x)=x\sigma(x)$ for all $x\in\mathbb{D}$, where $\sigma=\operatorname{\operatorname{id}}_{\mathbb{D}}$ is the sole involution of the $\mathbb{F}$-algebra $\mathbb{D}$.
Now, we deal with the case where $q$ is nondegenerate. We introduce the trace map $\operatorname{\operatorname{tr}}:x\in\mathbb{D}\mapsto b(1,x)\in\mathbb{F}$ and note that $\operatorname{\operatorname{tr}}(1)=2$. The key point consists in proving the quadratic identity
$\forall x\in\mathbb{D},\;x^{2}-\operatorname{\operatorname{tr}}(x)x+q(x)=0.$ (3)
To do so, we take arbitrary elements $y$ and $z$ in $\mathbb{D}$ and explore the identity $q(x+y)q(x+z)=q((x+y)(x+z))=q(x^{2}+yx+xz+yz)$. Thanks to $b$, we expand
$q(x^{2}+yx+xz+yz)=\underbrace{q(x^{2})}_{a_{4}}+\underbrace{b(x^{2},yx+xz)}_{a_{3}}+\underbrace{b(x^{2},yz)+q(yx+xz)}_{a_{2}}\\ +\underbrace{b(yx+xz,yz)}_{a_{1}}+\underbrace{q(yz)}_{a_{0}}$
and
$q(x+y)q(x+z)=\underbrace{q(x)^{2}}_{b_{4}}+\underbrace{q(x)(b(x,y)+b(x,z))}_{b_{3}}+\underbrace{q(x)(q(y)+q(z))+b(x,y)b(x,z)}_{b_{2}}\\ +\underbrace{q(z)b(x,y)+q(y)b(x,z)}_{b_{1}}+\underbrace{q(y)q(z)}_{b_{0}}.$
It is easily checked, by using either the multiplicativity of $q$ or (2), that $b_{i}=a_{i}$ for all $i\in\{0,1,3,4\}$, which leaves us with
$q(x)(q(y)+q(z))+b(x,y)\,b(x,z)=b(x^{2},yz)+q(yx+xz),$
and since q(yx+xz)=q(yx)+q(xz)+b(yx,xz)=q(y)q(x)+q(x)q(z)+b(yx,xz)q(yx+xz)=q(yx)+q(xz)+b(yx,xz)=q(y)q(x)+q(x)q(z)+b(yx,xz), this is further simplified as
$b(x,y)\,b(x,z)=b(x^{2},yz)+b(yx,xz).$
From there, we apply this identity to $y=1$ and an arbitrary $z$ to find $b(q(x),z)=q(x)\,b(1,z)=b(x,xz)=b(yx,xz)=b(x,1)\,b(x,z)-b(x^{2},z)=b(\operatorname{\operatorname{tr}}(x)x-x^{2},z)$, which yields $q(x)=\operatorname{\operatorname{tr}}(x)x-x^{2}$ because $q$ is nondegenerate. Next, we introduce the adjunction mapping
$x\in\mathbb{D}\mapsto x^{\star}:=\operatorname{\operatorname{tr}}(x)-x\in\mathbb{D},$
which is $\mathbb{F}$-linear and clearly satisfies $\operatorname{\operatorname{tr}}(x^{\star})=\operatorname{\operatorname{tr}}(x)$ and $(x^{\star})^{\star}=x$ for all $x\in\mathbb{D}$. We observe that (3) can be reformulated into
$\forall x\in\mathbb{D},\;q(x)=xx^{\star}=x^{\star}x.$
In particular, for all $x,y$ in $\mathbb{D}^{\times}$, we have $(xy)(xy)^{\star}=q(xy)=q(x)q(y)=xq(y)x^{\star}=xyy^{\star}x^{\star}$ and hence $(xy)^{\star}=y^{\star}x^{\star}$. It follows that $(-)^{\star}$ is an involution of the $\mathbb{F}$-algebra $\mathbb{D}$.
Assume now that $d>1$. Since $\mathbb{D}$ is a division ring, we deduce from (3) that it has a $2$-dimensional subfield, and hence $d$ is even. Assume that $d=2$. Since $q$ is nondegenerate, the trace does not vanish, so by linearity it does not vanish on $\mathbb{D}\smallsetminus\mathbb{F}$, and we deduce from (3) that $\mathbb{D}$ is a separable extension of $\mathbb{F}$ (by choosing $x\in\mathbb{D}\smallsetminus\mathbb{F}$ such that $\operatorname{\operatorname{tr}}(x)\neq 0$). And clearly $(-)^{\star}$ is not the identity, so it is the standard involution of $\mathbb{D}$ over $\mathbb{F}$.
Assume finally that $d>2$, and set p:=d2p:=d-2. Since $q$ is nondegenerate, the radical of its restriction to $H:=\operatorname{\operatorname{Ker}}(\operatorname{\operatorname{tr}})$ has dimension at most $1$, so we can find in $H$ an $\mathbb{F}$-linear subspace $V$ of dimension $p$ on which $q$ remains nondegenerate. Then the identity $\forall x\in V,\;x^{2}=-q(x)$ yields a (unique) homomorphism of $\mathbb{F}$-algebras $\Phi:\mathcal{C}(V,-q_{|V})\rightarrow\mathbb{D}$ such that $\forall x\in V,\;\Phi(x)=x$, where $\mathcal{C}(V,-q_{|V})$ stands for the Clifford algebra of the regular quadratic space $(V,-q_{|V})$. Classically, $\mathcal{C}(V,-q_{|V})$ is a simple algebra with dimension $2^{p}$ over $\mathbb{F}$, because $p$ is even. Hence $\Phi$ is injective, which leads to $2^{p}\leq d$ and hence to $d=4$. Then $\Phi$ is an isomorphism and $\mathbb{F}\cap V=\{0\}$, and we have recognized from $\Phi$ that $\mathbb{D}$ is a quaternion algebra over $\mathbb{F}$ and that the quaternionic conjugation $x\mapsto\overline{x}$ inherited from $\Phi$ satisfies $x\overline{x}=q(x)$ for all $x\in V$, which yields that it coincides with $(-)^{\star}$ on $V$. Hence $x\mapsto x^{\star}$ and $x\mapsto\overline{x}$ coincide on the $\mathbb{F}$-subalgebra generated by $\mathbb{F}\oplus V$, which must be $\mathbb{D}$ because it is a division algebra over $\mathbb{F}$ with dimension $3$ or $4$ that contains a subfield with dimension $2$ over $\mathbb{F}$. Therefore $x\mapsto x^{\star}$ is the quaternionic conjugation of $\mathbb{D}$, and our proof is completed. ∎
3 The greatest possible dimension for a trivial spectrum subspace
This short section is devoted to the proof of Theorem 1.6. We have seen right at the beginning of the article that for every integer $n\geq 1$ there exists an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $\alpha(n,d)$. Hence, in order to prove Theorem 1.6 it suffices to prove the following result:
Theorem 3.1.
Let $\mathbb{D}$ be a division ring with finite degree $d$ over a central subfield $\mathbb{F}$, and $V$ be a vector space over $\mathbb{D}$ with finite dimension $n$. Assume that $|\mathbb{F}|\geq dn$. Let $\mathcal{S}$ be an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then
$\dim\mathcal{S}\leq\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$.
The main key to this result is the use of the dual operator space of $\mathcal{S}$ (see Section 1.5) and what we call the Tangent Space Lemma. The latter is recalled in the next section.
3.1 The tangent space lemma, and the Flanders-Atkinson lemma
One of the main keys to Theorem 3.1 is the following lemma (see [2] and [20] for various proofs). The special case $k=0$ in the second set of identities is due to Flanders [6], and the general case is due to Atkinson.
Lemma 3.2 (Flanders-Atkinson lemma).
Let $n,p,r$ be integers with $0<r\leq\min(n,p)$. Assume that $|\mathbb{F}|>r$. Let $J_{r}:=\begin{bmatrix}I_{r}&0\\ 0&0\end{bmatrix}$ and $M=\begin{bmatrix}A&C\\ B&D\end{bmatrix}$ belong to $\operatorname{\operatorname{M}}_{n,p}(\mathbb{F})$, with $A\in\operatorname{\operatorname{M}}_{r}(\mathbb{F})$ and so on. Assume that $\operatorname{\operatorname{rk}}(sJ_{r}+tM)\leq r$ for all $(s,t)\in\mathbb{F}^{2}$. Then
$D=0\quad\text{and}\quad\forall k\geq 0,\;BA^{k}C=0.$
The following corollary is simply obtained by exploiting the identity $D=0$ in the Flanders-Atkinson lemma:
Lemma 3.3 (Tangent Space Lemma).
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{F}$, and $\mathcal{S}$ be a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$. In $\mathcal{S}$, take an element $u_{0}$ of maximal rank $r$, and assume that $|\mathbb{F}|>r$. Then every element of $\mathcal{S}$ maps $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
The reference to tangent spaces comes from the observation that if $\mathbb{F}$ is the field of real numbers then the manifold of all rank $r$ operators of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ has its tangent space at $u_{0}$ equal to the space of all operators in $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ that map $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
3.2 Proof of Theorem 3.1
We prove the result by induction on $n$. The case $n=0$ is trivial, so we assume that $n\geq 1$. Let us consider the dual operator space $\widehat{\mathcal{S}}$, consisting of the $\mathbb{F}$-linear mappings
$\widehat{x}:u\in\mathcal{S}\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
Note that $\widehat{\mathcal{S}}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$. Let $x\in V\smallsetminus\{0\}$. Since $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum, there is no $u\in\mathcal{S}$ such that $u(x)=x$; hence $\operatorname{\operatorname{rk}}(\widehat{x})\leq dn-1$.
Let us now take $x\in V\smallsetminus\{0\}$ such that $\widehat{x}$ has the greatest possible rank $r\leq dn-1$ in $\widehat{\mathcal{S}}$. Since $|\mathbb{F}|>r$, the Tangent Space Lemma (Lemma 3.3) yields
$\forall u\in\operatorname{\operatorname{Ker}}\widehat{x},\;\forall y\in V,\quad\widehat{y}(u)\in\operatorname{\operatorname{Im}}\widehat{x}$,
or in other words
$\forall u\in\mathcal{S},\;u(x)=0\Rightarrow\operatorname{\operatorname{Im}}u\subseteq\mathcal{S}x.$
Setting
$\mathcal{S}^{\prime}:=\operatorname{\operatorname{Ker}}\widehat{x}=\{u\in\mathcal{S}:\;u(x)=0\},$
we can apply the rank theorem to get
$\dim(\mathcal{S})=\dim(\mathcal{S}^{\prime})+\dim_{\mathbb{F}}(\mathcal{S}x)\leq\dim(\mathcal{S}^{\prime})+(dn-1).$ (4)
Now, the operators in $\mathcal{S}^{\prime}$ are actually $\mathbb{D}$-linear so their ranges are $\mathbb{D}$-linear subspaces of $V$, all included in $\mathcal{S}x$ and hence all included in the socle (see Definition 2.1) of $\mathcal{S}x$. This socle does not contain $x$ (because $x\not\in\mathcal{S}x$), and hence we can embed it into a $\mathbb{D}$-hyperplane $H$ of $V$ such that $x\mathbb{D}\oplus H=V$, and we note that $\operatorname{\operatorname{Im}}u\subseteq H$ for every $u\in\mathcal{S}^{\prime}$.
Hence, all the elements of $\mathcal{S}^{\prime}$ leave $H$ invariant, inducing $\mathbb{D}$-linear endomorphisms of $H$, all of them with $\mathbb{F}$-trivial spectrum. Since $x\mathbb{D}\oplus H=V$ and the elements of $\mathcal{S}^{\prime}$ vanish on $x\mathbb{D}$, the $\mathbb{F}$-linear mapping
$u\in\mathcal{S}^{\prime}\mapsto u_{H}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$
is injective. Its range is an $\mathbb{F}$-trivial spectrum subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$. Since $|\mathbb{F}|\geq(n-1)d$, we obtain by induction that
$\dim(\mathcal{S}^{\prime})=\dim(\mathcal{T})\leq d\dbinom{n-1}{2}+(n-1)(d-1).$
By combining this with (4), we conclude that
$\dim(\mathcal{S})\leq(n-1)(d-1)+d\dbinom{n-1}{2}+((d-1)+d(n-1))=n(d-1)+d\dbinom{n}{2}$,
which completes the proof.
3.3 Adaptation to deeply intransitive operator spaces
We shall immediately prove the first part of Theorem 1.20 by using essentially the same technique as in the previous proof. We restate the result below:
Proposition 3.4.
Let $U$ and $V$ be $\mathbb{D}$-vector spaces, with $n=\dim_{\mathbb{D}}V\geq 1$. Assume that $|\mathbb{F}|\geq nd$. Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Then $\dim\mathcal{S}\leq\alpha(n,d)$.
Proof.
We denote by $r$ the transitive rank of $\mathcal{S}$. If $r=0$ then $\mathcal{S}=\{0\}$ and we are done. Assume that $r>0$, and choose $x\in U$ such that $\dim_{\mathbb{F}}\mathcal{S}x=r$. Hence $\widehat{x}$ has the greatest possible rank in $\widehat{\mathcal{S}}$, and as $|\mathbb{F}|>r$ we can apply the tangent space lemma. This yields that all the operators in $\mathcal{S}^{\prime}:=\{f\in\mathcal{S}:f(x)=0\}$ map into $\mathcal{S}x$. As in the previous proof, we can find a $\mathbb{D}$-linear hyperplane $H$ of $V$ into which all the elements of $\mathcal{S}^{\prime}$ map. The rank theorem yields
$\dim\mathcal{S}=r+\dim\mathcal{S}^{\prime}.$
Then there are two cases:
  • If $n=1$ then $\mathcal{S}^{\prime}=\{0\}$ and we conclude that $\dim\mathcal{S}\leq d-1$, as claimed;
  • If $n>1$ then $\mathcal{S}^{\prime}$ is a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, by induction we find that
    $\dim\mathcal{S}\leq(nd-1)+\alpha(n-1,d)=\alpha(n,d),$
    and the proof is complete.
Point (b) of Theorem 1.20 is much more difficult when $d>1$, and we wait until Section 5 to prove it. The next point is an easy consequence of Proposition 3.4.
Corollary 3.5.
Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(3-d,1)$ and $n>0$. Then $\mathcal{S}$ is target-reduced.
Remark 3.1.
In the proof and in future use of the result, it is crucial to note that the fact that $\mathcal{S}$ is not target-reduced is equivalent to the existence of a $\mathbb{D}$-linear hyperplane $H$ of $V$ (and not simply of an $\mathbb{F}$-linear hyperplane) that includes the range of every element of $\mathcal{S}$. The converse implication is obvious. For the direct one, assume that we have an $\mathbb{F}$-linear hyperplane $V^{\prime}$ of $V$ that includes the range of every operator in $\mathcal{S}$. Then all the operators in $\mathcal{S}$ map into the socle of $V^{\prime}$, which is a $\mathbb{D}$-linear hyperplane of $V$.
Proof of Corollary 3.5.
Assume that $\mathcal{S}$ is not target-reduced. Then there exists a $\mathbb{D}$-linear hyperplane $H$ of $V$ such that $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
Because $\mathcal{S}$ is deeply intransitive we can view it as a deeply intransitive subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, and then Proposition 3.4 yields $\dim\mathcal{S}\leq\alpha(n-1,d)=\alpha(n,d)-(nd-1)$. Now (nd1)n=n(d1)1(nd-1)-n=n(d-1)-1, and n(d1)1>2n(d-1)-1>-2 if $d=1$, whereas n(d1)1>1n(d-1)-1>-1 whenever $d\geq 2$ and $n>0$. Hence, in any case we contradict the asumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(3-d,1)$. ∎
4 A review of the method of generic matrices
Here we review the technique of generic matrices that was introduced by Atkinson and Lloyd [1, 2] (see also [21]).
4.1 Generic matrices
Let $E$ be an $\mathbb{F}$-vector space. We consider its dual vector space $E^{\star}:=\operatorname{\operatorname{Hom}}_{\mathbb{F}}(E,\mathbb{F})$ and the (graded) symmetric algebra $R=S(E^{\star})$ over $\mathbb{F}$ (which is the direct sum of the symmetric tensor spaces $S^{n}(E^{\star})$ for $n\geq 0$) and its field of fractions $\mathbb{L}$. We naturally view $\mathbb{F}$ as a subfield of $\mathbb{L}$.
The $1$-homogeneous elements of $R$ constitute the space $S^{1}(E^{\star})=E^{\star}$ of symmetric $1$-tensors. We have a canonical ring homomorphism $S(E^{\star})\rightarrow\mathrm{Pol}(E,\mathbb{F})$ to the ring of polynomial functions on $E$ (it is an isomorphism if $\mathbb{F}$ is infinite). Hence for every $\mathbf{p}\in S(E^{\star})$ and every $z\in E$ we can consider the specialization $\mathbf{p}(z)\in\mathbb{F}$. More generally, given a matrix $\mathbf{M}\in\operatorname{\operatorname{M}}_{m,p}(R)$ and a vector $z\in E$, we denote by $\mathbf{M}(z)$ the matrix obtained by specializing all the entries of $\mathbf{M}$ at $z$.
Definition 4.1.
Let $m,p$ be positive integers, and let $\mathcal{M}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{m,p}(\mathbb{F})$. A generic matrix [^5] Beware that in [20], such a matrix is called a semi-generic matrix. for $\mathcal{M}$ (with respect to $R$) consists of a matrix $\mathbf{M}$ whose entries are $1$-homogeneous elements of $R$ and such that
$\mathcal{M}=\{\mathbf{M}(z)\mid z\in E\}.$
Remark 4.1.
In [1, 2, 21], the authors only dealt with polynomial rings that are formally of the form $\mathbb{F}[\mathbf{t_{1}},\dots,\mathbf{t_{r}}]$. Our viewpoint here will turn out to be more convenient, although it is not fundamentally more general than the traditional one. Indeed, if we have a basis $(e_{1},\dots,e_{r})$ of the $\mathbb{F}$-vector space $E$, we have an isomorphism of $\mathbb{F}$-algebras from $\mathbb{F}[\mathbf{t_{1}},\dots,\mathbf{t_{r}}]$ to $S(E^{\star})$ that maps each indeterminate $\mathbf{t_{i}}$ to the $1$-symmetric tensor $e_{i}^{\star}$, the $i$-th vector of the dual basis of $(e_{1},\dots,e_{r})$.
A critical property of generic matrices, when seen as matrices with entries in the fraction field $\mathbb{L}$, is that their rank coincides with the maximal rank of $\mathcal{M}$ when the field $\mathbb{F}$ is large enough.
Notation 4.2.
For a nonempty subset $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{m,p}(\mathbb{F})$, we set
$\operatorname{\operatorname{maxrk}}(\mathcal{M}):=\max\{\operatorname{\operatorname{rk}}M\mid M\in\operatorname{\operatorname{M}}_{m,p}(\mathbb{F})\}.$
The following classical lemma will be used later, and it also has a direct application to generic matrices:
Lemma 4.1.
Let $\mathbb{F}^{\prime}$ be an arbitrary field extension of $\mathbb{F}$, and $\mathcal{M}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{m,p}(\mathbb{F})$. Assume that $|\mathbb{F}|>\operatorname{\operatorname{maxrk}}(\mathcal{M})$. Then
$\operatorname{\operatorname{maxrk}}(\operatorname{\operatorname{span}}_{\mathbb{F}^{\prime}}(\mathcal{M}))=\operatorname{\operatorname{maxrk}}(\mathcal{M}).$
By working with $\mathbb{F}^{\prime}=\mathbb{L}$, it is easy to derive the following corollary:
Corollary 4.2 (See lemma 2.1 in [20]).
Let $\mathcal{M}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{m,p}(\mathbb{F})$, and $\mathbf{M}$ be a generic matrix of it. Assume that $|\mathbb{F}|>\operatorname{\operatorname{maxrk}}(\mathcal{M})$. Then
$\operatorname{\operatorname{rk}}_{\mathbb{L}}(\mathbf{M})=\operatorname{\operatorname{maxrk}}(\mathcal{M}).$
In connection with this corollary, we remark that the generic matrix $\mathbf{M}$ belongs to $\operatorname{\operatorname{span}}_{\mathbb{L}}(\mathcal{M})$, a fact that we will reuse later. To see this, we take a basis $(e_{1},\dots,e_{p})$ of $E$, consider the dual basis $(e_{1}^{\star},\dots,e_{p}^{\star})$, and note that $\mathbf{M}=\sum_{k=1}^{p}e_{k}^{\star}\,\mathbf{M}(e_{k})$.
4.2 Notions of ranks for generic row matrices
Let us now discuss row matrices over the ring $R$.
Definition 4.3.
Let $\mathbf{L}\in\operatorname{\operatorname{M}}_{1,n}(R)$ be a row matrix whose entries are $1$-homogeneous elements of $R$. The spanning rank of $\mathbf{L}$, denoted by $\operatorname{\operatorname{sprk}}(\mathbf{L})$, is defined as the dimension (over $\mathbb{F}$) of the specialized linear subspace $\{\mathbf{L}(z)\mid z\in E\}$ of $\operatorname{\operatorname{M}}_{1,n}(\mathbb{F})$.
Of course, one should not confuse the spanning rank of $\mathbf{L}$ with its (traditional) rank, i.e., its rank as a matrix of $\operatorname{\operatorname{M}}_{1,n}(\mathbb{L})$, which equals 0 or $1$ (whereas the spanning rank is an integer between 0 and $n$).
The following lemma (lemma 5.2 in [21]) will be particularly important in our study:
Lemma 4.3 (Factorization Lemma).
Let $n\geq 1$ and $\delta\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $\operatorname{\operatorname{M}}_{1,n}(R)$ and assume that:
  1. (i)
    The entries of $\mathbf{X}$ are $1$-homogeneous and the ones of $\mathbf{Y}$ are $\delta$-homogeneous.
  2. (ii)
    The spanning rank of $\mathbf{X}$ is greater than $1$.
  3. (iii)
    The vectors $\mathbf{X}$ and $\mathbf{Y}$ are linearly dependent in the $\mathbb{L}$-vector space $\operatorname{\operatorname{M}}_{1,n}(\mathbb{L})$.
Then $\mathbf{Y}=\mathbf{p}\,\mathbf{X}$ for some $(\delta-1)$-homogeneous $\mathbf{p}\in R$.
4.3 Recognizing alternators in generic matrix identities
We finish this introduction to generic matrices by discussing left annihilators of generic matrices and their relationship with alternators.
Definition 4.4.
Let $\mathbf{M}\in\operatorname{\operatorname{M}}_{m,p}(R)$. A left-annihilator of $\mathbf{M}$ is a row matrix $\mathbf{L}\in\operatorname{\operatorname{M}}_{1,m}(R)$ such that that $\mathbf{L}\mathbf{M}=0$. A catcher of $\mathbf{M}$ is a left-annihilator of $\mathbf{M}$ whose entries are $1$-homogeneous. We denote by $\operatorname{\operatorname{catch}}(\mathbf{M})$ the $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{1,m}(R)$ that consists of all the catchers of $\mathbf{M}$, and call it the catcher space of $\mathbf{M}$.
Catchers are intimately connected with alternators, as we shall now see.
Let us take finite-dimensional vector spaces $U$ and $V$ over $\mathbb{F}$, a linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, and let $\mathbf{B}_{\mathcal{S}}=(v_{1},\dots,v_{n})$ and $\mathbf{B}_{V}=(e_{1},\dots,e_{p})$ be respective bases of $\mathcal{S}$ and $V$.
Throughout, we assume that $\mathcal{S}$ is target-reduced (see the definition in Section 1.5), to the effect that $\widehat{\mathcal{S}}$ is also target-reduced. We consider the matrix space $\mathcal{M}$ that represents the dual operator space $\widehat{\mathcal{S}}$ in $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$, and we take an arbitrary generic matrix $\mathbf{M}$ that is associated with $\mathcal{M}$, with parameter space denoted by $E$. We define the following linear mappings:
  • $\Phi$ assigns to every $z\in E$ the operator in $\widehat{\mathcal{S}}$ whose matrix in the bases $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$ is $\mathbf{M}(z)$;
  • $\widehat{(-)}$ assigns to every $y\in U$ the evaluation mapping $\widehat{y}\in\widehat{\mathcal{S}}$.
These mappings are represented in the following diagram
[…]
Now, we consider an alternator $b$ of $\mathcal{S}$. The left mapping $L_{b}:y\in U\mapsto b(y,-)\in V^{\star}$ can then be inserted into the following diagram:
[…]
Next, we prove that $L_{b}$ factors into the composite of $\widehat{(-)}$ with a unique linear mapping $\Theta:\widehat{\mathcal{S}}\rightarrow V^{\star}$. To see this, we apply the factorization lemma for linear maps: all we need is to check that $L_{b}$ vanishes at every vector $y\in U$ such that $\widehat{y}=0$. So, let $y\in U$ satisfy $\widehat{y}=0$. Hence, for all $f\in\mathcal{S}$, we have b(y,f(z))=b(z,f(y))=0b(y,f(z))=-b(z,f(y))=0 because $f$ is $b$-alternating. Because $\mathcal{S}$ is target-reduced, this yields $y\in\operatorname{\operatorname{Ker}}L_{b}$, as claimed.
The previous diagram is then completed as:
[…]
where $\Psi:=\Theta\circ\Phi$. Finally, we consider the matrix $\mathbf{L}\in\operatorname{\operatorname{M}}_{1,p}(R)$ whose entries are the composite linear forms $z\in E\mapsto\Psi[z](e_{i})$, with $i\in\mathopen{[\![}1,p\mathclose{]\!]}$. We claim that $\mathbf{L}\mathbf{M}=0$. Since the entries of $\mathbf{L}\mathbf{M}$ are homogeneous of degree $2$ in $R$, it will suffice to prove that $\forall z\in E,\;\mathbf{L}(z)\mathbf{M}(z)=0$. Letting $z\in E$ and choosing $y\in U$ such that $\Phi(z)=\widehat{y}$, we see from the definition that, for all $j\in\mathopen{[\![}1,n\mathclose{]\!]}$, the $j$-th entry of $\mathbf{L}(z)\mathbf{M}(z)$ equals $\sum_{i=1}^{p}\Psi[z](e_{i})\,e_{i}^{\star}(v_{j}(y))$, where $(e_{1}^{\star},\dots,e_{p}^{\star})$ stands for the dual basis of $(e_{1},\dots,e_{p})$. Hence
$\forall j\in\mathopen{[\![}1,n\mathclose{]\!]},\;\bigl{(}\mathbf{L}(z)\mathbf{M}(z)\bigr{)}_{j}=\Psi[z][v_{j}(y)]=\Theta(\widehat{y})[v_{j}(y)]=b(y,v_{j}(y))=0.$
Moreover, if $\mathbf{L}=0$ then we successively find $\Psi=0$, $\Theta=0$ and finally $b=0$. Assigning $\mathbf{L}$ to $b$ thus creates an injective linear mapping
$\Lambda:\operatorname{\operatorname{Alt}}(\mathcal{S})\hookrightarrow\operatorname{\operatorname{catch}}(\mathbf{M}).$
Conversely, let $\mathbf{L}$ be a catcher of $\mathbf{M}$. Let $\Psi:E\rightarrow V^{\star}$ assign to every $z\in E$ the linear form on $V$ whose matrix in the basis $\mathbf{B}_{V}$ is $\mathbf{L}(z)$. We claim that there is a (unique) factorization $\Psi=\Theta\circ\Phi$ for some linear mapping $\Theta:\widehat{\mathcal{S}}\rightarrow V^{\star}$. Again, we use the factorization lemma. Let $z_{0}\in\operatorname{\operatorname{Ker}}\Phi$. Let us polarize the quadratic identity $\mathbf{L}\mathbf{M}=0$ at $z_{0}$. Because $\mathbf{M}(z_{0})=0$ by assumption, this yields $\mathbf{L}(z_{0})M=0$ for all $M\in\mathcal{M}$. In other words, every operator in $\widehat{\mathcal{S}}$ maps into $\operatorname{\operatorname{Ker}}\Psi(z_{0})$. Since $\widehat{\mathcal{S}}$ is target-reduced, we deduce that $\Psi(z_{0})=0$. Hence the existence of $\Theta$, the uniqueness of which is obvious because $\Phi$ is onto. Finally, we consider the composite mapping
$y\in U\longmapsto\Theta(\widehat{y})\in V^{\star}$
and the associated bilinear form
$b:(y,z)\in U\times V\longmapsto\Theta(\widehat{y})[z].$
Then, by the same computation as in the above we deduce from the identity $\mathbf{L}\mathbf{M}=0$ that $b$ is an alternator of $\mathcal{V}$. And finally it is clear that $\Lambda(b)=\mathbf{L}$.
We conclude that $\Lambda:\operatorname{\operatorname{Alt}}(\mathcal{V})\overset{\simeq}{\longrightarrow}\operatorname{\operatorname{catch}}(\mathbf{M})$ is a vector space isomorphism.
4.4 Simple applications to deeply intransitive operator spaces
We immediately give a simple application of the previous considerations to the structure of the alternator space of an intransitive operator space.
Proposition 4.4.
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$. Let $\mathcal{S}$ be a target-reduced $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Assume that $\operatorname{\operatorname{trk}}\mathcal{S}=nd-1$ and $|\mathbb{F}|\geq nd$. Then $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$.
Proof.
We choose respective bases $\mathbf{B}_{V}=(e_{1},\dots,e_{nd})$ and $\mathbf{B}_{\mathcal{S}}=(f_{1},\dots,f_{N})$ of $V$ and $\mathcal{S}$ seen as $\mathbb{F}$-vector spaces, and we consider the generic matrix $\mathbf{M}$ associated with $x\in U\mapsto\widehat{x}\in\widehat{\mathcal{S}}$ in these bases. The condition that $\operatorname{\operatorname{trk}}\mathcal{S}=nd-1$ yields $\operatorname{\operatorname{rk}}\mathbf{M}=nd-1$, by Corollary 4.2.
Let $b_{1}$ and $b_{2}$ be nonzero alternators of $\mathcal{S}$. We shall prove that $b_{1}$ and $b_{2}$ are linearly dependent over $\mathbb{F}$.
We apply the procedure explained in the previous paragraph to obtain associated catchers $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ of $\mathbf{M}$. Since $\operatorname{\operatorname{rk}}\mathbf{M}=nd-1$, these catchers are linearly dependent over the field $\mathbb{L}$ of fractions of the ring $R$ of symmetric tensors of the dual space $U^{\star}$.
Now, assume that $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}=1$. Then $\mathbf{B_{1}}=\mathbf{b_{1}}B_{1}$ for some nonzero row $B_{1}\in\operatorname{\operatorname{M}}_{1,nd}(\mathbb{F})\smallsetminus\{0\}$ and some $1$-homogeneous element $\mathbf{b_{1}}$ of $R\smallsetminus\{0\}$. Then the identity $\mathbf{B_{1}}\mathbf{M}=0$ leads to $\mathbf{b_{1}}(B_{1}\mathbf{M})=0$, and hence to $B_{1}\mathbf{M}=0$. The matrix $B_{1}$ represents a linear form $\varphi$ in the dual basis of $\mathbf{B}_{V}$, and the statement that $B_{1}\mathbf{M}=0$ means that all the operators in $\widehat{\mathcal{S}}$ map into the kernel of $\varphi$. Hence so do all the operators in $\mathcal{S}$, and we contradict the assumption that $\mathcal{S}$ is target-reduced.
Hence $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}>1$ and we deduce from the Factorization Lemma that $\mathbf{B_{2}}=\alpha\mathbf{B}_{1}$ for some scalar $\alpha\in\mathbb{F}$. This yields $b_{2}=\alpha b_{1}$. We conclude that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension at most $1$. ∎
Proposition 4.5.
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$. Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that $|\mathbb{F}|\geq nd$.
  1. (a)
    If $\mathcal{S}$ is weakly primitively intransitive, then it is deeply intransitive.
  2. (b)
    If $\mathcal{S}$ is primitively intransitive, then it is deeply intransitive.
Proof.
Assume that $\mathcal{S}$ is not deeply intransitive. Then we can find a nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that $\mathcal{S}^{V^{\prime}}$ is not intransitive. Denote by $\pi:V\twoheadrightarrow V/V^{\prime}$ the standard projection, and consider the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$.
Take a basis $(v_{1},\dots,v_{s})$ of $\mathcal{S}^{V^{\prime}}$; and lift a basis of $\pi\mathcal{S}$ to a family $(u_{1},\dots,u_{r})$ in $\mathcal{S}$, so that $(u_{1},\dots,u_{r},v_{1},\dots,v_{s})$ is a basis $\mathbf{B}_{\mathcal{S}}$ of $\mathcal{S}$. Take an arbitrary $\mathbb{F}$-basis $\mathbf{B}_{V}$ of $V$ that is adapted to $V^{\prime}$ and in which the first $d$ vectors are pairwise linearly dependent over $\mathbb{D}$, and consider a generic matrix $\mathbf{M}$ of $\widehat{\mathcal{S}}$ in the bases $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$. Then this matrix takes the form
$\mathbf{M}=\begin{bmatrix}[?]_{p\times r}&\mathbf{C}\\ \mathbf{B}&[0]\end{bmatrix}\quad\text{where $p:=\dim_{\mathbb{F}}(V^{\prime})$.}$
Note that $\mathbf{C}$ is a generic matrix of $\widehat{\mathcal{S}^{V^{\prime}}}$, whereas $\mathbf{B}$ is a generic matrix of $\widehat{\pi\mathcal{S}}$. Hence $\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{trk}}(\mathcal{S}^{V^{\prime}})=\dim_{\mathbb{F}}V^{\prime}=p$. Besides $\operatorname{\operatorname{rk}}\mathbf{C}+\operatorname{\operatorname{rk}}\mathbf{B}\leq\operatorname{\operatorname{rk}}\mathbf{M}$.
Consider the $\mathbb{D}$-linear subspace $L$ of $V$ spanned by the first $d$ vectors of $\mathbf{B}_{V}$, and the matrix $\mathbf{M^{\prime}}$ obtained by deleting the first $d$ rows of $\mathbf{M}$. Then $\mathbf{M^{\prime}}$ is a generic matrix of $\pi^{\prime}\mathcal{S}$ where $\pi^{\prime}:V\twoheadrightarrow V/L$ is the standard projection. By deleting the first $d$ rows we see that
$\operatorname{\operatorname{rk}}\mathbf{M^{\prime}}\leq d(\dim_{\mathbb{D}}V^{\prime}-1)+\operatorname{\operatorname{rk}}\mathbf{B}\leq d(\dim_{\mathbb{D}}V^{\prime}-1)+\operatorname{\operatorname{rk}}\mathbf{M}-\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{rk}}\mathbf{M}-d.$
Hence $\operatorname{\operatorname{trk}}(\pi^{\prime}\mathcal{S})\leq\operatorname{\operatorname{trk}}(\mathcal{S})-d$. We conclude that $\mathcal{S}$ is not weakly primitively intransitive.
The second point is deduced from the first one and from Proposition 1.23. ∎
5 Nondegenerate alternators of large deeply intransitive operator spaces
Our aim here is to complete the proof of Theorem 1.20. We have already proved point (a).
Our proof mainly has two steps. The first and most difficult step consists, along with the proof of point (b) of Theorem 1.20, in proving the following weak form of point (c):
Proposition 5.1.
Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim V$. Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
Then $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is $1$-dimensional.
Once this result is obtained, we will directly consider operator spaces that have a $1$-dimensional alternator space, and point (c) of Theorem 1.20 will be obtained along with Theorem 1.21 thanks to the classification of multiplicative quadratic forms on division algebras (Theorem 2.3).
5.1 Setup
Throughout, we let $n\geq 1$ be an integer, and $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}V=n$. We assume that $|\mathbb{F}|\geq nd$. For simplicity, we set $\varepsilon=1$ if $d=1$, and $\varepsilon=0$ otherwise. Finally, we take a deeply intransitive $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that
$\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon.$
We will prove the following two points by induction on $n$:
  • In any case $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$.
  • If $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ then $\mathcal{S}$ has a nonzero alternator.
Since $n>0$, we immediately retrieve from Corollary 3.5 that $\mathcal{S}$ is target-reduced.
Throughout, we set
$r:=\operatorname{\operatorname{trk}}(\mathcal{S})\in\mathopen{[\![}0,nd-1\mathclose{]\!]}.$
Since $\mathcal{S}$ is target-reduced we have $\mathcal{S}\neq\{0\}$ and hence $r>0$. We fix a vector $x\in U\smallsetminus\{0\}$ such that $\operatorname{\operatorname{rk}}\widehat{x}=r$. Then we apply the Tangent Space Lemma to recover that all the operators in
$\mathcal{S}^{\prime}:=\{u\in\mathcal{S}:\;u(x)=0\}$
map into $\mathcal{S}x$. We denote by $H$ the socle of $\mathcal{S}x$ and obtain that all the operators in $\mathcal{S}^{\prime}$ map into $H$ (despite the notation, at this point $H$ is not necessarily a linear hyperplane of the $\mathbb{D}$-vector space $V$). We set
$s:=\dim_{\mathbb{F}}H.$
Note that $\mathcal{S}^{\prime}$ is included in $\mathcal{S}^{H}$, which is deeply intransitive, and hence $\mathcal{S}^{\prime}$ is also a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
Now it is time to represent the dual operator space $\widehat{\mathcal{S}}$ in appropriate bases of $\mathcal{S}$ and $V$. We will choose such bases carefully with respect to the vector $x$. For the basis of $V$:
  • We start from a basis $(e_{1},\dots,e_{r})$ of $\mathcal{S}x$ in which the first $s$ vectors span $H$ (as an $\mathbb{F}$-linear subspace), and we extend it arbitrarily into a basis $\mathbf{E}$ of $V$.
  • We lift the first vectors $e_{1},\dots,e_{r}$ to obtain a family $(u_{1},\dots,u_{r})$ of operators in $\mathcal{S}$, so that $u_{i}(x)=e_{i}$ for all $i\in\mathopen{[\![}1,r\mathclose{]\!]}$. Then $(u_{1},\dots,u_{r})$ is linearly independent and we can choose an arbitrary basis $(u_{r+1},\dots,u_{\dim\mathcal{S}})$ of $\mathcal{S}^{\prime}$. Then our basis of $\mathcal{S}$ is $\mathbf{F}:=(u_{1},\dots,u_{r},u_{r+1},\dots,u_{\dim\mathcal{S}})$.
We consider the matrix space $\mathcal{M}$ that represents $\widehat{\mathcal{S}}$ in the bases $\mathbf{F}$ and $\mathbf{E}$. The matrix in $\mathcal{M}$ that corresponds to the operator $\widehat{x}$ is the reduced matrix
$J_{r}=\begin{bmatrix}I_{r}&[0]_{r\times(\dim\mathcal{S}-r)}\\ [0]_{(nd-r)\times r}&[0]_{(nd-r)\times(\dim\mathcal{S}-r)}\end{bmatrix}$.
Now, we take $E:=U$ as our parameter space, $R$ and $\mathbb{L}$ as the associated ring and field, and we construct a generic matrix $\mathbf{M}$ of $\mathcal{M}$ as follows: for all $(i,j)\in\mathopen{[\![}1,nd\mathclose{]\!]}\times\mathopen{[\![}1,\dim\mathcal{S}\mathclose{]\!]}$, we define $\mathbf{m}_{i,j}$ as the linear form on $U$ that takes every vector $z\in U$ to the $(i,j)$-entry of the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$. Hence, the specialization $\mathbf{M}(z)$ is simply the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
As we already know that all the operators in $\mathcal{S}^{\prime}$ map into $H$, the matrix $\mathbf{M}$ actually takes the refined form
$\mathbf{M}=\begin{bmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{B}&[0]_{(nd-r)\times(\dim\mathcal{S}-r)}\end{bmatrix}$
for some matrices $\mathbf{A}\in\operatorname{\operatorname{M}}_{r}(R)$, $\mathbf{B}\in\operatorname{\operatorname{M}}_{nd-r,r}(R)$ and $\mathbf{C}\in\operatorname{\operatorname{M}}_{r,\dim\mathcal{S}-r}(R)$ with $1$-homogeneous entries, and we can further refine
$\mathbf{C}=\begin{bmatrix}\mathbf{C_{1}}\\ [0]_{(r-s)\times(\dim S-r)}\end{bmatrix}\qquad\text{with $\mathbf{C_{1}}\in\operatorname{\operatorname{M}}_{s,\dim\mathcal{S}-r}(R)$.}$
Accordingly, we also refine
$\mathbf{B}=\begin{bmatrix}\mathbf{B_{1}}&\mathbf{B_{2}}\end{bmatrix}\qquad\text{with $\mathbf{B_{1}}\in\operatorname{\operatorname{M}}_{nd-r,s}(R)$ and $\mathbf{B_{2}}\in\operatorname{\operatorname{M}}_{nd-r,r-s}(R)$.}$
Finally, we write
$\mathbf{A}=\begin{bmatrix}\mathbf{A_{1}}&\mathbf{A_{3}}\\ \mathbf{A_{2}}&\mathbf{A_{4}}\end{bmatrix}\qquad\text{with $\mathbf{A_{1}}\in\operatorname{\operatorname{M}}_{s}(R)$, $\mathbf{A_{2}}\in\operatorname{\operatorname{M}}_{r-s,s}(R)$ and so on.}$
Critically, $\mathbf{C_{1}}$ is a generic matrix for the dual operator space $\widehat{\mathcal{S}^{\prime}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S}^{\prime},H)$ in the basis $(u_{r+1},\dots,u_{\dim\mathcal{S}})$ of $\mathcal{S}^{\prime}$ and the basis $(e_{1},\dots,e_{s})$ of the $\mathbb{F}$-vector space $H$. Here, the parameter space for this generic matrix remains $U$.
By the very definition of the transitive rank, the maximal rank in $\mathcal{M}$ is $r$. Since both matrices $J_{r}$ and $\mathbf{M}$ belong to $\operatorname{\operatorname{span}}_{\mathbb{L}}(\mathcal{M})$ (see the remark following Corollary 4.2), so does every linear combination of them over $\mathbb{L}$. Hence by Lemma 4.1 every linear combination of $J_{r}$ and $\mathbf{M}$ over $\mathbb{L}$ has rank at most $r$. Thanks to the infiniteness of $\mathbb{L}$ we can apply the Flanders-Atkinson Lemma (Lemma 3.2), yielding the identity
$\forall k\geq 0,\;\mathbf{B}\mathbf{A}^{k}\mathbf{C}=0.$ (5)
The case $k=0$ yields
$\mathbf{B_{1}}\mathbf{C_{1}}=0.$
Before we proceed, it is time to complete the special case $n=1$ for the proof that $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$. So, assume for a moment that $n=1$. In that case $\mathcal{S}^{\prime}=\{0\}$ and hence $u\in\mathcal{S}\mapsto u(x)\in\mathcal{S}x$ is an $\mathbb{F}$-linear isomorphism. It follows that $r=\dim\mathcal{S}$. The assumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon$ requires here that $\dim\mathcal{S}\geq nd-1$ and hence r=nd1r=nd-1.
From now on, we assume that $n\geq 2$. We will now draw some consequences of the assumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon$.
Claim 5.1.
Assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1. Then the space $H$ has dimension n1n-1 over $\mathbb{D}$, and the space $\mathcal{S}^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$ is target-reduced.
Proof.
Assume that $\dim_{\mathbb{D}}H<n-1$ or that $\mathcal{S}^{\prime}$ is not target-reduced. Recall from Remark 3.1 that this yields a $\mathbb{D}$-linear subspace $G$ of $V$ with dimension n2n-2 such that $\mathcal{S}^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,G)$. Since $\mathcal{S}^{\prime}$ is deeply intransitive we deduce from point (a) of Theorem 1.20 that $\dim\mathcal{S}^{\prime}\leq\alpha(n-2,d)$. Hence $\dim\mathcal{S}-r\leq\alpha(n-2,d)$. Yet the starting assumption shows that $\dim\mathcal{S}-r\geq\alpha(n-1,d)-n+2+\varepsilon$, and hence $\alpha(n-2,d)\geq\alpha(n-1,d)-n+2+\varepsilon$, to the effect that $(n-1)d-1=\alpha(n-1,d)-\alpha(n-2,d)\leq n-2-\varepsilon$. Hence $(n-1)(d-1)\leq-\varepsilon$, which contradicts the definition of $\varepsilon$ because $n\geq 2$. ∎
Claim 5.2.
There is no row $L\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$ such that $L\mathbf{B_{1}}=0$. In particular $\mathbf{B_{1}}\neq 0$.
Proof.
Assume that the contrary holds. Then there is a linear hyperplane $\mathcal{H}$ of the $\mathbb{F}$-vector space $V$ that includes $H$ and such that all the operators $u_{1},\dots,u_{(n-1)d}$ map into $\mathcal{H}$. Denote by $H^{\prime}$ the socle of $\mathcal{H}$, and note that $H\subseteq H^{\prime}$ and $H^{\prime}$ is a $\mathbb{D}$-linear hyperplane of $V$.
The $\mathbb{D}$-linear mappings $u_{1},\dots,u_{(n-1)d}$ map into $H^{\prime}$, and so do all the elements of $\mathcal{S}^{\prime}$. It follows that $\dim(\mathcal{S}^{H^{\prime}})\geq\dim\mathcal{S}^{\prime}+(n-1)d$. Yet $\mathcal{S}^{H^{\prime}}$ is deeply intransitive, so we can apply point (a) of Theorem 1.20 once more to find $\dim\mathcal{S}^{\prime}+(n-1)d\leq\alpha(n-1,d)$, leading to $\dim\mathcal{S}\leq\alpha(n-1,d)-(n-1)d+r=\alpha(n-2,d)+r-1$. A contradiction is then found like in the proof of Claim 5.1, using only $\dim\mathcal{S}-r\geq\alpha(n-1,d)-n+1+\varepsilon$ here. ∎
Claim 5.3.
Assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1. Let $X\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$. Then $\operatorname{\operatorname{sprk}}(X\mathbf{B_{1}})>1$.
Proof.
Assume that the contrary holds. By Claim 5.2, we have $X\mathbf{B_{1}}\neq 0$, and then $X\mathbf{B_{1}}=\mathbf{a}R_{1}$ for some row $R_{1}\in\operatorname{\operatorname{M}}_{1,(n-1)d}(\mathbb{F})\smallsetminus\{0\}$ and some $\mathbf{a}\in R\smallsetminus\{0\}$. The identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ then yields $\mathbf{a}(R_{1}\mathbf{C_{1}})=0$, whence $R_{1}\mathbf{C_{1}}=0$.
This contradicts the fact that $\mathcal{S}^{\prime}$ is target-reduced (see Claim 5.1). ∎
5.2 The transitive rank
Claim 5.4.
One has $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$.
Proof.
We assume that $r\leq nd-2$ and seek to find a contradiction.
Denote by $r^{\prime}$ the rank of $\mathbf{C_{1}}$: it is also the transitive rank of $\mathcal{S}^{\prime}$. If $r^{\prime}<(n-1)d-1$, then by induction $\dim\mathcal{S}^{\prime}<\alpha(n-1,d)-(n-1)+1+\varepsilon$ and hence
$\dim\mathcal{S}=r+\dim\mathcal{S}^{\prime}<nd-2+\alpha(n-1,d)-(n-1)+1+\varepsilon=\alpha(n,d)-n+1+\varepsilon,$
which contradicts our starting assumption.
Hence $r^{\prime}=(n-1)d-1$ (remember that $\mathcal{S}^{\prime}$ is intransitive). Let us pick $\mathbb{F}$-linearly independent rows $L_{1}$ and $L_{2}$ in $\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})$. Set $\mathbf{R_{i}}:=L_{i}\mathbf{B_{1}}$ for $i\in\{1,2\}$. We have $\mathbf{R_{i}}\mathbf{C_{1}}=0$ with $\operatorname{\operatorname{rk}}\mathbf{C_{1}}=(n-1)d-1$ and the number of rows of $\mathbf{C_{1}}$ is at most (n1)d(n-1)d. Hence $\mathbf{R_{1}}$ and $\mathbf{R_{2}}$ are linearly dependent over $\mathbb{L}$. By Claim 5.3 (remembering that we assume r<nd1r<nd-1) we know that $\operatorname{\operatorname{sprk}}(\mathbf{R_{1}})>1$, and hence the Factorization Lemma yields a scalar $\alpha\in\mathbb{F}$ such that $\mathbf{R_{2}}=\alpha\mathbf{R_{1}}$. Thus $(L_{2}-\alpha L_{1})\mathbf{B_{1}}=0$ with $L_{2}-\alpha L_{1}\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$, thereby contradicting Claim 5.2.
We conclude that r=nd1r=nd-1. ∎
5.3 Deciphering the large spaces further
Now we proceed to prove the last point. So, now we assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$, and we prove that $\mathcal{S}$ has a nonzero alternator. Note that the assumptions of Claims 5.1 and 5.3 are satisfied. In particular $\mathcal{S}^{\prime}$ is target-reduced. Next,
$\dim\mathcal{S}^{\prime}=\dim\mathcal{S}-r\geq\alpha(n-1,d)-(n-1)+1+\varepsilon,$
and hence by induction we recover $\operatorname{\operatorname{trk}}(\mathcal{S}^{\prime})=(n-1)d-1$. It then follows from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S}^{\prime})\leq 1$.
Since r=nd1r=nd-1, we see that $\mathbf{B_{1}}$ and $\mathbf{B_{2}}$ are simply row matrices, and we know from Claim 5.3 that $\operatorname{\operatorname{sprk}}(\mathbf{B_{1}})>1$.
Since $\mathcal{S}^{\prime}$ is target-reduced, the identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ yields that $\mathbf{B_{1}}$ represents a nonzero alternator $b:U\times H\rightarrow\mathbb{F}$ of $\mathcal{S}^{\prime}$ (seen as a subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$). We deduce that $\operatorname{\operatorname{Alt}}(\mathcal{S}^{\prime})=\mathbb{F}b$, and hence the left radical
$\mathcal{L}:=\operatorname{\operatorname{Lrad}}(b^{\prime})$
does not depend on the choice of $b^{\prime}$ in $\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$ (in particular it is the left radical of $b$).
For all $y\in U$, the specialized matrix $\mathbf{B_{1}}(y)$ represents the $\mathbb{F}$-linear form b(y,)b(y,-) in the basis $(e_{1},\dots,e_{(n-1)d})$ of $H$.
Then for all $y\in U\smallsetminus\mathcal{L}$ we have $\mathbf{B_{1}}(y)\neq 0$, which is interpreted as follows in geometric terms:
Claim 5.5.
Let $y\in U\smallsetminus\mathcal{L}$. The mappings $u\in\mathcal{S}$ such that $u(x)\in H$ do not all map $y$ into $\mathcal{S}x$.
We now move on to what is arguably the most subtle item in our proof:
Claim 5.6.
The compound matrix
$\mathbf{N}:=\begin{bmatrix}\mathbf{A_{2}}\\ \mathbf{B_{1}}\end{bmatrix}\in\operatorname{\operatorname{M}}_{d,(n-1)d}(R)$
has rank $d$.
Proof.
We take an arbitrary vector $y\in U\smallsetminus\mathcal{L}$ (which exists!). We shall prove that $\operatorname{\operatorname{rk}}\mathbf{N}(y)\geq d$: since $(n-1)d\geq d$ this will be enough to obtain the conclusion.
Assume on the contrary that $\operatorname{\operatorname{rk}}\mathbf{N}(y)<d$. Then there is a nonzero row $L=\begin{bmatrix}l_{1}\cdots l_{d}\end{bmatrix}\in\operatorname{\operatorname{M}}_{1,d}(\mathbb{F})$ such that $L\mathbf{N}(y)=0$. Denote by $\ell$ the $\mathbb{F}$-linear form on $V$ such that $\ell(e_{1})=\dots=\ell(e_{(n-1)d})=0$ and $\ell(e_{(n-1)d+i})=l_{i}$ for all $i\in\mathopen{[\![}1,d\mathclose{]\!]}$. This linear form is nonzero and its kernel includes $H$. The assumption $L\mathbf{N}(y)=0$ then shows that all the operators $u_{1},\dots,u_{(n-1)d}$ map $y$ into $\operatorname{\operatorname{Ker}}\ell$, and hence all the operators of $\mathcal{S}$ that map $x$ into $H$ also map $y$ into $\operatorname{\operatorname{Ker}}\ell$.
Next, we prove that $\operatorname{\operatorname{Ker}}\ell=(\mathcal{S}x)a^{-1}$ for some $a\in\mathbb{D}\smallsetminus\{0\}$. Choose indeed an $\mathbb{F}$-linear form $\ell^{\prime}$ on $V$ with kernel $\mathcal{S}x$. The linear forms $z\in V\mapsto\ell^{\prime}(za)$, with $a\in\mathbb{D}$, constitute a $d$-dimensional $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(V,\mathbb{F})$, and they all vanish on $H$. By orthogonality theory, it follows that they are exactly the $\mathbb{F}$-linear forms on $V$ whose kernel include $H$. In particular there exists $a\in\mathbb{D}$ such that $\forall z\in V,\;\ell(z)=\ell^{\prime}(za)$, and obviously $a\neq 0$. It follows that
$\operatorname{\operatorname{Ker}}\ell=(\operatorname{\operatorname{Ker}}\ell^{\prime})a^{-1}=(\mathcal{S}x)a^{-1}=\mathcal{S}(xa^{-1}).$
Now the trick is to use the fact that $xa^{-1}$ satisfies exactly the same assumptions as the vector $x$ we started from. Indeed $\operatorname{\operatorname{rk}}(\widehat{xa^{-1}})=\dim_{\mathbb{F}}\bigl{(}\mathcal{S}(xa^{-1})\bigr{)}=nd-1$, the operators in $\mathcal{S}$ that vanish at $x$ are also the ones that vanish at $xa^{-1}$, and $H$ is the $\mathbb{D}$-linear hyperplane of $V$ that is included in $\mathcal{S}(xa^{-1})$. Hence Claim 5.5 applies when $x$ is replaced with $xa^{-1}$. Every $u\in\mathcal{S}$ that maps $x$ into $H$ also maps $xa^{-1}$ into $H$ because $u$ is $\mathbb{D}$-linear, and hence there is at least one such operator that does not map $y$ into $\mathcal{S}(xa^{-1})$. This contradicts the first step of the present proof. We conclude that $\operatorname{\operatorname{rk}}\mathbf{N}(y)\geq d$, which completes the proof. ∎
5.4 Constructing a nonzero alternator of $\mathcal{S}$
Now everything is in place for us to produce a nonzero alternator of $\mathcal{S}$.
We start by applying identity (5) once more. This time around, by taking $k=1$ we find $\mathbf{B}\mathbf{A}\mathbf{C}=0$. Writing
$\mathbf{B}\mathbf{A}=\begin{bmatrix}\mathbf{B^{\prime}_{1}}&\mathbf{B^{\prime}_{2}}\end{bmatrix}\quad\text{with $\mathbf{B^{\prime}_{1}}\in\operatorname{\operatorname{M}}_{1,(n-1)d}(R)$ and $\mathbf{B^{\prime}_{2}}\in\operatorname{\operatorname{M}}_{1,d-1}(R)$,}$
we deduce that $\mathbf{B^{\prime}_{1}}$ is a left-annihilator of $\mathbf{C}_{1}$. Note that all the entries of $\mathbf{B^{\prime}_{1}}$ are $2$-homogeneous. Yet $\mathbf{B_{1}}$ is also a nonzero left-annihilator of $\mathbf{C}_{1}$. Since $\operatorname{\operatorname{rk}}\mathbf{C}_{1}=(n-1)d-1$, we deduce that $\mathbf{B^{\prime}_{1}}$ is a scalar multiple of $\mathbf{B_{1}}$ over the field $\mathbb{L}$. Thanks to $\operatorname{\operatorname{sprk}}\mathbf{B_{1}}>1$ (see Claim 5.3) we deduce from the Factorization Lemma (Lemma 4.3) that
$\mathbf{B^{\prime}_{1}}=\mathbf{a}\mathbf{B_{1}}\quad\text{for some $1$-homogeneous $\mathbf{a}\in R$.}$
Now, we consider the new row $\mathbf{R}:=\mathbf{B}(\mathbf{A}-\mathbf{a}I_{nd-1})$, which can be written $\begin{bmatrix}[0]_{1\times(n-1)d}&\mathbf{R_{2}}\end{bmatrix}$. We also write
$\mathbf{R}\mathbf{A}=\begin{bmatrix}\mathbf{R^{\prime}_{1}}&\mathbf{R^{\prime}_{2}}\end{bmatrix}$
with $\mathbf{R^{\prime}_{1}}\in\operatorname{\operatorname{M}}_{1,(n-1)d}(R)$ and $\mathbf{R^{\prime}_{2}}\in\operatorname{\operatorname{M}}_{1,d-1}(R)$, both with $3$-homogeneous entries. Using identity (5) with $k=1$ and $k=2$, we now obtain
$\mathbf{R}\mathbf{A}\mathbf{C}=0.$
Applying the Factorization Lemma once more, we obtain
$\mathbf{R^{\prime}_{1}}=\mathbf{p}\mathbf{B_{1}}\quad\text{for some $2$-homogeneous $\mathbf{p}\in R$,}$
which yields
$\begin{bmatrix}\mathbf{R_{2}}&-\mathbf{p}\end{bmatrix}\begin{bmatrix}\mathbf{A_{2}}\\ \mathbf{B_{1}}\end{bmatrix}=0.$
Using Claim 5.6, we deduce that $\begin{bmatrix}\mathbf{R_{2}}&-\mathbf{p}\end{bmatrix}=0$. In particular $\mathbf{R_{2}}=0$, and we conclude that
$\mathbf{B}(\mathbf{A}-\mathbf{a}I_{nd-1})=0.$
By combining this with $\mathbf{B}\mathbf{C}=0$, we conclude that $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$ is a (nonzero) catcher of $\mathbf{M}$.
Since it was known from the start that $\mathcal{S}$ is target-reduced, the process explained in the end of Section 4 yields a nonzero alternator $\widetilde{b}:U\times V\rightarrow\mathbb{F}$ of $\mathcal{S}$, associated with the nonzero catcher $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$.
Finally, since $\mathcal{S}$ has transitive rank nd1nd-1 and is target-reduced, we know from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
This completes the proof of Proposition 5.1.
5.5 Operator spaces with $1$-dimensional alternator spaces
We will now complete the proof of Theorem 1.20, by proving point (c).
We will prove the following general result, which avoids any dimensional assumption on $\mathcal{S}$ and focuses on the condition regarding the structure of the alternator space.
Proposition 5.2.
Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$, and let $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$. Then:
  1. (a)
    $\mathbb{D}$ has quadratic type over $\mathbb{F}$.
  2. (b)
    If we take a pair $(\sigma,e)$ that is associated with $(\mathbb{D},\mathbb{F})$, then there is a unique $\sigma$-sesquilinear mapping $B:U\times V\rightarrow\mathbb{D}$ such that $b(x,y)=e(B(x,y))$ for all $(x,y)\in U\times V$.
Proof.
The key is to consider the “action” of $\mathbb{D}$ on the set $\mathcal{B}(U\times V,\mathbb{F})$ of all bilinear forms on $U\times V$. For such a form $c$ and for $a\in\mathbb{D}$, we set
$c^{a}:(x,y)\in U\times V\mapsto c(xa,ya)\in\mathbb{F}.$
Then $a\mapsto[c\mapsto c^{a}]$ defines an $\mathbb{F}$-quadratic mapping $P:\mathbb{D}\rightarrow\operatorname{\operatorname{End}}_{\mathbb{F}}(\mathcal{B}(U\times V,\mathbb{F}))$, with $P(aa^{\prime})=P(a)\circ P(a^{\prime})$ for all $(a,a^{\prime})\in\mathbb{D}^{2}$, i.e., $P$ is multiplicative. Moreover, $P(a)$ leaves $\operatorname{\operatorname{Alt}}(\mathcal{S})$ invariant for all $a\in\mathbb{D}$, as for all $c\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ we observe that
$\forall u\in\mathcal{S},\;\forall x\in U,\;c(xa,u(x)a)=c(xa,u(xa))=0.$
Since $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$, we deduce that $b$ is an eigenvector of $P(a)$ for all $a\in\mathbb{D}$, and by denoting by $q(a)\in\mathbb{F}$ the corresponding eigenvalue we obtain an $\mathbb{F}$-quadratic form $q:\mathbb{D}\rightarrow\mathbb{F}$. Obviously $\operatorname{\operatorname{rk}}b^{a}=\operatorname{\operatorname{rk}}b$ for all $a\in\mathbb{D}^{\times}$ (because $\mathbb{D}$ is a division ring), and in particular $q$ is nonisotropic. Finally, since $P$ is multiplicative we obtain that $q$ is multiplicative.
Therefore, we are in the position to use Theorem 2.3: the algebra $\mathbb{D}$ has quadratic type over $\mathbb{F}$. Let us now choose a pair $(\sigma,e)$ that is associated with $(\mathbb{D},\mathbb{F})$. Then we also know that $q(a)=a\sigma(a)$ for all $a\in\mathbb{D}$. Moreover $e$ satisfies the identity $\forall(x,y)\in\mathbb{D}^{2},\;e(yx)=e(xy)$, which will ease things up from now on. As seen in Section 2.1, since $e$ is nonzero we recover a unique $\mathbb{F}$-bilinear and right-$\mathbb{D}$-linear form $B$ on $U\times V$ such that
$\forall(x,y)\in U\times V,\;b(x,y)=e(B(x,y)).$
Let $a\in\mathbb{D}$. Then
$\forall(x,y)\in U\times V,\;b^{a}(x,y)=e(B(xa,ya))=e(B(xa,y)a)=e(aB(xa,y))$
and we deduce that
$\forall(x,y)\in U\times V,\;aB(xa,y)=q(a)B(x,y)=a\sigma(a)B(x,y).$
If $a\neq 0$ we derive that
$\forall(x,y)\in U\times V,\;B(xa,y)=\sigma(a)B(x,y),$
which is of course obvious if $a=0$. Hence $B$ is $\sigma$-sesquilinear, as claimed. ∎
Now we can complete the proof of point (c) of Theorem 1.20. Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ with $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ and $|\mathbb{F}|\geq nd$. By Proposition 5.1 we have $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
Choose $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$. We need to prove that $b$ is right-nondegenerate. Set $R:=\operatorname{\operatorname{Rrad}}(b)$, which by point (c) of Proposition 5.2 is a $\mathbb{D}$-linear subspace of $V$. Set $t:=\dim_{\mathbb{D}}R\in\mathopen{[\![}0,n-1\mathclose{]\!]}$, and denote by $\pi:V\twoheadrightarrow V/R$ the standard projection, which is $\mathbb{D}$-linear. The bilinear form $b$ induces a right-nondegenerate $\mathbb{F}$-bilinear form $\overline{b}$ on $U\times(V/R)$. The space $\pi\mathcal{S}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/R)$, and it is obvious that $\overline{b}$ is an alternator of it. Since $\overline{b}$ is right-nondegenerate, we know from Proposition 1.18 that $\pi\mathcal{S}$ is deeply intransitive. Besides, the kernel of $u\in\mathcal{S}\mapsto\pi\circ u$ is precisely $\mathcal{S}^{R}$, which is deeply intransitive. Hence, by combining the rank theorem with point (a) of Theorem 1.20, we find
$\dim(\mathcal{S})=\dim(\pi\mathcal{S})+\dim(\mathcal{S}^{R})\leq\alpha(n-t,d)+\alpha(t,d)=\alpha(n,d)-dt(n-t).$
It $t>0$ then $dt(n-t)\geq d(n-1)$, and we observe that $d(n-1)>n-\max(4-d,2)$ because obviously $(d-1)(n-1)\geq 0>1-\max(4-d,2)$. This runs counter to our starting assumption on the dimension of $\mathcal{S}$.
Therefore $t=0$, to the effect that $b$ is right-nondegenerate. This completes the proof of point (c) in Theorem 1.20.
Then, by combining this very point with Proposition 5.2, we derive Theorem 1.21.
5.6 The converse statement
It remains to check that, in the case of a division algebra of quadratic type, the spaces $\mathcal{A}_{b,\mathbb{D}}$ truly have the claimed properties (and most critically that their dimension is the bound $\alpha(n,d)$ we have found). Interestingly, we can prove this without any assumption on the underlying field $\mathbb{F}$. This requires the next lemma, which for fields with large cardinality is a straightforward consequence of Corollary 3.5.
Lemma 5.3.
Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V>0$, and set $n:=\dim_{\mathbb{D}}U$. Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ for all $x\in U\smallsetminus\{0\}$. Assume that $\mathcal{S}$ is target-reduced. Then all the nonzero alternators of $\mathcal{S}$ are nondegenerate, and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension at most $1$.
Proof.
Let $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ be nonzero and degenerate. It follows that $nd=\dim_{\mathbb{F}}V>1$. Every element of $\mathcal{S}$ maps the left radical $L$ of $b$ into its right radical $R$, and here $\dim_{\mathbb{F}}L=\dim_{\mathbb{F}}R$ since $\dim_{\mathbb{F}}U=\dim_{\mathbb{F}}V$.
We choose $x\in L\smallsetminus\{0\}$ and deduce that $\dim_{\mathbb{F}}R\geq\dim_{\mathbb{F}}\mathcal{S}x=nd-1$. Hence $\operatorname{\operatorname{rk}}(b)=1$. Then each bilinear form $(x,y)\in U^{2}\mapsto b(x,u(y))$, with $u\in\mathcal{S}$, is alternating with rank at most $1$, and hence is zero. Hence $\operatorname{\operatorname{Im}}u\subseteq R$ for all $u\in\mathcal{S}$, which contradicts the assumption that $\mathcal{S}$ is target-reduced. Hence, every nonzero element of $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is nondegenerate.
Next, let $b$ and $b^{\prime}$ belong to $\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$. Choose $x\in U\smallsetminus\{0\}$. By the first step, $b(x,-)\neq 0$. The nonzero linear forms $b^{\prime}(x,-)$ and b(x,)b(x,-) vanish on the $\mathbb{F}$-linear hyperplane $\mathcal{S}x$, and hence $b^{\prime}(x,-)=\lambda b(x,-)$ for some $\lambda\in\mathbb{F}$. Then $b^{\prime}-\lambda b$ is a degenerate element in $\operatorname{\operatorname{Alt}}(\mathcal{S})$, and by the first step $b^{\prime}-\lambda b=0$. We conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$. ∎
Proposition 5.4.
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, and denote by $(\sigma,e)$ an associated pair. Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $n:=\dim_{\mathbb{D}}V>0$. Let $B:U\times V\rightarrow\mathbb{D}$ be a right-nondegenerate $\sigma$-sesquilinear form, and set $b:(x,y)\in U\times V\longmapsto e(B(x,y))$. Then:
  1. (a)
    $\dim\mathcal{A}_{b,\mathbb{D}}=\alpha(n,d)$.
  2. (b)
    If $\dim_{\mathbb{D}}U=n$ then $\operatorname{\operatorname{Alt}}(\mathcal{A}_{b,\mathbb{D}})=\mathbb{F}b$ and $\dim(\mathcal{A}_{b,\mathbb{D}}\,x)=nd-1$ for all $x\in U\smallsetminus\{0\}$.
  3. (c)
    If $U=V$, then $\mathcal{A}_{b,\mathbb{D}}$ has trivial spectrum if and only if $B$ is $e$-nonisotropic, in which case it is irreducible.
Note that point (a) is enough to yield Proposition 1.22. Indeed, in the conclusion, the deep intransitivity of $\mathcal{A}_{b,\mathbb{D}}$ is a straightforward consequence of point (b) in Proposition 1.18 (remember from Lemma 2.2 that the right-nondegeneracy of $b$ is equivalent to the one of $B$).
Proof.
Set $\mathcal{S}:=\mathcal{A}_{b,\mathbb{D}}$. To start with, to prove the first point we reduce the situation to the one where $B$ is nondegenerate. To see this, we consider the left radical $L$ of $B$, which equals the one of $b$, and the induced form $\overline{B}:(U/L)\times V\rightarrow\mathbb{D}$, and the $\mathbb{F}$-bilinear form $\overline{b}:(x,y)\in(U/L)\times V\mapsto e(\overline{B}(x,y))$. All the operators in $\mathcal{S}$ map $L$ into the right radical of $b$, which is zero, so they all vanish on $L$. Hence every $u\in\mathcal{S}$ induces a $\mathbb{D}$-linear mapping $\overline{u}:U/L\rightarrow V$, and we have $\overline{b}(\overline{x},\overline{u}(\overline{x}))=0$ for all $x\in U/L$. We denote by $\overline{\mathcal{S}}$ the space of all mappings $\overline{u}$ with $u\in\mathcal{S}$.
Conversely, for all $u\in\mathcal{A}_{\overline{b},\mathbb{D}}$ it is clear that we have $u\circ\pi\in\mathcal{S}$ for the standard projection $\pi:U\twoheadrightarrow U/L$. Therefore $\overline{\mathcal{S}}=\mathcal{A}_{\overline{b},\mathbb{D}}$. Since $u\in\mathcal{S}\mapsto\overline{u}\in\overline{\mathcal{S}}$ is clearly injective, we deduce that $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{A}_{\overline{b},\mathbb{D}}$.
Hence, in the remainder of the proof we need only consider the situation where $\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V=n$. We shall use a matrix interpretation. Let $x\in U\smallsetminus\{0\}$, and extend it into a $\mathbb{D}$-basis $\mathbf{B}_{U}$ of $U$ with $x$ as first vector, and consider a $\mathbb{D}$-basis $\mathbf{B}_{V}$ of $V$.
Denote by $P$ the Gram matrix of $B$ in the bases $U$ and $V$, and set $\mathcal{M}=\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$. Then the elements of $\mathcal{S}$ are represented by the elements of $P^{-1}\mathcal{M}$ in the bases $\mathbf{B}_{U}$ and $\mathbf{B}_{V}$, and in particular $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{M}$. Then, it is easily checked in each case that
$\dim\mathcal{M}=n(d-1)+d\dbinom{n}{2}=\alpha(n,d);$
indeed, each diagonal entry varies independently in $\operatorname{\operatorname{Ker}}e$, the strictly upper-diagonal entries vary independently in $\mathbb{D}$ and the strictly lower-diagonal entries are the opposites of the conjugates of the strictly upper-diagonal ones under $\sigma$.
Next, denote by $X$ the first vector of the standard basis of $\mathbb{D}^{n}$. Then we observe that $\mathcal{M}X=(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-1}$ and hence $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$. Moreover if $n\geq 2$ and we take $X^{\prime}$ as the second vector of the standard basis we find $\mathcal{M}X^{\prime}=\mathbb{D}\times(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-2}$, and it follows that $\mathcal{S}x+\mathcal{S}x^{\prime}=V$ for some $x^{\prime}\in U$. If $d>1$ or $n>1$ we have collected enough information to see that $\mathcal{S}$ is target-reduced, and then we conclude from Lemma 5.3 that $\operatorname{\operatorname{Alt}}(\mathcal{S})=\mathbb{F}b$. If $d=1$ and $n=1$ then $\mathcal{S}=\{0\}$ and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$ (it is the space of all $\mathbb{F}$-bilinear forms on $U\times V$).
Assume finally that $U=V$. By definition, $B$ is $e$-nonisotropic if and only if $b$ is nonisotropic. We already know that $\mathcal{A}_{b,\mathbb{D}}$ has trivial spectrum if $b$ is nonisotropic.
Assume for the remainder of the proof that $\mathcal{A}_{b,\mathbb{D}}$ has trivial spectrum.
  • Let $x\in V\smallsetminus\{0\}$. Then as $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ and $\mathcal{S}x$ is included in the right-orthogonal of $x$ under $b$, it equals this right-orthogonal, which yields $b(x,x)\neq 0$ since $x\not\in\mathcal{S}x$. Hence $b$ is nonisotropic.
  • Let us finally prove that $\mathcal{S}$ is irreducible. Let $x\in V\smallsetminus\{0\}$. Combining $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ with $x\not\in\mathcal{S}x$, we find $\mathbb{F}x\oplus\mathcal{S}x=V$. Therefore, $\mathcal{S}$ is irreducible.
6 Trivial spectrum subspaces
In this section, we will quickly derive the main results on irreducible optimal trivial spectrum subspaces from their analogues on deeply intransitive operator spaces.
6.1 The Invariant Subspace Lemma
We start with a simple lemma, which will allow us to deal with the case of non-primitively intransitive spaces.
Lemma 6.1 (Invariant Subspace Lemma).
Let $\mathcal{S}$ be an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Assume that there exists a $\mathbb{D}$-linear subspace $W$ of $V$ such that $\mathcal{S}$ contains every operator $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that satisfies $\operatorname{\operatorname{Im}}u\subseteq W\subseteq\operatorname{\operatorname{Ker}}u$. Then $W$ is invariant under $\mathcal{S}$.
Proof.
Let $u\in\mathcal{S}$. Assume that there is a (nonzero) vector $x\in W$ such that $u(x)\not\in W$. The vector u(x)xu(x)-x is not in $W$, and hence it belongs to a direct summand $W^{\prime}$ of $W$ in $V$ (i.e., a $\mathbb{D}$-linear subspace of $V$ such that $V=W\oplus W^{\prime}$). Denote by $\pi$ the projection of $V$ onto $W$ along $W^{\prime}$, and by $v\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ the linear mapping that vanishes on $W$ and satisfies $v(y)=\pi(u(y))$ for all $y\in W^{\prime}$. Then $v\in\mathcal{S}$ by our assumptions, and hence $u-v\in\mathcal{S}$. Note that uvu-v leaves $W^{\prime}$ invariant. Moreover $(u-v)(x)=u(x)\equiv x$ mod $W^{\prime}$. The endomorphism of $V/W^{\prime}$ induced by uvu-v then has a nonzero fixed vector, which contradicts the assumption that it should have $\mathbb{F}$-trivial spectrum. Therefore $W$ is invariant under $\mathcal{S}$. ∎
6.2 The partial structure theorem for optimal spaces
Now, we shall complete the proof of Theorem 1.7. The case $n=1$ is obvious. So, we let $n\geq 2$ be an integer, $V$ be an $n$-dimensional $\mathbb{D}$-vector space, and $\mathcal{S}$ be an irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$. We also assume that $|\mathbb{F}|\geq nd$.
We immediately note that $\mathcal{S}$ is intransitive.
Case 1. The space $\mathcal{S}$ is primitively intransitive.
Then $\mathcal{S}$ is deeply intransitive, by Proposition 4.5.
  • Assume first that $(n,d)\neq(2,1)$. Then $n\geq\max(4-d,2)$, and hence $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$. By point (c) of Theorem 1.20, we find that $\mathcal{S}$ has a nondegenerate alternator $b$.
  • Assume now that $(n,d)=(2,1)$. Then we simply note that $\dim\mathcal{S}=1$, so we can write $\mathcal{S}=\mathbb{F}u_{0}$ for some operator $u_{0}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$; then $u_{0}$ is invertible because of the irreducibility assumption, and hence by picking an arbitrary symplectic form $s$ on the $\mathbb{F}$-vector space $V$ (which exists because $n=2$ and $d=1$), we see that $b:(x,y)\mapsto s(x,u_{0}^{-1}(y))$ is a nondegenerate alternator of $\mathcal{S}$.
In any case, we have found a nondegenerate alternator $b$ of $\mathcal{S}$, to the effect that $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$. Yet $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive (by point (b) of Proposition 1.18), and hence $\dim\mathcal{A}_{b,\mathbb{D}}\leq\alpha(n,d)$ by point (a) of Theorem 1.20. Therefore $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$.
Case 2. The space $\mathcal{S}$ is not primitively intransitive.
Then we take a maximal $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the projected space $\pi\mathcal{S}$ is intransitive. Note that $\{0\}\subset V^{\prime}\subset V$. By the maximality of $V^{\prime}$, the space $\pi\mathcal{S}$ is primitively intransitive, and hence deeply intransitive (see Proposition 4.5). Set then $t:=\dim_{\mathbb{D}}V^{\prime}$. Applying point (a) of Theorem 1.20, we find $\dim(\pi\mathcal{S})\leq\alpha(n-t,d)$. Now, we look at the space $\mathcal{S}^{V^{\prime}}$. First of all, the rank theorem yields
$\dim(\mathcal{S}^{V^{\prime}})=\dim(\mathcal{S})-\dim(\pi\mathcal{S})\geq\alpha(n,d)-\alpha(n-t,d)=dt(n-t)+\alpha(t,d).$
Assigning to every $f\in\mathcal{S}^{V^{\prime}}$ its restriction $f_{V^{\prime}}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$, we create an $\mathbb{F}$-linear mapping $\Phi$ whose range is an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$ and whose kernel is the subspace $\mathcal{D}$ of all $f\in\mathcal{S}^{V^{\prime}}$ that vanish on $V^{\prime}$. By Theorem 3.1 we have $\operatorname{\operatorname{rk}}\Phi\leq\alpha(t,d)$, whence the rank theorem yields
$\dim\mathcal{D}\geq\dim(\mathcal{S}^{V^{\prime}})-\alpha(t,d)\geq dt(n-t).$
Yet $\mathcal{D}$ is included in the space of all $\mathbb{D}$-endomorphisms of $V$ that vanish on $V^{\prime}$ and map into $V^{\prime}$, a space which has dimension dt(nt)dt(n-t) over $\mathbb{F}$. Hence $\mathcal{D}$ equals that space, to the effect that all such endomorphisms belong to $\mathcal{S}$.
Then the Invariant Subspace Lemma yields that $V^{\prime}$ is invariant under $\mathcal{S}$. However this contradicts the assumption that $\mathcal{S}$ is irreducible, and we conclude that Case 2 is actually impossible.
Hence, Theorem 1.7 is now proved.
6.3 The full structure theorem for optimal spaces
Now, we turn to Theorem 1.8.
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$ such that $\mathcal{S}:=\mathcal{A}_{b,\mathbb{D}}$ is an optimal irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
We shall prove that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$ in any case.
  • Assume first that $n\geq\max(4-d,2)$. Then $\dim\mathcal{S}=\alpha(n,d)\geq\alpha(n,d)-n+\max(4-d,2)$ and the claimed statement is deduced from point (c) of Theorem 1.20.
  • Assume now that $n=2$ and $d=1$. Then $\dim\mathcal{S}=1$, and we take $u_{0}\in\mathcal{S}\smallsetminus\{0\}$. Since $\mathcal{S}$ is irreducible, $u_{0}$ must be invertible, and we deduce that $\dim\mathcal{S}x=1=nd-1$ for all $x\in V\smallsetminus\{0\}$. Moreover $\mathcal{S}$ is target-reduced because $u_{0}$ is surjective. Hence Lemma 5.3 shows that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$ because $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})$.
  • If $n=1$ and $d>1$ then $\dim_{\mathbb{F}}(\mathcal{S}x)=\dim\mathcal{S}=nd-1$ for all $x\in V\smallsetminus\{0\}$, and obviously $\mathcal{S}$ is target-reduced, so once more we find by Proposition 5.3 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
  • If $n=1$ and $d=1$ then $\mathcal{S}=\{0\}$ and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is the space of all $\mathbb{F}$-bilinear forms on $V$, which has dimension $1$.
Then, we apply Proposition 5.2 and find that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and that, for an associated pair $(\sigma,e)$, there exists a $\sigma$-sesquilinear form $B:V^{2}\rightarrow\mathbb{D}$ such that $b:(x,y)\mapsto e(B(x,y))$. Besides, it has already been observed in Section 2.1 that $B$ is uniquely determined by $b$.
Finally, the converse statement in Theorem 1.8 has already been proved through Proposition 5.4.
6.4 The quadratic type case
We finish this section by proving the results that are specific to the division algebras of quadratic type: Theorems 1.10 and 1.11.
So, assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and denote by $(\sigma,e)$ an associated pair. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $n\geq 1$. Let $\mathcal{S}$ be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ such that $|\mathbb{F}|\geq nd$.
If $n\geq 2$ then we apply Theorems 1.7 and 1.8 to find a $\sigma$-sesquilinear nondegenerate form $B:V^{2}\rightarrow\mathbb{D}$ such that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$ for $b:(x,y)\mapsto e(B(x,y))$.
Assume now that $n=1$. In that case we need only consider the special case where $V=\mathbb{D}$. By the isomorphism $a\in\mathbb{D}\mapsto[x\mapsto ax]\in\operatorname{\operatorname{End}}_{\mathbb{D}}(\mathbb{D})$ of $\mathbb{F}$-vector spaces, we associate to $\mathcal{S}$ an $\mathbb{F}$-linear hyperplane $H$ of $\mathbb{D}$. Thus $H$ is the kernel of $x\mapsto e(\alpha x)$ for some $\alpha\in\mathbb{D}^{\times}$. Then, for all $a\in H$, we have
$\forall x\in\mathbb{D},\;e(x^{\sigma}\alpha(ax))=e(\alpha a\underbrace{xx^{\sigma}}_{\in\mathbb{F}})=(xx^{\sigma})\,e(\alpha a)=0$
This shows that the nondegenerate bilinear form $b:(x,y)\in\mathbb{D}^{2}\mapsto e(x^{\sigma}\alpha y)$ is an alternator of $\mathcal{S}$. Hence $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$. Yet $\dim\mathcal{A}_{b,\mathbb{D}}\leq d-1$ because $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive, and we conclude that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$. Note here that $b:(x,y)\mapsto e(B(x,y))$ for the $\sigma$-sesquilinear form $B:(x,y)\mapsto x^{\sigma}\alpha y$. By point (c) of Proposition 5.4, we see that $b$ is nonisotropic.
From here, we deduce from point (b) of Proposition 5.4 that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$: it follows that for every nondegenerate bilinear form $b^{\prime}$ on $V$ such that $\mathcal{S}=\mathcal{A}_{b^{\prime},\mathbb{D}}$, we have $b^{\prime}\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ and hence $b^{\prime}\in\mathbb{F}b$. This completes the proof of Theorem 1.10.
Let us finally derive Theorem 1.11. So, let $B$ and $B^{\prime}$ be nondegenerate $\sigma$-sesquilinear forms on $V$. Let $\varphi\in\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$ and set $b:(x,y)\in V^{2}\mapsto e(B(x,y))$ and $b^{\prime}:(x,y)\in V^{2}\mapsto e(B^{\prime}(x,y))$.
Then, for all $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$,
$\varphi\circ u\circ\varphi^{-1}\in\mathcal{A}_{b}$ $\Leftrightarrow\forall x\in V,\;e(B(x,\varphi(u(\varphi^{-1}(x)))=0$
$\Leftrightarrow\forall y\in V,\;e(B(\varphi(y),\varphi(u(y)))=0$
$\Leftrightarrow u\in\mathcal{A}_{b^{\prime\prime}}$
where $b^{\prime\prime}:(x,y)\mapsto e(B^{\prime\prime}(x,y))$ for the $\sigma$-sesquilinear form $B^{\prime\prime}:(x,y)\mapsto B(\varphi(x),\varphi(y))$. Hence $\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is the conjugate of $\mathcal{A}_{b,\mathbb{D}}$ under $\varphi^{-1}$, and more generally $\mathcal{A}_{\alpha b^{\prime\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is conjugated to $\mathcal{A}_{b,\mathbb{D}}$ for all $\alpha\in\mathbb{F}^{\times}$.
Conversely, assume that $\mathcal{A}_{b,\mathbb{D}}$ is conjugated to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ under an element of $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$. Then by the above $\mathcal{A}_{b^{\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ for some $b^{\prime\prime}$ of the form $(x,y)\mapsto e(B^{\prime\prime}(x,y))$, where $B^{\prime\prime}$ is equivalent to $B$. The uniqueness statement in Theorem 1.10 yields $b^{\prime}=\alpha b^{\prime\prime}$ for some $\alpha\in\mathbb{F}^{\times}$, and hence $B^{\prime}=\alpha B^{\prime\prime}$. Hence $B^{\prime}$ is equivalent to $\alpha B$ for some $\alpha\in\mathbb{F}^{\times}$. This completes the proof of Theorem 1.11.
7 Semiaffine spaces of matrices with rank bounded below
In this short section, we apply the results on trivial spectrum subspaces to obtain significant ones on $\mathbb{F}$-affine subspaces of matrices with rank bounded below. This uses a strategy devised by Meshulam [10].
7.1 Semiaffine spaces of matrices with rank bounded below
Theorem 7.1.
Let $n,p,r$ be integers with $\min(n,p)\geq r>0$. Let $\mathcal{S}$ be an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which every element has rank at least $r$. Assume that $|\mathbb{F}|\geq rd$. Then
$\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq r+d\,\dbinom{r}{2}$,
where the codimension is to be understood as a codimension over $\mathbb{F}$. Moreover, this inequality is optimal, and equality holds only if $\mathcal{S}$ contains a rank $r$ matrix.
The proof requires the following weak version of Theorem 1.6:
Proposition 7.2.
Let $r\geq 1$ be such that $|\mathbb{F}|\geq rd$, and let $n\geq r$. Let $S$ be an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq r+d\dbinom{r}{2}$, with equality attained only if $n=r$.
For large values of $n$ with respect to $r$, this inequality can be considerably improved with limited effort, but we will not need such refinements.
Proof of Proposition 7.2.
We prove the result by induction on $n$. The case $n=r$ is a reformulation of Theorem 1.6. Assume now that $n>r$. Denote by $x$ the first vector of the standard basis of $\mathbb{D}^{n}$. Then $\dim_{\mathbb{F}}(Sx)\leq nd-1$ because $S$ is intransitive. Denote by $S^{\prime}$ the set of all matrices $N\in\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$ such that $\begin{bmatrix}0&[0]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&N\end{bmatrix}$ belongs to $S$, and note that $S^{\prime}$ is an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$. By the rank theorem
$\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq\operatorname{\operatorname{codim}}_{\mathbb{D}^{n}}(Sx)+\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})}S^{\prime},$
and by induction $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S^{\prime}\geq r+d\dbinom{r}{2}$. Hence $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq 1+r+d\dbinom{r}{2}$, which completes the proof. ∎
We can now prove Theorem 7.1.
Proof of Theorem 7.1.
The optimality of the inequality is exemplified by taking the space of all matrices of the form
$\begin{bmatrix}I_{r}+M&[?]_{r\times(p-r)}\\ [?]_{(n-r)\times r}&[?]_{(n-r)\times(p-r)}\end{bmatrix}$
where $M$ ranges over an $\mathbb{F}$-trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{r}(\mathbb{F})$ with dimension $\alpha(r,d)$ (say, one that is described on page 1.1). It is clear that the codimension of this space in $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ is the one of $\mathcal{M}$ in $\operatorname{\operatorname{M}}_{r}(\mathbb{D})$, that is $r+d\,\dbinom{r}{2}$. Every matrix in that space has its upper-left $r\times r$ block invertible, and hence its rank is at least $r$.
Now, denote by $s$ the least possible rank in $\mathcal{S}$ (at this point it may very well be that $s>r$). Replacing $\mathcal{S}$ with an equivalent space if necessary, we can then assume that $\mathcal{S}$ contains $J_{s}=\begin{bmatrix}I_{s}&[0]_{s\times(p-s)}\\ [0]_{(n-s)\times s}&[0]_{(n-s)\times(p-s)}\end{bmatrix}$. Denote by $\overrightarrow{\mathcal{S}}$ the translation vector space of $\mathcal{S}$, and write every matrix $M$ of it as $M=\begin{bmatrix}A(M)&C(M)\\ B(M)&D(M)\end{bmatrix}$ with $A(M)\in\operatorname{\operatorname{M}}_{s}(\mathbb{F})$ and so on. Consider the $\mathbb{F}$-linear subspace $W$ of $\overrightarrow{\mathcal{S}}$ consisting of all the matrices $M\in\overrightarrow{\mathcal{S}}$ such that $B(M)=0$, $C(M)=0$ and $D(M)=0$. Hence $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{s}(\mathbb{D})}A(W)$. For all $M\in W$, the matrix
$J_{s}+M=\begin{bmatrix}I_{s}+A(M)&[0]_{s\times(p-s)}\\ [0]_{(n-s)\times s}&[0]_{(n-s)\times(p-s)}\end{bmatrix}$
has rank at least $s$ because it belongs to $\mathcal{S}$, and hence $I_{s}+A(M)$ is invertible. It follows that $A(W)$ is an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{s}(\mathbb{D})$. Since $|\mathbb{F}|\geq rd\geq sd$, we deduce from Proposition 7.2 that $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{s}(\mathbb{D})}A(W)\geq r+d\dbinom{r}{2}$, with strict equality if $s>r$, and we conclude that $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq r+d\dbinom{r}{2}$, with strict inequality if $s>r$. This proves all the claimed statements. ∎
Remark 7.1.
If we had used Theorem 1.6 instead of Proposition 7.2, we would have needed the inequality $|\mathbb{F}|\geq d\max(n,p)$ because we have no prior control on the least possible rank in $\mathcal{S}$. Clearly, this requirement is not very satisfying, hence our insistence in requiring only that $|\mathbb{F}|\geq rd$.
7.2 Optimal semiaffine spaces of nonsingular matrices
At this point, it is natural that we wish to obtain a description of the optimal $\mathbb{F}$-affine subspaces of matrices with rank at least $r$, meaning the ones with the minimal possible codimension in $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$. Judging from the length of [16], which covers only the case $\mathbb{D}=\mathbb{F}$, it is safer to leave this problem for future work. Nevertheless, we will tackle here the limit case $n=p=r$ as it is can be obtained with few technical difficulties from our results on trivial spectrum linear subspaces.
Let us start with some terminology and notation.
Definition 7.1.
Two subsets $\mathcal{X}$ and $\mathcal{Y}$ of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ are called equivalent , and we write $\mathcal{X}\sim\mathcal{Y}$, when there exists a pair $(P,Q)\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})\times\operatorname{\operatorname{GL}}_{p}(\mathbb{D})$ such that $\mathcal{Y}=P\mathcal{X}Q$, i.e., when $\mathcal{X}$ and $\mathcal{Y}$ represent the same set of linear mappings from a $p$-dimensional $\mathbb{D}$-vector space to an $n$-dimensional $\mathbb{D}$-vector space, in a different choice of bases of these spaces.
Theorem 7.3.
Let $n\geq 1$. Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$, and that $|\mathbb{F}|\geq nd$. Then, up to equivalence, there is exactly one optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. One such space is the $n$-fold joint $(1_{\mathbb{D}}+H)\vee\cdots\vee(1_{\mathbb{D}}+H)$, where $H$ is an arbitrary $\mathbb{F}$-linear hyperplane of $\mathbb{D}$ that does not contain $1_{\mathbb{D}}$.
Proof.
First of all it is clear that the $n$-fold joint $(1_{\mathbb{D}}+H)\vee\cdots\vee(1_{\mathbb{D}}+H)$ is an $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, with dimension $n(d-1)+d\dbinom{n}{2}$.
[…]
Finally, $n$-fold joints are clearly compatible with equivalence, to the effect that for any list $\mathcal{H}_{1},\dots,\mathcal{H}_{n}$ of $\mathbb{F}$-affine hyperplanes of $\mathbb{D}$ that do not contain $0_{\mathbb{D}}$, we have $\mathcal{M}\sim\mathcal{H}_{1}\vee\cdots\vee\mathcal{H}_{n}$. The claimed result ensues. ∎
The situation is both more complicated and more interesting when $\mathbb{D}$ has quadratic type over $\mathbb{F}$.
Theorem 7.4 (Classification of optimal $\mathbb{F}$-affine subspaces of invertible matrices).
Let $n\geq 1$. Assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, with $|\mathbb{F}|\geq nd$. Let $(\sigma,e)$ be an associated pair.
Let $\mathcal{M}$ be an optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then there exists a partition $n=n_{1}+\cdots+n_{p}$ and a list $(P_{1},\dots,P_{p})$ of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$, such that
$\mathcal{M}\simeq(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D})).$
Conversely, let $n=n_{1}+\cdots+n_{p}$ be a partition, and $(P_{1},\dots,P_{p})$ be a list of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$. Then $(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D}))\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
Finally, let $m=m_{1}+\cdots+m_{q}$ be a partition, and $(P^{\prime}_{1},\dots,P^{\prime}_{q})$ be a list of $e$-nonisotropic matrices, with $P^{\prime}_{i}\in\operatorname{\operatorname{GL}}_{m_{i}}(\mathbb{D})$. Then $(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D}))\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is equivalent to $(P^{\prime}_{1}+\mathcal{S}\mathcal{H}_{m_{1}}(\mathbb{D}))\vee\cdots\vee(P^{\prime}_{q}+\mathcal{S}\mathcal{H}_{m_{q}}(\mathbb{D}))$ if and only if $(n_{1},\dots,n_{p})=(m_{1},\dots,m_{q})$ and, for all $k\in\mathopen{[\![}1,p\mathclose{]\!]}$, there exists an invertible matrix $Q_{k}\in\operatorname{\operatorname{GL}}_{n_{k}}(\mathbb{D})$ and a scalar $\alpha\in\mathbb{F}^{\times}$ such that $\forall X\in\mathbb{D}^{n_{k}},\;e(X^{\star}P^{\prime}_{k}X)=\alpha\,e((Q_{k}X)^{\star}P_{k}(Q_{k}X))$.
The proof involves the following lemma:
Lemma 7.5.
Let $\mathcal{S}$ be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, where $V$ is an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Then $\mathbb{F}x+\mathcal{S}x=V$ for all $x\in V\smallsetminus\{0\}$.
Proof.
We combine Theorems 1.8 and 1.10 to find that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$ for some nondegenerate $\mathbb{F}$-bilinear form $b$ of type $(x,y)\mapsto e(B(x,y))$, where $B$ is a nondegenerate $\sigma$-sesquilinear form on $V$. Let $x\in V\smallsetminus\{0\}$. By point (b) of Proposition 5.4 we find $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$. The conclusion is obtained by noting that $x\not\in\mathcal{S}x$ because $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum. ∎
Proof of Theorem 7.4.
To simplify the writing of the proof, we set $\mathcal{M}_{k}:=\mathcal{S}\mathcal{H}_{k}(\mathbb{D})$ for every integer $k\geq 1$.
[…]
$\overrightarrow{\mathcal{M}}\simeq(P_{1}^{-1}\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}^{-1}\mathcal{M}_{n_{p}}).$
Hence
$\mathcal{M}\simeq I_{n}+\bigl{(}(P_{1}^{-1}\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}^{-1}\mathcal{M}_{n_{p}})\bigr{)}=(I_{n_{1}}+P_{1}^{-1}\mathcal{M}_{n_{1}})\vee\cdots\vee(I_{n_{p}}+P_{p}^{-1}\mathcal{M}_{n_{p}}),$
leading to
$\mathcal{M}\sim(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}}).$
The second statement is obvious because, for every $k\in\mathopen{[\![}1,p\mathclose{]\!]}$ and every $e$-nonisotropic $P\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ we have $P+\mathcal{M}_{k}\sim I_{k}+P^{-1}\mathcal{M}_{k}$, and $P^{-1}\mathcal{M}_{k}$ has $\mathbb{F}$-trivial spectrum and dimension $\alpha(k,d)$ over $\mathbb{F}$; note also, for the remainder of the proof, that $\operatorname{\operatorname{span}}_{\mathbb{F}}(P+\mathcal{M}_{k})X=\mathbb{D}^{k}$ for all $X\in\mathbb{D}^{k}\smallsetminus\{0\}$, as a consequence of Lemma 7.5.
It remains to prove the uniqueness statement. It is better to think in geometrical terms here. So, let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, and let $\mathcal{S}$ be an optimal $\mathbb{F}$-affine subspace of automorphisms of $V$. Assume that we have bases $(e_{1},\dots,e_{n})$ and $(f_{1},\dots,f_{n})$ of $V$ such that the matrices that represent the elements of $\mathcal{S}$ form the space $(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})$. We consider the integers $N_{i}:=\sum_{k=1}^{i}n_{k}$ and the spaces $U_{i}:=\operatorname{\operatorname{span}}(e_{1},\dots,e_{N_{i}})$ and $V_{i}:=\operatorname{\operatorname{span}}(f_{1},\dots,f_{N_{i}})$ for $i\in\mathopen{[\![}0,p\mathclose{]\!]}$. We shall prove that any other choice of bases will yield the same spaces. To see this, we look at the dimension $\dim_{\mathbb{F}}(\mathcal{S}x)$ when $x$ ranges over $V$. Using the previous remark, we see that for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ and all $x\in U_{i}\smallsetminus U_{i-1}$, we have $\operatorname{\operatorname{span}}_{\mathbb{F}}(\mathcal{S}x)=V_{i}$. Hence the spaces $V_{0},\dots,V_{p}$ are uniquely determined by $\mathcal{S}$ only, as the possible $\mathcal{S}x$ spaces when $x$ ranges over $V$; and then $U_{i}=\{x\in V:\;\mathcal{S}x\subseteq V_{i}\}$ is also uniquely determined by $\mathcal{S}$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$.
As a consequence, if we have invertible matrices $R,S$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that
$(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})=R\left((P^{\prime}_{1}+\mathcal{M}_{n^{\prime}_{1}})\vee\cdots\vee(P^{\prime}_{q}+\mathcal{M}_{n^{\prime}_{q}})\right)S$
then $(n_{1},\dots,n_{p})=(n^{\prime}_{1},\dots,n^{\prime}_{q})$ and both matrices $R$ and $S$ are block upper-triangular with respect to the partition $(n_{1},\dots,n_{p})$. Their respective diagonal blocks are invertible, and it follows that
$\forall i\in\mathopen{[\![}1,p\mathclose{]\!]},\;P_{i}+\mathcal{M}_{n_{i}}\sim P^{\prime}_{i}+\mathcal{M}_{n_{i}}.$
Conversely, it is obvious that if $(n_{1},\dots,n_{p})=(n^{\prime}_{1},\dots,n^{\prime}_{q})$ and $\forall i\in\mathopen{[\![}1,p\mathclose{]\!]},\;P_{i}+\mathcal{M}_{n_{i}}\sim P^{\prime}_{i}+\mathcal{M}_{n_{i}}$ then $(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})\sim(P^{\prime}_{1}+\mathcal{M}_{n^{\prime}_{1}})\vee\cdots\vee(P^{\prime}_{q}+\mathcal{M}_{n^{\prime}_{q}})$.
In order to conclude, we go back to an arbitrary integer $k\in\mathopen{[\![}1,n\mathclose{]\!]}$, we take two $e$-nonisotropic matrices $P$ and $Q$ of $\operatorname{\operatorname{M}}_{k}(\mathbb{D})$, and we characterize the equivalence $P+\mathcal{M}_{k}\sim Q+\mathcal{M}_{k}$. First of all, note that $R^{\star}\mathcal{M}_{k}R=\mathcal{M}_{k}$ for all $R\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ (which easily follows from Proposition 1.12). Now, say that we have $R,S$ in $\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ such that $P+\mathcal{M}_{k}=R(Q+\mathcal{M}_{k})S$. Then $P+\mathcal{M}_{k}=RQS+R\mathcal{M}_{k}S$. Comparing the translation vector spaces yields $R\mathcal{M}_{k}S=\mathcal{M}_{k}$ i.e., $R(S^{\star})^{-1}(S^{\star}\mathcal{M}_{k}S)=\mathcal{M}_{k}$, leading to $R(S^{\star})^{-1}\mathcal{M}_{k}=\mathcal{M}_{k}$. We observe that $(X,Y)\mapsto e(X^{\star}S^{\star}R^{-1}Y)$ is a nondegenerate alternator of $R(S^{\star})^{-1}\mathcal{M}_{k}$, that $(X,Y)\mapsto e(X^{\star}Y)$ is a nonzero alternator of $\mathcal{M}_{k}$, and finally that $\dim_{\mathbb{F}}(\mathcal{M}_{k}X)=kd-1$ for all $X\in\mathbb{D}^{k}\smallsetminus\{0\}$. Hence by Proposition 5.3 those alternators are $\mathbb{F}$-linearly dependent, yielding $S^{\star}R^{-1}=\alpha I_{k}$ for some $\alpha\in\mathbb{F}^{\times}$. Hence $P+\mathcal{M}_{k}=P^{\prime}+\mathcal{M}_{k}$ for $P^{\prime}:=\alpha^{-1}S^{\star}QS$. Finally, $P=P^{\prime}+M$ for some $M\in\mathcal{M}_{k}$, which yields $\forall X\in\mathbb{D}^{n},\;e(X^{\star}PX)=e(X^{\star}P^{\prime}X)=\alpha^{-1}\,e((SX)^{\star}Q(SX))$.
Conversely. assume that $\forall X\in\mathbb{D}^{k},\;e(X^{\star}PX)=\alpha\,e((SX)^{\star}Q(SX))$ for some $\alpha\in\mathbb{F}^{\times}$ and some $S\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$. Then $P-\alpha S^{\star}QS$ belongs to $\mathcal{M}_{k}$, whence
$P+\mathcal{M}_{k}=\alpha S^{\star}QS+\mathcal{M}_{k}\sim Q+(S^{\star})^{-1}\mathcal{M}_{k}S^{-1}=Q+\mathcal{M}_{k}.$
This completes our proof. ∎
It remains to understand the nature of the mappings $X\in\mathbb{D}^{n}\mapsto e(X^{\star}PX)$ when $n\geq 1$ and $P$ ranges over the $e$-nonisotropic matrices of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$. In the special case $d=1$, they are simply the nonisotropic quadratic forms on $\mathbb{D}^{n}$. Hence, as seen in [18], the optimal affine subspaces of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are classified up to equivalence by lists of congruence classes of nonisotropic quadratic forms. In particular, if $\mathbb{F}=\mathbb{R}$ then there is only one such congruence class of $k$-dimensional forms for each $k\geq 1$, and we obtain that every optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ is equivalent to $(I_{n_{1}}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{R}))\vee\cdots\vee(I_{n_{p}}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{R}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely every space of this form is an optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$.
The situation is not as clear if $d>1$. But say that $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$, and keep the assumption that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$. Then $d\in\{2,4\}$. Let $P\in\operatorname{\operatorname{M}}_{n}(\mathbb{F})$. Then there is a unique splitting $P=Q+R$ where $Q$ is Hermitian (that is $Q^{\star}=Q$) and $R$ is skew-Hermitian. We can then replace $P+\mathcal{M}_{n}$ with $Q+\mathcal{M}_{n}$, and now $X\mapsto X^{\star}QX$ takes all its values in $\mathbb{F}$, so $\forall X\in\mathbb{D}^{n},\;e(X^{\star}QX)=2\,X^{\star}QX$. Moreover, it is known by polarizing that $X\mapsto X^{\star}QX$ determines $Q$. Hence the equivalence class of $P+\mathcal{M}_{n}$ determines $Q$ up to star-congruence and multiplication with an element of $\mathbb{F}^{\times}$. As a consequence, optimal affine subspaces of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are classified, up to equivalence, by lists $([Q_{1}],\dots,[Q_{p}])$ of classes of nonisotropic Hermitian matrices over $\mathbb{D}$, where each class is for the equivalence relation where $Q$ and $Q^{\prime}$ are equivalent if and only if $Q^{\prime}=\alpha R^{\star}QR$ for some invertible $R$ and some $\alpha\in\mathbb{F}^{\times}$. In particular, if $\mathbb{D}\in\{\mathbb{C},\mathbb{H}\}$ and $\mathbb{F}=\mathbb{R}$, then we recover from the classification of Hermitian forms that every optimal $\mathbb{R}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is equivalent to $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal $\mathbb{R}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for every partition $n=n_{1}+\cdots+n_{p}$.
For fields of characteristic $2$, Hermitian matrices are skew-Hermitian and hence there is no obvious way to simplify our vision on the classification of the forms $X\mapsto e(X^{\star}PX)$. We leave the problem for future work.
8 Linear subspaces of $\mathbb{F}$-diagonalisable matrices
In this final section, we give a series of applications of the structure theorem for optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
8.1 $\mathbb{F}$-diagonalisable subspaces
Let $V$ be a (finite-dimensional) vector space over $\mathbb{D}$. An endomorphism of the $\mathbb{D}$-vector space $V$ is called $\mathbb{F}$-diagonalisable when it is representable over $\mathbb{D}$ by a diagonal matrix whose diagonal entries all belong to $\mathbb{F}$, or equivalently when it is annihilated by a polynomial $p(t)\in\mathbb{F}[t]$ that is split and with simple roots. An $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is called $\mathbb{F}$-diagonalisable when all its elements are $\mathbb{F}$-diagonalisable. In some sense, being $\mathbb{F}$-diagonalisable is sort of the opposite of having $\mathbb{F}$-trivial spectrum: if an endomorphism $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is $\mathbb{F}$-diagonalisable and has $\mathbb{F}$-trivial spectrum, then it must be zero (because in a diagonalisation, the diagonal entries must equal zero). We define $\mathbb{F}$-diagonalisable matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ and $\mathbb{F}$-diagonalisable linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ accordingly.
Now we introduce three large spaces that are known to have this property. The space $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ of all real symmetric matrices; the space $\mathcal{H}_{n}(\mathbb{C})$ of all Hermitian matrices over the complex numbers; and finally the space $\mathcal{H}_{n}(\mathbb{H})$ of all quaternionic Hermitian matrices (see e.g. [5] for a proof that the latter is $\mathbb{R}$-diagonalisable). More generally, assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ and that $\mathbb{D}$ is not a hyper-radicial extension of $\mathbb{F}$. Consider the associated pair $(\sigma,e)$ (so here $e:=\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$) and set
$\mathcal{H}_{n}(\mathbb{D}):=\{M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D}):\;M^{\star}=M\},$
where $M^{\star}$ is defined with respect to $\sigma$. Then $\mathcal{H}_{n}(\mathbb{D})$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ and its dimension is $n+d\binom{n}{2}$. However, it might not be $\mathbb{F}$-diagonalisable (see the case of the symmetric matrices over the complex numbers). Nevertheless, we easily have the following result:
Theorem 8.1.
Let $\mathcal{V}$ be an $\mathbb{F}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then
$\dim\mathcal{V}\leq n+d\binom{n}{2}$.
Proof.
We know that $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ contains an $\mathbb{F}$-trivial spectrum linear subspace $\mathcal{S}$ with dimension $\alpha(n,d)$. By a previous remark we must have $\mathcal{V}\cap\mathcal{S}=\{0\}$ and hence
$\dim\mathcal{V}\leq\dim_{\mathbb{F}}\operatorname{\operatorname{M}}_{n}(\mathbb{D})-\dim\mathcal{S}=n+d\binom{n}{2}$.
In most cases, the inequality from Theorem 8.1 is very far from optimality. Indeed if $\mathbb{D}=\mathbb{F}=\mathbb{C}$ it is known by the Motzkin-Taussky theorem [11] that the greatest possible dimension for a $\mathbb{C}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ is $n$ (every such subspace is conjugated to the space of all diagonal matrices), and strangely this holds for finite fields as well. It would be amazing if the latter result has not been found earlier, but we were unable to find it in the literature on the topic, so we include a quick proof.
Theorem 8.2 (Motzkin-Taussky theorem for finite fields).
Assume that $\mathbb{F}$ is finite, let $V$ be a finite-dimensional $\mathbb{F}$-vector space, and $u$ and $v$ be endomorphisms of $V$. Assume that every linear combination of $u$ and $v$ with coefficients in $\mathbb{F}$ is diagonalisable. Then $u$ and $v$ commute.
Proof.
Denote by $q$ the cardinality of $\mathbb{F}$. An endomorphism of $V$ is diagonalisable if and only if it is annihilated by the polynomial $P(t):=\prod_{\lambda\in\mathbb{F}}(t-\lambda)=t^{q}-t$. Hence $(\lambda u+v)^{q}=\lambda u+v$ for all $\lambda\in\mathbb{F}$. Yet there is a list $(a_{1},\dots,a_{p-1})$ of endomorphisms of $V$ such that $(\lambda u+v)^{q}=(\lambda u)^{q}+v^{q}+\sum_{k=1}^{q-1}\lambda^{k}a_{k}$ for all $\lambda\in\mathbb{F}$, where in particular $a_{1}=\sum_{k=0}^{q-1}v^{k}uv^{q-1-k}$. Note that $\forall\lambda\in\mathbb{F},\;(\lambda u)^{q}=\lambda u^{q}$. Hence $\sum_{k=1}^{q-1}\lambda^{k}a_{k}=0$ for all $\lambda\in\mathbb{F}$. Since $|\mathbb{F}|=q$ all the $a_{k}$’s vanish and in particular $a_{1}=0$. As a consequence, $\operatorname{\operatorname{Ker}}v$ is invariant under $u$: indeed for all $x\in\operatorname{\operatorname{Ker}}v$ we obtain $0=a_{1}(x)=v^{q-1}(u(x))$, and since $v$ is diagonalisable we classically have $\operatorname{\operatorname{Ker}}v^{q-1}=\operatorname{\operatorname{Ker}}v$, whence $v(u(x))=0$.
Now, we note that the assumptions are preserved in replacing $v$ with $v-\alpha\operatorname{\operatorname{id}}_{V}$ for an arbitrary $\alpha\in\mathbb{F}$, so the previous step shows that $u$ leaves all the eigenspaces of $v$ invariant, to the effect that $uv$ and $vu$ coincide on all these eigenspaces. Since $v$ is diagonalisable, it follows that $uv=vu$. ∎
Of course, in case $d=1$ the bound $n$ that is obtained for the field of complex numbers (and more generally for any algebraically closed fields of characteristic 0) and for finite fields is very far from the bound $\dbinom{n+1}{2}$ given in Theorem 8.1.
Here, we will consider the $\mathbb{F}$-diagonalisable subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ that have dimension $n+d\dbinom{n}{2}$. The case $n=1$ is trivial, so we only consider greater values of $n$.
8.2 Main results
We recall that the special case $d=1$ is already known (see [3, 13, 23] for various proofs):
Theorem 8.3.
Let $n\geq 2$. Let $\mathcal{S}$ be an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ such that $\dim\mathcal{S}=n+\binom{n}{2}$. Then $\mathbb{F}$ is formally real and $\mathcal{S}$ is similar to $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$.
To state the generalization, we need an additional notion:
Definition 8.1.
Assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and denote by $x\mapsto x^{\star}$ its standard involution over $\mathbb{F}$. We say that $x\in\mathbb{D}\mapsto xx^{\star}$, which is valued in $\mathbb{F}$, is the standard norm of $\mathbb{D}$ over $\mathbb{F}$.
Here is the generalization we are aiming at:
Theorem 8.4.
Let $n\geq 2$. Let $\mathcal{S}$ be an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ such that $\dim\mathcal{S}=n+d\binom{n}{2}$. Then $\mathbb{F}$ is formally real and Pythagorean [^6] The field $\mathbb{F}$ is Pythagorean when $\forall(x,y)\in\mathbb{F}^{2},\;\exists z\in\mathbb{F}:x^{2}+y^{2}=z^{2}$. (so it has characteristic 0), $\mathbb{D}$ has quadratic type over $\mathbb{F}$, its standard norm is square-valued in $\mathbb{F}$, and $\mathcal{S}$ is similar to $\mathcal{H}_{n}(\mathbb{D})$.
Hence, if there exists an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ such that $\dim\mathcal{S}=n+d\binom{n}{2}$, then $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and every Hermitian matrix of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable! Finally, we note that there is a known characterization of the fields $\mathbb{F}$ over which every symmetric matrix is diagonalisable [28] but there is no real hope of understanding, for a fixed $n\geq 3$, the fields $\mathbb{F}$ over which every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is diagonalisable. Nevertheless, if $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and characteristic 0, and if for some $n\geq 2$ every matrix of $\mathcal{H}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable, then so is every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$. Conversely, assume that, for some $n\geq 2$, every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is diagonalisable over $\mathbb{F}$, and that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and its standard norm is square-valued in $\mathbb{F}$. Then $\mathbb{F}$ is Pythagorean and formally real, and it is then easy to prove, by using the Gram-Schmidt orthonormalization process, that every matrix of $\mathcal{H}_{n}(\mathbb{D})$ is unitarily similar to a matrix of $\mathcal{H}_{n}(\mathbb{D})$ with entries in $\mathbb{F}$, i.e., to a matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$. Hence, $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is $\mathbb{F}$-diagonalisable if and only if $\mathcal{H}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable. As a consequence, we obtain:
Corollary 8.5.
Let $n\geq 2$. The following conditions are equivalent:
  1. (i)
    There exists an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\binom{n}{2}$.
  2. (ii)
    Every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is diagonalisable, $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ and its standard norm is square-valued in $\mathbb{F}$.
8.3 The orthogonality method
The strategy of proof of Theorem 8.4 is essentially similar to the one we used in [23]: We will make a direct connection between $\mathbb{F}$-diagonalisable subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\binom{n}{2}$ and $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. This connection is obtained thanks to a trace-orthogonality argument.
To obtain it, we consider the traditional trace $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$, defined as in Section 2.1. The trace of an element $a\in\mathbb{D}$ is the trace of the endomorphism $x\mapsto ax$ of the $\mathbb{F}$-vector space $\mathbb{D}$. It satisfies $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ab)=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ba)$ for all $(a,b)\in\mathbb{D}^{2}$. We will simply write $\operatorname{\operatorname{Tr}}$ for $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ in what follows.
One major difficulty is that $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}$, and in case $\operatorname{\operatorname{char}}(\mathbb{F})>0$ we might then have $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})=0$, and it might even hold that $\operatorname{\operatorname{Tr}}$ is the zero form.
In the remainder of this section, we assume that $\operatorname{\operatorname{char}}(\mathbb{F})$ does not divide $d$ (early on in our proof of Theorem 8.4, we will prove that $\operatorname{\operatorname{char}}(\mathbb{F})=0$ under specific assumptions, which will alleviate this difficulty).
The trace extends to square matrices because
$\forall(A,B)\in\operatorname{\operatorname{M}}_{n}(\mathbb{D})^{2},\;\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(AB))=\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(BA)),$
where $\operatorname{\operatorname{tr}}$ stands for the trace of matrices in $\mathbb{D}$ (i.e., the sum of the diagonal entries). Then
$\langle A,B\rangle:=\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(AB))$
defines a symmetric $\mathbb{F}$-bilinear form on $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and it satisfies the key identity
$\langle PAP^{-1},PBP^{-1}\rangle=\langle A,B\rangle.$
Thanks to $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})\neq 0$, this $\mathbb{F}$-bilinear form is nondegenerate. This way, we can define, on any $\mathbb{D}$-vector space $V$ of finite dimension, the $\mathbb{F}$-inner product of two $\mathbb{D}$-endomorphisms $u$ and $v$ of $V$, so that any matrix representation of $u$ and $v$ in a basis yields the same result.
Now, given an $\mathbb{F}$-linear subspace $\mathcal{V}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ that satisfies the conditions of Theorem 8.4, we shall consider the orthogonal complement $\mathcal{V}^{\bot}$ of $\mathcal{V}$ for the previous inner product: this gives an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $dn^{2}-n-d\binom{n}{2}=\alpha(n,d)$. And we will prove that $\mathcal{V}^{\bot}$ has $\mathbb{F}$-trivial spectrum!
There are two problems with this strategy: the first one is the assumption that $\operatorname{\operatorname{char}}(\mathbb{F})$ does not divide $d$, and the second one is the cardinality assumptions in the theorems that describe the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace and the structure of the optimal spaces. Fortunately, we will prove early on that $\operatorname{\operatorname{char}}(\mathbb{F})=0$, and in particular that $\mathbb{F}$ must be infinite (this could of course be immediately derived from Theorem 8.2). Hence, no major difficulty will be encountered. To obtain these results however, we must use an induction on $n$, where the case $n=2$ is critical. Before we go on, we prove a critical result that relates the properties of a space $\mathcal{V}$ to the ones of its orthogonal complement $\mathcal{V}^{\bot}$ when $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})\neq 0$. Then we will deal with the case $n=2$, and finally we will be able to prove Theorem 8.4 by using our results on trivial spectrum spaces.
8.4 Connecting trivial spectrality with a property on rank $1$ idempotents
Definition 8.2.
Let $V$ be a (finite-dimensional) $\mathbb{D}$-vector space. An $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have the full rank $1$ idempotent property whenever for every $1$-dimensional $\mathbb{D}$-linear subspace $E$ of $V$, the space $\mathcal{S}$ contains an idempotent endomorphism with range $E$.
Lemma 8.6.
Assume that $\operatorname{\operatorname{char}}(\mathbb{F})$ does not divide $d$. Let $V$ be a (finite-dimensional) $\mathbb{D}$-vector space. Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that has the full rank $1$ idempotent property. Then $\mathcal{S}^{\bot}$ has $\mathbb{F}$-trivial spectrum.
It can be shown that the converse holds true if $\mathbb{D}=\mathbb{F}$, but we will not need this result here.
Proof.
Let $u\in\mathcal{S}^{\bot}$. Assume that there is a nonzero vector $x\in V$ such that $u(x)=x$.
By assumption $\mathcal{S}$ contains an idempotent endomorphism $p$ with range $x\mathbb{D}$, and we have $\langle p,u\rangle=0$. For clarity, we extend $x$ to a basis $(x,e_{2},\dots,e_{n})$ of $V$, and we write the corresponding matrices of $u$ and $p$ as
$\begin{bmatrix}1_{\mathbb{D}}&[?]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&[?]_{(n-1)\times(n-1)}\end{bmatrix}\quad\text{and}\quad\begin{bmatrix}1_{\mathbb{D}}&[?]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&[0]_{(n-1)\times(n-1)}\end{bmatrix}$.
Then $\operatorname{\operatorname{tr}}(up)=\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}\neq 0$, which contradicts our assumptions. Hence $\mathcal{S}^{\bot}$ has $\mathbb{F}$-trivial spectrum. ∎
8.5 The case $n=2$
Lemma 8.7.
Assume that $\operatorname{\operatorname{M}}_{2}(\mathbb{D})$ contains an $\mathbb{F}$-diagonalisable subspace $\mathcal{V}$ with dimension at least $d+2$. Then $\mathbb{F}$ is formally real and Pythagorean.
Due to the potential non-commutativity of $\mathbb{D}$, the proof is surprisingly difficult because we are not allowed to use the trace and determinant to analyse $\mathbb{F}$-diagonalisable matrices.
Proof.
Without loss of generality, we assume that $I_{2}\in\mathcal{V}$ (as we can always replace $\mathcal{V}$ with $\mathcal{V}+\mathbb{F}I_{2}$).
We shall construct a linear form $t:\mathcal{V}\rightarrow\mathbb{F}$ and a quadratic form $q:\mathcal{V}\rightarrow\mathbb{F}$ such that
$\forall a\in\mathcal{V},\;a^{2}=t(a)\,a-q(a)\,I_{2}.$
If $\mathbb{D}$ is commutative it suffices to take for $t$ the trace, and for $q$ the determinant.
Assume now that $\mathbb{D}$ is noncommutative. Hence $\mathbb{F}$ is infinite by Wedderburn’s theorem. This way, we can use arithmetics on polynomial functions.
To start with, we take a linear hyperplane $H$ of $\mathcal{V}$ that does not contain $I_{2}$. For every $a\in H\smallsetminus\{0\}$, since $\mathcal{V}$ is $\mathbb{F}$-diagonalisable there are scalars $t(a)$ and $q(a)$ in $\mathbb{F}$ such that
$a^{2}=t(a)\,a-q(a)\,I_{2},$ (6)
and these scalars are uniquely determined by $a$. We also put $t(0)=0$ and $q(0)=0$. We shall see that $t(a)$ and $q(a)$ are homogeneous polynomial functions of $a$.
Clearly, as $I_{2}$ and $a$ are $\mathbb{F}$-linearly independent except for $a=0$, $t$ and $q$ are homogeneous rational functions on $H\smallsetminus\{0\}$ with respective degrees $1$ and $2$. We claim that they are actually polynomials. Split indeed $t=-\frac{r}{s}$ for mutually prime homogeneous polynomials $r$ and $s$, and assume that $s$ is nonconstant. Let $\varphi$ be an arbitrary nonzero $\mathbb{F}$-linear form on $\mathcal{V}$ that vanishes at $I_{2}$. Then we apply $\varphi$ to identity (6) to obtain $\forall a\in H,\;t(a)\,\varphi(a)=\varphi(a^{2})$.
Then $a\mapsto s(a)\,\varphi(a^{2})$ divides $r\varphi$, and hence $s$ divides $\varphi$. Because $\varphi$ has degree $1$ we deduce that $s=\lambda\varphi$ for some $\lambda\in\mathbb{F}^{\times}$. Hence, every $\mathbb{F}$-linear form on $\mathcal{V}$ that vanishes at $I_{2}$ must have its restriction to $H$ linearly dependent on $s$, to the effect that the space of all such linear forms has dimension (at most) $1$. Yet by orthogonality theory this space has dimension $\dim\mathcal{V}-1$, which is greater than $1$ because $\dim\mathcal{V}\geq d+2$. This contradiction shows that $s$ is constant.
Hence $t$ is actually a homogeneous polynomial function of degree $1$, and it follows that $q$ is a quadratic form on $H$. We can then extend $t$ and $q$ to $\mathcal{V}$ by putting $t(a+\lambda I_{2}):=t(a)+2\lambda$ and $q(a+\lambda I_{2}):=q(a)+\lambda^{2}+\lambda t(a)$ for all $a\in H$ and all $\lambda\in\mathbb{F}$, and now we have the formula
$\forall a\in\mathcal{V},\;a^{2}=t(a)\,a-q(a)\,I_{2}.$
Clearly $t$ is a linear form on $\mathcal{V}$, and $q$ is a quadratic form on $\mathcal{V}$.
Now, we return to the general case (we do not assume anymore that $\mathbb{D}$ is noncommutative). Let us take $W:=\operatorname{\operatorname{Ker}}t$, which has dimension at least $d+1\geq 2$. Let $a\in W$. Then $a^{2}=-q(a)I_{2}$. Since $a$ is similar to a diagonal matrix with entries in $\mathbb{F}$, we deduce that q(a)-q(a) is a square in $\mathbb{F}$, and $q(a)=0$ only if $a=0$. Hence q-q is a nonisotropic quadratic form on the $\mathbb{F}$-vector space $W$ whose values are all squares in $\mathbb{F}$. If either $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$ or $\dim W\geq 3$ then we can take $\mathbb{F}$-linearly independent vectors $a$ and $b$ in $W$ that are orthogonal for the polar form of $q$, and we deduce that $\mathbb{F}$ is Pythagorean and formally real. Otherwise $\operatorname{\operatorname{char}}(\mathbb{F})=2$ and $d=1$, but then the only trace zero matrices of $\operatorname{\operatorname{M}}_{2}(\mathbb{D})$ that are diagonalisable are the scalar multiples of $I_{2}$, which would lead to $\dim\mathcal{V}\leq 2$ and would thereby contradict our assumptions. ∎
8.6 The proof
Here is the key step in our proof.
Lemma 8.8.
Let $n\geq 2$. Let $\mathcal{V}$ be an $\mathbb{F}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\dbinom{n}{2}$. Then:
  1. (a)
    The space $\mathcal{V}$ has the full rank $1$ idempotent property.
  2. (b)
    There exists an $\mathbb{F}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$ with dimension $(n-1)+d\dbinom{n-1}{2}$.
Proof.
Let $E$ be a $1$-dimensional $\mathbb{D}$-linear subspace of $\mathbb{D}^{n}$. Conjugating $\mathcal{V}$, we lose no generality in assuming that $E=\mathbb{D}\times\{0_{n-1}\}$. For $M\in\mathcal{V}$, write
$M=\begin{bmatrix}a(M)&R(M)\\ C(M)&K(M)\end{bmatrix}\quad\text{where $a(M)\in\mathbb{D}$, $C(M)\in\mathbb{D}^{n-1}$ and so on.}$
Set
$\mathcal{W}:=\{M\in\mathcal{V}:C(M)=0\}.$
Note that
$\dim\mathcal{W}\geq\dim\mathcal{V}-(n-1)d.$
For every $M\in\mathcal{W}$, the scalar $a(M)$ must belong to $\mathbb{F}$ and the matrix $K(M)$ must be $\mathbb{F}$-diagonalisable (use, e.g., the characterization in terms of annihilating polynomials with coefficients in $\mathbb{F}$). Hence $K(\mathcal{W})$ is an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$, and $a(\mathcal{W})\subseteq\mathbb{F}$. By the former
$\dim K(\mathcal{W})\leq(n-1)+d\binom{n-1}{2}$.
Moreover, for all $M\in\mathcal{W}$, if $K(M)=0$ and $a(M)=0$ then $M$ is nilpotent and hence $R(M)=0$. It follows that the $\mathbb{F}$-linear mapping
$\varphi:M\in\mathcal{W}\mapsto(a(M),K(M))\in\mathbb{F}\times K(\mathcal{W})$
is injective, leading to
$\dim\mathcal{W}\leq 1+\dim K(\mathcal{W})\leq n+d\dbinom{n-1}{2}$.
Yet $\dim\mathcal{V}-(n-1)d=n+d\dbinom{n-1}{2}$. Hence $\dim\mathcal{W}=n+d\dbinom{n-1}{2}$ and we deduce:
  • That $\dim K(\mathcal{W})=(n-1)+d\dbinom{n-1}{2}$, which yields point (b);
  • That $\varphi$ is onto.
By the latter, we can find an $M\in\mathcal{W}$ such that $a(M)=1$ and $K(M)=0$, and then obviously $M\in\mathcal{V}$ is idempotent with range $E$. This yields point (a). ∎
As a consequence, by induction we deduce from point (b) and from Lemma 8.7:
Corollary 8.9.
Let $n\geq 2$. If there exists an $\mathbb{F}$-diagonalisable subspace $\mathcal{V}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\dbinom{n}{2}$, then $\mathbb{F}$ is formally real and consequently $\operatorname{\operatorname{char}}(\mathbb{F})=0$.
Now, we can finish the proof. So, let $n\geq 2$ and $\mathcal{V}$ be an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\dbinom{n}{2}$. By the previous corollary, we are on the safe side: the field $\mathbb{F}$ is infinite, and the trace map $\operatorname{\operatorname{Tr}}$ satisfies $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})\neq 0$.
Since $\mathbb{F}I_{n}+\mathcal{V}$ remains $\mathbb{F}$-diagonalisable, we deduce from Theorem 8.1 that $\mathcal{V}=\mathbb{F}I_{n}+\mathcal{V}$ and hence $I_{n}\in\mathcal{V}$. Now, we consider the standard inner product
$\langle A,B\rangle:=\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(AB))$
on $\mathcal{V}$ and the orthogonal complement $\mathcal{V}^{\bot}$ of $\mathcal{V}$, which is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n(d-1)+d\dbinom{n}{2}$.
By Lemma 8.8, $\mathcal{V}$ has the full rank $1$ idempotent property, and hence Lemma 8.6 yields that $\mathcal{V}^{\bot}$ has $\mathbb{F}$-trivial spectrum. And finally, $\mathcal{V}^{\bot}$ is optimal as an $\mathbb{F}$-trivial spectrum subspace, because $\mathbb{F}$ is infinite. Here is our next step:
Claim 8.1.
The space $\mathcal{V}^{\bot}$ is irreducible.
Proof.
Assume on the contrary that $\mathcal{V}^{\bot}$ is reducible. Because the inner product is invariant under conjugation, we can assume without loss of generality that $\mathbb{D}^{p}\times\{0_{n-p}\}$ is invariant under $\mathcal{V}^{\bot}$ for some $p\in\mathopen{[\![}1,n-1\mathclose{]\!]}$. Then clearly the unit matrix $E_{1,n}$ (with zero entries everywhere except on the first row and $n$-th column, where the entry equals $1$) belongs to $(\mathcal{V}^{\bot})^{\bot}=\mathcal{V}$. Yet $E_{1,n}$ is nonzero and nilpotent, so it is not $\mathbb{F}$-diagonalisable. This is a contradiction. Hence $\mathcal{V}^{\bot}$ is irreducible. ∎
Now, we can use Theorem 1.8. Since $n\geq 2$, the division ring $\mathbb{D}$ must be of quadratic type over $\mathbb{F}$. Since $\operatorname{\operatorname{char}}(\mathbb{F})=0$, the hyper-radicial case is discarded, so there are only three cases: either $d=1$, or $d=2$ and $\mathbb{D}$ is a (separable) quadratic extension of $\mathbb{F}$, or $\mathbb{D}$ is a quaternion algebra over $\mathbb{F}$ (in which case $\operatorname{\operatorname{Tr}}$ equals twice the quaternionic trace). Denote by $\sigma:x\mapsto x^{\star}$ the standard involution of $\mathbb{D}$ over $\mathbb{F}$, and recall that $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})=\{M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D}):M^{\star}=-M\}$ because $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$. By Theorem 1.8 there exists a matrix $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that
$\mathcal{V}^{\bot}=P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$
(here we do not care that $P$ should be $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$-nonisotropic). As obviously $\mathcal{H}_{n}(\mathbb{D})$ is orthogonal to $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ for the inner product, whereas $\dim_{\mathbb{F}}\mathcal{H}_{n}(\mathbb{D})+\dim_{\mathbb{F}}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})=dn^{2}$, we find
$\mathcal{V}=\mathcal{H}_{n}(\mathbb{D})P.$
The conclusion is near. Remembering that $I_{n}\in\mathcal{V}$, we get $P^{-1}\in\mathcal{H}_{n}(\mathbb{D})$, and hence $P\in\mathcal{H}_{n}(\mathbb{D})$. Now, note that we can safely replace $P$ with any star-congruent matrix: indeed if we take $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ then
$Q^{-1}\mathcal{V}Q=Q^{-1}\mathcal{H}_{n}(\mathbb{D})PQ=Q^{-1}\mathcal{H}_{n}(\mathbb{D})(Q^{-1})^{\star}Q^{\star}PQ=\mathcal{H}_{n}(\mathbb{D})(Q^{\star}PQ).$
In particular, by reduction theory for Hermitian forms (recall that $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$) we can now replace $P$ with a diagonal matrix $\Delta$ with diagonal entries $\delta_{1},\dots,\delta_{n}$ in $\mathbb{F}^{\times}$. Hence $\mathcal{H}_{n}(\mathbb{D})\Delta$ is $\mathbb{F}$-diagonalisable.
Let $i\in\mathopen{[\![}2,n\mathclose{]\!]}$. The matrix $M:=E_{1,i}+E_{i,1}$ belongs to $\mathcal{H}_{n}(\mathbb{D})$, so $M\Delta=\delta_{i}E_{1,i}+\delta_{1}E_{i,1}$ must be $\mathbb{F}$-diagonalisable, and since its entries are in $\mathbb{F}$ we can safely use its characteristic polynomial $t^{n-2}(t^{2}-\delta_{1}\delta_{i})$. It follows that $\delta_{1}\delta_{i}$ is a nonzero square in $\mathbb{F}$, and hence $\delta_{i}=(\delta_{i})^{2}\,(\delta_{1}\delta_{i})^{-1}\delta_{1}$ is the product of $\delta_{1}$ with a nonzero square in $\mathbb{F}$. We conclude that $\Delta$ is star-congruent to $\delta_{1}I_{n}$, and hence $\mathcal{V}$ is conjugated to $\mathcal{H}_{n}(\mathbb{D})(\delta_{1}I_{n})=\mathcal{H}_{n}(\mathbb{D})$. In particular $\mathcal{H}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable.
Finally, let $x\in\mathbb{D}$. Then the matrix $xE_{1,2}+x^{\star}E_{2,1}$ must be diagonalisable over $\mathbb{F}$. The square of this matrix is diagonal with diagonal entries $xx^{\star},x^{\star}x,0,\dots,0$, and hence all these values must be squares of elements of $\mathbb{F}$. In particular $xx^{\star}=\alpha^{2}$ for some $\alpha\in\mathbb{F}$. This completes the proof of Theorem 8.4.
8.7 Real subspaces of complex-diagonalisable matrices
Let us give another application of our results on trivial spectrum subspaces. This time around, we consider the problem of $\mathbb{R}$-linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ in which all the elements are diagonalisable over $\mathbb{C}$. An example of such a subspace if $\mathbb{R}I_{n}\oplus i\mathcal{H}_{n}(\mathbb{C})$, the space of all sums of a real scalar multiple of the identity matrix with a skew-Hermitian matrix. As will turn out, this space has the greatest possible dimension for the problem under consideration.
Theorem 8.10.
Let $\mathcal{S}$ be an $\mathbb{R}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ in which all the matrices are diagonalisable over $\mathbb{C}$. Then $\dim\mathcal{S}\leq 1+n^{2}$. Moreover, if $n\geq 3$ and equality holds there exist $\alpha\in\mathbb{C}^{\times}$ and $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{C})$ such that $\mathcal{S}$ is conjugated to $\alpha(\mathbb{R}I_{n}\oplus i\mathcal{H}_{n}(\mathbb{C}))$.
Interestingly, and in contrast with the earlier results on spaces of diagonalisable matrices, in the special case $n=2$ there are several orbits of such spaces with dimension $5$ under the action of $\mathbb{C}^{\times}\times\operatorname{\operatorname{GL}}_{n}(\mathbb{C})$ by $(\alpha,P).\mathcal{S}:=\alpha P\mathcal{S}P^{-1}$. This can be seen heuristically as follows: take a solution $\mathcal{S}$ with dimension $5$. Then as $\mathbb{C}I_{2}+\mathcal{S}$ is also a solution, we see that $\mathbb{C}I_{2}\subseteq\mathcal{S}$, and hence $\{M\in\mathcal{S}:\operatorname{\operatorname{tr}}(M)=0\}$ is a $3$-dimensional real linear subspace of $\mathfrak{sl}_{2}(\mathbb{C})$. Conversely, given a real linear subspace $\mathcal{M}$ of $\mathfrak{sl}_{2}(\mathbb{C})$, the condition that all the matrices in $\mathcal{M}$ are diagonalisable over $\mathbb{C}$ is equivalent to the condition that all the nonzero matrices in $\mathcal{M}$ are nonsingular. Hence the set of all $3$-dimensional real linear subspaces of $\mathfrak{sl}_{2}(\mathbb{C})$ that arise in this way is an open subset in the projective Grassmannian variety of all $3$-dimensional real subspaces of $\mathfrak{sl}_{2}(\mathbb{C})$, to the effect that its topological dimension is 3(63)=93(6-3)=9. However, under the said action the orbits are exactly the same ones as under the restricted action of the subgroup $\mathbb{C}^{\times}\times\mathrm{SL}_{2}(\mathbb{C})$, because every matrix of $\operatorname{\operatorname{GL}}_{2}(\mathbb{C})$ splits as $\alpha P$ for some $\alpha\in\mathbb{C}^{\times}$ and some $P\in\mathrm{SL}_{2}(\mathbb{C})$. Since the group $\mathbb{C}^{\times}\times\mathrm{SL}_{2}(\mathbb{C})$ has topological dimension $8$, this hints at the presence of several orbits (and even at infinitely many orbits).
We can actually give a very simple example of a solution that is not conjugated to $\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C})$ under the previous action.
Proposition 8.11.
Consider the space $W$ consisting of all the matrices of the form
$M(x,z)=\begin{bmatrix}ix&-\overline{z}\\ (1+i)z&-ix\end{bmatrix}\quad\text{with $(x,z)\in\mathbb{R}\times\mathbb{C}$.}$
Then $\mathbb{C}I_{2}\oplus W$ is a $5$-dimensional real linear subspace of $\operatorname{\operatorname{M}}_{2}(\mathbb{C})$ in which all the elements are diagonalisable (over $\mathbb{C}$). Moreover, there is no pair $(\alpha,P)\in\mathbb{C}^{\times}\times\operatorname{\operatorname{GL}}_{2}(\mathbb{C})$ such that $\mathbb{C}I_{2}\oplus W=\alpha P(\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C}))P^{-1}$.
Proof.
Clearly $W$ is a $3$-dimensional real linear subspace of $\mathfrak{sl}_{2}(\mathbb{C})$. For all $(x,z)\in\mathbb{R}\times\mathbb{C}$, we notice that $\det(M(x,z))=x^{2}+|z|^{2}(1+i)$, which vanishes only if $x=z=0$. Hence, all the matrices in $W$ are diagonalisable over $\mathbb{C}$. Then $\mathcal{S}:=\mathbb{C}I_{2}\oplus W$ is a $5$-dimensional real linear subspace of $\operatorname{\operatorname{M}}_{2}(\mathbb{C})$ in which all the matrices are diagonalisable over $\mathbb{C}$. Obviously $W=\mathcal{S}\cap\mathfrak{sl}_{2}(\mathbb{C})$.
Now, assume that there exists a pair $(\alpha,P)\in\mathbb{C}^{\times}\times\operatorname{\operatorname{GL}}_{2}(\mathbb{C})$ such that $\mathcal{S}=\alpha P(\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C}))P^{-1}$. All trace zero matrices in $\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C})$ actually belong to $i\mathcal{H}_{2}(\mathbb{C})$, and hence $W\subseteq\alpha P(i\mathcal{H}_{2}(\mathbb{C}))P^{-1}$. Hence the determinant of every element of $W$ is a real multiple of $\alpha^{2}$. However, it is clear from the previous computation that no $1$-dimensional real linear subspace of $\mathbb{C}$ contains the determinant of all the elements of $W$. Therefore, $\mathbb{C}I_{2}+W$ is not in the orbit of $\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C})$. ∎
We suspect that classifying the orbits of the optimal solutions for $n=2$ is very difficult, if not intractable.
Now, we set out to prove Theorem 8.10. We start by giving an inductive proof of the first statement, in the form of a theorem on operator spaces. The statement is already obvious for $n=1$. Let $n\geq 2$, let $V$ be an $n$-dimensional complex vector space, and $\mathcal{S}$ be a real linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{C}}(V)$ in which every element is diagonalisable (as a $\mathbb{C}$-endomorphism). Throughout, we will use the basic observation that the only nilpotent element of $\mathcal{S}$ is the zero endomorphism. We introduce the trace-orthogonal
$\mathcal{S}^{\bot}:=\{v\in\operatorname{\operatorname{End}}_{\mathbb{C}}(V):\forall u\in\mathcal{S},\;\operatorname{\operatorname{Re}}(\operatorname{\operatorname{tr}}(uv))=0\},$
which is a real linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{C}}(V)$ with dimension $2n^{2}-\dim\mathcal{S}$. Throughout the proof, we will need the following important observation. For a nonzero vector $x$ of $V$, we denote by $\operatorname{\operatorname{EV}}(\mathcal{S},x)$ the space of all $\lambda\in\mathbb{C}$ for which $\lambda x\in\mathcal{S}x$, or in other words the set of all complex numbers that are, for some $u\in\mathcal{S}$, the eigenvalue of $u$ for the eigenvector $x$. Obviously $\operatorname{\operatorname{EV}}(\mathcal{S},x)$ is a real linear subspace of $\mathbb{C}$.
Lemma 8.12.
Let $x$ be a nonzero vector of $V$. Then $\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)$ is orthogonal to $\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)$ for the $\mathbb{R}$-bilinear form $(y,z)\in\mathbb{C}^{2}\mapsto\operatorname{\operatorname{Re}}(yz)$.
Proof.
Let indeed $(\lambda,\mu)\in\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)\times\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)$. We can choose $u\in\mathcal{S}$ such that $\operatorname{\operatorname{Im}}u\subseteq\mathbb{C}x$ and $u(x)=\lambda x$, and choose $v\in\mathcal{S}^{\bot}$ such that $v(x)=\mu\,x$. Then $\operatorname{\operatorname{Im}}(uv)\subseteq\mathbb{C}x$ and $(uv)(x)=\lambda\mu x$. It follows that $\operatorname{\operatorname{tr}}(uv)=\lambda\mu$, and we conclude that $\operatorname{\operatorname{Re}}(\lambda\mu)=0$ because $u\in\mathcal{S}$ and $v\in\mathcal{S}^{\bot}$. ∎
Now we can move forward. Let $x\in V\smallsetminus\{0\}$ be arbitrary. We consider the real linear subspace $\mathcal{S}^{\prime}$ of all $u\in\mathcal{S}$ such that $u(x)\in\mathbb{C}x$. Every $u\in\mathcal{S}^{\prime}$ induces a complex diagonalisable endomorphism $\overline{u}$ of the quotient space $V/\mathbb{C}x$, and we denote by $\overline{\mathcal{S}}$ the (real) linear subspace of all such endomorphisms. By induction $\dim\overline{\mathcal{S}}\leq(n-1)^{2}+1$. By the rank theorem applied to $u\mapsto[u(x)]\in V/\mathbb{C}x$, we see that
$\dim\mathcal{S}\leq 2(n-1)+\dim\mathcal{S}^{\prime}.$
The kernel of $u\in\mathcal{S}^{\prime}\mapsto\overline{u}\in\overline{\mathcal{S}}$ equals $\mathcal{S}^{\mathbb{C}x}$. Every element in this kernel has $x$ as eigenvector. Assign to every $u\in\mathcal{S}^{\mathbb{C}x}$ the eigenvalue $\lambda(u)\in\mathbb{C}$ of $u$ associated with the eigenvector $x$, and note that if $\lambda(u)=0$ then $u$ is nilpotent, and hence $u=0$. Hence the $\mathbb{R}$-linear mapping $u\in\mathcal{S}^{\mathbb{C}x}\mapsto\lambda(u)$ is injective, and we deduce that
$\dim(\mathcal{S}^{\mathbb{C}x})=\dim_{\mathbb{R}}(\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x))\leq 2.$
Finally, the rank theorem yields $\dim\mathcal{S}^{\prime}=\dim(\mathcal{S}^{\mathbb{C}x})+\dim\overline{\mathcal{S}}$. This already yields
$\dim\mathcal{S}\leq n^{2}+2.$
Now, let us assume that $\dim\mathcal{S}=n^{2}+2$. Then $\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)=\mathbb{C}$, and we deduce from Lemma 8.12 that $\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)=\{0\}$. Varying $x$, we deduce that no operator in $\mathcal{S}^{\bot}$ has a nonzero (complex) eigenvalue.
Let us draw some consequences of this. First of all, we deduce that all the operators in $\mathcal{S}^{\bot}$ are nilpotent, and hence we are in the position to use the extension of Gerstenhaber’s theorem to skew fields. By theorem 2 of [17], applied with $\mathbb{K}=\mathbb{C}$ and $\mathbb{K}_{0}=\mathbb{R}$, we deduce that $\dim\mathcal{S}^{\bot}\leq 2\dfrac{n(n-1)}{2}=n(n-1)$, and if equality occurs then $\mathcal{S}^{\bot}$ is reducible (because $n>1$).
However, if $\mathcal{S}^{\bot}$ has a nontrivial invariant $\mathbb{C}$-linear subspace $W$, then by double-orthogonality $\mathcal{S}$ would contain all the $\mathbb{C}$-linear endomorphisms that vanish on $W$ and map into $W$; yet all such endomorphisms are nilpotent, and one of them is nonzero. Hence $\mathcal{S}^{\bot}$ is irreducible (as a subset of $\operatorname{\operatorname{End}}_{\mathbb{C}}(V)$). It follows that $\dim\mathcal{S}^{\bot}\leq n^{2}-n-1$, to the effect that $\dim\mathcal{S}\geq 2n^{2}-n(n-1)+1=n^{2}+n+1>n^{2}+2$. This contradicts an earlier result.
We conclude that $\dim\mathcal{S}\leq n^{2}+1$.
From now on, we assume that $\dim\mathcal{S}=n^{2}+1$ and $n\geq 3$.
Step 1. $\mathcal{S}^{\bot}$ is intransitive.
To see this, we consider an arbitrary vector $x\in V\smallsetminus\{0\}$. By coming back to the above proof, we note that $\dim\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)\geq 1$, for the contrary would yield $\dim\mathcal{S}\leq n^{2}$. By Lemma 8.12, we recover $\dim\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)\leq 1$. In particular $\mathbb{C}x\not\subset\mathcal{S}^{\bot}x$, and hence $V\neq\mathcal{S}^{\bot}x$.
Step 2. $\mathcal{S}^{\bot}$ is primitively intransitive.
Assume the contrary. Then we can take a maximal nonzero $\mathbb{C}$-linear subspace $W$ of $V$ such that, for the projection $\pi:V\twoheadrightarrow V/W$, the projected space $\pi\mathcal{S}^{\bot}$ is intransitive. By the maximality of $W$, the operator space $\pi\mathcal{S}^{\bot}$ is primitively intransitive. Denote by $t$ the dimension of $V/W$ over $\mathbb{C}$, and note that $1\leq t<n$. Consider the subspace $\mathcal{S}_{W}$ of $\mathcal{S}$ consisting of its operators that vanish on $W$. For every $u\in\mathcal{S}_{W}$, the induced endomorphism $\overline{u}$ of $V/W$ remains $\mathbb{C}$-diagonalisable, and the mapping $u\mapsto\overline{u}$ is injective because its kernel contains only nilpotent operators that belong to $\mathcal{S}$. Hence $\dim\mathcal{S}_{W}\leq 1+t^{2}$ by the first part of the proof. Besides, it is standard (see e.g. lemma 3.4 in [25], for the case of fields) that
$\dim\mathcal{S}_{W}+\dim(\pi\mathcal{S}^{\bot})=\dim_{\mathbb{R}}\operatorname{\operatorname{Hom}}_{\mathbb{C}}(V/W,V)=2nt.$
However, since $\pi\mathcal{S}^{\bot}$ is primitively intransitive it is also deeply intransitive (see Proposition 4.5; here $\mathbb{R}$ is infinite) and it follows that $\dim(\pi\mathcal{S}^{\bot})\leq\alpha(t,2)=t^{2}$. Therefore $2nt\leq 2t^{2}+1$, which yields $2t(n-t)\leq 1$ and obviously contradicts the fact that $t\in\mathopen{[\![}1,n-1\mathclose{]\!]}$.
Step 3. Applying the classification theorem for deeply intransitive spaces.
Since $\mathcal{S}^{\bot}$ is primitively intransitive, it is also deeply intransitive (again, by Proposition 4.5). Here $\dim\mathcal{S}^{\bot}=\alpha(n,2)-1$, and we observe that $n-\max(4-2,2)=n-2\geq 1$, to the effect that $\dim\mathcal{S}^{\bot}\geq\alpha(n,2)-n+\max(4-2,2)$. Hence we can apply Theorem 1.21, with $e=\operatorname{\operatorname{Re}}$ and $\sigma:z\mapsto\overline{z}$ the complex conjugation. This yields a nondegenerate sesquilinear form $B$ on $V$ such that, for $b:(x,y)\in V^{2}\mapsto\operatorname{\operatorname{Re}}(B(x,y))$, one has $\mathcal{S}^{\bot}\subseteq\mathcal{A}_{b,\mathbb{C}}$. Let us consider an arbitrary basis $\mathbf{B}$ of $V$, and consider the Gram matrix $P$ of $B$ in $\mathbf{B}$. Let $u\in\mathcal{S}^{\bot}$ have representing matrix $M$ in $\mathbf{B}$. Then $\forall X\in\mathbb{C}^{n},\;\operatorname{\operatorname{Re}}(X^{\star}PMX)=0$, and we deduce that $PM\in i\mathcal{H}_{n}(\mathbb{C})$.
Step 4. Back to $\mathcal{S}$.
Let us now denote by $\mathcal{M}$ the matrix space that represents $\mathcal{S}$ in the basis $\mathbf{B}$. Since $\mathcal{S}^{\bot}$ is represented in that basis by a real linear subspace of $iP^{-1}\mathcal{H}_{n}(\mathbb{C})$, we find by double-orthogonality that $\mathcal{M}\supseteq\mathcal{H}_{n}(\mathbb{C})P$. Note that $\dim_{\mathbb{R}}\mathcal{M}=n^{2}+1$, whence $\mathcal{H}_{n}(\mathbb{C})P$ is a (real) linear hyperplane of $\mathcal{M}$. Taking $Q:=P^{\star}$, we see that $\mathcal{H}_{n}(\mathbb{C})P=Q^{-1}\mathcal{H}_{n}(\mathbb{C})$.
It remains to analyze the matrix $Q$. Noting that $\mathcal{M}+\mathbb{C}I_{n}$ is a real linear subspace of $\mathbb{C}$-diagonalisable matrices, we see from the maximality of the dimension of $\mathcal{S}$ that $\mathbb{C}I_{n}\subseteq\mathcal{M}$. It follows in particular that $Q^{-1}\mathcal{H}_{n}(\mathbb{C})$ contains a nonzero scalar matrix. Hence $Q=\lambda Q^{\prime}$ for some $\lambda\in\mathbb{C}^{\times}$ and some $Q^{\prime}\in\mathcal{H}_{n}(\mathbb{C})\cap\operatorname{\operatorname{GL}}_{n}(\mathbb{C})$.
Multiplying $\mathcal{S}$ with $\lambda$, we then lose no generality in assuming that $Q$ is Hermitian. As in Section 8.6, we can replace $Q$ at will with a star-congruent matrix. Assume that $Q$ is isotropic. Then $Q$ is star-congruent to $Q_{1}:=\begin{bmatrix}K&[0]_{2\times(n-2)}\\ [0]_{(n-2)\times 2}&Q_{2}\end{bmatrix}$ for $K:=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$ and some invertible Hermitian matrix $Q_{2}$, and it follows that every matrix in $Q_{1}^{-1}\mathcal{H}_{n}(\mathbb{C})$ is diagonalisable. Yet $Q_{1}^{-1}\operatorname{\operatorname{Diag}}(1,0,\dots,0)$ is clearly nilpotent and nonzero, and hence it is not diagonalisable.
It follows that $Q$ is nonisotropic, and hence star-congruent to $\pm I_{n}$. Replacing $\mathcal{M}$ with a similar subspace, we can reduce the situation to the one where $Q=\pm I_{n}$, in which case we have $\mathcal{M}\supseteq\mathcal{H}_{n}(\mathbb{C})$. In this reduced situation, $i\mathcal{M}\supseteq\mathbb{C}I_{n}+i\mathcal{H}_{n}(\mathbb{C})$. Since the dimensions (over $\mathbb{R}$) are equal we conclude that
$i\mathcal{M}=\mathbb{C}I_{n}+i\mathcal{H}_{n}(\mathbb{C})=\mathbb{R}I_{n}\oplus i\mathcal{H}_{n}(\mathbb{C}),$
which completes the proof.
8.8 Subspaces of semi-simple endomorphisms
Here we study a variation of the previous problems. So far we have investigated the real linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ that consist only of diagonalisable matrices, the real linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ that consist only of diagonalisable matrices with real eigenvalues, and we have just investigated the real linear subspaces $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ that consist only of diagonalisable matrices. Hence it seems natural to study the (real) linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ that consist only of matrices that are diagonalisable over the complex numbers. The natural generalization of this question over an arbitrary field is to consider, for an $\mathbb{F}$-vector space $V$ with finite dimension $n\geq 1$, the linear subspaces of $\operatorname{\operatorname{End}}(V)$ in which all the elements are semi-simple. Recall that an endomorphism $u$ of $V$ is semi-simple if every $u$-invariant subspace $W$ of $V$ has a complementary subspace that is $u$-invariant. Classically, this is equivalent to the property that all the irreducible divisors of the minimal polynomial of $u$ have multiplicity $1$, and this property is inherited by induced endomorphisms, meaning that if $u$ is semi-simple and $W$ is a $u$-invariant subspace then the endomorphisms of $W$ and $V/W$ induced by $u$ are also semi-simple. Finally, no nonzero nilpotent endomorphism is semi-simple. In particular, if $\mathcal{S}$ is a linear subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple, then by taking the intersection of $\mathcal{S}$ with a subspace of nilpotent endomorphisms with dimension $\dbinom{\dim V}{2}$ we see that $\dim\mathcal{S}\leq\dbinom{\dim V+1}{2}$.
Now, we give a simple example of a large subspace of semi-simple endomorphisms. Assume that $V$ has a nonisotropic symmetric bilinear form $b$. Consider the space $\mathcal{S}_{b}$ of all $b$-symmetric endomorphisms, i.e., the endomorphisms $u$ of $V$ such that $(x,y)\mapsto b(x,u(y))$ is symmetric (or, equivalently, the $b$-selfadjoint endomorphisms). For every such endomorphism $u$ and every $u$-invariant subspace $W$, it is classical that $W^{\bot_{b}}$ is $u$-invariant, while $W\oplus W^{\bot_{b}}=V$ since $b$ is nonisotropic. It follows that every element of $\mathcal{S}_{b}$ is semi-simple. Moreover $\dim\mathcal{S}_{b}=\dbinom{n+1}{2}$ since $\mathcal{S}_{b}$ is isomorphic to the space of all symmetric bilinear forms on $V$. The next theorem shows that these are the only optimal solutions:
Theorem 8.13.
Let $\mathbb{F}$ be an arbitrary field, and $V$ be an $n$-dimensional vector space over $\mathbb{F}$, with $n\geq 1$. Let $\mathcal{S}$ be a linear subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple. Then:
  1. (a)
    $\dim\mathcal{S}\leq\dbinom{n+1}{2}$;
  2. (b)
    If $\dim\mathcal{S}=\dbinom{n+1}{2}$ then $\mathcal{S}=\mathcal{S}_{b}$ for some nonisotropic symmetric bilinear form $b$ on $V$.
Note that here no cardinality assumption is made on the field $\mathbb{F}$ under consideration. Hence, for fields with characteristic other than $2$, the existence of a subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple, and with dimension $\dbinom{n+1}{2}$, is connected to the so-called u-invariant of $\mathbb{F}$ (the supremum of the dimensions of the nonisotropic quadratic forms over $\mathbb{F}$).
We have already proved point (a). The proof of point (b) is an easy adaptation of the one given previously for diagonalisable subspaces, and here we explain the adaptation. First of all, since here $\mathbb{D}=\mathbb{F}$ there is no issue of characteristic when defining the orthogonal complement. Thus, let us take $\mathcal{S}$ as a linear subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple, and with dimension $\dbinom{n+1}{2}$. Then, exactly as previously we obtain that $\mathcal{S}^{\bot}$ has trivial spectrum and is irreducible (this relies upon the observation that the only nilpotent and semi-simple endomorphism is the zero endomorphism, and that semi-simplicity is inherited by induced endomorphisms).
There is however a difficulty related to the cardinality condition in Theorem 1.7, and it can be circumvented as follows if $n>|\mathbb{F}|$, which requires that $n>3$ if $|\mathbb{F}|>2$, and $n>2$ otherwise. Set $k:=3$ if $|\mathbb{F}|>2$, and $k=2$ if $|\mathbb{F}|=2$. In that case, we proceed as in Lemma 8.8 to find, by induction, that there is a $k$-dimensional vector space $W$ and a linear subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}(W)$ with dimension $\dbinom{k+1}{2}$ and in which all the elements are semi-simple. Hence by induction $\mathcal{T}=\mathcal{S}_{c}$ for some nonisotropic symmetric bilinear form $c$ on $W$. But here $\mathbb{F}$ is finite and:
  • Either $k=3$, in which case there is no nonisotropic symmetric bilinear form on $W$, because it is known that every quadratic form on $W$ is isotropic;
  • Or $k=2$, in which case $|\mathbb{F}|=2$ and every symmetric bilinear form on $W$ is isotropic (because the associated quadratic form is actually linear).
Hence, in any case we find a contradiction.
It follows that $n\leq|\mathbb{F}|$. Hence, we can safely use Theorem 1.7 (actually, we could also have used theorem 3 of [18] for the case $|\mathbb{F}|>2$, but this would still leave out the case of fields with cardinality $2$) and recover that, in some basis $\mathbf{B}$, the space $\mathcal{S}^{\bot}$ is represented by the matrix space $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ for some nonisotropic matrix $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$. Then $\mathcal{S}$ is represented by $\operatorname{\operatorname{S}}_{n}(\mathbb{F})P$ in the same basis. Once more, we use the maximality of the dimension of $\mathcal{S}$ to observe that $\operatorname{\operatorname{id}}_{V}\in\mathcal{S}$, and we recover that $P^{-1}$ is symmetric, and hence $P$ is symmetric. Then $\operatorname{\operatorname{S}}_{n}(\mathbb{F})P=(P^{T})^{-1}(P^{T}\operatorname{\operatorname{S}}_{n}(\mathbb{F})P)=(P^{T})^{-1}\operatorname{\operatorname{S}}_{n}(\mathbb{F})$, and since $P^{T}$ is obviously nonisotropic we conclude that $\mathcal{S}=\mathcal{S}_{b}$ for the nonisotropic symmetric bilinear form $b$ that is represented by $P^{T}$ in $\mathbf{B}$.
In the special case $\mathbb{F}=\mathbb{R}$, we obtain exactly the same optimal spaces as for the diagonalisability assumption. In other words, up to conjugation $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ is the only linear subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ with dimension $\dbinom{n+1}{2}$ and in which every element is diagonalisable over $\mathbb{C}$.
Judging from the difficulties encountered in Section 8.7, we suspect that generalizing the previous approach to semilinear subspaces of semi-simple endomorphisms over general division rings could be very difficult.
8.9 Towards linear subspaces of matrices with no nonzero nilpotent element?
Let us finish with another variation of the previous problem. Instead of considering linear subspaces of semi-simple endomorphisms, why not consider an even looser property, which is the absence of nonzero nilpotent elements in the subspace, still for $\mathbb{F}$-linear endomorphisms of an $\mathbb{F}$-vector space $V$ with finite dimension $n$? For such a subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}(V)$, we still have $\dim\mathcal{S}\leq\dbinom{n+1}{2}$ with the same proof as for spaces of diagonalisable endomorphisms (for algebraically closed fields, the basic theory of projective algebraic varieties yields the upper bound $\dim\mathcal{S}\leq n$: apply [27] corollary 1.7 to the coefficients of the characteristic polynomial). However, for the analysis of the case of equality, the previous method fails because the property is not inherited by induced endomorphisms anymore (if we have a non-nilpotent endomorphism $u$ that leaves invariant a linear subspace $W$, it is possible that $u_{W}$ be nilpotent and nonzero). Hence, it seems difficult to classify the spaces that reach the critical dimension $\dbinom{n+1}{2}$. First of all, if we take a nonisotropic bilinear form $b$ on $V$, then the space $\mathcal{S}_{b}$ of all $b$-symmetric endomorphisms contains no nonzero nilpotent endomorphism. Indeed, for $u\in\mathcal{S}_{b}$, the kernel and the range of $b$ have no nonzero common vector (for $x\in\operatorname{\operatorname{Ker}}u\cap\operatorname{\operatorname{Im}}u$, we write $x=u(y)$ to get $b(x,x)=b(x,u(y))=b(y,u(x))=0$, and deduce from the nonisotropy of $b$ that $x=0$). In contrast with the previous case of spaces of semi-simple endomorphisms, here we do not require that $b$ be symmetric.
A naive conjecture is that every space with no nonzero nilpotent element and dimension $\dbinom{n+1}{2}$ equals $\mathcal{S}_{b}$ for some nonisotropic bilinear form $b$ on $V$. Yet this is disproved in the case of the field of reals by appealing to geometric considerations. Indeed, in that case, the set of all $\dbinom{n+1}{2}-$dimensional linear subspaces that contain no nonzero nilpotent matrix is clearly open in the Grassmanian manifold of all $\dbinom{n+1}{2}-$dimensional linear subspaces. Yet it is easily proved that the set of all $\mathcal{S}_{b}$ spaces, when $b$ ranges over the nonisotropic bilinear forms on $V$, is a manifold with dimension $n^{2}-1$ (where the “minus 11” comes from the observation that $\mathcal{S}_{b}=\mathcal{S}_{b^{\prime}}$ whenever $b$ and $b^{\prime}$ are linearly dependent). Yet the Grassmannian manifold under consideration has dimension $n^{2}\dbinom{n+1}{2}-\dbinom{n+1}{2}^{2}=\dfrac{n^{2}(n^{2}-1)}{4}$, which largely exceeds $n^{2}-1$ when $n>2$. In contrast, for $n=2$ it is an easy exercise to validate the conjecture.
Index

### v2

@@
-Semilinear spaces of matrices with transitivity, rank or spectral constraints
+Affine subspaces of units in simple algebras
 Abstract
-Let $\mathbb{D}$ be a division ring with finite degree $d$ over a central subfield $\mathbb{F}$.
-Under mild cardinality assumptions on $\mathbb{F}$, we determine the greatest possible dimension for an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ in which no matrix has a nonzero fixed vector, and we elucidate the structure of the spaces that have the greatest dimension.
-The classification involves the associative composition algebras over $\mathbb{F}$ and requires the study of intransitive spaces of operators between finite-dimensional vector spaces.
-These results are applied to solve various problems of the same flavor, including the determination of the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which all the matrices have rank at least $r$ (for some fixed $r$), as well as the structure of the spaces that have the greatest possible dimension if $n=p=r$; and finally the structure of large $\mathbb{F}$-linear subspaces of $\mathbb{F}$-diagonalisable matrices is elucidated.
+Let $\mathcal{A}$ be a simple algebra over a field $\mathbb{F}$.
+Under a mild cardinality assumption on $\mathbb{F}$, we determine the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\mathcal{A}$ that is included in the group of units $\mathcal{A}^{\times}$, and we describe the spaces that have the greatest possible dimension.
+This is equivalent to the problem of determining the greatest possible dimension for an $\mathbb{F}$-linear subspace $S$ of $\mathcal{A}$ in which $x-1_{\mathcal{A}}$ is a unit for all $x\in S$, and we elucidate the structure of these linear subspaces up to conjugation when their dimension reaches the greatest possible one.
+These classifications involve the associative composition algebras over $\mathbb{F}$.
+Over fields of characteristic other than $2$, the first problem is essentially reduced to the classification of nonisotropic quadratic forms over $\mathbb{F}$ and of nonisotropic Hermitian forms over quadratic and quaternionic extensions of $\mathbb{F}$.
+These results are intimately connected with the problem of intransitive operator spaces between finite-dimensional vector spaces over division rings, which we study in depth: in particular, we generalize a dual version of Atkinson’s theorem on primitive spaces of bounded rank matrices.
 AMS MSC: 15A30, 16K20
-Keywords: affine space, rank, matrices over division rings, intransitivity, trivial spectrum spaces, simultaneous reduction, associative composition algebras
-List of notation
-[…]
+Keywords: simple algebras, affine spaces, matrices over division rings, intransitivity, trivial spectrum spaces, associative composition algebras
 1 Introduction
 1.1 The problem
-Throughout, we let $\mathbb{D}$ be a (unital and associative) division ring, $\mathbb{F}$ be a subfield of the center of $\mathbb{D}$, and we assume that $\mathbb{D}$ has finite degree $d\geq 1$ over $\mathbb{F}$.
-All the vector spaces over $\mathbb{D}$ we manipulate will be finite-dimensional and right vector spaces, with the notable exception of dual spaces which are naturally left vector spaces.
-This way, we can safely practice matrix computations in the usual sense, with vectors represented by columns.
-The space $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ of all endomorphisms of the $\mathbb{D}$-vector space $V$ has no natural structure of $\mathbb{D}$-vector space (because we are avoiding any commutativity assumption), but it can nevertheless be seen as a vector space over $\mathbb{F}$.
-An element $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have $\mathbb{F}$-trivial spectrum when it has no nonzero (right) eigenvalue in $\mathbb{F}$ (we adopt a similar definition for matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$).
-A subset of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have $\mathbb{F}$-trivial spectrum whenever all its elements have $\mathbb{F}$-trivial spectrum.
-For an $\mathbb{F}$-linear subspace $S$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, to have $\mathbb{F}$-trivial spectrum is equivalent to the condition that no element of $S$ has $1$ as eigenvalue, i.e., no element of $S$ has a nonzero fixed vector.
-In matrix terms, an example of large $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with $\mathbb{F}$-trivial spectrum is the set of all upper-triangular matrices
-$\begin{bmatrix}x_{1}&&*\\ &\ddots&\\ (0)&&x_{n}\end{bmatrix}$
-in which the $i$-th diagonal entry $x_{i}$ is required to belong to a given $\mathbb{F}$-linear hyperplane $H_{i}$ of $\mathbb{D}$ such that $\mathbb{F}\oplus H_{i}=\mathbb{D}$.
-In case $\mathbb{D}=\mathbb{F}$, all the hyperplanes $H_{i}$ are zero and we have simply described the strictly upper-triangular matrices over $\mathbb{D}$.
-Note that the dimension over $\mathbb{F}$ of the above space is
-$\alpha(n,d):=n(d-1)+d\,\frac{n(n-1)}{2},$
-and it will be regularly useful to observe that
-$\forall m,\;\forall n,\quad\alpha(m+n,d)=\alpha(m,d)+\alpha(n,d)+mnd.$
-Trivial spectrum subspaces have been studied in the last decade [15, 18], but only in the special case $\mathbb{D}=\mathbb{F}$.
-One motivation for considering such spaces is that they are a natural variation of linear subspaces of nilpotent operators: it is unavoidable to cite the celebrated Gerstenhaber theorem [7], which states that the greatest possible dimension for a linear subspace of nilpotent $n$-by-$n$ matrices over a field is $\frac{n(n-1)}{2}$, with equality attained only for the conjugates of the subspace of all strictly upper-triangular matrices.
-See [17] for a generalization to arbitrary division rings, and [9, 26] for alternative proofs.
-There is a more profound justification for studying trivial spectrum subspaces: they are intimately connected with $\mathbb{F}$-affine subspaces of invertible elements of the ring $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
-Let indeed $\mathcal{S}$ be an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, with translation vector space denoted by $\overrightarrow{S}$ (hence $\overrightarrow{S}$ is a $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$).
-Assume that $\mathcal{S}$ contains $\operatorname{\operatorname{id}}_{V}$.
-Then $\mathcal{S}\subseteq\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$ if and only if $\operatorname{\operatorname{id}}_{V}-u$ is invertible for all $u\in\overrightarrow{S}$, i.e., no element of $\overrightarrow{S}$ has a nonzero fixed vector.
-As noted in the above, this is equivalent to the condition that $\overrightarrow{S}$ has $\mathbb{F}$-trivial spectrum.
-It follows that the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is exactly the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that is included in $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$.
-And by a method due to Meshulam [10], this can also be used to obtain the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which all the elements have their rank greater than or equal to some fixed $r\in\mathopen{[\![}1,\min(n,p)\mathclose{]\!]}$.
-In [23], we have also discovered a surprising connection between $\mathbb{F}$-trivial spectrum subspaces (when $\mathbb{D}=\mathbb{F}$) and the apparently opposite notion of an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, meaning an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ in which every element is diagonalisable over $\mathbb{F}$ (i.e., every element is represented in some basis by a diagonal matrix with all diagonal entries in $\mathbb{F}$, or alternatively is annihilated by a polynomial of $\mathbb{F}[t]$ that is split with simple roots over $\mathbb{F}$).
-In particular, it was shown that the classification of the trivial spectrum subspaces over $\mathbb{R}$ (there, $\mathbb{D}=\mathbb{F}=\mathbb{R}$) with the greatest possible dimension yields the one of the $\mathbb{R}$-diagonalisable linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ with maximal dimension, which simply turn out to be the conjugates of the subspace $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ of all $n$-by-$n$ symmetric matrices.
-As a special case of the results on $\mathbb{F}$-trivial spectrum subspaces, we will obtain both the Hermitian and quaternionic cases, and in particular we will prove that every $\mathbb{R}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ has dimension over $\mathbb{R}$ at most $n^{2}$, with equality attained only for the conjugates of the subspace $\mathcal{H}_{n}(\mathbb{C})$ of all $n$-by-$n$ (complex) Hermitian matrices.
-1.2 Reduction to the irreducible case
-An $\mathbb{F}$-trivial spectrum (linear) subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is called optimal when it has the greatest possible dimension among such spaces.
-A subset $\mathcal{X}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is called irreducible when $V$ is nonzero and $\mathcal{X}$ has no nontrivial invariant $\mathbb{D}$-linear subspace, and is called reducible otherwise.
-The study of optimal $\mathbb{F}$-trivial spectrum subspaces can entirely be reduced to the one of irreducible optimal subspaces.
-The reduction process is not specific to $\mathbb{F}$-trivial spectrum subspaces of endomorphisms, and is actually shared by many properties on matrices: an account of this is given in the appendix of [25], and here we will simply retrieve the results.
-Key to this is the observation that the $\mathbb{F}$-trivial spectrum property satisfies the following two conditions:
-  • It is invariant under similarity, i.e., if $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ has this property then so does $\varphi\circ u\circ\varphi^{-1}$ for every isomorphism $\varphi:V\overset{\simeq}{\longrightarrow}V^{\prime}$ of $\mathbb{D}$-vector spaces;
-  • It satisfies the strong inheritance condition, meaning that for every $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ and every nontrivial $\mathbb{D}$-linear subspace $W$ of $V$ that is invariant under $u$, the endomorphism $u$ has $\mathbb{F}$-trivial spectrum if and only if the endomorphisms of $W$ and $V/W$ induced by $u$ have $\mathbb{F}$-trivial spectrum (this follows from the observation that $u$ can be seen as an endomorphism of the $\mathbb{F}$-vector space $V$, which is finite-dimensional).
-By lemma A.1 in [25], we deduce the following result:
-Proposition 1.1.
-Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
-Then the $\mathcal{S}$-invariant subspaces of $V$ are totally ordered under inclusion.
-A flag of ($\mathbb{D}$-linear subspaces of) $V$ is an increasing list $V_{0}\subset\cdots\subset V_{p}$ of $\mathbb{D}$-linear subspaces with $V_{0}=\{0\}$ and $V_{p}=V$.
-Beware that we do not require that $\dim_{\mathbb{D}}V_{i}=i$ for all $i\in\mathopen{[\![}0,p\mathclose{]\!]}$ (in which case we would say that this flag is complete).
-The previous result can then be restated as saying that when $\mathcal{S}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, the $\mathcal{S}$-invariant subspaces form a flag.
-Next, let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, and consider the $\mathcal{S}$-invariant subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$.
-Given respective subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{1}/V_{0}),\dots,\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{p}/V_{p-1})$, one denotes by $\mathcal{X}_{1}\vee\cdots\vee\mathcal{X}_{p}$ the set of all $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that leave all the $V_{i}$’s invariant and induce, for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an element of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that belongs to $\mathcal{X}_{i}$.
-Then, as a consequence of theorem A.2 in [25], we obtain:
+Let $\mathbb{F}$ be a field.
+In the seminal [5], Murray Gerstenhaber proved that every linear subspace of nilpotent $n$-by-$n$ matrices (over $\mathbb{F}$) has its dimension no larger than $\frac{n(n-1)}{2}$, and equality is attained only for the conjugates of the space $\operatorname{\operatorname{NT}}_{n}(\mathbb{F})$ of all strictly upper-triangular matrices (Gerstenhaber actually required that $|\mathbb{F}|\geq n$, but his result has been later revealed to hold over an field [20]).
+[…]
+Meshulam’s problem was solved in the early 2010’s.
+In [9] and [10], the bound $\frac{n(n-1)}{2}$ from Gerstenhaber’s theorem was generalized to trivial spectrum subspaces over any field, and then a complete classification of the trivial spectrum subspaces with dimension $\frac{n(n-1)}{2}$, called the optimal ones, was achieved [12] over any field with cardinality greater than $2$.
+Spectacularly, the classification is intimately connected with the quadratic structure of the field $\mathbb{F}$: the classification of the optimal trivial spectrum subspaces is reduced to the classification of nonisotropic bilinear forms over finite-dimensional $\mathbb{F}$-vector spaces, up to congruence (i.e., equivalence and multiplication with a nonzero scalar).
+In particular, for quadratically closed fields the trivial spectrum subspaces of maximal dimension all consist of nilpotent operators.
+As a consequence, all the affine subspaces of invertible elements of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ have dimension at most $\frac{n(n-1)}{2}$, and (for fields with more than $2$ elements) the classification up to equivalence of the ones of maximal dimension, called the optimal ones, has been reduced to the classification of nonisotropic quadratic forms over $\mathbb{F}$ up to congruence.
+To be more precise, to each equivalence class of optimal affine subspace of invertible elements of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ can be attached a list $(q_{1},\dots,q_{p})$ of nonisotropic quadratic forms on finite-dimensional vector spaces over $\mathbb{F}$ such that $n=\sum_{k=1}^{p}\dim q_{k}$, and the equivalence class of the said matrix space determines the congruence class of each $q_{i}$.
+We must of course insist that this result fails for fields with $2$ elements.
+For example, for $\mathbb{F}=\mathbb{R}$ this shows that there are as many equivalence classes of optimal affine spaces as there are ordered partitions of $n$.
+Finally, these results have been applied to the study of affine subspaces of matrices with rank bounded below [11] and to the one of linear subspaces of diagonalisable matrices [18].
+Here we consider the more general problem of affine subspaces of units in simple algebras.
+A simple algebra over $\mathbb{F}$ is a nontrivial unital associative $\mathbb{F}$-algebra $\mathcal{A}$ with finite dimension over $\mathbb{F}$ and which has no nontrivial two-sided ideal.
+Two subsets $\mathcal{S}$ and $\mathcal{T}$ of $\mathcal{A}$ are called equivalent whenever there exist units $a,b$ in $\mathcal{A}^{\times}$ such that $\mathcal{T}=a\mathcal{S}b$, in which case we write $\mathcal{S}\sim\mathcal{T}$, and they are called similar whenever there exists a unit $a\in\mathcal{A}^{\times}$ such that $\mathcal{T}=a\mathcal{S}a^{-1}$, in which case we write $\mathcal{S}\simeq\mathcal{T}$.
+By a classical theorem of Wedderburn, the simple $\mathbb{F}$-algebras are the isomorphs of the matrix algebras of the form $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, where $\mathbb{D}$ is a division ring which contains $\mathbb{F}$ as a central subfield (i.e., $\mathbb{F}$ is included in the center of $\mathbb{D}$) and has finite dimension over $\mathbb{F}$, and $n$ is a positive integer.
+Here, we expressly do not require that $\mathcal{A}$ is central as an $\mathbb{F}$-algebra, for this would leave out many interesting cases, e.g., the one of the $\mathbb{R}$-algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$.
+Our main problem is the extension of Meshulam’s: Given a simple $\mathbb{F}$-algebra $\mathcal{A}$, what is the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\mathcal{A}$ that is included in the unit group $\mathcal{A}^{\times}$?
+And what are the ones that have the greatest possible dimension?
+We can easily translate Meshulam’s observation in this general situation: every $\mathbb{F}$-affine subspace of $\mathcal{A}$ is equivalent to one which contains the unity $1_{\mathcal{A}}$, and hence we can largely bound the discussion by studying the $\mathbb{F}$-linear subspaces $S$ of $\mathcal{A}$ for which $1_{A}-x$ is a unit for all $x\in S$.
+Definition 1.1.
+A subset of $\mathcal{A}$ is called unital when it is included in the group $\mathcal{A}^{\times}$ of units of $\mathcal{A}$.
+A unital $\mathbb{F}$-affine subspace is called optimal when it has the greatest possible dimension among such subspaces.
+An $\mathbb{F}$-linear subspace $S$ of $\mathcal{A}$ is said to have trivial spectrum whenever $1_{\mathcal{A}}+S$ is unital, i.e., $1_{\mathcal{A}}-x\in\mathcal{A}^{\times}$ for all $x\in S$.
+It is called an optimal trivial spectrum subspace whenever it has the greatest possible dimension among such subspaces.
+We note that, at the other end of the spectrum, the analogue problem for the $\mathbb{F}$-affine subspaces of elements of $\mathcal{A}$ that contain no unit has been entirely solved in [16] (the greatest possible dimension has been obtained, as well as the structure of the spaces with greatest possible dimension).
+It can hardly be disputed that this problem is considerably simpler than the one we are tackling here.
+1.2 Reduction to the irreducible spaces
+In order to state our results more clearly, it is unavoidable to use Wedderburn’s structure theorem for simple algebras and to frame the problem into one on spaces of linear operators or of matrices.
+So, we now let $\mathbb{D}$ be an arbitrary division ring which contains $\mathbb{F}$ as a central subfield, and with finite dimension $d$ over $\mathbb{F}$.
+We will consider right vector spaces over $\mathbb{D}$ throughout, enabling us to represent $\mathbb{D}$-linear mappings between vector spaces by matrices with entries in $\mathbb{D}$ in the usual way (so that vector are represented by columns and applying an operator corresponds to multiplying the coordinate vector on the left by the matrix of the operator).
+For such vector spaces $U$ and $V$, we get the $\mathbb{F}$-vector space $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ of all $\mathbb{D}$-linear mappings from $U$ to $V$, and the simple $\mathbb{F}$-algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(U)$ of all endomorphisms of the $\mathbb{D}$-vector space $U$.
+The latter is isomorphic to the matrix $\mathbb{F}$-algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ where $n:=\dim_{\mathbb{D}}U$ through a choice a basis of $U$.
+Throughout, we consider $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ as an $\mathbb{F}$-vector space and will not consider potential enriched structures (for example, when $\mathbb{D}=\mathbb{C}$ and $\mathbb{F}=\mathbb{R}$, we only view $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ as an $\mathbb{R}$-vector space, and never as a $\mathbb{C}$-vector space).
+We can immediately give an example of a very large trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, which generalizes Gerstenhaber’s nilpotent spaces: we can choose various $\mathbb{F}$-linear hyperplanes $H_{1},\dots,H_{n}$ of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$, and form the set $H_{1}\vee\cdots\vee H_{n}$ of all matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ of the form
+$\begin{bmatrix}h_{1}&&[*]\\ &\ddots&\\ (0)&&h_{n}\end{bmatrix}$
+where $h_{i}\in H_{i}$ for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, and the entries above the diagonal are arbitrary.
+It is clear that this is a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
+We note that this space has dimension
+$\alpha(n,d):=n(d-1)+d\binom{n}{2}$.
+More generally, we can create new trivial spectrum subspaces of matrices from known ones thanks to the junction operator $\vee$.
+For nonempty subsets $\mathcal{M}_{1}\subseteq\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\mathcal{M}_{p}\subseteq\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, we define their joint $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ as the set of all block upper-triangular matrices of the form
+$\begin{bmatrix}M_{1}&&[*]\\ &\ddots&\\ [0]&&M_{p}\end{bmatrix}$
+where $M_{1}\in\mathcal{M}_{1},\dots,M_{p}\in\mathcal{M}_{p}$ and the blocks above the diagonal are arbitrary.
+Since $\mathbb{D}$ is finite-dimensional over $\mathbb{F}$, for any block upper-triangular matrix $M=\begin{bmatrix}A&B\\ [0]&C\end{bmatrix}$ with entries in $\mathbb{D}$ the invertibility of $M$ is equivalent to the one of $A$ and $C$.
+As a consequence, $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ is a unital $\mathbb{F}$-affine subspace if and only if each $\mathcal{M}_{i}$ is a unital $\mathbb{F}$-affine subspace, and $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ is a trivial spectrum linear subspace if and only if each $\mathcal{M}_{i}$ is a trivial spectrum linear subspace.
+The above construction is of course intimately connected with the notion of irreducibility for subsets of endomorphisms.
+Given a finite-dimensional vector space $V$ over $\mathbb{D}$ and a subset $\mathcal{X}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, a $\mathbb{D}$-linear subspace $V_{0}$ of $V$ is called $\mathcal{X}$-invariant whenever $\forall f\in\mathcal{X},\;\forall x\in V_{0},\;f(x)\in V_{0}$, and we say that $\mathcal{X}$ is irreducible whenever $V\neq\{0\}$ and the only $\mathcal{X}$-invariant subspaces of $V$ are $\{0\}$ and $V$.
+If $S$ is a trivial spectrum subspace and $V_{0}$ is an $S$-invariant subspace of $V$, then $S$ induces respective linear subspaces of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V_{0})$ and $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(V/V_{0})$, and both are trivial spectrum subspaces.
+Then, we have the equivalent of the junction operator for subsets of endomorphisms: given a flag $(V_{0},\dots,V_{p})$ of $\mathbb{D}$-linear subspaces of $V$, and given (non-empty) subsets $\mathcal{S}_{1}\subseteq\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{1}/V_{0}),\dots,\mathcal{S}_{p}\subseteq\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{p}/V_{p-1})$, we denote by $\mathcal{S}_{1}\vee\cdots\vee\mathcal{S}_{p}$ the set of all $\mathbb{D}$-linear endomorphisms of $V$ that leave each space $V_{i}$ invariant and induce, for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an endomorphism of $V_{i}/V_{i-1}$ that belongs to $\mathcal{S}_{i}$.
+We can now invoke a result that was obtained in the appendix of [19] as a special case of a more general one:
 Theorem 1.1.
-Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum (linear) subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, with invariant subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$.
-Denote, for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, by $\mathcal{S}_{i}$ the set of all endomorphisms of $V_{i}/V_{i-1}$ induced by the elements of $\mathcal{S}$.
-Then
-$\mathcal{S}=\mathcal{S}_{1}\vee\cdots\vee\mathcal{S}_{p},$
-and for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the set $\mathcal{S}_{i}$ is an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$.
-Let us now give a matrix interpretation of the above.
-Two subsets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are called similar, and we write $\mathcal{X}\simeq\mathcal{X}^{\prime}$, when there exists $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that $\mathcal{X}^{\prime}=P\mathcal{X}P^{-1}$.
-Let $n_{1},\dots,n_{p}$ be positive integers.
-For respective subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$ of $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, we denote by $\mathcal{X}_{1}\vee\cdots\vee\mathcal{X}_{p}$ the set of all matrices of the form
-$M=\begin{bmatrix}M_{1}&&[?]\\ &\ddots&\\ [0]&&M_{p}\end{bmatrix}\quad\text{where $M_{1}\in\mathcal{X}_{1},\dots,M_{p}\in\mathcal{X}_{p}$.}$
-We call it the joint of the sets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$.
-Let $\mathcal{M}$ be an optimal $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
-Then there exists a unique partition $n=n_{1}+\cdots+n_{p}$ into positive integers, and, for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an irreducible optimal $\mathbb{F}$-linear subspace $\mathcal{M}_{i}$ of $\operatorname{\operatorname{M}}_{n_{i}}(\mathbb{D})$ such that
-$\mathcal{M}\simeq\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}.$
-Moreover, each $\mathcal{M}_{i}$ is uniquely determined by $\mathcal{M}$ up to similarity.
-As a conclusion, we find that understanding the optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ essentially amounts to understanding the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{k}(\mathbb{D})$ for all integers $k\in\mathopen{[\![}1,n\mathclose{]\!]}$.
-There is a small difficulty however: given a partition $n=n_{1}+\cdots+n_{p}$ and respective irreducible optimal $\mathbb{F}$-trivial spectrum subspaces $\mathcal{M}_{1},\dots,\mathcal{M}_{p}$ of $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, it is possible in theory that $\mathcal{M}_{1}\vee\dots\vee\mathcal{M}_{p}$, which is an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is not an optimal one.
-We call this the composition problem.
-Fortunately, in the situations we will consider, it will turn out that the joint of irreducible optimal $\mathbb{F}$-trivial spectrum subspaces is always optimal, and this will stem from the observation that
-$\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}=\alpha(n,d).$
-1.3 Basic considerations on trivial spectrum subspaces, and a review of known results
-Now, we recall the main known results on trivial spectrum subspaces in the case $\mathbb{D}=\mathbb{F}$.
-Theorem 1.3 (See [12, 15]).
-Let $V$ be a vector space over $\mathbb{F}$ with finite dimension $n$.
-Then the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}(V)$ is $\dbinom{n}{2}=\alpha(n,1)$.
-In particular, the composition problem stated in the previous section has a positive answer when $\mathbb{D}=\mathbb{F}$.
-As far as the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces are concerned, their structure is described by the main theorem of [18] when $\mathbb{D}=\mathbb{F}$, which excludes the case $|\mathbb{F}|=2$ (in that case there are enough counterexamples to discourage anyone to try formulating a plausible conjecture [22]).
-Before we restate this theorem, some additional definitions and notation are required.
-Beware that throughout we manipulate bilinear forms that are possibly nonsymmetric.
-Definition 1.1.
-Let $b$ be a bilinear form on an $\mathbb{F}$-vector space $V$. We say that $b$ is nonisotropic whenever
-$\forall x\in V,\;b(x,x)\neq 0.$
-An endomorphism $u\in\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ is called $b$-alternating (on the right) whenever
-$\forall x\in V,\;b(x,u(x))=0,$
-i.e., the bilinear form $(x,y)\in V^{2}\mapsto b(x,u(y))$ is alternating.
-We denote by $\mathcal{A}_{b}$ the set of all $b$-alternating $\mathbb{F}$-endomorphisms of $V$ (it is a linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$).
-If $V$ has an additional structure of $\mathbb{D}$-vector space (which extends its structure of $\mathbb{F}$-vector space), we set
-$\mathcal{A}_{b,\mathbb{D}}:=\mathcal{A}_{b}\cap\operatorname{\operatorname{End}}_{\mathbb{D}}(V).$
-If $V$ has dimension $n$ over $\mathbb{D}$ and $b$ is nondegenerate, then $\mathcal{A}_{b}$ is naturally isomorphic to the space of all alternating bilinear forms on $V$, and hence has dimension $\frac{nd(nd-1)}{2}$ over $\mathbb{F}$.
-Moreover, if $b$ is nonisotropic then the linear subspace $\mathcal{A}_{b}$ has $\mathbb{F}$-trivial spectrum, as for all $x\in V$ and $u\in\mathcal{A}_{b}$ the equality $u(x)=x$ implies $b(x,x)=b(x,u(x))=0$ and hence $x=0$.
-Now, we can state the main result of [18], reformulated in geometric terms.
-Theorem 1.4 (See theorems 3 and 4 in [18]).
-Let $V$ be a nonzero (finite-dimensional) $\mathbb{F}$-vector space, with $|\mathbb{F}|>2$. Then:
+Let $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ for some vector space $V$. Then:
   1. (a)
-    The irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are the spaces of the form $\mathcal{A}_{b}$, where $b$ is a nonisotropic bilinear form on $V$.
+    The $S$-invariant subspaces of $V$ are totally ordered under inclusion, thereby forming a (potentially incomplete) flag $(V_{0},\dots,V_{p})$ of linear subspaces of $V$.
   2. (b)
-    Given two nonisotropic bilinear forms $b$ and $b^{\prime}$ on $V$, the equality $\mathcal{A}_{b}=\mathcal{A}_{b^{\prime}}$ holds if and only if $b^{\prime}$ is a scalar multiple of $b$, and $\mathcal{A}_{b}$ is similar [^3] Two subsets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are called similar when there exists a vector space automorphism $\varphi$ of $V$ such that $\mathcal{X}^{\prime}=\{\varphi\circ u\circ\varphi^{-1}\mid u\in\mathcal{X}\}$.
-to $\mathcal{A}_{b^{\prime}}$ if and only if $b^{\prime}$ is congruent to $b$ (i.e., equivalent to a scalar multiple of $b$).
-Denote by $\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ the space of all alternating matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ (a matrix $M$ is alternating if and only if it represents an alternating bilinear form in a basis, or in other words $M^{T}=-M$ and all the diagonal entries of $M$ are equal).
-In matrix terms, point (a) of Theorem 1.4 means that the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are the spaces of the form $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ for the nonisotropic matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$.
-And the second part of point (b) means that two such spaces $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ and $Q^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ are similar if and only if $P$ is congruent to a scalar multiple of $Q$ (i.e., $P=\lambda\,RQR^{T}$ for some $\lambda\in\mathbb{F}^{\times}$ and some $R\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$).
-Moreover, for the more general case of (potentially reducible) optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$, we have the following corollary:
+    For all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, the $\mathbb{F}$-linear subspace $S_{i}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ induced by $S$ is an irreducible optimal trivial spectrum subspace.
+  3. (c)
+    The space $S$ equals $S_{1}\vee\cdots\vee S_{p}$.
+In matrix terms, this translates as follows:
 Theorem 1.2.
-Let $V$ be an $\mathbb{F}$-vector space, where $|\mathbb{F}|>2$, and $\mathcal{S}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$.
-Then the $\mathcal{S}$-invariant linear subspaces form a flag of subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and there exists a list $(b_{1},\dots,b_{p})$ of nonisotropic bilinear forms, with $b_{i}$ defined on $(V_{i}/V_{i-1})^{2}$, such that $\mathcal{S}$ is the space of all $u\in\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ that leave each subspace $V_{i}$ invariant and induce a $b_{i}$-alternating endomorphism $u_{i}$ of $V_{i}/V_{i-1}$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$.
-Moreover each form $b_{i}$ is uniquely determined up to multiplication by a nonzero scalar.
-In matrix terms this yields that for every optimal trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$, there is a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of nonisotropic matrices such that $\mathcal{M}$ is conjugated to the space of all matrices of the form
-$\begin{bmatrix}P_{1}^{-1}A_{1}&[?]&&[?]\\ [0]&P_{2}^{-1}A_{2}&\ddots&\\ \vdots&\ddots&\ddots&[?]\ \\ [0]&\cdots&[0]&P_{p}^{-1}A_{p}\end{bmatrix}$
-where $A_{1}\in\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}),\dots,A_{p}\in\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F})$ are variable alternating matrices, and the question marks represent arbitrary blocks.
-Moreover, the list $(n_{1},\dots,n_{p})$ is uniquely determined by $\mathcal{M}$, and each $P_{i}$ is uniquely determined by $\mathcal{M}$ up to congruence and multiplication with a nonzero scalar.
-Conversely, every space of the above form is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
-Hence the structure of the optimal trivial spectrum subspaces is intimately connected with the quadratic structure of $\mathbb{F}$.
-If $\mathbb{F}$ is quadratically closed, irreducible optimal subspaces only exist in dimension $1$, and we obtain the same result as in Gerstenhaber’s theorem (simultaneous triangularization).
-If $\mathbb{F}$ is finite then irreducible optimal subspaces only exist in dimensions $1$ and $2$, and there is a very limited number of similarity classes of such spaces.
-If $\mathbb{F}$ is the field of reals, there are irreducible optimal subspaces in all dimensions, and their classification up to conjugation is rather easy (it can be reduced to the determination of the orbits, under orthogonal conjugation, of the skewsymmetric matrices).
-1.4 Main new results on trivial spectrum subspaces
-We will now state a generalization of the previous results, however with a cardinality assumption on $\mathbb{F}$.
+Let $\mathcal{M}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for some integer $n>0$.
+Then there is a partition $n=n_{1}+\cdots+n_{p}$ into positive integers, and a list $(\mathcal{M}_{1},\dots,\mathcal{M}_{p})$ in which:
+  1. (i)
+    For each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, the set $\mathcal{M}_{i}$ is an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n_{i}}(\mathbb{D})$.
+  2. (ii)
+    The space $\mathcal{M}$ is similar to $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$.
+  3. (iii)
+    The similarity class of each $\mathcal{M}_{i}$ is uniquely determined by the one of $\mathcal{M}$.
+This shows that the problem of classifying, up to conjugation, the optimal trivial spectrum subspaces of matrices over $\mathbb{D}$ can be reduced to the subproblem of classifying the irreducible ones.
+One difficulty must be pointed out though, and we call it the reconstruction problem.
+If we have irreducible optimal trivial spectrum subspaces of matrices $\mathcal{M}_{1},\dots,\mathcal{M}_{p}$, then although their joint $\mathcal{M}_{1}\vee\dots\vee\mathcal{M}_{p}$ is surely a trivial spectrum subspace, there can be no simple way to prove that it is optimal.
+To see this, consider the mundane case where we take $\mathbb{F}$-linear hyperplanes $H_{1},\dots,H_{n}$ of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$, and we form the joint $H_{1}\vee\cdots\vee H_{n}$.
+Clearly each $H_{i}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{1}(\mathbb{D})$, and their joint can be optimal only if we prove that the greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is $\alpha(n,d)$.
+Yet, as we shall see this result is not known when $\mathbb{F}$ is an arbitrary finite field (unless $d=1$).
+However, if we have an integer $n>0$ and we know that $\alpha(k,d)$ is the greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{k}(\mathbb{D})$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$ then for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(\mathcal{M}_{1},\dots,\mathcal{M}_{p})$ of optimal trivial spectrum subspaces of, respectively, $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, the space $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ is optimal since it is easily observed that
+$n(d-1)+d\dbinom{n}{2}=\sum_{i=1}^{p}\left(n_{i}(d-1)+d\dbinom{n_{i}}{2}\right)+\sum_{1\leq i<j\leq p}dn_{i}n_{j}.$
+1.3 Review of older results
+Now that the main concepts are in place, we can review the known results on our problem. As said in the introduction, these results deal with the special case $\mathbb{D}=\mathbb{F}$ only:
+Theorem 1.3 (Quinlan [9], de Seguins Pazzis [10]).
+Let $n>0$. The greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ is $\dbinom{n}{2}$.
+The structure of the spaces of greatest possible dimension is connected with bilinear forms.
+Let $V$ be an $\mathbb{F}$-vector space and $b$ be a bilinear form on $V$.
+An endomorphism $u$ of $V$ is called $b$-alternating (on the right) whenever $\forall x\in V,\;b(x,u(x))=0$, i.e.
+the bilinear form $(x,y)\in V^{2}\mapsto b(x,u(y))$ is alternating.
+Notation 1.2.
+We denote by $\mathcal{A}_{b}$ the set consisting of all the $b$-alternating endomorphisms of $V$. This is a linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$.
+Assume now that $b$ is nonisotropic, that is $\forall x\in V,\;b(x,x)\neq 0$.
+Then $\mathcal{A}_{b}$ has dimension $\frac{n(n-1)}{2}$ (by the non-degeneracy of $b$), and it is a trivial spectrum subspace (indeed, for all $u\in\mathcal{A}_{b}$, having $u(x)=x$ for some $x\in V\smallsetminus\{0\}$ implies that $0=b(x,u(x))=b(x,x)$).
+It is not difficult to prove that it is also an irreducible one if $n>0$.
+Theorem 1.4 (de Seguins Pazzis [12]).
+Assume that $|\mathbb{F}|>2$.
+Let $V$ be an $n$-dimensional vector space over $\mathbb{F}$, with $n>0$.
+Then the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are the spaces of the form $\mathcal{A}_{b}$ for some nonisotropic bilinear form $b$ on $V$.
+For two such forms $b$ and $b^{\prime}$, the spaces $\mathcal{A}_{b}$ and $\mathcal{A}_{b^{\prime}}$ are equal if and only if $b^{\prime}=\lambda b$ for some $\lambda\in\mathbb{F}^{\times}$, and they are conjugated in $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ if and only if $b^{\prime}$ is congruent to $b$ (i.e., there exists an automorphism $\varphi$ of $V$ and a scalar $\alpha\in\mathbb{F}^{\times}$ such that $b^{\prime}(x,y)=\alpha\,b(\varphi(x),\varphi(y))$ for all $x,y$ in $V$).
+As a consequence, we get the full picture for optimal trivial spectrum subspaces:
 Theorem 1.5.
-Let $\mathbb{D}$ be a division ring with degree $d$ over a field $\mathbb{F}$, and $V$ be a (right) vector space over $\mathbb{D}$ with finite dimension $n>0$.
+Assume that $|\mathbb{F}|>2$.
+Let $V$ be an $n$-dimensional vector space over $\mathbb{F}$, with $n>0$.
+Let $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$, with flag of invariant subspaces denoted by $(V_{0},\dots,V_{p})$.
+Then for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there is a nonisotropic bilinear form $b_{i}$ on $V_{i}/V_{i-1}$, uniquely determined by $S$ up to multiplication with a nonzero scalar, such that $S=\mathcal{A}_{b_{1}}\vee\cdots\vee\mathcal{A}_{b_{p}}$.
+Conversely, given a flag $(V_{0},\dots,V_{p})$ of linear subspaces of $V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ a nonisotropic bilinear form $b_{i}$ on $V_{i}/V_{i-1}$, the set $S:=\mathcal{A}_{b_{1}}\vee\cdots\vee\mathcal{A}_{b_{p}}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ with associated flag $(V_{0},\dots,V_{p})$ of invariant subspaces.
+We can also state the matrix version of the problem, where a square matrix $P\in\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ is called nonisotropic whenever $\forall X\in\mathbb{F}^{n}\smallsetminus\{0\},\;X^{T}PX\neq 0$, and two square matrices $M$ and $M^{\prime}$ are called homocongruent whenever there exist a scalar $\alpha\in\mathbb{F}^{\times}$ and an invertible matrix $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ such that $M^{\prime}=\alpha\,QMQ^{T}$.
+Also, for $n\in\mathbb{N}^{*}$ we denote by $\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ the space of all alternating $n$-by-$n$ matrices (i.e., skew-symmetric matrices with all diagonal entries zero).
+Theorem 1.6.
+Assume that $|\mathbb{F}|>2$.
+  1. (a)
+    Let $n>0$ and $\mathcal{M}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
+Then there exist a partition $n=n_{1}+\cdots+n_{p}$ and nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ such that M(P_1^-1 A_n_1(F)) ∨⋯∨(P_p^-1 A_n_p(F)).
+The partition $(n_{1},\dots,n_{p})$ is then uniquely determined by $n$, as well as the homocongruence classes of the matrices $P_{1},\dots,P_{p}$.
+  2. (b)
+    Conversely, for each partition $n=n_{1}+\cdots+n_{p}$ and each list of nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$, the set $(P_{1}^{-1}\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}))\vee\cdots\vee(P_{p}^{-1}\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F}))$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
+Finally, we can state the solution of the initial problem of classifying the optimal unital affine subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
+Here, the matrix formulation turns out to be the clearer one:
+Theorem 1.7.
+Assume that $|\mathbb{F}|>2$.
+  1. (a)
+    Let $n>0$ and $\mathcal{M}$ be an optimal unital affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
+Then there exist a partition $n=n_{1}+\cdots+n_{p}$ and nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ such that M(P_1+ A_n_1(F)) ∨⋯∨(P_p+ A_n_p(F)).
+The partition $(n_{1},\dots,n_{p})$ is then uniquely determined by $n$, as well as the congruence class of each quadratic form $X\mapsto X^{T}P_{i}X$.
+  2. (b)
+    Conversely for each partition $n=n_{1}+\cdots+n_{p}$ and each list of nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$, the set $(P_{1}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}))\vee\cdots\vee(P_{p}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F}))$ is an optimal unital affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
+Hence, this theorem gives a one-to-one correspondence between the equivalence classes of the optimal unital affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ on the one hand, and the lists of congruence classes of nonisotropic quadratic forms of respective dimensions $n_{1},\dots,n_{p}$ such that $n=\sum_{k=1}^{p}n_{k}$ on the other hand.
+Remark 1.1.
+In the previous classification results, the case of fields with two elements is systematically omitted.
+Indeed, it has been shown through various examples (see section 5.2 of [17]) that a classification theorem for $\mathbb{F}_{2}$ is hopeless.
+1.4 The new results on trivial spectrum subspaces
+We are now ready to state the results on the more general case of a simple $\mathbb{F}$-algebra.
+From now on, we consider a division ring $\mathbb{D}$ with finite dimension $d$ over a central subfield $\mathbb{F}$.
+We start with the statement on the greatest possible dimension for a trivial spectrum (linear) subspace. This generalizes Theorem 1.3, however with a cardinality assumption on $\mathbb{F}$.
+Theorem 1.8.
+Let $\mathbb{D}$ be a division ring with dimension $d$ over a central subfield $\mathbb{F}$, and $V$ be a (right) vector space over $\mathbb{D}$ with finite dimension $n>0$.
 Assume that $|\mathbb{F}|\geq dn$.
-Then the greatest possible dimension for an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or, equivalently, of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is
-$\alpha(n,d)=d\,\dbinom{n}{2}+n(d-1).$
-An immediate consequence is that when $|\mathbb{F}|\geq dn$ the composition problem stated in the end of Section 1.2 has a positive answer in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$: Indeed in that case $|\mathbb{F}|\geq dk$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$, and for every partition $n=n_{1}+\cdots+n_{p}$ the number $\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}$ equals $\alpha(n,d)$.
-Under the same assumption as in Theorem 1.6, the classification of the irreducible optimal spaces is partly given by the following theorem:
-Theorem 1.7 (Partial classification theorem for irreducible optimal $\mathbb{F}$-trivial spectrum subspaces).
+Then the greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or, equivalently, of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is
+$\alpha(n,d)=n(d-1)+d\,\dbinom{n}{2}$.
+An immediate consequence is that when $|\mathbb{F}|\geq dn$ the reconstruction problem stated in the end of Section 1.2 has a positive answer in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$: Indeed in that case $|\mathbb{F}|\geq dk$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$, and for every partition $n=n_{1}+\cdots+n_{p}$ the number $\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}$ equals $\alpha(n,d)$.
+Next, we turn to the structure of the optimal spaces. Here, we consider an $\mathbb{F}$-bilinear form $b:V^{2}\rightarrow\mathbb{F}$ for a $\mathbb{D}$-vector space $V$, and set
+$\mathcal{A}_{b,\mathbb{D}}:=\mathcal{A}_{b}\cap\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$
+(which is much smaller than the full space $\mathcal{A}_{b}$ as we are only considering $\mathbb{D}$-linear endomorphisms).
+Theorem 1.9 (Partial classification theorem for irreducible optimal trivial spectrum subspaces).
 Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$.
   1. (i)
-    If $n=1$ then the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ are the $\mathbb{F}$-linear hyperplanes of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that do not contain $\operatorname{\operatorname{id}}_{V}$.
+    If $n=1$ then the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ are the $\mathbb{F}$-linear hyperplanes of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that do not contain $\operatorname{\operatorname{id}}_{V}$.
   2. (ii)
@@
 Of course, the above two theorems look on the surface as plain generalizations of the case $\mathbb{D}=\mathbb{F}$, and in some sense they are.
-But there is a subtlety: although $\mathcal{A}_{b,\mathbb{D}}$ is always an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, in general it is not an optimal one!
+But there is a subtlety: although $\mathcal{A}_{b,\mathbb{D}}$ is always a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, in general it is not an optimal one!
 Hence it is necessary to understand when optimality happens.
@@
 Denote by $\sigma$ either the identity of $\mathbb{F}$ in cases (i) and (iv), the non-identity automorphism in case (ii), and the quaternionic conjugation in case (iii); we say that $\sigma$ is the standard involution on $\mathbb{D}$.
-In cases (i) and (ii), we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the standard trace of $\mathbb{D}$ over $\mathbb{F}$ (see Section 2.1 for details), and in case (iii) we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the reduced trace $x\mapsto x+\sigma(x)$ (the reduced trace of $x$ is either $2x$ if $x$ belongs to $\mathbb{F}1_{\mathbb{D}}$, otherwise it equals the opposite of the coefficient on $t$ of the minimal polynomial of $x$ over $\mathbb{F}$).
+In cases (i) and (ii), we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the Galois trace of $\mathbb{D}$ over $\mathbb{F}$, and in case (iii) we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the reduced trace $x\mapsto x+\sigma(x)$ (the reduced trace of $x$ is either $2x$ if $x$ belongs to $\mathbb{F}1_{\mathbb{D}}$, otherwise it equals the opposite of the coefficient on $t$ of the minimal polynomial of $x$ over $\mathbb{F}$).
 We say that $(\sigma,e)$ is an associated pair for $\mathbb{D}$ over $\mathbb{F}$ when $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, $\sigma$ is its standard involution, and $e$ is either $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ when $\mathbb{D}$ is of separable type, or an arbitrary nonzero $\mathbb{F}$-linear form on $\mathbb{D}$ otherwise.
-Finally, if $\sigma$ is an antiautomorphism of the $\mathbb{F}$-algebra $\mathbb{D}$, we define a sesquilinear form on $V$ as a mapping $B:V^{2}\rightarrow\mathbb{D}$ that is right-linear and left-$\sigma$-quasilinear, meaning that $B(\alpha x+y,-)=\sigma(\alpha)B(x,-)+B(y,-)$ and $B(-,\alpha x+y)=\alpha B(-,x)+B(-,y)$ for all $\alpha\in\mathbb{D}$ and $x,y$ in $V$.
+Finally, if $\sigma$ is an antiautomorphism of the $\mathbb{F}$-algebra $\mathbb{D}$, we define a $\sigma$-sesquilinear on $V$ as a mapping $B:V^{2}\rightarrow\mathbb{D}$ that is right-linear and left-$\sigma$-quasilinear, meaning that $B(x\alpha+y,-)=\sigma(\alpha)B(x,-)+B(y,-)$ and $B(-,x\alpha+y)=B(-,x)\alpha+B(-,y)$ for all $\alpha\in\mathbb{D}$ and $x,y$ in $V$.
 Two such forms $B$ and $B^{\prime}$ are called equivalent if there exists an automorphism $\varphi$ of the $\mathbb{D}$-vector space $V$ such that $B^{\prime}(\varphi(x),\varphi(y))=B(x,y)$ for all $(x,y)\in V^{2}$.
-Theorem 1.8 (Full classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
+Theorem 1.10 (Full classification theorem for optimal trivial spectrum subspaces).
 Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$.
 Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$.
-For $\mathcal{A}_{b,\mathbb{D}}$ to be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, it is necessary and sufficient that all the following conditions hold:
+For $\mathcal{A}_{b,\mathbb{D}}$ to be an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, it is necessary and sufficient that both of the following conditions hold:
   1. (i)
@@
   2. (ii)
-    Given an associated pair $(\sigma,e)$ for $\mathbb{D}$ over $\mathbb{F}$, there is a $\sigma$-sesquilinear form $B$ on the $\mathbb{D}$-vector space $V$ such that
-    $\forall(x,y)\in V^{2},\;b(x,y)=e(B(x,y)).$
+    Given an associated pair $(\sigma,e)$ for $\mathbb{D}$ over $\mathbb{F}$, there is a $\sigma$-sesquilinear form $B$ on the $\mathbb{D}$-vector space $V$ such that (x,y)V^2,   b(x,y)=e(B(x,y)).
 In that case, $B$ is uniquely determined by $b$.
-Hence, if $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ with associated pair $(\sigma,e)$, and we take any $\sigma$-sesquilinear form $B$ on $V$, the space $\mathcal{A}_{b,\mathbb{D}}$ is an irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$ provided that $B$ is $e$-nonisotropic , meaning that $\forall x\in V,\;e(B(x,x))\neq 0$ (for this very result the provision $|\mathbb{F}|\geq nd$ is not needed, see the proof in Section 5.6).
-As a consequence of Theorems 1.7 and 1.8, and still under the cardinality assumption from Theorem 1.6, we can refine our understanding of the irreducible optimal trivial spectrum subspaces:
+Hence, if $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ with associated pair $(\sigma,e)$, and we take any $\sigma$-sesquilinear form $B$ on $V$, the space $\mathcal{A}_{b,\mathbb{D}}$ is an irreducible trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$ provided that $B$ is $e$-nonisotropic, meaning that $\forall x\in V,\;e(B(x,x))\neq 0$ (for this very result the provision $|\mathbb{F}|\geq nd$ is not needed, see the proof in Section 5.6).
+As a consequence of Theorems 1.9 and 1.10, and still under the cardinality assumption from Theorem 1.8, we can refine our understanding of the irreducible optimal trivial spectrum subspaces:
 Corollary 1.11.
 Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$.
-Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$.
-The space $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ contains an irreducible optimal $\mathbb{F}$-trivial spectrum subspace only if $n=1$.
+The space $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ contains an irreducible optimal trivial spectrum subspace only if $n=1$.
 In case $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, the following result slightly reinforces Theorem 1.9 by taking the $1$-dimensional spaces into account.
@@
 Moreover, for two such forms $b$ and $b^{\prime}$, the equality $\mathcal{A}_{b,\mathbb{D}}=\mathcal{A}_{b^{\prime},\mathbb{D}}$ is satisfied if and only if $b^{\prime}$ is a scalar multiple of $b$.
-The next result deals with the conjugacy class of $\mathcal{A}_{b,\mathbb{D}}$ inside the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+The next result deals with the similarity class of $\mathcal{A}_{b,\mathbb{D}}$ inside the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
 Theorem 1.13.
@@
 Let $B$ and $B^{\prime}$ be $e$-nonisotropic $\sigma$-sesquilinear forms on $V$, and set $b:(x,y)\mapsto e(B(x,y))$ and $b^{\prime}:(x,y)\mapsto e(B^{\prime}(x,y))$.
-Then $\mathcal{A}_{b,\mathbb{D}}$ is conjugated to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ in the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ if and only if $B^{\prime}$ is equivalent to $\alpha B$ as a $\sigma$-sesquilinear form for some $\alpha\in\mathbb{F}^{\times}$.
+Then $\mathcal{A}_{b,\mathbb{D}}$ is similar to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ in the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ if and only if $B^{\prime}$ is equivalent to $\alpha B$ as a $\sigma$-sesquilinear form for some $\alpha\in\mathbb{F}^{\times}$.
 It is a good idea now to express the above results in matrix terms.
@@
 Conversely, let $M\in\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$.
-By the same computation as in the above, we see that the identity $M^{\star}=-M$ leads to $X\in(\mathbb{D}^{n})\mapsto e(X^{\star}MX)$ being additive (its polar form vanishes).
+By the same computation as in the above, we see that the identity $M^{\star}=-M$ leads to $X\in\mathbb{D}^{n}\mapsto e(X^{\star}MX)$ being additive (its polar form vanishes).
 As a consequence, for all $X=\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$, we find
@@
   • If $d\geq 2$ and $\mathbb{D}$ is inseparable over $\mathbb{F}$ then $\sigma=\operatorname{\operatorname{id}}$ and $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is the space of all symmetric matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ whose diagonal entries belong to $\operatorname{\operatorname{Ker}}e$ (in that case the constraint $\forall i\in\mathopen{[\![}1,n\mathclose{]\!]},\;e(m_{i,i})=0$ is unavoidable).
+Finally, two matrices $A$ and $B$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ will be called homo-star-congruent whenever $B=\alpha\,PAP^{\star}$ for some $\alpha\in\mathbb{F}^{\times}$ and some $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$.
 Hence, according to the type of $\mathbb{D}$ over $\mathbb{F}$, we can partially translate the previous results in matrix terms:
@@
   1. (a)
-    The irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces of the form $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ where $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ is $e$-nonisotropic.
+    The irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces of the form $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ where $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ is $e$-nonisotropic.
   2. (b)
-    Given $e$-nonisotropic matrices $P_{1}$ and $P_{2}$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$, the spaces $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ and $P_{2}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ are conjugated under an element of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ if and only if $P_{2}$ is star-congruent to $\alpha P_{1}$ for some $\alpha\in\mathbb{F}^{\times}$ (two matrices $M$ and $M^{\prime}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are star-congruent if and only if there exists $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that $M^{\prime}=QMQ^{\star}$).
+    Given $e$-nonisotropic matrices $P_{1}$ and $P_{2}$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$, the spaces $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ and $P_{2}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ are similar if and only if $P_{2}$ is homo-star-congruent to $P_{1}$.
 At this point, the reader might be puzzled by the hyper-radicial case and doubt the existence of such kinds of optimal trivial spectrum spaces.
@@
 Hence $P$ is $e$-nonisotropic.
-Therefore $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and both $n$ and $p$ can be taken as large as we wish.
-Let us conclude by combining Theorem 1.2 with the previous results, so as to give a full description of the optimal $\mathbb{F}$-trivial spectrum spaces.
+Therefore $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and both $n$ and $p$ can be taken as large as we wish.
+Let us conclude by combining Theorem 1.1 with the previous results, so as to give a full description of the optimal trivial spectrum spaces.
 We distinguish between two cases, whether $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ or not.
@@
 Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$.
-Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
-Then the $\mathcal{S}$-invariant $\mathbb{D}$-linear subspaces form a complete flag $(V_{0},\dots,V_{n})$ of $V$, and $\mathcal{S}=H_{1}\vee\cdots\vee H_{n}$, where for each $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, $H_{i}$ is an $\mathbb{F}$-linear hyperplane of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that does not contain the identity of $V_{i}/V_{i-1}$.
-Conversely, given a complete flag $(V_{0},\dots,V_{n})$ of $V$, and for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$ an $\mathbb{F}$-linear hyperplane $H_{i}$ of $V_{i}/V_{i-1}$ that does not contain the identity of $V_{i}/V_{i-1}$, the space $H_{1}\vee\cdots\vee H_{n}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+Let $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+Then the $S$-invariant $\mathbb{D}$-linear subspaces form a complete flag $(V_{0},\dots,V_{n})$ of $V$, and $S=H_{1}\vee\cdots\vee H_{n}$, where for each $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, $H_{i}$ is an $\mathbb{F}$-linear hyperplane of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that does not contain the identity of $V_{i}/V_{i-1}$.
+Conversely, given a complete flag $(V_{0},\dots,V_{n})$ of $V$, and for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$ an $\mathbb{F}$-linear hyperplane $H_{i}$ of $V_{i}/V_{i-1}$ that does not contain the identity of $V_{i}/V_{i-1}$, the space $H_{1}\vee\cdots\vee H_{n}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
 Here is a restatement in matrix terms:
@@
 Let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$, and assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$.
-The optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces that are conjugated to a space of the form
-$H_{1}\vee\cdots\vee H_{n}=\left\{\begin{bmatrix}\alpha_{1}&&[*]\\ &\ddots&\\ [0]&&\alpha_{n}\end{bmatrix}\mid(\alpha_{1},\dots,\alpha_{n})\in H_{1}\times\cdots\times H_{n}\right\}$
-where $H_{1},\dots,H_{n}$ are $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$.
-Two such spaces $H_{1}\vee\cdots\vee H_{n}$ and $H^{\prime}_{1}\vee\cdots\vee H^{\prime}_{n}$ are conjugated in the algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ if and only if, for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, the hyperplanes $H_{i}$ and $H^{\prime}_{i}$ are conjugated in $\mathbb{D}$.
-Theorem 1.16 (Full classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
+The optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces that are similar to a space of the form $H_{1}\vee\cdots\vee H_{n}$ where $H_{1},\dots,H_{n}$ are $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$.
+Two such spaces $H_{1}\vee\cdots\vee H_{n}$ and $H^{\prime}_{1}\vee\cdots\vee H^{\prime}_{n}$ are similar in the algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ if and only if, for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, the hyperplanes $H_{i}$ and $H^{\prime}_{i}$ are conjugated in $\mathbb{D}$.
+Hence, if $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$ (and the cardinality of $\mathbb{F}$ is large enough) the classification of optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is reduced to the classification of the $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$, up to conjugation.
+The latter classification is directly connected with the classification of conjugacy classes of elements in $\mathbb{D}$ by duality.
+To see this, consider the composite $\operatorname{\operatorname{Tr}}:\mathbb{D}\rightarrow\mathbb{F}$ of the reduced trace of $\mathbb{D}$ over its center $C$ and of the Galois trace $\operatorname{\operatorname{Tr}}_{C/\mathbb{F}}$: this is a nonzero $\mathbb{F}$-linear mapping that satisfies $\operatorname{\operatorname{Tr}}(xy)=\operatorname{\operatorname{Tr}}(yx)$ for all $x,y$ in $\mathbb{D}$.
+Then $a\mapsto[b\mapsto\operatorname{\operatorname{Tr}}(ab)]$ induces an $\mathbb{F}$-linear isomorphism from $\mathbb{D}$ to the dual space of the $\mathbb{F}$-vector space $\mathbb{D}$.
+It is then easily checked that this yields a bijective correspondence between the $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ and the equivalence classes in $\mathbb{D}^{\times}$ for the relation $a\cong b$ defined by $\exists\lambda\in\mathbb{F}^{\times}:\exists u\in\mathbb{D}^{\times}:b=\lambda\,uau^{-1}$.
+The $\mathbb{F}$-linear hyperplanes that do not contain $1_{\mathbb{D}}$ correspond to the elements of $\mathbb{D}$ with non-vanishing trace.
+Finally, by the Skolem-Noether theorem, two elements of $\mathbb{D}$ are conjugated if and only if they have the same minimal polynomial over the center $C$.
+We finish with the case where $\mathbb{D}$ is of quadratic type over $\mathbb{F}$.
+Theorem 1.18.
 Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$.
-Let $V$ be a $\mathbb{D}$-vector space with finite dimension $n>0$, where $|\mathbb{F}|\geq nd$, and $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
-Then the $\mathcal{S}$-invariant $\mathbb{D}$-linear subspaces constitute a (potentially incomplete) flag $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there exists an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$ such that, for $b_{i}:=e\circ B_{i}$,
-$\mathcal{S}=\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}.$
-Moreover, each form $B_{i}$ is uniquely determined by $\mathcal{S}$ up to multiplication by an element of $\mathbb{F}^{\times}$, and the equivalence class of $B_{i}$ is uniquely determined by the conjugacy class of $\mathcal{S}$ in $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ up to multiplication by an element of $\mathbb{F}^{\times}$.
-Conversely, given a (potentially incomplete) flag $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$, with associated $\mathbb{F}$-bilinear form $b_{i}:=e\circ B_{i}$, the space $\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+Let $V$ be a $\mathbb{D}$-vector space with finite dimension $n>0$, where $|\mathbb{F}|\geq nd$, and $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+Then the $S$-invariant $\mathbb{D}$-linear subspaces constitute a (potentially incomplete) flag $(V_{0},V_{1},\dots,V_{p})$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there exists an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$ such that, for $b_{i}:=e\circ B_{i}$,
+$S=\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}.$
+Moreover, each form $B_{i}$ is uniquely determined by $S$ up to multiplication by an element of $\mathbb{F}^{\times}$, and the equivalence class of $B_{i}$ is uniquely determined by the similarity class of $S$ in $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ up to multiplication by an element of $\mathbb{F}^{\times}$.
+Conversely, given a (potentially incomplete) flag $(V_{0},V_{1},\dots,V_{p})$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$, with associated $\mathbb{F}$-bilinear form $b_{i}:=e\circ B_{i}$, the space $\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
 We finish with a restatement in matrix terms:
-Theorem 1.17 (Classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
+Theorem 1.19.
 Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$.
 Let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$.
-For every optimal $\mathbb{F}$-trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, there exists a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{D})$ of $e$-nonisotropic matrices such that
+For every optimal trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, there exists a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{D})$ of $e$-nonisotropic matrices such that
 $\mathcal{M}\simeq P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}).$
-The integers $n_{1},\dots,n_{p}$ are uniquely determined by $n$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the matrix $P_{i}$ is uniquely determined by $\mathcal{M}$ up to star-congruence and multiplication by an element of $\mathbb{F}^{\times}$.
-Conversely, for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of $e$-nonisotropic matrices, the space $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D})$, as well as all its conjugates in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
+The integers $n_{1},\dots,n_{p}$ are uniquely determined by $n$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the matrix $P_{i}$ is uniquely determined by $\mathcal{M}$ up to homo-star-congruence.
+Conversely, for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of $e$-nonisotropic matrices, the space $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D})$, as well as all its conjugates in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
+The corresponding results for the initial problem of optimal unital affine subspaces will be stated and proved in Section 7.
 1.5 The strategy, and intransitive operator spaces
-At this point, we have stated the main results of the present article, which are the ones on trivial spectrum spaces.
-They are far from the only results here, and we will expand on the other results in a moment.
-But before we come to the other results, let us explain the main strategy for proving those on trivial spectrum spaces.
+Let us now explain the main strategy for proving our new results on trivial spectrum spaces.
 So far, we have not succeeded in generalizing the techniques of [10, 12] to the case $d>1$.
 This has led us to reconsider an alternative proof of Theorem 1.4, limited to fields with more than n1n-1 elements, which we discovered ten years ago [13].
-Critical to the proof is the observation that the $\mathbb{F}$-trivial spectrum property implies a lack of transitivity from the said spaces of operators, meaning that if $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum then $x\not\in\mathcal{S}x$ for every nonzero vector $x\in V\smallsetminus\{0\}$, and in particular $\mathcal{S}x\subsetneq V$.
-The key insight is then to introduce the dual operator space $\widehat{\mathcal{S}}$ of $\mathcal{S}$, which consists of all the evaluation mappings
-$\widehat{x}:u\in\mathcal{S}\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
-Then $\widehat{\mathcal{S}}$ is a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$, and every element of $\widehat{\mathcal{S}}$ is nonsurjective and hence has its rank less than $\dim_{\mathbb{F}}V$.
-The key in [19] was to rely upon the theory of primitive spaces of bounded rank matrices, studied first by Atkinson and Lloyd [1, 2] and rediscovered a few years later by Eisenbud and Harris [4].
+Critical to the proof is the observation that the trivial spectrum property implies a lack of transitivity from the said spaces of operators, meaning that if $S$ has trivial spectrum then $x\not\in Sx$ for every nonzero vector $x\in V\smallsetminus\{0\}$, and in particular $Sx\subsetneq V$.
+The key insight is then to introduce the dual operator space $\widehat{S}$ of $S$, which consists of all the evaluation mappings
+$\widehat{x}:u\in S\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
+Then $\widehat{S}$ is a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S,V)$, and every element of $\widehat{S}$ is nonsurjective and hence has its rank less than $\dim_{\mathbb{F}}V$.
+The key in [13] was to rely upon the theory of primitive spaces of bounded rank matrices, studied first by Atkinson and Lloyd [1, 2] and rediscovered a few years later by Eisenbud and Harris [3] (see also the recent [6]).
 Thanks to Atkinson’s classification of primitive matrix spaces that satisfy an extremal condition between the numbers of rows and columns (theorem C of [2]), and more precisely to our own generalization to semi-primitive spaces (see section 5 of [15] and a restatement as theorem 3.1 in [13]), it was then possible to find an easy derivation of Theorem 1.4 in case $|\mathbb{F}|\geq n$ (an unavoidable assumption if one wants to use Atkinson’s theorem).
-Yet, for the case $d>1$ a major issue is that the space $\widehat{\mathcal{S}}$ is far from satisfying Atkinson’s extremal condition between the respective dimensions of the source and target spaces, even if $\mathcal{S}$ is optimal; hence nothing can be obtained from the theory of primitive spaces of bounded rank matrices.
-Nevertheless, a version of Atkinson’s theorem (theorem 5.3 of [21]) can be entirely reformulated in the viewpoint of spaces of operators that lack transitivity, as we shall now see.
+Yet, for the case $d>1$ a major issue is that the space $\widehat{S}$ is far from satisfying Atkinson’s extremal condition between the respective dimensions of the source and target spaces, even if $S$ is optimal; hence nothing can be obtained from the theory of primitive spaces of bounded rank matrices.
+Nevertheless, we discovered that a version of Atkinson’s theorem (theorem 5.3 of [15]) can be entirely reformulated in the viewpoint of spaces of operators that lack transitivity, as we shall now see.
 A few basic definitions and notation will help:
-Notation 1.4.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $V^{\prime}$ be a $\mathbb{D}$-linear subspace of $V$.
-Given a subset $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, we denote by $\mathcal{S}^{V^{\prime}}:=\mathcal{S}\cap\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$ the subset consisting of the elements of $\mathcal{S}$ that map into $V^{\prime}$.
+Notation 1.5.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $V^{\prime}$ be a $\mathbb{D}$-linear subspace of $V$.
+Given a subset $S$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, we denote by $S^{V^{\prime}}:=S\cap\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$ the subset consisting of the elements of $S$ that map into $V^{\prime}$.
 Definition 1.6.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is:
-  • intransitive whenever $\forall x\in U,\;\mathcal{S}x\neq V$ ;
-  • deeply intransitive whenever it is intransitive and, for every nonzero linear subspace $V^{\prime}$ of $V$, the set $\mathcal{S}^{V^{\prime}}$ is intransitive as an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$.
-In any case, the transitive rank $\operatorname{\operatorname{trk}}(\mathcal{S})$ of $\mathcal{S}$ is defined as the greatest possible dimension $\dim_{\mathbb{F}}(\mathcal{S}x)$ when $x$ ranges over $U$.
-Note that $\mathcal{S}$ is intransitive only if $V\neq\{0\}$, and in the definition of deep intransitivity we can replace the condition that $\mathcal{S}$ be intransitive with the one that $V\neq\{0\}$.
-It is obvious that if $\mathcal{S}$ is deeply intransitive then so is $\mathcal{S}^{V^{\prime}}$ for every nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $S$ is:
+  • intransitive whenever $\forall x\in U,\;Sx\neq V$ ;
+  • deeply intransitive whenever it is intransitive and, for every nonzero linear subspace $V^{\prime}$ of $V$, the set $S^{V^{\prime}}$ is intransitive as an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$.
+In any case, the transitive rank $\operatorname{\operatorname{trk}}(S)$ of $S$ is defined as the greatest possible dimension $\dim_{\mathbb{F}}(Sx)$ when $x$ ranges over $U$.
+Note that $S$ is intransitive only if $V\neq\{0\}$, and in the definition of deep intransitivity we can replace the condition that $S$ be intransitive with the one that $V\neq\{0\}$.
+It is obvious that if $S$ is deeply intransitive then so is $S^{V^{\prime}}$ for every nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$.
 Constructing large deeply intransitive spaces can be achieved by using nonzero bilinear forms, as we shall now explain.
 Definition 1.7.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
-An $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$ is called an alternator of $\mathcal{S}$ whenever
-$\forall x\in U,\;\forall f\in\mathcal{S},\;b(x,f(x))=0.$
-The set of all such forms is denoted by $\operatorname{\operatorname{Alt}}(\mathcal{S})$ (it is obviously a linear subspace of the set of all $\mathbb{F}$-bilinear forms on $U\times V$).
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+An $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$ is called an alternator of $S$ whenever
+$\forall x\in U,\;\forall f\in S,\;b(x,f(x))=0.$
+The set of all such forms is denoted by $\operatorname{\operatorname{Alt}}(S)$ (it is obviously a linear subspace of the set of all $\mathbb{F}$-bilinear forms on $U\times V$).
 For an $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$, recall that the left radical of $b$ is the $\mathbb{F}$-linear subspace
@@
 Proposition 1.20.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
   1. (a)
-    If $\mathcal{S}$ has a nonzero alternator then $\mathcal{S}$ is intransitive.
+    If $S$ has a nonzero alternator then $S$ is intransitive.
   2. (b)
-    If $\mathcal{S}$ has a right-nondegenerate alternator and $V\neq\{0\}$ then $\mathcal{S}$ is deeply intransitive.
+    If $S$ has a right-nondegenerate alternator and $V\neq\{0\}$ then $S$ is deeply intransitive.
 Proof.
-Assume that $\mathcal{S}$ has a nonzero alternator $b$.
-It follows by polarizing that $\forall(x,x^{\prime})\in U^{2},\;\forall f\in\mathcal{S},\;b(x,f(x^{\prime}))+b(x^{\prime},f(x))=0$, from which we observe that all the elements of $\mathcal{S}$ map $\operatorname{\operatorname{Lrad}}(b)$ into $\operatorname{\operatorname{Rrad}}(b)$.
+Assume that $S$ has a nonzero alternator $b$.
+It follows by polarizing that $\forall(x,x^{\prime})\in U^{2},\;\forall f\in S,\;b(x,f(x^{\prime}))+b(x^{\prime},f(x))=0$, from which we observe that all the elements of $S$ map $\operatorname{\operatorname{Lrad}}(b)$ into $\operatorname{\operatorname{Rrad}}(b)$.
 The latter is not the full space $V$ because $b$ is nonzero.
 Now, let $x\in U$.
-If $x\in\operatorname{\operatorname{Lrad}}(b)$ then $\mathcal{S}x\subseteq\operatorname{\operatorname{Rrad}}(b)\subset V$.
-Otherwise $\mathcal{S}x$ is included in $\operatorname{\operatorname{Ker}}b(x,-)$, an $\mathbb{F}$-linear hyperplane of $V$.
-Hence $\mathcal{S}$ is intransitive.
-Assume now that $\mathcal{S}$ has a right-nondegenerate alternator $b$ and $V\neq\{0\}$.
+If $x\in\operatorname{\operatorname{Lrad}}(b)$ then $Sx\subseteq\operatorname{\operatorname{Rrad}}(b)\subset V$.
+Otherwise $Sx$ is included in $\operatorname{\operatorname{Ker}}b(x,-)$, an $\mathbb{F}$-linear hyperplane of $V$.
+Hence $S$ is intransitive.
+Assume now that $S$ has a right-nondegenerate alternator $b$ and $V\neq\{0\}$.
 Let $V^{\prime}$ be a nonzero $\mathbb{D}$-linear subspace of $V$.
-The restriction $b^{\prime}$ of $b$ to $U\times V^{\prime}$ is nonzero because $b$ is right-nondegenerate, and obviously $b^{\prime}$ is an alternator of $\mathcal{S}^{V^{\prime}}$.
-Hence $\mathcal{S}^{V^{\prime}}$ is intransitive, by point (a).
+The restriction $b^{\prime}$ of $b$ to $U\times V^{\prime}$ is nonzero because $b$ is right-nondegenerate, and obviously $b^{\prime}$ is an alternator of $S^{V^{\prime}}$.
+Hence $S^{V^{\prime}}$ is intransitive, by point (a).
 ∎
 Hence, whenever we take a right-nondegenerate bilinear form $b:U\times V\rightarrow\mathbb{F}$ and $V\neq\{0\}$, every $\mathbb{F}$-linear subspace of $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive.
-Atkinson’s theorem can be essentially summed up in saying that all large enough operator spaces are of this kind.
+When $\mathbb{D}=\mathbb{F}$, Atkinson’s theorem can be essentially summed up in saying that all large enough operator spaces are of this kind.
 Let us now be more precise:
 Theorem 1.21 (Atkinson’s theorem for deeply intransitive operator spaces).
-Let $U$ and $V$ be finite-dimensional $\mathbb{F}$-vector spaces and $\mathcal{S}$ be a deeply intransitive linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$.
+Let $U$ and $V$ be finite-dimensional $\mathbb{F}$-vector spaces and $S$ be a deeply intransitive linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$.
 Set $n:=\dim V$ and assume that $|\mathbb{F}|\geq n$.
@@
   1. (a)
-    $\dim\mathcal{S}\leq\frac{n(n-1)}{2}$, and even $\dim\mathcal{S}\leq\frac{(n-1)(n-2)}{2}$ if $\operatorname{\operatorname{trk}}(\mathcal{S})<n-1$.
+    $\dim S\leq\frac{n(n-1)}{2}$, and even $\dim S\leq\frac{(n-1)(n-2)}{2}$ if $\operatorname{\operatorname{trk}}(S)<n-1$.
   2. (b)
-    If $\dim\mathcal{S}\geq\frac{n(n-1)}{2}-(n-3)$ then $\mathcal{S}$ has a right-nondegenerate alternator $b$, and its alternator space $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$.
-As an extension of point (b), we note, with the same polarization argument as in the proof of Proposition 1.18, that all the operators in $\mathcal{S}$ then vanish on the left radical $\operatorname{\operatorname{Lrad}}(b)$, and hence $\mathcal{S}$ can naturally be identified with a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U/\operatorname{\operatorname{Lrad}}(b),V)$, for which the form induced by $b$ on $(U/\operatorname{\operatorname{Lrad}}(b))\times V$ is a nondegenerate alternator.
+    If $\dim S\geq\frac{n(n-1)}{2}-(n-3)$ then $S$ has a right-nondegenerate alternator $b$, and $\operatorname{\operatorname{Alt}}(S)=\mathbb{F}b$.
+As an extension of point (b), we note, with the same polarization argument as in the proof of Proposition 1.20, that all the operators in $S$ then vanish on the left radical $\operatorname{\operatorname{Lrad}}(b)$, and hence $S$ can naturally be identified with a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U/\operatorname{\operatorname{Lrad}}(b),V)$, for which the form induced by $b$ on $(U/\operatorname{\operatorname{Lrad}}(b))\times V$ is a nondegenerate alternator.
 This fundamentally reduces the situation to the one where $\dim U=\dim V$, in which case we could go further and simply consider spaces of endomorphisms of $V$.
@@
 Primitivity consists of as many as four conditions that roughly mean that the rank constraint cannot be easily localized.
-In contrast, here we replace all four assumptions in the definition of primitivity by a single assumption, which is deep intransitivity (plain intransitivity is of course a far too weak hypothesis, allowing, e.g., $\mathcal{S}$ to be the space of all operators from $U$ to a given linear hyperplane of $V$).
-The deep intransitivity assumption is a dual counterpart of what Atkinson and Lloyd [1] called the row condition for matrix spaces with bounded rank.
-The row condition for such a matrix space $\mathcal{M}\subseteq\operatorname{\operatorname{M}}_{n,p}(\mathbb{F})$ means that, whenever we have invertible matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ and $Q\in\operatorname{\operatorname{GL}}_{p}(\mathbb{F})$, and integers $r\in\mathopen{[\![}1,n\mathclose{]\!]}$ and $s\in\mathopen{[\![}1,p\mathclose{]\!]}$ such that every matrix $M$ of $\mathcal{M}$ looks like
+In contrast, here we replace all four assumptions in the definition of primitivity by a single assumption, which is deep intransitivity (plain intransitivity is of course a far too weak hypothesis, allowing, e.g., $S$ to be the space of all operators from $U$ to a given linear hyperplane of $V$).
+The deep intransitivity assumption is a dual counterpart of what Atkinson and Lloyd [1] called the row property for matrix spaces with bounded rank.
+The row property for such a matrix space $\mathcal{M}\subseteq\operatorname{\operatorname{M}}_{n,p}(\mathbb{F})$ means that, whenever we have invertible matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ and $Q\in\operatorname{\operatorname{GL}}_{p}(\mathbb{F})$, and integers $r\in\mathopen{[\![}1,n\mathclose{]\!]}$ and $s\in\mathopen{[\![}1,p\mathclose{]\!]}$ such that every matrix $M$ of $\mathcal{M}$ looks like
 $M=P\begin{bmatrix}[?]_{r\times s}&K(M)\\ [?]_{(n-r)\times s}&[0]_{(n-r)\times(p-s)}\end{bmatrix}Q,$
 all the matrices in the extracted matrix space $K(\mathcal{M})$ have their rank less than $r$ (the number of rows).
-Now, if we take an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, consider its dual operator space $\widehat{\mathcal{S}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$, and represent the latter by a matrix space $\mathcal{M}$ in bases of $\mathcal{S}$ and $V$, the row condition on $\mathcal{M}$ is equivalent to the fact that whenever we take a nonzero linear subspace $\mathcal{S}^{\prime}$ of $\mathcal{S}$ and we know that all the operators in $\mathcal{S}^{\prime}$ map into some nonzero $\mathbb{F}$-linear subspace $V^{\prime}$ of $V$, then the rank of $s^{\prime}\in\mathcal{S}^{\prime}\mapsto s^{\prime}(x)\in V^{\prime}$ is less than the dimension of $V^{\prime}$ for all $x\in U$, which is exactly saying that $\mathcal{S}^{\prime}$ is intransitive as a linear subspace of mappings from $U$ to $V^{\prime}$.
-Hence, $\mathcal{M}$ satisfies the row condition if and only if $\mathcal{S}$ is deeply intransitive.
-Among the four conditions that define primitivity, two are worth mentioning:
+Now, if we take an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, consider its dual operator space $\widehat{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S,V)$, and represent the latter by a matrix space $\mathcal{M}$ in bases of $S$ and $V$, the row condition on $\mathcal{M}$ is equivalent to the fact that whenever we take a nonzero linear subspace $S^{\prime}$ of $S$ and we know that all the operators in $S^{\prime}$ map into some nonzero $\mathbb{F}$-linear subspace $V^{\prime}$ of $V$, then the rank of $s^{\prime}\in S^{\prime}\mapsto s^{\prime}(x)\in V^{\prime}$ is less than the dimension of $V^{\prime}$ for all $x\in U$, which is exactly saying that $S^{\prime}$ is intransitive as a linear subspace of mappings from $U$ to $V^{\prime}$.
+Hence, $\mathcal{M}$ has the row property if and only if $S$ is deeply intransitive.
+Among the four conditions that define primitivity under the Atkinson-Lloyd definition, two are worth mentioning:
 Definition 1.8.
-An operator space $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ between $\mathbb{F}$-vector spaces $U$ and $V$ is called:
-  • source-reduced when $\underset{f\in\mathcal{S}}{\bigcap}\operatorname{\operatorname{Ker}}f=\{0\}$, i.e., no nonzero vector is annihilated by all the operators in $\mathcal{S}$;
-  • target-reduced when $\underset{f\in\mathcal{S}}{\sum}\operatorname{\operatorname{Im}}f=V$, i.e., no proper linear subspace of $V$ includes the range of all the operators in $\mathcal{S}$.
-Finally, $\mathcal{S}$ is called reduced when it is both source-reduced and target-reduced.
-Atkinson’s theorem is frequently formulated by adding the provision that the operator space $\mathcal{S}$ under consideration is reduced.
+An operator space $S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ between $\mathbb{F}$-vector spaces $U$ and $V$ is called:
+  • source-reduced when $\underset{f\in S}{\bigcap}\operatorname{\operatorname{Ker}}f=\{0\}$, i.e., no nonzero vector is annihilated by all the operators in $S$;
+  • target-reduced when $\underset{f\in S}{\sum}\operatorname{\operatorname{Im}}f=V$, i.e., no proper linear subspace of $V$ includes the range of all the operators in $S$.
+Finally, $S$ is called reduced when it is both source-reduced and target-reduced.
+Atkinson’s theorem is frequently formulated by adding the provision that the operator space $S$ under consideration is reduced (see e.g., [15]).
 However, this assumption is not necessary for the formulation in terms of intransitive spaces.
-Indeed, on the one hand the dual operator space $\widehat{\mathcal{S}}$ of an operator space $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}(U,V)$ is automatically source-reduced (as only the zero operator can be annihilated by all the evaluation mappings).
-On the other hand, if $\mathcal{S}$ is not target-reduced we can take the sum $V^{\prime}$ of all the ranges of the operators in $\mathcal{S}$, then observe that $\widehat{\mathcal{S}}$ can be viewed as a reduced subspace of $\operatorname{\operatorname{Hom}}(\mathcal{S},V^{\prime})$; applying point (a) in Atkinson’s theorem then readily shows that $\dim\mathcal{S}\leq\frac{\dim V^{\prime}(\dim V^{\prime}-1)}{2}\cdot$
+Indeed, on the one hand the dual operator space $\widehat{S}$ of an operator space $S\subseteq\operatorname{\operatorname{Hom}}(U,V)$ is automatically source-reduced (as only the zero operator can be annihilated by all the evaluation mappings).
+On the other hand, if $S$ is not target-reduced we can take the sum $V^{\prime}$ of all the ranges of the operators in $S$, then observe that $\widehat{S}$ can be viewed as a reduced subspace of $\operatorname{\operatorname{Hom}}(S,V^{\prime})$; applying point (a) in Atkinson’s theorem then readily shows that $\dim S\leq\frac{\dim V^{\prime}(\dim V^{\prime}-1)}{2}\cdot$
 Atkinson’s theorem deals with vector spaces over $\mathbb{F}$, but here we need to extend the situation to $\mathbb{F}$-linear subspaces of $\mathbb{D}$-linear mappings.
-In that context, it makes no sense to try to extend Atkinson’s formulation, as by representing the operators of $\widehat{\mathcal{S}}$ by matrices with entries in $\mathbb{F}$, we forget that the operators in $\mathcal{S}$ are $\mathbb{D}$-linear, and not only $\mathbb{F}$-linear.
-The correct way to generalize the results seems to be as a theorem on intransitive operator spaces.
+In that context, it makes no sense to try to extend Atkinson’s formulation, as by representing the operators of $\widehat{S}$ by matrices with entries in $\mathbb{F}$, we miss the critical fact that the operators in $S$ are $\mathbb{D}$-linear, and not only $\mathbb{F}$-linear.
+The correct way to generalize the result seems to be as a theorem on intransitive operator spaces.
 We now state the generalization:
 Theorem 1.22 (Main theorem on deeply intransitive operator spaces).
-Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces and $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces and $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
 Set $n:=\dim_{\mathbb{D}}V$ and assume that $|\mathbb{F}|\geq nd$.
@@
   1. (a)
-    $\dim\mathcal{S}\leq\alpha(n,d)$.
+    $\dim S\leq\alpha(n,d)$.
   2. (b)
-    $\dim\mathcal{S}\leq\alpha(n,d)-n+\max(2-d,0)$ if $\operatorname{\operatorname{trk}}(\mathcal{S})<nd-1$.
+    $\dim S\leq\alpha(n,d)-n+\max(2-d,0)$ if $\operatorname{\operatorname{trk}}(S)<nd-1$.
   3. (c)
-    If $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ then $\mathcal{S}$ has a right-nondegenerate alternator $b$, and its alternator space is $1$-dimensional.
+    If $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$ then $S$ has a right-nondegenerate alternator $b$, and $\operatorname{\operatorname{Alt}}(S)=\mathbb{F}b$.
 Remark 1.2.
-Due to point (a), the inequality $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ of point (c) requires that $n\geq 2$.
+Due to point (a), the inequality $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$ of point (c) requires that $n\geq 2$, and even that $n\geq 3$ if $d=1$.
 The next theorem further elucidates the structure of the alternators:
 Theorem 1.23.
-Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
-Set $n:=\dim_{\mathbb{D}}V$, assume that $|\mathbb{F}|\geq nd$, that $\mathcal{S}$ has a right-nondegenerate alternator $b$ and that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
-Then $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and for any associated pair $(\sigma,e)$ the form $b$ equals $(x,y)\mapsto e(B(x,y))$ for a unique right-nondegenerate $\sigma$-sesquilinear form $B$ on $U\times V$, and of course $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+Set $n:=\dim_{\mathbb{D}}V$, assume that $|\mathbb{F}|\geq nd$, that $S$ has a right-nondegenerate alternator $b$ and that $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$.
+Then $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and for any associated pair $(\sigma,e)$ the form $b$ equals $(x,y)\mapsto e(B(x,y))$ for a unique right-nondegenerate $\sigma$-sesquilinear form $B$ on $U\times V$, and of course $S\subseteq\mathcal{A}_{b,\mathbb{D}}$.
 Here is a converse statement:
@@
 Having these results, it will be fairly easy to derive the ones on optimal irreducible trivial spectrum subspaces, by adapting the method of [13].
-The main difficulty is that although a trivial spectrum subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (with $V$ nonzero) is obviously intransitive (since $\mathcal{S}x$ can never contain $x$ when $x$ is a nonzero vector of $V$), it is however not deeply intransitive in general.
+The main difficulty is that although a trivial spectrum subspace $S$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (with $V$ nonzero) is obviously intransitive (since $Sx$ can never contain $x$ when $x$ is a nonzero vector of $V$), it is however not deeply intransitive in general.
 A classical way of obtaining the deep intransitivity is through the following property, which is essentially the analogue of Atkinson and Lloyd’s condition (C) in the definition of a primitive space:
 Definition 1.9.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
-We say that $\mathcal{S}$ is weakly primitively intransitive whenever there is no $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi\mathcal{S})\leq\operatorname{\operatorname{tr}}(\mathcal{S})-d$.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+We say that $S$ is weakly primitively intransitive whenever there is no $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi S)\leq\operatorname{\operatorname{tr}}(S)-d$.
 In fact, for most applications we will prefer the following stronger property, which is easier to work with:
 Definition 1.10.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
-We say that $\mathcal{S}$ is primitively intransitive whenever it is intransitive and there is no nonzero linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ is intransitive.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+We say that $S$ is primitively intransitive whenever it is intransitive and there is no nonzero linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ is intransitive.
 The following implication is easy to see:
@@
 Proof.
-Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces, $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that there exists a $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi\mathcal{S})\leq\operatorname{\operatorname{tr}}(\mathcal{S})-d$.
-If $\mathcal{S}$ is intransitive, then we deduce that $\operatorname{\operatorname{trk}}(\pi\mathcal{S})<\dim_{\mathbb{F}}(V/V^{\prime})$ and hence $\pi\mathcal{S}$ is intransitive as well: this shows that $\mathcal{S}$ is not primitively intransitive.
+Let $U$ and $V$ be $\mathbb{D}$-vector spaces, $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that there exists a $1$-dimensional $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$ satisfies $\operatorname{\operatorname{trk}}(\pi S)\leq\operatorname{\operatorname{tr}}(S)-d$.
+If $S$ is intransitive, then we deduce that $\operatorname{\operatorname{trk}}(\pi S)<\dim_{\mathbb{F}}(V/V^{\prime})$ and hence $\pi S$ is intransitive as well: this shows that $S$ is not primitively intransitive.
 ∎
-A key insight of Atkinson and Lloyd can be rephrased in saying that, provided that the field $\mathbb{F}$ has enough elements, every weakly primitively intransitive space is deeply intransitive.
-The primitively intransitivity condition has the shortcoming that it is generally not inherited by $\mathcal{S}^{V^{\prime}}$ when $V^{\prime}$ is a linear subspace of $V$, in contrast with the deep intransitivity condition.
+A key insight of Atkinson and Lloyd can be rephrased in saying that, provided that the field $\mathbb{F}$ has enough elements, every weakly primitively intransitive space is deeply intransitive (in their terminology, this is the statement that every matrix space that satisfies condition (C) has the row property).
+The primitive intransitivity condition has the shortcoming that it is generally not inherited by $S^{V^{\prime}}$ when $V^{\prime}$ is a linear subspace of $V$, in contrast with the deep intransitivity condition.
 However, it is the most convenient condition when applying the previous theorem.
@@
 1.6 Structure of the article
-As hinted at in the previous paragraph, our focus is now shifting from the trivial spectrum subspaces, which are the prime motivation for this article, to the deeply intransitive operator spaces, which can be viewed as a more general topic.
-We expect that studying the later will yield many applications in the future, just like Atkinson’s theorem has found many unexpected applications in the recent past (see e.g.
-[19, 25]).
+As hinted at in the previous paragraph, our focus, which had already shifted from unital affine subspaces to trivial spectrum linear subspaces, is now shifting from trivial spectrum subspaces to deeply intransitive operator spaces, a more profound and general problem.
+We expect that studying the later will yield many applications in the future, just like Atkinson’s theorem has found many unexpected applications in the recent past (see e.g., [13, 19]).
 In an early version of the present article, we wrote a proof of the result on trivial spectrum subspaces alone.
-Surprising as it may seem, we found later that the result on deeply intransitive operator spaces is easier to prove.
-The icing on the cake is that it also holds more potential for future applications.
+Surprising as it may seem, we found later that the result on deeply intransitive operator spaces was much easier to prove.
+The icing on the cake is that it also holds far more potential for future applications.
 Let us now discuss the structure of the remainder of the article.
@@
 The next section (Section 4) introduces the key technical tools to study the intransitive operator spaces: there we make no claims of originality, as we will use the generic matrix machinery that allowed Atkinson to prove his theorem, a machinery which has already demonstrated its remarkable efficiency for such problems.
-The main results on intransitive operator spaces are obtained in Section 5: there, the technical details for a general division ring $\mathbb{D}$ are substantially more difficult than in Atkinson’s article.
+The main results on intransitive operator spaces are obtained in Section 5: there, the technical details for a general division ring $\mathbb{D}$ are much more difficult than in Atkinson’s article.
 In Section 6 we derive all the main results on optimal trivial spectrum subspaces from the ones on intransitive operator spaces.
-The remaining sections are devoted to applications of trivial spectrum subspaces.
-In Section 7, we obtain the greatest possible dimension for an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ with rank at least $r$ (Section 7.1), and in Section 7.2 we characterize the spaces with the greatest possible dimension when $n=p=r$ (i.e., affine subspaces of nonsingular matrices).
-The characterization for a general $r$ is very difficult (see [16] for the case $d=1$, which takes no less than twenty-three pages of proof), so we believe the reader will be relieved to learn that this issue is postponed to a subsequent article.
-The last section is devoted to the applications of $\mathbb{F}$-trivial spectrum subspaces to subspaces of $\mathbb{F}$-diagonalisable matrices.
-There, we successfully adapt the method featured in [23], and we also give a brand new result for real vector spaces of complex diagonalisable matrices.
-This new application involves the main theorem on deeply intransitive operator spaces in a case where the space under consideration has its dimension slightly lower than the critical one.
-We finish with an adaptation to spaces of semi-simple endomorphisms and with considerations on spaces of endomorphisms in which only the zero endomorphism is nilpotent.
+Finally, in Section 7 we conclude the study by solving the initial problem of classifying the optimal unital $\mathbb{F}$-affine subspaces of a simple $\mathbb{F}$-algebra, up to equivalence.
+1.7 Potential applications
+There are many applications of the results featured in this article, but due to a lack of space we have decided to keep them out of this article. However, we can quickly discuss two of them.
+The first application is towards affine subspaces of operators with rank bounded below.
+In this, one takes finite-dimensional $\mathbb{D}$-vector spaces $U$ and $V$, together with a positive integer $r\leq\min(\dim_{\mathbb{D}}U,\dim_{\mathbb{D}}V)$, and one asks about the greatest possible dimension for an $\mathbb{F}$-affine subspace of elements of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ in which every operator has rank at least $r$, and how to classify the spaces of greatest possible dimension.
+The special case $r=\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V$ is essentially the problem of unital affine subspaces in the simple $\mathbb{F}$-algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(U)$.
+Deriving the greatest possible dimension in the general case is not difficult, but classifying the optimal spaces is highly non-trivial: see [11] for the special case where $\mathbb{D}=\mathbb{F}$.
+A second potential application is towards $\mathbb{F}$-linear subspaces of $\mathbb{F}$-diagonalisable elements of a simple $\mathbb{F}$-algebra $\mathcal{A}$.
+There, an element $a\in\mathcal{A}$ is called $\mathbb{F}$-diagonalisable whenever its minimal polynomial over $\mathbb{F}$ splits with simple roots, or equivalently when $x\mapsto ax$ is a diagonalisable endomorphism of the $\mathbb{F}$-vector space $\mathcal{A}$.
+When $\mathcal{A}=\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ this problem has already been studied, in particular over the field of real numbers: for an arbitrary field $\mathbb{F}$, any $\mathbb{F}$-linear subspace of $\mathbb{F}$-diagonalisable elements of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ has dimension at most $\frac{n(n+1)}{2}$, with equality attained only for conjugates of the space $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ of all symmetric matrices (and if $n\geq 2$ such spaces exist only if $\mathbb{F}$ is formally real and Pythagorean).
+A connection with the problem of trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ was discovered and exploited in [18]: In a sequel to this article we will use this connection to generalize the study of linear subspaces of $\mathbb{F}$-diagonalisable elements in any simple $\mathbb{F}$-algebra, and in particular we will obtain the structure of the $\mathbb{R}$-linear subspaces of $\mathbb{R}$-diagonalisable matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$: it will turn out that the optimal spaces are the conjugates of the space $\operatorname{\operatorname{H}}_{n}(\mathbb{C})$ of all Hermitian matrices.
+There is also a quaternionic analogue of this result.
 2 On $\mathbb{D}$-vector spaces seen as $\mathbb{F}$-vector spaces
-2.1 Traces
-Throughout, we will need to consider $\mathbb{F}$-linear forms on $\mathbb{D}$.
-There is a canonical choice of such a linear form: the trace of $\mathbb{D}$ over $\mathbb{F}$, denote by $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ (with capital T), maps every $a\in\mathbb{D}$ to the trace of the $\mathbb{F}$-endomorphism $x\mapsto ax$ of $\mathbb{D}$.
-Consequently, it is an $\mathbb{F}$-linear form that enjoys the property that $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ab)=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ba)$ for all $a,b$ in $\mathbb{D}$.
-Let $a\in\mathbb{D}$.
-Denoting by $\pi=t^{n}-\alpha t^{n-1}+\cdots$ the minimal polynomial of $a$ in the $\mathbb{F}$-algebra $\mathbb{D}$, it is easily seen that $n$ divides $d$ and that $x\mapsto ax$ is the direct sum of $\frac{d}{n}$ cyclic endomorphisms of $\mathbb{F}$-vector spaces with minimal polynomial $\pi$, which yields $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)=\frac{d}{n}\,\alpha.$ Applying the same line of reasoning to $x\mapsto xa$ yields that $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)$ is also the trace of the $\mathbb{F}$-linear endomorphism $x\mapsto xa$.
-Moreover $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(\varphi(a))=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(a)$ for every automorphism $\varphi$ of the $\mathbb{F}$-algebra $\mathbb{D}$.
-Alas, in some cases $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ might be useless.
-This is the case for example if $\mathbb{D}$ is a totally radicial extension of $\mathbb{F}$: in that case $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ is identically zero.
-The same holds if $\mathbb{D}$ is a quaternion algebra and $\operatorname{\operatorname{char}}(\mathbb{F})=2$: in that case the minimal polynomial of every element of $\mathbb{D}$ has degree $1$ or $2$, and hence the trace of such an element is either $2$ times or $4$ times the trace of this polynomial, so it is zero in any case.
-If however $\operatorname{\operatorname{char}}(\mathbb{F})=0$ then $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}$ is nonzero and we can safely use $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ as a nonzero linear form on $\mathbb{D}$.
+From now on, $\mathbb{D}$ denotes a division algebra that contains $\mathbb{F}$ as a central subfield, and we denote by $d$ the dimension of $\mathbb{D}$ over $\mathbb{F}$.
 2.1 The socle of an $\mathbb{F}$-linear subspace
@@
 2.3 Multiplicative quadratic forms on division algebras
-Our study will require the following result.
-It can be extracted from the general theory of composition algebras, but since we are in the very specific context of a division ring $\mathbb{D}$, the proof can be considerably shortened.
-We include a proof below for non-specialists, but advise the reader to skip it at first reading.
+Our study will require the following classical result from the theory of composition algebras.
 Theorem 2.3.
@@
 Then $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, and $q:x\mapsto x\sigma(x)$ for the standard involution $\sigma$ of $\mathbb{D}$ over $\mathbb{F}$.
-Proof.
-Throughout we consider the polar form $b:(x,y)\in\mathbb{D}^{2}\mapsto q(x+y)-q(x)-q(y)$ of $q$, and note that the multiplicativity of $q$ yields
-$\forall(a,x,y)\in\mathbb{D}^{3},\;b(ax,ay)=q(a)\,b(x,y)=b(xa,ya).$ (2)
-Assume first that $q$ is degenerate.
-Then $\operatorname{\operatorname{char}}(\mathbb{F})=2$.
-We start by proving that $q$ is totally degenerate (i.e., additive).
-Let us consider the radical $R$ of $b$.
-We have $R\neq\{0\}$ by assumption.
-Let us choose $x\in R\smallsetminus\{0\}$.
-For all $a\in\mathbb{D}\smallsetminus\{0\}$ we have
-$\forall y\in\mathbb{D},\;b(ax,ay)=q(a)\,b(x,y)=0,$
-and since $\{ay\mid y\in\mathbb{D}\}=\mathbb{D}$ we deduce that $ax\in R$. Hence $R=\mathbb{D}$, meaning that $q$ is additive.
-Next, for all $x,y$ in $\mathbb{D}$ we find
-q(xyyx)=q(xy)q(yx)=q(x)q(y)q(y)q(x)=0,q(xy-yx)=q(xy)-q(yx)=q(x)q(y)-q(y)q(x)=0,
-and hence xyyx=0xy-yx=0 because $q$ is nonisotropic. Hence $\mathbb{D}$ is commutative.
-Let $x\in\mathbb{D}$. Then
-$q(x^{2}-q(x))=q(x^{2})-q(q(x).1_{\mathbb{D}})=q(x)^{2}-q(x)^{2}q(1_{\mathbb{D}})=0,$
-and hence $x^{2}=q(x)\in\mathbb{F}$ because $q$ is nonisotropic.
-Therefore $\mathbb{D}$ is a hyper-radicial extension of $\mathbb{F}$, and $q(x)=x\sigma(x)$ for all $x\in\mathbb{D}$, where $\sigma=\operatorname{\operatorname{id}}_{\mathbb{D}}$ is the sole involution of the $\mathbb{F}$-algebra $\mathbb{D}$.
-Now, we deal with the case where $q$ is nondegenerate.
-We introduce the trace map $\operatorname{\operatorname{tr}}:x\in\mathbb{D}\mapsto b(1,x)\in\mathbb{F}$ and note that $\operatorname{\operatorname{tr}}(1)=2$.
-The key point consists in proving the quadratic identity
-$\forall x\in\mathbb{D},\;x^{2}-\operatorname{\operatorname{tr}}(x)x+q(x)=0.$ (3)
-To do so, we take arbitrary elements $y$ and $z$ in $\mathbb{D}$ and explore the identity $q(x+y)q(x+z)=q((x+y)(x+z))=q(x^{2}+yx+xz+yz)$. Thanks to $b$, we expand
-$q(x^{2}+yx+xz+yz)=\underbrace{q(x^{2})}_{a_{4}}+\underbrace{b(x^{2},yx+xz)}_{a_{3}}+\underbrace{b(x^{2},yz)+q(yx+xz)}_{a_{2}}\\ +\underbrace{b(yx+xz,yz)}_{a_{1}}+\underbrace{q(yz)}_{a_{0}}$
-and
-$q(x+y)q(x+z)=\underbrace{q(x)^{2}}_{b_{4}}+\underbrace{q(x)(b(x,y)+b(x,z))}_{b_{3}}+\underbrace{q(x)(q(y)+q(z))+b(x,y)b(x,z)}_{b_{2}}\\ +\underbrace{q(z)b(x,y)+q(y)b(x,z)}_{b_{1}}+\underbrace{q(y)q(z)}_{b_{0}}.$
-It is easily checked, by using either the multiplicativity of $q$ or (2), that $b_{i}=a_{i}$ for all $i\in\{0,1,3,4\}$, which leaves us with
-$q(x)(q(y)+q(z))+b(x,y)\,b(x,z)=b(x^{2},yz)+q(yx+xz),$
-and since q(yx+xz)=q(yx)+q(xz)+b(yx,xz)=q(y)q(x)+q(x)q(z)+b(yx,xz)q(yx+xz)=q(yx)+q(xz)+b(yx,xz)=q(y)q(x)+q(x)q(z)+b(yx,xz), this is further simplified as
-$b(x,y)\,b(x,z)=b(x^{2},yz)+b(yx,xz).$
-From there, we apply this identity to $y=1$ and an arbitrary $z$ to find $b(q(x),z)=q(x)\,b(1,z)=b(x,xz)=b(yx,xz)=b(x,1)\,b(x,z)-b(x^{2},z)=b(\operatorname{\operatorname{tr}}(x)x-x^{2},z)$, which yields $q(x)=\operatorname{\operatorname{tr}}(x)x-x^{2}$ because $q$ is nondegenerate.
-Next, we introduce the adjunction mapping
-$x\in\mathbb{D}\mapsto x^{\star}:=\operatorname{\operatorname{tr}}(x)-x\in\mathbb{D},$
-which is $\mathbb{F}$-linear and clearly satisfies $\operatorname{\operatorname{tr}}(x^{\star})=\operatorname{\operatorname{tr}}(x)$ and $(x^{\star})^{\star}=x$ for all $x\in\mathbb{D}$.
-We observe that (3) can be reformulated into
-$\forall x\in\mathbb{D},\;q(x)=xx^{\star}=x^{\star}x.$
-In particular, for all $x,y$ in $\mathbb{D}^{\times}$, we have $(xy)(xy)^{\star}=q(xy)=q(x)q(y)=xq(y)x^{\star}=xyy^{\star}x^{\star}$ and hence $(xy)^{\star}=y^{\star}x^{\star}$.
-It follows that $(-)^{\star}$ is an involution of the $\mathbb{F}$-algebra $\mathbb{D}$.
-Assume now that $d>1$.
-Since $\mathbb{D}$ is a division ring, we deduce from (3) that it has a $2$-dimensional subfield, and hence $d$ is even.
-Assume that $d=2$.
-Since $q$ is nondegenerate, the trace does not vanish, so by linearity it does not vanish on $\mathbb{D}\smallsetminus\mathbb{F}$, and we deduce from (3) that $\mathbb{D}$ is a separable extension of $\mathbb{F}$ (by choosing $x\in\mathbb{D}\smallsetminus\mathbb{F}$ such that $\operatorname{\operatorname{tr}}(x)\neq 0$).
-And clearly $(-)^{\star}$ is not the identity, so it is the standard involution of $\mathbb{D}$ over $\mathbb{F}$.
-Assume finally that $d>2$, and set p:=d2p:=d-2.
-Since $q$ is nondegenerate, the radical of its restriction to $H:=\operatorname{\operatorname{Ker}}(\operatorname{\operatorname{tr}})$ has dimension at most $1$, so we can find in $H$ an $\mathbb{F}$-linear subspace $V$ of dimension $p$ on which $q$ remains nondegenerate.
-Then the identity $\forall x\in V,\;x^{2}=-q(x)$ yields a (unique) homomorphism of $\mathbb{F}$-algebras $\Phi:\mathcal{C}(V,-q_{|V})\rightarrow\mathbb{D}$ such that $\forall x\in V,\;\Phi(x)=x$, where $\mathcal{C}(V,-q_{|V})$ stands for the Clifford algebra of the regular quadratic space $(V,-q_{|V})$.
-Classically, $\mathcal{C}(V,-q_{|V})$ is a simple algebra with dimension $2^{p}$ over $\mathbb{F}$, because $p$ is even.
-Hence $\Phi$ is injective, which leads to $2^{p}\leq d$ and hence to $d=4$.
-Then $\Phi$ is an isomorphism and $\mathbb{F}\cap V=\{0\}$, and we have recognized from $\Phi$ that $\mathbb{D}$ is a quaternion algebra over $\mathbb{F}$ and that the quaternionic conjugation $x\mapsto\overline{x}$ inherited from $\Phi$ satisfies $x\overline{x}=q(x)$ for all $x\in V$, which yields that it coincides with $(-)^{\star}$ on $V$.
-Hence $x\mapsto x^{\star}$ and $x\mapsto\overline{x}$ coincide on the $\mathbb{F}$-subalgebra generated by $\mathbb{F}\oplus V$, which must be $\mathbb{D}$ because it is a division algebra over $\mathbb{F}$ with dimension $3$ or $4$ that contains a subfield with dimension $2$ over $\mathbb{F}$.
-Therefore $x\mapsto x^{\star}$ is the quaternionic conjugation of $\mathbb{D}$, and our proof is completed.
-∎
+The result is rarely presented in this way in the literature, in particular because most references on the topic do not restrict the scope to nonisotropic quadratic forms.
+Hence, some comments are necessary.
+If $q$ is non-degenerate then the theory of unital composition algebras yields that either $\mathbb{D}$ is a Cayley algebra over $\mathbb{F}$ or $\mathbb{D}$ is of separable quadratic type over $\mathbb{F}$ and $q:x\mapsto x\sigma(x)$ for the standard involution $\sigma$ of $\mathbb{D}$ over $\mathbb{F}$ (see theorem 33.17 in [7]).
+Yet, here the multiplication map is associative, so Cayley algebras are discarded.
+Finally, if $q$ is degenerate then $\operatorname{\operatorname{char}}(\mathbb{F})=2$ and it is known in that case that the radical of $q$ equals the full space $\mathbb{D}$ (see proposition 33.4 in [7]), and from there it is known that $\mathbb{D}$ is a hyper-radicial extension of $\mathbb{F}$ and that $q:x\mapsto x^{2}$ (see corollary 33.6 in [7]).
 3 The greatest possible dimension for a trivial spectrum subspace
 This short section is devoted to the proof of Theorem 1.8.
-We have seen right at the beginning of the article that for every integer $n\geq 1$ there exists an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $\alpha(n,d)$.
+We have seen right at the beginning of Section 1.2 that for every integer $n\geq 1$ there exists a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$.
 Hence, in order to prove Theorem 1.8 it suffices to prove the following result:
 Theorem 3.1.
-Let $\mathbb{D}$ be a division ring with finite degree $d$ over a central subfield $\mathbb{F}$, and $V$ be a vector space over $\mathbb{D}$ with finite dimension $n$.
+Let $\mathbb{D}$ be a division ring with finite dimension $d$ over a central subfield $\mathbb{F}$, and $V$ be a vector space over $\mathbb{D}$ with finite dimension $n$.
 Assume that $|\mathbb{F}|\geq dn$.
-Let $\mathcal{S}$ be an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+Let $S$ be a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
 Then
-$\dim\mathcal{S}\leq\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$.
-The main key to this result is the use of the dual operator space of $\mathcal{S}$ (see Section 1.5) and what we call the Tangent Space Lemma. The latter is recalled in the next section.
+$\dim S\leq\alpha(n,d)=n(d-1)+d\dbinom{n}{2}$.
+The main key to this result is the use of the dual operator space of $S$ (see Section 1.5) and what we call the Tangent Space Lemma. The latter is recalled in the next section.
 3.1 The Tangent Space Lemma, and the Flanders-Atkinson lemma
@@
 Lemma 3.3 (Tangent Space Lemma).
-Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{F}$, and $\mathcal{S}$ be a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$.
-In $\mathcal{S}$, take an element $u_{0}$ of maximal rank $r$, and assume that $|\mathbb{F}|>r$.
-Then every element of $\mathcal{S}$ maps $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
+Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{F}$, and $S$ be a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$.
+In $S$, take an element $u_{0}$ of maximal rank $r$, and assume that $|\mathbb{F}|>r$.
+Then every element of $S$ maps $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
 The reference to tangent spaces comes from the observation that if $\mathbb{F}$ is the field of real numbers then the manifold of all rank $r$ operators of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ has its tangent space at $u_{0}$ equal to the space of all operators in $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ that map $\operatorname{\operatorname{Ker}}u_{0}$ into $\operatorname{\operatorname{Im}}u_{0}$.
 3.2 Proof of Theorem 3.1
-We prove the result by induction on $n$.
-The case $n=0$ is trivial, so we assume that $n\geq 1$.
-Let us consider the dual operator space $\widehat{\mathcal{S}}$, consisting of the $\mathbb{F}$-linear mappings
-$\widehat{x}:u\in\mathcal{S}\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
-Note that $\widehat{\mathcal{S}}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$.
+We prove the result by induction on $n$. The case $n=0$ is trivial, so we assume that $n\geq 1$. Let us consider the dual operator space $\widehat{S}$, consisting of the $\mathbb{F}$-linear mappings
+$\widehat{x}:u\in S\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
+Note that $\widehat{S}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S,V)$.
 Let $x\in V\smallsetminus\{0\}$.
-Since $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum, there is no $u\in\mathcal{S}$ such that $u(x)=x$; hence $\operatorname{\operatorname{rk}}(\widehat{x})\leq dn-1$.
-Let us now take $x\in V\smallsetminus\{0\}$ such that $\widehat{x}$ has the greatest possible rank $r\leq dn-1$ in $\widehat{\mathcal{S}}$.
-Since $|\mathbb{F}|>r$, the Tangent Space Lemma (Lemma 3.3) yields
+Since $S$ has trivial spectrum, there is no $u\in S$ such that $u(x)=x$; hence $\operatorname{\operatorname{rk}}(\widehat{x})\leq dn-1$.
+Let us now take $x\in V\smallsetminus\{0\}$ such that $\widehat{x}$ has the greatest possible rank $r\leq dn-1$ in $\widehat{S}$. Since $|\mathbb{F}|>r$, the Tangent Space Lemma (Lemma 3.3) yields
 $\forall u\in\operatorname{\operatorname{Ker}}\widehat{x},\;\forall y\in V,\quad\widehat{y}(u)\in\operatorname{\operatorname{Im}}\widehat{x}$,
 or in other words
-$\forall u\in\mathcal{S},\;u(x)=0\Rightarrow\operatorname{\operatorname{Im}}u\subseteq\mathcal{S}x.$
+$\forall u\in S,\;u(x)=0\Rightarrow\operatorname{\operatorname{Im}}u\subseteq Sx.$
 Setting
-$\mathcal{S}^{\prime}:=\operatorname{\operatorname{Ker}}\widehat{x}=\{u\in\mathcal{S}:\;u(x)=0\},$
+$S^{\prime}:=\operatorname{\operatorname{Ker}}\widehat{x}=\{u\in S:\;u(x)=0\},$
 we can apply the rank theorem to get
-$\dim(\mathcal{S})=\dim(\mathcal{S}^{\prime})+\dim_{\mathbb{F}}(\mathcal{S}x)\leq\dim(\mathcal{S}^{\prime})+(dn-1).$ (4)
-Now, the operators in $\mathcal{S}^{\prime}$ are actually $\mathbb{D}$-linear so their ranges are $\mathbb{D}$-linear subspaces of $V$, all included in $\mathcal{S}x$ and hence all included in the socle (see Definition 2.1) of $\mathcal{S}x$.
-This socle does not contain $x$ (because $x\not\in\mathcal{S}x$), and hence we can embed it into a $\mathbb{D}$-hyperplane $H$ of $V$ such that $x\mathbb{D}\oplus H=V$, and we note that $\operatorname{\operatorname{Im}}u\subseteq H$ for every $u\in\mathcal{S}^{\prime}$.
-Hence, all the elements of $\mathcal{S}^{\prime}$ leave $H$ invariant, inducing $\mathbb{D}$-linear endomorphisms of $H$, all of them with $\mathbb{F}$-trivial spectrum.
-Since $x\mathbb{D}\oplus H=V$ and the elements of $\mathcal{S}^{\prime}$ vanish on $x\mathbb{D}$, the $\mathbb{F}$-linear mapping
-$u\in\mathcal{S}^{\prime}\mapsto u_{H}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$
-is injective. Its range is an $\mathbb{F}$-trivial spectrum subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$. Since $|\mathbb{F}|\geq(n-1)d$, we obtain by induction that
-$\dim(\mathcal{S}^{\prime})=\dim(\mathcal{T})\leq d\dbinom{n-1}{2}+(n-1)(d-1).$
+$\dim(S)=\dim(S^{\prime})+\dim_{\mathbb{F}}(Sx)\leq\dim(S^{\prime})+(dn-1).$ (2)
+Now, the operators in $S^{\prime}$ are actually $\mathbb{D}$-linear so their ranges are $\mathbb{D}$-linear subspaces of $V$, all included in $Sx$ and hence all included in the socle (see Definition 2.1) of $Sx$.
+This socle does not contain $x$ (because $x\not\in Sx$), and hence we can embed it into a $\mathbb{D}$-hyperplane $H$ of $V$ such that $x\mathbb{D}\oplus H=V$, and we note that $\operatorname{\operatorname{Im}}u\subseteq H$ for every $u\in S^{\prime}$.
+Hence, all the elements of $S^{\prime}$ leave $H$ invariant, inducing $\mathbb{D}$-linear endomorphisms of $H$, all of them with trivial spectrum.
+Since $x\mathbb{D}\oplus H=V$ and the elements of $S^{\prime}$ vanish on $x\mathbb{D}$, the $\mathbb{F}$-linear mapping
+$u\in S^{\prime}\mapsto u_{H}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$
+is injective. Its range is a trivial spectrum subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(H)$. Since $|\mathbb{F}|\geq(n-1)d$, we obtain by induction that
+$\dim(S^{\prime})=\dim(\mathcal{T})\leq d\dbinom{n-1}{2}+(n-1)(d-1).$
 By combining this with (2), we conclude that
-$\dim(\mathcal{S})\leq(n-1)(d-1)+d\dbinom{n-1}{2}+((d-1)+d(n-1))=n(d-1)+d\dbinom{n}{2}$,
+$\dim(S)\leq(n-1)(d-1)+d\dbinom{n-1}{2}+((d-1)+d(n-1))=n(d-1)+d\dbinom{n}{2}$,
 which completes the proof.
@@
 Assume that $|\mathbb{F}|\geq nd$.
-Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
-Then $\dim\mathcal{S}\leq\alpha(n,d)$.
+Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+Then $\dim S\leq\alpha(n,d)$.
 Proof.
-We denote by $r$ the transitive rank of $\mathcal{S}$.
-If $r=0$ then $\mathcal{S}=\{0\}$ and we are done.
-Assume that $r>0$, and choose $x\in U$ such that $\dim_{\mathbb{F}}\mathcal{S}x=r$.
-Hence $\widehat{x}$ has the greatest possible rank in $\widehat{\mathcal{S}}$, and as $|\mathbb{F}|>r$ we can apply the tangent space lemma.
-This yields that all the operators in $\mathcal{S}^{\prime}:=\{f\in\mathcal{S}:f(x)=0\}$ map into $\mathcal{S}x$.
-As in the previous proof, we can find a $\mathbb{D}$-linear hyperplane $H$ of $V$ into which all the elements of $\mathcal{S}^{\prime}$ map.
+We denote by $r$ the transitive rank of $S$.
+If $r=0$ then $S=\{0\}$ and we are done.
+Assume that $r>0$, and choose $x\in U$ such that $\dim_{\mathbb{F}}Sx=r$.
+Hence $\widehat{x}$ has the greatest possible rank in $\widehat{S}$, and as $|\mathbb{F}|>r$ we can apply the tangent space lemma.
+This yields that all the operators in $S^{\prime}:=\{f\in S:f(x)=0\}$ map into $Sx$.
+As in the previous proof, we can find a $\mathbb{D}$-linear hyperplane $H$ of $V$ into which all the elements of $S^{\prime}$ map.
 The rank theorem yields
-$\dim\mathcal{S}=r+\dim\mathcal{S}^{\prime}.$
+$\dim S=r+\dim S^{\prime}.$
 Then there are two cases:
-  • If $n=1$ then $\mathcal{S}^{\prime}=\{0\}$ and we conclude that $\dim\mathcal{S}\leq d-1$, as claimed;
-  • If $n>1$ then $\mathcal{S}^{\prime}$ is a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, by induction we find that
-    $\dim\mathcal{S}\leq(nd-1)+\alpha(n-1,d)=\alpha(n,d),$
+  • If $n=1$ then $S^{\prime}=\{0\}$ and we conclude that $\dim S\leq d-1$, as claimed;
+  • If $n>1$ then $S^{\prime}$ is a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, by induction we find that
+    $\dim S\leq(nd-1)+\alpha(n-1,d)=\alpha(n,d),$
     and the proof is complete.
@@
 Corollary 3.5.
-Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(3-d,1)$ and $n>0$.
-Then $\mathcal{S}$ is target-reduced.
+Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim S\geq\alpha(n,d)-n+\max(3-d,1)$ and $n>0$.
+Then $S$ is target-reduced.
 Remark 3.1.
-In the proof and in future use of the result, it is crucial to note that the fact that $\mathcal{S}$ is not target-reduced is equivalent to the existence of a $\mathbb{D}$-linear hyperplane $H$ of $V$ (and not simply of an $\mathbb{F}$-linear hyperplane) that includes the range of every element of $\mathcal{S}$.
+In the proof and in future use of the result, it is crucial to note that the fact that $S$ is not target-reduced is equivalent to the existence of a $\mathbb{D}$-linear hyperplane $H$ of $V$ (and not simply of an $\mathbb{F}$-linear hyperplane) that includes the range of every element of $S$.
 The converse implication is obvious.
-For the direct one, assume that we have an $\mathbb{F}$-linear hyperplane $V^{\prime}$ of $V$ that includes the range of every operator in $\mathcal{S}$.
-Then all the operators in $\mathcal{S}$ map into the socle of $V^{\prime}$, which is a $\mathbb{D}$-linear hyperplane of $V$.
+For the direct one, assume that we have an $\mathbb{F}$-linear hyperplane $V^{\prime}$ of $V$ that includes the range of every operator in $S$.
+Then all the operators in $S$ map into the socle of $V^{\prime}$, which is a $\mathbb{D}$-linear hyperplane of $V$.
 Proof of Corollary 3.5.
-Assume that $\mathcal{S}$ is not target-reduced. Then there exists a $\mathbb{D}$-linear hyperplane $H$ of $V$ such that $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
-Because $\mathcal{S}$ is deeply intransitive we can view it as a deeply intransitive subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, and then Proposition 3.4 yields $\dim\mathcal{S}\leq\alpha(n-1,d)=\alpha(n,d)-(nd-1)$.
+Assume that $S$ is not target-reduced. Then there exists a $\mathbb{D}$-linear hyperplane $H$ of $V$ such that $S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
+Because $S$ is deeply intransitive we can view it as a deeply intransitive subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$, and then Proposition 3.4 yields $\dim S\leq\alpha(n-1,d)=\alpha(n,d)-(nd-1)$.
 Now (nd1)n=n(d1)1(nd-1)-n=n(d-1)-1, and n(d1)1>2n(d-1)-1>-2 if $d=1$, whereas n(d1)1>1n(d-1)-1>-1 whenever $d\geq 2$ and $n>0$.
-Hence, in any case we contradict the asumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(3-d,1)$.
+Hence, in any case we contradict the asumption that $\dim S\geq\alpha(n,d)-n+\max(3-d,1)$.
 ∎
@@
 Catchers are intimately connected with alternators, as we shall now see.
-Let us take finite-dimensional vector spaces $U$ and $V$ over $\mathbb{F}$, a linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, and let $\mathbf{B}_{\mathcal{S}}=(v_{1},\dots,v_{n})$ and $\mathbf{B}_{V}=(e_{1},\dots,e_{p})$ be respective bases of $\mathcal{S}$ and $V$.
-Throughout, we assume that $\mathcal{S}$ is target-reduced (see the definition in Section 1.5), to the effect that $\widehat{\mathcal{S}}$ is also target-reduced.
-We consider the matrix space $\mathcal{M}$ that represents the dual operator space $\widehat{\mathcal{S}}$ in $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$, and we take an arbitrary generic matrix $\mathbf{M}$ that is associated with $\mathcal{M}$, with parameter space denoted by $E$.
+Let us take finite-dimensional vector spaces $U$ and $V$ over $\mathbb{F}$, a linear subspace $S$ of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, and let $\mathbf{B}_{S}=(v_{1},\dots,v_{n})$ and $\mathbf{B}_{V}=(e_{1},\dots,e_{p})$ be respective bases of $S$ and $V$.
+Throughout, we assume that $S$ is target-reduced (see the definition in Section 1.5), to the effect that $\widehat{S}$ is also target-reduced.
+We consider the matrix space $\mathcal{M}$ that represents the dual operator space $\widehat{S}$ in $\mathbf{B}_{S}$ and $\mathbf{B}_{V}$, and we take an arbitrary generic matrix $\mathbf{M}$ that is associated with $\mathcal{M}$, with parameter space denoted by $E$.
 We define the following linear mappings:
-  • $\Phi$ assigns to every $z\in E$ the operator in $\widehat{\mathcal{S}}$ whose matrix in the bases $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$ is $\mathbf{M}(z)$;
-  • $\widehat{(-)}$ assigns to every $y\in U$ the evaluation mapping $\widehat{y}\in\widehat{\mathcal{S}}$.
+  • $\Phi$ assigns to every $z\in E$ the operator in $\widehat{S}$ whose matrix in the bases $\mathbf{B}_{S}$ and $\mathbf{B}_{V}$ is $\mathbf{M}(z)$;
+  • $\widehat{(-)}$ assigns to every $y\in U$ the evaluation mapping $\widehat{y}\in\widehat{S}$.
 These mappings are represented in the following diagram
 […]
-Now, we consider an alternator $b$ of $\mathcal{S}$. The left mapping $L_{b}:y\in U\mapsto b(y,-)\in V^{\star}$ can then be inserted into the following diagram:
+Now, we consider an alternator $b$ of $S$. The left mapping $L_{b}:y\in U\mapsto b(y,-)\in V^{\star}$ can then be inserted into the following diagram:
 […]
-Next, we prove that $L_{b}$ factors into the composite of $\widehat{(-)}$ with a unique linear mapping $\Theta:\widehat{\mathcal{S}}\rightarrow V^{\star}$.
+Next, we prove that $L_{b}$ factors into the composite of $\widehat{(-)}$ with a unique linear mapping $\Theta:\widehat{S}\rightarrow V^{\star}$.
 To see this, we apply the factorization lemma for linear maps: all we need is to check that $L_{b}$ vanishes at every vector $y\in U$ such that $\widehat{y}=0$.
 So, let $y\in U$ satisfy $\widehat{y}=0$.
-Hence, for all $f\in\mathcal{S}$, we have b(y,f(z))=b(z,f(y))=0b(y,f(z))=-b(z,f(y))=0 because $f$ is $b$-alternating.
-Because $\mathcal{S}$ is target-reduced, this yields $y\in\operatorname{\operatorname{Ker}}L_{b}$, as claimed.
+Hence, for all $f\in S$, we have b(y,f(z))=b(z,f(y))=0b(y,f(z))=-b(z,f(y))=0 because $f$ is $b$-alternating.
+Because $S$ is target-reduced, this yields $y\in\operatorname{\operatorname{Ker}}L_{b}$, as claimed.
 The previous diagram is then completed as:
@@
 Moreover, if $\mathbf{L}=0$ then we successively find $\Psi=0$, $\Theta=0$ and finally $b=0$. Assigning $\mathbf{L}$ to $b$ thus creates an injective linear mapping
-$\Lambda:\operatorname{\operatorname{Alt}}(\mathcal{S})\hookrightarrow\operatorname{\operatorname{catch}}(\mathbf{M}).$
+$\Lambda:\operatorname{\operatorname{Alt}}(S)\hookrightarrow\operatorname{\operatorname{catch}}(\mathbf{M}).$
 Conversely, let $\mathbf{L}$ be a catcher of $\mathbf{M}$.
 Let $\Psi:E\rightarrow V^{\star}$ assign to every $z\in E$ the linear form on $V$ whose matrix in the basis $\mathbf{B}_{V}$ is $\mathbf{L}(z)$.
-We claim that there is a (unique) factorization $\Psi=\Theta\circ\Phi$ for some linear mapping $\Theta:\widehat{\mathcal{S}}\rightarrow V^{\star}$.
+We claim that there is a (unique) factorization $\Psi=\Theta\circ\Phi$ for some linear mapping $\Theta:\widehat{S}\rightarrow V^{\star}$.
 Again, we use the factorization lemma.
@@
 Because $\mathbf{M}(z_{0})=0$ by assumption, this yields $\mathbf{L}(z_{0})M=0$ for all $M\in\mathcal{M}$.
-In other words, every operator in $\widehat{\mathcal{S}}$ maps into $\operatorname{\operatorname{Ker}}\Psi(z_{0})$.
-Since $\widehat{\mathcal{S}}$ is target-reduced, we deduce that $\Psi(z_{0})=0$.
+In other words, every operator in $\widehat{S}$ maps into $\operatorname{\operatorname{Ker}}\Psi(z_{0})$.
+Since $\widehat{S}$ is target-reduced, we deduce that $\Psi(z_{0})=0$.
 Hence the existence of $\Theta$, the uniqueness of which is obvious because $\Phi$ is onto.
@@
 Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$.
-Let $\mathcal{S}$ be a target-reduced $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
-Assume that $\operatorname{\operatorname{trk}}\mathcal{S}=nd-1$ and $|\mathbb{F}|\geq nd$.
-Then $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$.
+Let $S$ be a target-reduced $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
+Assume that $\operatorname{\operatorname{trk}}S=nd-1$ and $|\mathbb{F}|\geq nd$.
+Then $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$.
 Proof.
-We choose respective bases $\mathbf{B}_{V}=(e_{1},\dots,e_{nd})$ and $\mathbf{B}_{\mathcal{S}}=(f_{1},\dots,f_{N})$ of $V$ and $\mathcal{S}$ seen as $\mathbb{F}$-vector spaces, and we consider the generic matrix $\mathbf{M}$ associated with $x\in U\mapsto\widehat{x}\in\widehat{\mathcal{S}}$ in these bases.
-The condition that $\operatorname{\operatorname{trk}}\mathcal{S}=nd-1$ yields $\operatorname{\operatorname{rk}}\mathbf{M}=nd-1$, by Corollary 4.2.
-Let $b_{1}$ and $b_{2}$ be nonzero alternators of $\mathcal{S}$. We shall prove that $b_{1}$ and $b_{2}$ are linearly dependent over $\mathbb{F}$.
+We choose respective bases $\mathbf{B}_{V}=(e_{1},\dots,e_{nd})$ and $\mathbf{B}_{S}=(f_{1},\dots,f_{N})$ of $V$ and $S$ seen as $\mathbb{F}$-vector spaces, and we consider the generic matrix $\mathbf{M}$ associated with $x\in U\mapsto\widehat{x}\in\widehat{S}$ in these bases.
+The condition that $\operatorname{\operatorname{trk}}S=nd-1$ yields $\operatorname{\operatorname{rk}}\mathbf{M}=nd-1$, by Corollary 4.2.
+Let $b_{1}$ and $b_{2}$ be nonzero alternators of $S$. We shall prove that $b_{1}$ and $b_{2}$ are linearly dependent over $\mathbb{F}$.
 We apply the procedure explained in the previous paragraph to obtain associated catchers $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ of $\mathbf{M}$.
@@
 Then the identity $\mathbf{B_{1}}\mathbf{M}=0$ leads to $\mathbf{b_{1}}(B_{1}\mathbf{M})=0$, and hence to $B_{1}\mathbf{M}=0$.
-The matrix $B_{1}$ represents a linear form $\varphi$ in the dual basis of $\mathbf{B}_{V}$, and the statement that $B_{1}\mathbf{M}=0$ means that all the operators in $\widehat{\mathcal{S}}$ map into the kernel of $\varphi$.
-Hence so do all the operators in $\mathcal{S}$, and we contradict the assumption that $\mathcal{S}$ is target-reduced.
+The matrix $B_{1}$ represents a linear form $\varphi$ in the dual basis of $\mathbf{B}_{V}$, and the statement that $B_{1}\mathbf{M}=0$ means that all the operators in $\widehat{S}$ map into the kernel of $\varphi$.
+Hence so do all the operators in $S$, and we contradict the assumption that $S$ is target-reduced.
 Hence $\operatorname{\operatorname{sprk}}\mathbf{B}_{1}>1$ and we deduce from the Factorization Lemma that $\mathbf{B_{2}}=\alpha\mathbf{B}_{1}$ for some scalar $\alpha\in\mathbb{F}$.
 This yields $b_{2}=\alpha b_{1}$.
-We conclude that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension at most $1$.
+We conclude that $\operatorname{\operatorname{Alt}}(S)$ has dimension at most $1$.
 ∎
@@
 Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim_{\mathbb{D}}V$.
-Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that $|\mathbb{F}|\geq nd$.
+Let $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, and assume that $|\mathbb{F}|\geq nd$.
   1. (a)
-    If $\mathcal{S}$ is weakly primitively intransitive, then it is deeply intransitive.
+    If $S$ is weakly primitively intransitive, then it is deeply intransitive.
   2. (b)
-    If $\mathcal{S}$ is primitively intransitive, then it is deeply intransitive.
+    If $S$ is primitively intransitive, then it is deeply intransitive.
 Proof.
-Assume that $\mathcal{S}$ is not deeply intransitive.
-Then we can find a nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that $\mathcal{S}^{V^{\prime}}$ is not intransitive.
-Denote by $\pi:V\twoheadrightarrow V/V^{\prime}$ the standard projection, and consider the operator space $\pi\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$.
-Take a basis $(v_{1},\dots,v_{s})$ of $\mathcal{S}^{V^{\prime}}$; and lift a basis of $\pi\mathcal{S}$ to a family $(u_{1},\dots,u_{r})$ in $\mathcal{S}$, so that $(u_{1},\dots,u_{r},v_{1},\dots,v_{s})$ is a basis $\mathbf{B}_{\mathcal{S}}$ of $\mathcal{S}$.
-Take an arbitrary $\mathbb{F}$-basis $\mathbf{B}_{V}$ of $V$ that is adapted to $V^{\prime}$ and in which the first $d$ vectors are pairwise linearly dependent over $\mathbb{D}$, and consider a generic matrix $\mathbf{M}$ of $\widehat{\mathcal{S}}$ in the bases $\mathbf{B}_{\mathcal{S}}$ and $\mathbf{B}_{V}$.
+Assume that $S$ is not deeply intransitive.
+Then we can find a nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that $S^{V^{\prime}}$ is not intransitive.
+Denote by $\pi:V\twoheadrightarrow V/V^{\prime}$ the standard projection, and consider the operator space $\pi S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/V^{\prime})$.
+Take a basis $(v_{1},\dots,v_{s})$ of $S^{V^{\prime}}$; and lift a basis of $\pi S$ to a family $(u_{1},\dots,u_{r})$ in $S$, so that $(u_{1},\dots,u_{r},v_{1},\dots,v_{s})$ is a basis $\mathbf{B}_{S}$ of $S$.
+Take an arbitrary $\mathbb{F}$-basis $\mathbf{B}_{V}$ of $V$ that is adapted to $V^{\prime}$ and in which the first $d$ vectors are pairwise linearly dependent over $\mathbb{D}$, and consider a generic matrix $\mathbf{M}$ of $\widehat{S}$ in the bases $\mathbf{B}_{S}$ and $\mathbf{B}_{V}$.
 Then this matrix takes the form
 $\mathbf{M}=\begin{bmatrix}[?]_{p\times r}&\mathbf{C}\\ \mathbf{B}&[0]\end{bmatrix}\quad\text{where $p:=\dim_{\mathbb{F}}(V^{\prime})$.}$
-Note that $\mathbf{C}$ is a generic matrix of $\widehat{\mathcal{S}^{V^{\prime}}}$, whereas $\mathbf{B}$ is a generic matrix of $\widehat{\pi\mathcal{S}}$.
-Hence $\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{trk}}(\mathcal{S}^{V^{\prime}})=\dim_{\mathbb{F}}V^{\prime}=p$.
+Note that $\mathbf{C}$ is a generic matrix of $\widehat{S^{V^{\prime}}}$, whereas $\mathbf{B}$ is a generic matrix of $\widehat{\pi S}$.
+Hence $\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{trk}}(S^{V^{\prime}})=\dim_{\mathbb{F}}V^{\prime}=p$.
 Besides $\operatorname{\operatorname{rk}}\mathbf{C}+\operatorname{\operatorname{rk}}\mathbf{B}\leq\operatorname{\operatorname{rk}}\mathbf{M}$.
 Consider the $\mathbb{D}$-linear subspace $L$ of $V$ spanned by the first $d$ vectors of $\mathbf{B}_{V}$, and the matrix $\mathbf{M^{\prime}}$ obtained by deleting the first $d$ rows of $\mathbf{M}$.
-Then $\mathbf{M^{\prime}}$ is a generic matrix of $\pi^{\prime}\mathcal{S}$ where $\pi^{\prime}:V\twoheadrightarrow V/L$ is the standard projection.
+Then $\mathbf{M^{\prime}}$ is a generic matrix of $\pi^{\prime}S$ where $\pi^{\prime}:V\twoheadrightarrow V/L$ is the standard projection.
 By deleting the first $d$ rows we see that
 $\operatorname{\operatorname{rk}}\mathbf{M^{\prime}}\leq d(\dim_{\mathbb{D}}V^{\prime}-1)+\operatorname{\operatorname{rk}}\mathbf{B}\leq d(\dim_{\mathbb{D}}V^{\prime}-1)+\operatorname{\operatorname{rk}}\mathbf{M}-\operatorname{\operatorname{rk}}\mathbf{C}=\operatorname{\operatorname{rk}}\mathbf{M}-d.$
-Hence $\operatorname{\operatorname{trk}}(\pi^{\prime}\mathcal{S})\leq\operatorname{\operatorname{trk}}(\mathcal{S})-d$. We conclude that $\mathcal{S}$ is not weakly primitively intransitive.
+Hence $\operatorname{\operatorname{trk}}(\pi^{\prime}S)\leq\operatorname{\operatorname{trk}}(S)-d$. We conclude that $S$ is not weakly primitively intransitive.
 The second point is deduced from the first one and from Proposition 1.25. ∎
-5 Nondegenerate alternators of large deeply intransitive operator spaces
+5 Nondegenerate alternators of deeply intransitive operator spaces
 Our aim here is to complete the proof of Theorem 1.22. We have already proved point (a).
@@
 Let $U$ and $V$ be finite-dimensional vector spaces over $\mathbb{D}$, with $n:=\dim V$.
-Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
-Then $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is $1$-dimensional.
+Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$.
+Then $\operatorname{\operatorname{Alt}}(S)$ is $1$-dimensional.
 Once this result is obtained, we will directly consider operator spaces that have a $1$-dimensional alternator space, and point (c) of Theorem 1.22 will be obtained along with Theorem 1.23 thanks to the classification of multiplicative quadratic forms on division algebras (Theorem 2.3).
@@
 For simplicity, we set $\varepsilon=1$ if $d=1$, and $\varepsilon=0$ otherwise.
-Finally, we take a deeply intransitive $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that
-$\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon.$
+Finally, we take a deeply intransitive $\mathbb{F}$-linear subspace $S$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that
+$\dim S\geq\alpha(n,d)-n+1+\varepsilon.$
 We will prove the following two points by induction on $n$:
-  • In any case $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$.
-  • If $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ then $\mathcal{S}$ has a nonzero alternator.
-Since $n>0$, we immediately retrieve from Corollary 3.5 that $\mathcal{S}$ is target-reduced.
+  • In any case $\operatorname{\operatorname{trk}}(S)=nd-1$.
+  • If $\dim S\geq\alpha(n,d)-n+2+\varepsilon$ then $S$ has a nonzero alternator.
+Since $n>0$, we immediately retrieve from Corollary 3.5 that $S$ is target-reduced.
 Throughout, we set
-$r:=\operatorname{\operatorname{trk}}(\mathcal{S})\in\mathopen{[\![}0,nd-1\mathclose{]\!]}.$
-Since $\mathcal{S}$ is target-reduced we have $\mathcal{S}\neq\{0\}$ and hence $r>0$.
+$r:=\operatorname{\operatorname{trk}}(S)\in\mathopen{[\![}0,nd-1\mathclose{]\!]}.$
+Since $S$ is target-reduced we have $S\neq\{0\}$ and hence $r>0$.
 We fix a vector $x\in U\smallsetminus\{0\}$ such that $\operatorname{\operatorname{rk}}\widehat{x}=r$.
 Then we apply the Tangent Space Lemma to recover that all the operators in
-$\mathcal{S}^{\prime}:=\{u\in\mathcal{S}:\;u(x)=0\}$
-map into $\mathcal{S}x$.
-We denote by $H$ the socle of $\mathcal{S}x$ and obtain that all the operators in $\mathcal{S}^{\prime}$ map into $H$ (despite the notation, at this point $H$ is not necessarily a linear hyperplane of the $\mathbb{D}$-vector space $V$).
+$S^{\prime}:=\{u\in S:\;u(x)=0\}$
+map into $Sx$.
+We denote by $H$ the socle of $Sx$ and obtain that all the operators in $S^{\prime}$ map into $H$ (despite the notation, at this point $H$ is not necessarily a linear hyperplane of the $\mathbb{D}$-vector space $V$).
 We set
 $s:=\dim_{\mathbb{F}}H.$
-Note that $\mathcal{S}^{\prime}$ is included in $\mathcal{S}^{H}$, which is deeply intransitive, and hence $\mathcal{S}^{\prime}$ is also a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
-Now it is time to represent the dual operator space $\widehat{\mathcal{S}}$ in appropriate bases of $\mathcal{S}$ and $V$.
-We will choose such bases carefully with respect to the vector $x$.
-For the basis of $V$:
-  • We start from a basis $(e_{1},\dots,e_{r})$ of $\mathcal{S}x$ in which the first $s$ vectors span $H$ (as an $\mathbb{F}$-linear subspace), and we extend it arbitrarily into a basis $\mathbf{E}$ of $V$.
-  • We lift the first vectors $e_{1},\dots,e_{r}$ to obtain a family $(u_{1},\dots,u_{r})$ of operators in $\mathcal{S}$, so that $u_{i}(x)=e_{i}$ for all $i\in\mathopen{[\![}1,r\mathclose{]\!]}$.
-Then $(u_{1},\dots,u_{r})$ is linearly independent and we can choose an arbitrary basis $(u_{r+1},\dots,u_{\dim\mathcal{S}})$ of $\mathcal{S}^{\prime}$.
-Then our basis of $\mathcal{S}$ is $\mathbf{F}:=(u_{1},\dots,u_{r},u_{r+1},\dots,u_{\dim\mathcal{S}})$.
-We consider the matrix space $\mathcal{M}$ that represents $\widehat{\mathcal{S}}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
+Note that $S^{\prime}$ is included in $S^{H}$, which is deeply intransitive, and hence $S^{\prime}$ is also a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$.
+Now it is time to represent the dual operator space $\widehat{S}$ in appropriate bases of $S$ and $V$. We will choose such bases carefully with respect to the vector $x$. For the basis of $V$:
+  • We start from a basis $(e_{1},\dots,e_{r})$ of $Sx$ in which the first $s$ vectors span $H$ (as an $\mathbb{F}$-linear subspace), and we extend it arbitrarily into a basis $\mathbf{E}$ of $V$.
+  • We lift the first vectors $e_{1},\dots,e_{r}$ to obtain a family $(u_{1},\dots,u_{r})$ of operators in $S$, so that $u_{i}(x)=e_{i}$ for all $i\in\mathopen{[\![}1,r\mathclose{]\!]}$.
+Then $(u_{1},\dots,u_{r})$ is linearly independent and we can choose an arbitrary basis $(u_{r+1},\dots,u_{\dim S})$ of $S^{\prime}$.
+Then our basis of $S$ is $\mathbf{F}:=(u_{1},\dots,u_{r},u_{r+1},\dots,u_{\dim S})$.
+We consider the matrix space $\mathcal{M}$ that represents $\widehat{S}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
 The matrix in $\mathcal{M}$ that corresponds to the operator $\widehat{x}$ is the reduced matrix
-$J_{r}=\begin{bmatrix}I_{r}&[0]_{r\times(\dim\mathcal{S}-r)}\\ [0]_{(nd-r)\times r}&[0]_{(nd-r)\times(\dim\mathcal{S}-r)}\end{bmatrix}$.
-Now, we take $E:=U$ as our parameter space, $R$ and $\mathbb{L}$ as the associated ring and field, and we construct a generic matrix $\mathbf{M}$ of $\mathcal{M}$ as follows: for all $(i,j)\in\mathopen{[\![}1,nd\mathclose{]\!]}\times\mathopen{[\![}1,\dim\mathcal{S}\mathclose{]\!]}$, we define $\mathbf{m}_{i,j}$ as the linear form on $U$ that takes every vector $z\in U$ to the $(i,j)$-entry of the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
+$J_{r}=\begin{bmatrix}I_{r}&[0]_{r\times(\dim S-r)}\\ [0]_{(nd-r)\times r}&[0]_{(nd-r)\times(\dim S-r)}\end{bmatrix}$.
+Now, we take $E:=U$ as our parameter space, $R$ and $\mathbb{L}$ as the associated ring and field, and we construct a generic matrix $\mathbf{M}$ of $\mathcal{M}$ as follows: for all $(i,j)\in\mathopen{[\![}1,nd\mathclose{]\!]}\times\mathopen{[\![}1,\dim S\mathclose{]\!]}$, we define $\mathbf{m}_{i,j}$ as the linear form on $U$ that takes every vector $z\in U$ to the $(i,j)$-entry of the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
 Hence, the specialization $\mathbf{M}(z)$ is simply the matrix of $\widehat{z}$ in the bases $\mathbf{F}$ and $\mathbf{E}$.
-As we already know that all the operators in $\mathcal{S}^{\prime}$ map into $H$, the matrix $\mathbf{M}$ actually takes the refined form
-$\mathbf{M}=\begin{bmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{B}&[0]_{(nd-r)\times(\dim\mathcal{S}-r)}\end{bmatrix}$
-for some matrices $\mathbf{A}\in\operatorname{\operatorname{M}}_{r}(R)$, $\mathbf{B}\in\operatorname{\operatorname{M}}_{nd-r,r}(R)$ and $\mathbf{C}\in\operatorname{\operatorname{M}}_{r,\dim\mathcal{S}-r}(R)$ with $1$-homogeneous entries, and we can further refine
-$\mathbf{C}=\begin{bmatrix}\mathbf{C_{1}}\\ [0]_{(r-s)\times(\dim S-r)}\end{bmatrix}\qquad\text{with $\mathbf{C_{1}}\in\operatorname{\operatorname{M}}_{s,\dim\mathcal{S}-r}(R)$.}$
+As we already know that all the operators in $S^{\prime}$ map into $H$, the matrix $\mathbf{M}$ actually takes the refined form
+$\mathbf{M}=\begin{bmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{B}&[0]_{(nd-r)\times(\dim S-r)}\end{bmatrix}$
+for some matrices $\mathbf{A}\in\operatorname{\operatorname{M}}_{r}(R)$, $\mathbf{B}\in\operatorname{\operatorname{M}}_{nd-r,r}(R)$ and $\mathbf{C}\in\operatorname{\operatorname{M}}_{r,\dim S-r}(R)$ with $1$-homogeneous entries, and we can further refine
+$\mathbf{C}=\begin{bmatrix}\mathbf{C_{1}}\\ [0]_{(r-s)\times(\dim S-r)}\end{bmatrix}\qquad\text{with $\mathbf{C_{1}}\in\operatorname{\operatorname{M}}_{s,\dim S-r}(R)$.}$
 Accordingly, we also refine
@@
 $\mathbf{A}=\begin{bmatrix}\mathbf{A_{1}}&\mathbf{A_{3}}\\ \mathbf{A_{2}}&\mathbf{A_{4}}\end{bmatrix}\qquad\text{with $\mathbf{A_{1}}\in\operatorname{\operatorname{M}}_{s}(R)$, $\mathbf{A_{2}}\in\operatorname{\operatorname{M}}_{r-s,s}(R)$ and so on.}$
-Critically, $\mathbf{C_{1}}$ is a generic matrix for the dual operator space $\widehat{\mathcal{S}^{\prime}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S}^{\prime},H)$ in the basis $(u_{r+1},\dots,u_{\dim\mathcal{S}})$ of $\mathcal{S}^{\prime}$ and the basis $(e_{1},\dots,e_{s})$ of the $\mathbb{F}$-vector space $H$.
+Critically, $\mathbf{C_{1}}$ is a generic matrix for the dual operator space $\widehat{S^{\prime}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S^{\prime},H)$ in the basis $(u_{r+1},\dots,u_{\dim S})$ of $S^{\prime}$ and the basis $(e_{1},\dots,e_{s})$ of the $\mathbb{F}$-vector space $H$.
 Here, the parameter space for this generic matrix remains $U$.
@@
 $\mathbf{B_{1}}\mathbf{C_{1}}=0.$
-Before we proceed, it is time to complete the special case $n=1$ for the proof that $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$.
+Before we proceed, it is time to complete the special case $n=1$ for the proof that $\operatorname{\operatorname{trk}}(S)=nd-1$.
 So, assume for a moment that $n=1$.
-In that case $\mathcal{S}^{\prime}=\{0\}$ and hence $u\in\mathcal{S}\mapsto u(x)\in\mathcal{S}x$ is an $\mathbb{F}$-linear isomorphism.
-It follows that $r=\dim\mathcal{S}$.
-The assumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon$ requires here that $\dim\mathcal{S}\geq nd-1$ and hence r=nd1r=nd-1.
-From now on, we assume that $n\geq 2$. We will now draw some consequences of the assumption that $\dim\mathcal{S}\geq\alpha(n,d)-n+1+\varepsilon$.
+In that case $S^{\prime}=\{0\}$ and hence $u\in S\mapsto u(x)\in Sx$ is an $\mathbb{F}$-linear isomorphism.
+It follows that $r=\dim S$.
+The assumption that $\dim S\geq\alpha(n,d)-n+1+\varepsilon$ requires here that $\dim S\geq nd-1$ and hence r=nd1r=nd-1.
+From now on, we assume that $n\geq 2$. We will now draw some consequences of the assumption that $\dim S\geq\alpha(n,d)-n+1+\varepsilon$.
 Claim 5.1.
-Assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1.
-Then the space $H$ has dimension n1n-1 over $\mathbb{D}$, and the space $\mathcal{S}^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$ is target-reduced.
+Assume that $\dim S\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1.
+Then the space $H$ has dimension n1n-1 over $\mathbb{D}$, and the space $S^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$ is target-reduced.
 Proof.
-Assume that $\dim_{\mathbb{D}}H<n-1$ or that $\mathcal{S}^{\prime}$ is not target-reduced.
-Recall from Remark 3.1 that this yields a $\mathbb{D}$-linear subspace $G$ of $V$ with dimension n2n-2 such that $\mathcal{S}^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,G)$.
-Since $\mathcal{S}^{\prime}$ is deeply intransitive we deduce from point (a) of Theorem 1.20 that $\dim\mathcal{S}^{\prime}\leq\alpha(n-2,d)$.
-Hence $\dim\mathcal{S}-r\leq\alpha(n-2,d)$.
-Yet the starting assumption shows that $\dim\mathcal{S}-r\geq\alpha(n-1,d)-n+2+\varepsilon$, and hence $\alpha(n-2,d)\geq\alpha(n-1,d)-n+2+\varepsilon$, to the effect that $(n-1)d-1=\alpha(n-1,d)-\alpha(n-2,d)\leq n-2-\varepsilon$.
+Assume that $\dim_{\mathbb{D}}H<n-1$ or that $S^{\prime}$ is not target-reduced.
+Recall from Remark 3.1 that this yields a $\mathbb{D}$-linear subspace $G$ of $V$ with dimension n2n-2 such that $S^{\prime}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,G)$.
+Since $S^{\prime}$ is deeply intransitive we deduce from point (a) of Theorem 1.22 that $\dim S^{\prime}\leq\alpha(n-2,d)$.
+Hence $\dim S-r\leq\alpha(n-2,d)$.
+Yet the starting assumption shows that $\dim S-r\geq\alpha(n-1,d)-n+2+\varepsilon$, and hence $\alpha(n-2,d)\geq\alpha(n-1,d)-n+2+\varepsilon$, to the effect that $(n-1)d-1=\alpha(n-1,d)-\alpha(n-2,d)\leq n-2-\varepsilon$.
 Hence $(n-1)(d-1)\leq-\varepsilon$, which contradicts the definition of $\varepsilon$ because $n\geq 2$.
@@
 Denote by $H^{\prime}$ the socle of $\mathcal{H}$, and note that $H\subseteq H^{\prime}$ and $H^{\prime}$ is a $\mathbb{D}$-linear hyperplane of $V$.
-The $\mathbb{D}$-linear mappings $u_{1},\dots,u_{(n-1)d}$ map into $H^{\prime}$, and so do all the elements of $\mathcal{S}^{\prime}$.
-It follows that $\dim(\mathcal{S}^{H^{\prime}})\geq\dim\mathcal{S}^{\prime}+(n-1)d$.
-Yet $\mathcal{S}^{H^{\prime}}$ is deeply intransitive, so we can apply point (a) of Theorem 1.20 once more to find $\dim\mathcal{S}^{\prime}+(n-1)d\leq\alpha(n-1,d)$, leading to $\dim\mathcal{S}\leq\alpha(n-1,d)-(n-1)d+r=\alpha(n-2,d)+r-1$.
-A contradiction is then found like in the proof of Claim 5.1, using only $\dim\mathcal{S}-r\geq\alpha(n-1,d)-n+1+\varepsilon$ here.
+The $\mathbb{D}$-linear mappings $u_{1},\dots,u_{(n-1)d}$ map into $H^{\prime}$, and so do all the elements of $S^{\prime}$.
+It follows that $\dim(S^{H^{\prime}})\geq\dim S^{\prime}+(n-1)d$.
+Yet $S^{H^{\prime}}$ is deeply intransitive, so we can apply point (a) of Theorem 1.22 once more to find $\dim S^{\prime}+(n-1)d\leq\alpha(n-1,d)$, leading to $\dim S\leq\alpha(n-1,d)-(n-1)d+r=\alpha(n-2,d)+r-1$.
+A contradiction is then found like in the proof of Claim 5.1, using only $\dim S-r\geq\alpha(n-1,d)-n+1+\varepsilon$ here.
 ∎
 Claim 5.3.
-Assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1.
+Assume that $\dim S\geq\alpha(n,d)-n+2+\varepsilon$ or r<nd1r<nd-1.
 Let $X\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$.
@@
 The identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ then yields $\mathbf{a}(R_{1}\mathbf{C_{1}})=0$, whence $R_{1}\mathbf{C_{1}}=0$.
-This contradicts the fact that $\mathcal{S}^{\prime}$ is target-reduced (see Claim 5.1). ∎
+This contradicts the fact that $S^{\prime}$ is target-reduced (see Claim 5.1). ∎
 5.2 The transitive rank
 Claim 5.4.
-One has $\operatorname{\operatorname{trk}}(\mathcal{S})=nd-1$.
+One has $\operatorname{\operatorname{trk}}(S)=nd-1$.
 Proof.
 We assume that $r\leq nd-2$ and seek to find a contradiction.
-Denote by $r^{\prime}$ the rank of $\mathbf{C_{1}}$: it is also the transitive rank of $\mathcal{S}^{\prime}$.
-If $r^{\prime}<(n-1)d-1$, then by induction $\dim\mathcal{S}^{\prime}<\alpha(n-1,d)-(n-1)+1+\varepsilon$ and hence
-$\dim\mathcal{S}=r+\dim\mathcal{S}^{\prime}<nd-2+\alpha(n-1,d)-(n-1)+1+\varepsilon=\alpha(n,d)-n+1+\varepsilon,$
+Denote by $r^{\prime}$ the rank of $\mathbf{C_{1}}$: it is also the transitive rank of $S^{\prime}$.
+If $r^{\prime}<(n-1)d-1$, then by induction $\dim S^{\prime}<\alpha(n-1,d)-(n-1)+1+\varepsilon$ and hence
+$\dim S=r+\dim S^{\prime}<nd-2+\alpha(n-1,d)-(n-1)+1+\varepsilon=\alpha(n,d)-n+1+\varepsilon,$
 which contradicts our starting assumption.
-Hence $r^{\prime}=(n-1)d-1$ (remember that $\mathcal{S}^{\prime}$ is intransitive).
+Hence $r^{\prime}=(n-1)d-1$ (remember that $S^{\prime}$ is intransitive).
 Let us pick $\mathbb{F}$-linearly independent rows $L_{1}$ and $L_{2}$ in $\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})$.
@@
 Now we proceed to prove the last point.
-So, now we assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$, and we prove that $\mathcal{S}$ has a nonzero alternator.
+So, now we assume that $\dim S\geq\alpha(n,d)-n+2+\varepsilon$, and we prove that $S$ has a nonzero alternator.
 Note that the assumptions of Claims 5.1 and 5.3 are satisfied.
-In particular $\mathcal{S}^{\prime}$ is target-reduced.
+In particular $S^{\prime}$ is target-reduced.
 Next,
-$\dim\mathcal{S}^{\prime}=\dim\mathcal{S}-r\geq\alpha(n-1,d)-(n-1)+1+\varepsilon,$
-and hence by induction we recover $\operatorname{\operatorname{trk}}(\mathcal{S}^{\prime})=(n-1)d-1$.
-It then follows from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S}^{\prime})\leq 1$.
+$\dim S^{\prime}=\dim S-r\geq\alpha(n-1,d)-(n-1)+1+\varepsilon,$
+and hence by induction we recover $\operatorname{\operatorname{trk}}(S^{\prime})=(n-1)d-1$. It then follows from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(S^{\prime})\leq 1$.
 Since r=nd1r=nd-1, we see that $\mathbf{B_{1}}$ and $\mathbf{B_{2}}$ are simply row matrices, and we know from Claim 5.3 that $\operatorname{\operatorname{sprk}}(\mathbf{B_{1}})>1$.
-Since $\mathcal{S}^{\prime}$ is target-reduced, the identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ yields that $\mathbf{B_{1}}$ represents a nonzero alternator $b:U\times H\rightarrow\mathbb{F}$ of $\mathcal{S}^{\prime}$ (seen as a subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$).
-We deduce that $\operatorname{\operatorname{Alt}}(\mathcal{S}^{\prime})=\mathbb{F}b$, and hence the left radical
+Since $S^{\prime}$ is target-reduced, the identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ yields that $\mathbf{B_{1}}$ represents a nonzero alternator $b:U\times H\rightarrow\mathbb{F}$ of $S^{\prime}$ (seen as a subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$).
+We deduce that $\operatorname{\operatorname{Alt}}(S^{\prime})=\mathbb{F}b$, and hence the left radical
 $\mathcal{L}:=\operatorname{\operatorname{Lrad}}(b^{\prime})$
-does not depend on the choice of $b^{\prime}$ in $\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$ (in particular it is the left radical of $b$).
+does not depend on the choice of $b^{\prime}$ in $\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$ (in particular it is the left radical of $b$).
 For all $y\in U$, the specialized matrix $\mathbf{B_{1}}(y)$ represents the $\mathbb{F}$-linear form b(y,)b(y,-) in the basis $(e_{1},\dots,e_{(n-1)d})$ of $H$.
@@
 Claim 5.5.
-Let $y\in U\smallsetminus\mathcal{L}$. The mappings $u\in\mathcal{S}$ such that $u(x)\in H$ do not all map $y$ into $\mathcal{S}x$.
+Let $y\in U\smallsetminus\mathcal{L}$. The mappings $u\in S$ such that $u(x)\in H$ do not all map $y$ into $Sx$.
 We now move on to what is arguably the most subtle item in our proof:
@@
 This linear form is nonzero and its kernel includes $H$.
-The assumption $L\mathbf{N}(y)=0$ then shows that all the operators $u_{1},\dots,u_{(n-1)d}$ map $y$ into $\operatorname{\operatorname{Ker}}\ell$, and hence all the operators of $\mathcal{S}$ that map $x$ into $H$ also map $y$ into $\operatorname{\operatorname{Ker}}\ell$.
-Next, we prove that $\operatorname{\operatorname{Ker}}\ell=(\mathcal{S}x)a^{-1}$ for some $a\in\mathbb{D}\smallsetminus\{0\}$.
-Choose indeed an $\mathbb{F}$-linear form $\ell^{\prime}$ on $V$ with kernel $\mathcal{S}x$.
+The assumption $L\mathbf{N}(y)=0$ then shows that all the operators $u_{1},\dots,u_{(n-1)d}$ map $y$ into $\operatorname{\operatorname{Ker}}\ell$, and hence all the operators of $S$ that map $x$ into $H$ also map $y$ into $\operatorname{\operatorname{Ker}}\ell$.
+Next, we prove that $\operatorname{\operatorname{Ker}}\ell=(Sx)a^{-1}$ for some $a\in\mathbb{D}\smallsetminus\{0\}$.
+Choose indeed an $\mathbb{F}$-linear form $\ell^{\prime}$ on $V$ with kernel $Sx$.
 The linear forms $z\in V\mapsto\ell^{\prime}(za)$, with $a\in\mathbb{D}$, constitute a $d$-dimensional $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(V,\mathbb{F})$, and they all vanish on $H$.
@@
 It follows that
-$\operatorname{\operatorname{Ker}}\ell=(\operatorname{\operatorname{Ker}}\ell^{\prime})a^{-1}=(\mathcal{S}x)a^{-1}=\mathcal{S}(xa^{-1}).$
+$\operatorname{\operatorname{Ker}}\ell=(\operatorname{\operatorname{Ker}}\ell^{\prime})a^{-1}=(Sx)a^{-1}=S(xa^{-1}).$
 Now the trick is to use the fact that $xa^{-1}$ satisfies exactly the same assumptions as the vector $x$ we started from.
-Indeed $\operatorname{\operatorname{rk}}(\widehat{xa^{-1}})=\dim_{\mathbb{F}}\bigl{(}\mathcal{S}(xa^{-1})\bigr{)}=nd-1$, the operators in $\mathcal{S}$ that vanish at $x$ are also the ones that vanish at $xa^{-1}$, and $H$ is the $\mathbb{D}$-linear hyperplane of $V$ that is included in $\mathcal{S}(xa^{-1})$.
+Indeed $\operatorname{\operatorname{rk}}(\widehat{xa^{-1}})=\dim_{\mathbb{F}}\bigl(S(xa^{-1})\bigr)=nd-1$, the operators in $S$ that vanish at $x$ are also the ones that vanish at $xa^{-1}$, and $H$ is the $\mathbb{D}$-linear hyperplane of $V$ that is included in $S(xa^{-1})$.
 Hence Claim 5.5 applies when $x$ is replaced with $xa^{-1}$.
-Every $u\in\mathcal{S}$ that maps $x$ into $H$ also maps $xa^{-1}$ into $H$ because $u$ is $\mathbb{D}$-linear, and hence there is at least one such operator that does not map $y$ into $\mathcal{S}(xa^{-1})$.
+Every $u\in S$ that maps $x$ into $H$ also maps $xa^{-1}$ into $H$ because $u$ is $\mathbb{D}$-linear, and hence there is at least one such operator that does not map $y$ into $S(xa^{-1})$.
 This contradicts the first step of the present proof.
@@
 ∎
-5.4 Constructing a nonzero alternator of $\mathcal{S}$
-Now everything is in place for us to produce a nonzero alternator of $\mathcal{S}$.
+5.4 Constructing a nonzero alternator of $S$
+Now everything is in place for us to produce a nonzero alternator of $S$.
 We start by applying identity (3) once more. This time around, by taking $k=1$ we find $\mathbf{B}\mathbf{A}\mathbf{C}=0$. Writing
@@
 $\mathbf{B^{\prime}_{1}}=\mathbf{a}\mathbf{B_{1}}\quad\text{for some $1$-homogeneous $\mathbf{a}\in R$.}$
-Now, we consider the new row $\mathbf{R}:=\mathbf{B}(\mathbf{A}-\mathbf{a}I_{nd-1})$, which can be written $\begin{bmatrix}[0]_{1\times(n-1)d}&\mathbf{R_{2}}\end{bmatrix}$. We also write
+Now, we consider the new row $\mathbf{R}:=\mathbf{B}(\mathbf{A}-\mathbf{a}I_{nd-1})=\begin{bmatrix}[0]_{1\times(n-1)d}&\mathbf{R_{2}}\end{bmatrix}$. We also write
 $\mathbf{R}\mathbf{A}=\begin{bmatrix}\mathbf{R^{\prime}_{1}}&\mathbf{R^{\prime}_{2}}\end{bmatrix}$
@@
 By combining this with $\mathbf{B}\mathbf{C}=0$, we conclude that $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$ is a (nonzero) catcher of $\mathbf{M}$.
-Since it was known from the start that $\mathcal{S}$ is target-reduced, the process explained in the end of Section 4 yields a nonzero alternator $\widetilde{b}:U\times V\rightarrow\mathbb{F}$ of $\mathcal{S}$, associated with the nonzero catcher $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$.
-Finally, since $\mathcal{S}$ has transitive rank nd1nd-1 and is target-reduced, we know from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
+Since it was known from the start that $S$ is target-reduced, the process explained in the end of Section 4 yields a nonzero alternator $\widetilde{b}:U\times V\rightarrow\mathbb{F}$ of $S$, associated with the nonzero catcher $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$.
+Finally, since $S$ has transitive rank nd1nd-1 and is target-reduced, we know from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(S)=1$.
 This completes the proof of Proposition 5.1.
@@
 We will now complete the proof of Theorem 1.22, by proving point (c).
-We will prove the following general result, which avoids any dimensional assumption on $\mathcal{S}$ and focuses on the condition regarding the structure of the alternator space.
+We will prove the following general result, which avoids any dimensional assumption on $S$ and focuses on the condition regarding the structure of the alternator space.
 Proposition 5.2.
-Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$, and let $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$.
+Let $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$, and let $b\in\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$.
 Then:
@@
 Then $a\mapsto[c\mapsto c^{a}]$ defines an $\mathbb{F}$-quadratic mapping $P:\mathbb{D}\rightarrow\operatorname{\operatorname{End}}_{\mathbb{F}}(\mathcal{B}(U\times V,\mathbb{F}))$, with $P(aa^{\prime})=P(a)\circ P(a^{\prime})$ for all $(a,a^{\prime})\in\mathbb{D}^{2}$, i.e., $P$ is multiplicative.
-Moreover, $P(a)$ leaves $\operatorname{\operatorname{Alt}}(\mathcal{S})$ invariant for all $a\in\mathbb{D}$, as for all $c\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ we observe that
-$\forall u\in\mathcal{S},\;\forall x\in U,\;c(xa,u(x)a)=c(xa,u(xa))=0.$
-Since $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$, we deduce that $b$ is an eigenvector of $P(a)$ for all $a\in\mathbb{D}$, and by denoting by $q(a)\in\mathbb{F}$ the corresponding eigenvalue we obtain an $\mathbb{F}$-quadratic form $q:\mathbb{D}\rightarrow\mathbb{F}$.
+Moreover, $P(a)$ leaves $\operatorname{\operatorname{Alt}}(S)$ invariant for all $a\in\mathbb{D}$, as for all $c\in\operatorname{\operatorname{Alt}}(S)$ we observe that
+$\forall u\in S,\;\forall x\in U,\;c(xa,u(x)a)=c(xa,u(xa))=0.$
+Since $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$, we deduce that $b$ is an eigenvector of $P(a)$ for all $a\in\mathbb{D}$, and by denoting by $q(a)\in\mathbb{F}$ the corresponding eigenvalue we obtain an $\mathbb{F}$-quadratic form $q:\mathbb{D}\rightarrow\mathbb{F}$.
 Obviously $\operatorname{\operatorname{rk}}b^{a}=\operatorname{\operatorname{rk}}b$ for all $a\in\mathbb{D}^{\times}$ (because $\mathbb{D}$ is a division ring), and in particular $q$ is nonisotropic.
@@
 Moreover $e$ satisfies the identity $\forall(x,y)\in\mathbb{D}^{2},\;e(yx)=e(xy)$, which will ease things up from now on.
-As seen in Section 2.1, since $e$ is nonzero we recover a unique $\mathbb{F}$-bilinear and right-$\mathbb{D}$-linear form $B$ on $U\times V$ such that
+As seen in Section 2, since $e$ is nonzero we recover a unique $\mathbb{F}$-bilinear and right-$\mathbb{D}$-linear form $B$ on $U\times V$ such that
 $\forall(x,y)\in U\times V,\;b(x,y)=e(B(x,y)).$
@@
 Now we can complete the proof of point (c) of Theorem 1.22.
-Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ with $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ and $|\mathbb{F}|\geq nd$.
-By Proposition 5.1 we have $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
-Choose $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$.
+Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ with $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$ and $|\mathbb{F}|\geq nd$.
+By Proposition 5.1 we have $\dim\operatorname{\operatorname{Alt}}(S)=1$.
+Choose $b\in\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$.
 We need to prove that $b$ is right-nondegenerate.
@@
 The bilinear form $b$ induces a right-nondegenerate $\mathbb{F}$-bilinear form $\overline{b}$ on $U\times(V/R)$.
-The space $\pi\mathcal{S}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/R)$, and it is obvious that $\overline{b}$ is an alternator of it.
-Since $\overline{b}$ is right-nondegenerate, we know from Proposition 1.18 that $\pi\mathcal{S}$ is deeply intransitive.
-Besides, the kernel of $u\in\mathcal{S}\mapsto\pi\circ u$ is precisely $\mathcal{S}^{R}$, which is deeply intransitive.
+The space $\pi S$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/R)$, and it is obvious that $\overline{b}$ is an alternator of it.
+Since $\overline{b}$ is right-nondegenerate, we know from Proposition 1.20 that $\pi S$ is deeply intransitive.
+Besides, the kernel of $u\in S\mapsto\pi\circ u$ is precisely $S^{R}$, which is deeply intransitive.
 Hence, by combining the rank theorem with point (a) of Theorem 1.22, we find
-$\dim(\mathcal{S})=\dim(\pi\mathcal{S})+\dim(\mathcal{S}^{R})\leq\alpha(n-t,d)+\alpha(t,d)=\alpha(n,d)-dt(n-t).$
-It $t>0$ then $dt(n-t)\geq d(n-1)$, and we observe that $d(n-1)>n-\max(4-d,2)$ because obviously $(d-1)(n-1)\geq 0>1-\max(4-d,2)$.
-This runs counter to our starting assumption on the dimension of $\mathcal{S}$.
+$\dim(S)=\dim(\pi S)+\dim(S^{R})\leq\alpha(n-t,d)+\alpha(t,d)=\alpha(n,d)-dt(n-t).$
+It $t>0$ then $dt(n-t)\geq d(n-1)$, and we observe that $d(n-1)>n-\max(4-d,2)$ because obviously $(d-1)(n-1)\geq 0>1-\max(4-d,2)$. This runs counter to our starting assumption on the dimension of $S$.
 Therefore $t=0$, to the effect that $b$ is right-nondegenerate. This completes the proof of point (c) in Theorem 1.22.
@@
 Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V>0$, and set $n:=\dim_{\mathbb{D}}U$.
-Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ for all $x\in U\smallsetminus\{0\}$.
-Assume that $\mathcal{S}$ is target-reduced.
-Then all the nonzero alternators of $\mathcal{S}$ are nondegenerate, and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension at most $1$.
+Let $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim_{\mathbb{F}}(Sx)=nd-1$ for all $x\in U\smallsetminus\{0\}$.
+Assume that $S$ is target-reduced.
+Then all the nonzero alternators of $S$ are nondegenerate, and $\operatorname{\operatorname{Alt}}(S)$ has dimension at most $1$.
 Proof.
-Let $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ be nonzero and degenerate.
+Let $b\in\operatorname{\operatorname{Alt}}(S)$ be nonzero and degenerate.
 It follows that $nd=\dim_{\mathbb{F}}V>1$.
-Every element of $\mathcal{S}$ maps the left radical $L$ of $b$ into its right radical $R$, and here $\dim_{\mathbb{F}}L=\dim_{\mathbb{F}}R$ since $\dim_{\mathbb{F}}U=\dim_{\mathbb{F}}V$.
-We choose $x\in L\smallsetminus\{0\}$ and deduce that $\dim_{\mathbb{F}}R\geq\dim_{\mathbb{F}}\mathcal{S}x=nd-1$.
+Every element of $S$ maps the left radical $L$ of $b$ into its right radical $R$, and here $\dim_{\mathbb{F}}L=\dim_{\mathbb{F}}R$ since $\dim_{\mathbb{F}}U=\dim_{\mathbb{F}}V$.
+We choose $x\in L\smallsetminus\{0\}$ and deduce that $\dim_{\mathbb{F}}R\geq\dim_{\mathbb{F}}Sx=nd-1$.
 Hence $\operatorname{\operatorname{rk}}(b)=1$.
-Then each bilinear form $(x,y)\in U^{2}\mapsto b(x,u(y))$, with $u\in\mathcal{S}$, is alternating with rank at most $1$, and hence is zero.
-Hence $\operatorname{\operatorname{Im}}u\subseteq R$ for all $u\in\mathcal{S}$, which contradicts the assumption that $\mathcal{S}$ is target-reduced.
-Hence, every nonzero element of $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is nondegenerate.
-Next, let $b$ and $b^{\prime}$ belong to $\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$.
+Then each bilinear form $(x,y)\in U^{2}\mapsto b(x,u(y))$, with $u\in S$, is alternating with rank at most $1$, and hence is zero.
+Hence $\operatorname{\operatorname{Im}}u\subseteq R$ for all $u\in S$, which contradicts the assumption that $S$ is target-reduced.
+Hence, every nonzero element of $\operatorname{\operatorname{Alt}}(S)$ is nondegenerate.
+Next, let $b$ and $b^{\prime}$ belong to $\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$.
 Choose $x\in U\smallsetminus\{0\}$.
 By the first step, $b(x,-)\neq 0$.
-The nonzero linear forms $b^{\prime}(x,-)$ and b(x,)b(x,-) vanish on the $\mathbb{F}$-linear hyperplane $\mathcal{S}x$, and hence $b^{\prime}(x,-)=\lambda b(x,-)$ for some $\lambda\in\mathbb{F}$.
-Then $b^{\prime}-\lambda b$ is a degenerate element in $\operatorname{\operatorname{Alt}}(\mathcal{S})$, and by the first step $b^{\prime}-\lambda b=0$.
-We conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$.
+The nonzero linear forms $b^{\prime}(x,-)$ and b(x,)b(x,-) vanish on the $\mathbb{F}$-linear hyperplane $Sx$, and hence $b^{\prime}(x,-)=\lambda b(x,-)$ for some $\lambda\in\mathbb{F}$.
+Then $b^{\prime}-\lambda b$ is a degenerate element in $\operatorname{\operatorname{Alt}}(S)$, and by the first step $b^{\prime}-\lambda b=0$.
+We conclude that $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$.
 ∎
@@
 Proof.
-Set $\mathcal{S}:=\mathcal{A}_{b,\mathbb{D}}$.
+Set $S:=\mathcal{A}_{b,\mathbb{D}}$.
 To start with, to prove the first point we reduce the situation to the one where $B$ is nondegenerate.
 To see this, we consider the left radical $L$ of $B$, which equals the one of $b$, and the induced form $\overline{B}:(U/L)\times V\rightarrow\mathbb{D}$, and the $\mathbb{F}$-bilinear form $\overline{b}:(x,y)\in(U/L)\times V\mapsto e(\overline{B}(x,y))$.
-All the operators in $\mathcal{S}$ map $L$ into the right radical of $b$, which is zero, so they all vanish on $L$.
-Hence every $u\in\mathcal{S}$ induces a $\mathbb{D}$-linear mapping $\overline{u}:U/L\rightarrow V$, and we have $\overline{b}(\overline{x},\overline{u}(\overline{x}))=0$ for all $x\in U/L$.
-We denote by $\overline{\mathcal{S}}$ the space of all mappings $\overline{u}$ with $u\in\mathcal{S}$.
-Conversely, for all $u\in\mathcal{A}_{\overline{b},\mathbb{D}}$ it is clear that we have $u\circ\pi\in\mathcal{S}$ for the standard projection $\pi:U\twoheadrightarrow U/L$.
-Therefore $\overline{\mathcal{S}}=\mathcal{A}_{\overline{b},\mathbb{D}}$.
-Since $u\in\mathcal{S}\mapsto\overline{u}\in\overline{\mathcal{S}}$ is clearly injective, we deduce that $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{A}_{\overline{b},\mathbb{D}}$.
+All the operators in $S$ map $L$ into the right radical of $b$, which is zero, so they all vanish on $L$.
+Hence every $u\in S$ induces a $\mathbb{D}$-linear mapping $\overline{u}:U/L\rightarrow V$, and we have $\overline{b}(\overline{x},\overline{u}(\overline{x}))=0$ for all $x\in U/L$.
+We denote by $\overline{S}$ the space of all mappings $\overline{u}$ with $u\in S$.
+Conversely, for all $u\in\mathcal{A}_{\overline{b},\mathbb{D}}$ it is clear that we have $u\circ\pi\in S$ for the standard projection $\pi:U\twoheadrightarrow U/L$.
+Therefore $\overline{S}=\mathcal{A}_{\overline{b},\mathbb{D}}$.
+Since $u\in S\mapsto\overline{u}\in\overline{S}$ is clearly injective, we deduce that $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{A}_{\overline{b},\mathbb{D}}$.
 Hence, in the remainder of the proof we need only consider the situation where $\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V=n$.
@@
 Let $x\in U\smallsetminus\{0\}$, and extend it into a $\mathbb{D}$-basis $\mathbf{B}_{U}$ of $U$ with $x$ as first vector, and consider a $\mathbb{D}$-basis $\mathbf{B}_{V}$ of $V$.
-Denote by $P$ the Gram matrix of $B$ in the bases $U$ and $V$, and set $\mathcal{M}=\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$.
-Then the elements of $\mathcal{S}$ are represented by the elements of $P^{-1}\mathcal{M}$ in the bases $\mathbf{B}_{U}$ and $\mathbf{B}_{V}$, and in particular $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{M}$.
+Denote by $P$ the Gram matrix of $B$ in the bases $U$ and $V$, and set $\mathcal{M}=S\mathcal{H}_{n}(\mathbb{D})$.
+Then the elements of $S$ are represented by the elements of $P^{-1}\mathcal{M}$ in the bases $\mathbf{B}_{U}$ and $\mathbf{B}_{V}$, and in particular $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{M}$.
 Then, it is easily checked in each case that
@@
 Next, denote by $X$ the first vector of the standard basis of $\mathbb{D}^{n}$.
-Then we observe that $\mathcal{M}X=(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-1}$ and hence $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$.
-Moreover if $n\geq 2$ and we take $X^{\prime}$ as the second vector of the standard basis we find $\mathcal{M}X^{\prime}=\mathbb{D}\times(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-2}$, and it follows that $\mathcal{S}x+\mathcal{S}x^{\prime}=V$ for some $x^{\prime}\in U$.
-If $d>1$ or $n>1$ we have collected enough information to see that $\mathcal{S}$ is target-reduced, and then we conclude from Lemma 5.3 that $\operatorname{\operatorname{Alt}}(\mathcal{S})=\mathbb{F}b$.
-If $d=1$ and $n=1$ then $\mathcal{S}=\{0\}$ and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$ (it is the space of all $\mathbb{F}$-bilinear forms on $U\times V$).
+Then we observe that $\mathcal{M}X=(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-1}$ and hence $\dim_{\mathbb{F}}(Sx)=nd-1$.
+Moreover if $n\geq 2$ and we take $X^{\prime}$ as the second vector of the standard basis we find $\mathcal{M}X^{\prime}=\mathbb{D}\times(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-2}$, and it follows that $Sx+Sx^{\prime}=V$ for some $x^{\prime}\in U$.
+If $d>1$ or $n>1$ we have collected enough information to see that $S$ is target-reduced, and then we conclude from Lemma 5.3 that $\operatorname{\operatorname{Alt}}(S)=\mathbb{F}b$.
+If $d=1$ and $n=1$ then $S=\{0\}$ and $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$ (it is the space of all $\mathbb{F}$-bilinear forms on $U\times V$).
 Assume finally that $U=V$. By definition, $B$ is $e$-nonisotropic if and only if $b$ is nonisotropic. We already know that $\mathcal{A}_{b,\mathbb{D}}$ has trivial spectrum if $b$ is nonisotropic.
@@
   • Let $x\in V\smallsetminus\{0\}$.
-Then as $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ and $\mathcal{S}x$ is included in the right-orthogonal of $x$ under $b$, it equals this right-orthogonal, which yields $b(x,x)\neq 0$ since $x\not\in\mathcal{S}x$.
+Then as $\dim_{\mathbb{F}}(Sx)=nd-1$ and $Sx$ is included in the right-orthogonal of $x$ under $b$, it equals this right-orthogonal, which yields $b(x,x)\neq 0$ since $x\not\in Sx$.
 Hence $b$ is nonisotropic.
-  • Let us finally prove that $\mathcal{S}$ is irreducible.
+  • Let us finally prove that $S$ is irreducible.
 Let $x\in V\smallsetminus\{0\}$.
-Combining $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ with $x\not\in\mathcal{S}x$, we find $\mathbb{F}x\oplus\mathcal{S}x=V$.
-Therefore, $\mathcal{S}$ is irreducible.
-6 Trivial spectrum subspaces
+Combining $\dim_{\mathbb{F}}(Sx)=nd-1$ with $x\not\in Sx$, we find $\mathbb{F}x\oplus Sx=V$.
+Therefore, $S$ is irreducible.
+6 Optimal trivial spectrum subspaces
 In this section, we will quickly derive the main results on irreducible optimal trivial spectrum subspaces from their analogues on deeply intransitive operator spaces.
@@
 Lemma 6.1 (Invariant Subspace Lemma).
-Let $\mathcal{S}$ be an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
-Assume that there exists a $\mathbb{D}$-linear subspace $W$ of $V$ such that $\mathcal{S}$ contains every operator $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that satisfies $\operatorname{\operatorname{Im}}u\subseteq W\subseteq\operatorname{\operatorname{Ker}}u$.
-Then $W$ is invariant under $\mathcal{S}$.
+Let $S$ be a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+Assume that there exists a $\mathbb{D}$-linear subspace $W$ of $V$ such that $S$ contains every operator $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that satisfies $\operatorname{\operatorname{Im}}u\subseteq W\subseteq\operatorname{\operatorname{Ker}}u$.
+Then $W$ is invariant under $S$.
 Proof.
-Let $u\in\mathcal{S}$.
+Let $u\in S$.
 Assume that there is a (nonzero) vector $x\in W$ such that $u(x)\not\in W$.
@@
 Denote by $\pi$ the projection of $V$ onto $W$ along $W^{\prime}$, and by $v\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ the linear mapping that vanishes on $W$ and satisfies $v(y)=\pi(u(y))$ for all $y\in W^{\prime}$.
-Then $v\in\mathcal{S}$ by our assumptions, and hence $u-v\in\mathcal{S}$.
+Then $v\in S$ by our assumptions, and hence $u-v\in S$.
 Note that uvu-v leaves $W^{\prime}$ invariant.
 Moreover $(u-v)(x)=u(x)\equiv x$ mod $W^{\prime}$.
-The endomorphism of $V/W^{\prime}$ induced by uvu-v then has a nonzero fixed vector, which contradicts the assumption that it should have $\mathbb{F}$-trivial spectrum.
-Therefore $W$ is invariant under $\mathcal{S}$.
+The endomorphism of $V/W^{\prime}$ induced by uvu-v then has a nonzero fixed vector, which contradicts the assumption that it should have trivial spectrum.
+Therefore $W$ is invariant under $S$.
 ∎
@@
 The case $n=1$ is obvious.
-So, we let $n\geq 2$ be an integer, $V$ be an $n$-dimensional $\mathbb{D}$-vector space, and $\mathcal{S}$ be an irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$.
+So, we let $n\geq 2$ be an integer, $V$ be an $n$-dimensional $\mathbb{D}$-vector space, and $S$ be an irreducible trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$.
 We also assume that $|\mathbb{F}|\geq nd$.
-We immediately note that $\mathcal{S}$ is intransitive.
-Case 1. The space $\mathcal{S}$ is primitively intransitive.
-Then $\mathcal{S}$ is deeply intransitive, by Proposition 4.5.
-  • Assume first that $(n,d)\neq(2,1)$.
-Then $n\geq\max(4-d,2)$, and hence $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
-By point (c) of Theorem 1.20, we find that $\mathcal{S}$ has a nondegenerate alternator $b$.
+We immediately note that $S$ is intransitive.
+Case 1. The space $S$ is primitively intransitive.
+Then $S$ is deeply intransitive, by Proposition 4.5.
+  • Assume first that $(n,d)\neq(2,1)$. Then $n\geq\max(4-d,2)$, and hence $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$. By point (c) of Theorem 1.22, we find that $S$ has a nondegenerate alternator $b$.
   • Assume now that $(n,d)=(2,1)$.
-Then we simply note that $\dim\mathcal{S}=1$, so we can write $\mathcal{S}=\mathbb{F}u_{0}$ for some operator $u_{0}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$; then $u_{0}$ is invertible because of the irreducibility assumption, and hence by picking an arbitrary symplectic form $s$ on the $\mathbb{F}$-vector space $V$ (which exists because $n=2$ and $d=1$), we see that $b:(x,y)\mapsto s(x,u_{0}^{-1}(y))$ is a nondegenerate alternator of $\mathcal{S}$.
-In any case, we have found a nondegenerate alternator $b$ of $\mathcal{S}$, to the effect that $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$.
+Then we simply note that $\dim S=1$, so we can write $S=\mathbb{F}u_{0}$ for some operator $u_{0}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$; then $u_{0}$ is invertible because of the irreducibility assumption, and hence by picking an arbitrary symplectic form $s$ on the $\mathbb{F}$-vector space $V$ (which exists because $n=2$ and $d=1$), we see that $b:(x,y)\mapsto s(x,u_{0}^{-1}(y))$ is a nondegenerate alternator of $S$.
+In any case, we have found a nondegenerate alternator $b$ of $S$, to the effect that $S\subseteq\mathcal{A}_{b,\mathbb{D}}$.
 Yet $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive (by point (b) of Proposition 1.20), and hence $\dim\mathcal{A}_{b,\mathbb{D}}\leq\alpha(n,d)$ by point (a) of Theorem 1.22.
-Therefore $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$.
-Case 2. The space $\mathcal{S}$ is not primitively intransitive.
-Then we take a maximal $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the projected space $\pi\mathcal{S}$ is intransitive.
+Therefore $S=\mathcal{A}_{b,\mathbb{D}}$.
+Case 2. The space $S$ is not primitively intransitive.
+Then we take a maximal $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the projected space $\pi S$ is intransitive.
 Note that $\{0\}\subset V^{\prime}\subset V$.
-By the maximality of $V^{\prime}$, the space $\pi\mathcal{S}$ is primitively intransitive, and hence deeply intransitive (see Proposition 4.5).
+By the maximality of $V^{\prime}$, the space $\pi S$ is primitively intransitive, and hence deeply intransitive (see Proposition 4.5).
 Set then $t:=\dim_{\mathbb{D}}V^{\prime}$.
-Applying point (a) of Theorem 1.20, we find $\dim(\pi\mathcal{S})\leq\alpha(n-t,d)$.
-Now, we look at the space $\mathcal{S}^{V^{\prime}}$.
+Applying point (a) of Theorem 1.22, we find $\dim(\pi S)\leq\alpha(n-t,d)$.
+Now, we look at the space $S^{V^{\prime}}$.
 First of all, the rank theorem yields
-$\dim(\mathcal{S}^{V^{\prime}})=\dim(\mathcal{S})-\dim(\pi\mathcal{S})\geq\alpha(n,d)-\alpha(n-t,d)=dt(n-t)+\alpha(t,d).$
-Assigning to every $f\in\mathcal{S}^{V^{\prime}}$ its restriction $f_{V^{\prime}}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$, we create an $\mathbb{F}$-linear mapping $\Phi$ whose range is an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$ and whose kernel is the subspace $\mathcal{D}$ of all $f\in\mathcal{S}^{V^{\prime}}$ that vanish on $V^{\prime}$.
+$\dim(S^{V^{\prime}})=\dim(S)-\dim(\pi S)\geq\alpha(n,d)-\alpha(n-t,d)=dt(n-t)+\alpha(t,d).$
+Assigning to every $f\in S^{V^{\prime}}$ its restriction $f_{V^{\prime}}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$, we create an $\mathbb{F}$-linear mapping $\Phi$ whose range is a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$ and whose kernel is the subspace $\mathcal{D}$ of all $f\in S^{V^{\prime}}$ that vanish on $V^{\prime}$.
 By Theorem 3.1 we have $\operatorname{\operatorname{rk}}\Phi\leq\alpha(t,d)$, whence the rank theorem yields
-$\dim\mathcal{D}\geq\dim(\mathcal{S}^{V^{\prime}})-\alpha(t,d)\geq dt(n-t).$
+$\dim\mathcal{D}\geq\dim(S^{V^{\prime}})-\alpha(t,d)\geq dt(n-t).$
 Yet $\mathcal{D}$ is included in the space of all $\mathbb{D}$-endomorphisms of $V$ that vanish on $V^{\prime}$ and map into $V^{\prime}$, a space which has dimension dt(nt)dt(n-t) over $\mathbb{F}$.
-Hence $\mathcal{D}$ equals that space, to the effect that all such endomorphisms belong to $\mathcal{S}$.
-Then the Invariant Subspace Lemma yields that $V^{\prime}$ is invariant under $\mathcal{S}$.
-However this contradicts the assumption that $\mathcal{S}$ is irreducible, and we conclude that Case 2 is actually impossible.
+Hence $\mathcal{D}$ equals that space, to the effect that all such endomorphisms belong to $S$.
+Then the Invariant Subspace Lemma yields that $V^{\prime}$ is invariant under $S$. However this contradicts the assumption that $S$ is irreducible, and we conclude that Case 2 is actually impossible.
 Hence, Theorem 1.9 is now proved.
@@
 Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$.
-Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$ such that $\mathcal{S}:=\mathcal{A}_{b,\mathbb{D}}$ is an optimal irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
-We shall prove that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$ in any case.
-  • Assume first that $n\geq\max(4-d,2)$. Then $\dim\mathcal{S}=\alpha(n,d)\geq\alpha(n,d)-n+\max(4-d,2)$ and the claimed statement is deduced from point (c) of Theorem 1.20.
+Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$ such that $S:=\mathcal{A}_{b,\mathbb{D}}$ is an optimal irreducible trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
+We shall prove that $\dim\operatorname{\operatorname{Alt}}(S)=1$ in any case.
+  • Assume first that $n\geq\max(4-d,2)$. Then $\dim S=\alpha(n,d)\geq\alpha(n,d)-n+\max(4-d,2)$ and the claimed statement is deduced from point (c) of Theorem 1.22.
   • Assume now that $n=2$ and $d=1$.
-Then $\dim\mathcal{S}=1$, and we take $u_{0}\in\mathcal{S}\smallsetminus\{0\}$.
-Since $\mathcal{S}$ is irreducible, $u_{0}$ must be invertible, and we deduce that $\dim\mathcal{S}x=1=nd-1$ for all $x\in V\smallsetminus\{0\}$.
-Moreover $\mathcal{S}$ is target-reduced because $u_{0}$ is surjective.
-Hence Lemma 5.3 shows that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$ because $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})$.
-  • If $n=1$ and $d>1$ then $\dim_{\mathbb{F}}(\mathcal{S}x)=\dim\mathcal{S}=nd-1$ for all $x\in V\smallsetminus\{0\}$, and obviously $\mathcal{S}$ is target-reduced, so once more we find by Proposition 5.3 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
-  • If $n=1$ and $d=1$ then $\mathcal{S}=\{0\}$ and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is the space of all $\mathbb{F}$-bilinear forms on $V$, which has dimension $1$.
+Then $\dim S=1$, and we take $u_{0}\in S\smallsetminus\{0\}$.
+Since $S$ is irreducible, $u_{0}$ must be invertible, and we deduce that $\dim Sx=1=nd-1$ for all $x\in V\smallsetminus\{0\}$.
+Moreover $S$ is target-reduced because $u_{0}$ is surjective.
+Hence Lemma 5.3 shows that $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(S)=1$ because $b\in\operatorname{\operatorname{Alt}}(S)$.
+  • If $n=1$ and $d>1$ then $\dim_{\mathbb{F}}(Sx)=\dim S=nd-1$ for all $x\in V\smallsetminus\{0\}$, and obviously $S$ is target-reduced, so once more we find by Proposition 5.3 that $\dim\operatorname{\operatorname{Alt}}(S)=1$.
+  • If $n=1$ and $d=1$ then $S=\{0\}$ and $\operatorname{\operatorname{Alt}}(S)$ is the space of all $\mathbb{F}$-bilinear forms on $V$, which has dimension $1$.
 Then, we apply Proposition 5.2 and find that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and that, for an associated pair $(\sigma,e)$, there exists a $\sigma$-sesquilinear form $B:V^{2}\rightarrow\mathbb{D}$ such that $b:(x,y)\mapsto e(B(x,y))$.
-Besides, it has already been observed in Section 2.1 that $B$ is uniquely determined by $b$.
+Besides, it has already been observed in Section 2 that $B$ is uniquely determined by $b$.
 Finally, the converse statement in Theorem 1.10 has already been proved through Proposition 5.4.
 6.4 The quadratic type case
-We finish this section by proving the results that are specific to the division algebras of quadratic type: Theorems 1.10 and 1.11.
+We finish this section by proving the results that are specific to the division algebras of quadratic type: Theorems 1.12 and 1.13.
 So, assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and denote by $(\sigma,e)$ an associated pair.
 Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $n\geq 1$.
-Let $\mathcal{S}$ be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ such that $|\mathbb{F}|\geq nd$.
-If $n\geq 2$ then we apply Theorems 1.7 and 1.8 to find a $\sigma$-sesquilinear nondegenerate form $B:V^{2}\rightarrow\mathbb{D}$ such that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$ for $b:(x,y)\mapsto e(B(x,y))$.
+Let $S$ be an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ such that $|\mathbb{F}|\geq nd$.
+If $n\geq 2$ then we apply Theorems 1.9 and 1.10 to find a $\sigma$-sesquilinear nondegenerate form $B:V^{2}\rightarrow\mathbb{D}$ such that $S=\mathcal{A}_{b,\mathbb{D}}$ for $b:(x,y)\mapsto e(B(x,y))$.
 Assume now that $n=1$.
 In that case we need only consider the special case where $V=\mathbb{D}$.
-By the isomorphism $a\in\mathbb{D}\mapsto[x\mapsto ax]\in\operatorname{\operatorname{End}}_{\mathbb{D}}(\mathbb{D})$ of $\mathbb{F}$-vector spaces, we associate to $\mathcal{S}$ an $\mathbb{F}$-linear hyperplane $H$ of $\mathbb{D}$.
+By the isomorphism $a\in\mathbb{D}\mapsto[x\mapsto ax]\in\operatorname{\operatorname{End}}_{\mathbb{D}}(\mathbb{D})$ of $\mathbb{F}$-vector spaces, we associate to $S$ an $\mathbb{F}$-linear hyperplane $H$ of $\mathbb{D}$.
 Thus $H$ is the kernel of $x\mapsto e(\alpha x)$ for some $\alpha\in\mathbb{D}^{\times}$.
@@
 $\forall x\in\mathbb{D},\;e(x^{\sigma}\alpha(ax))=e(\alpha a\underbrace{xx^{\sigma}}_{\in\mathbb{F}})=(xx^{\sigma})\,e(\alpha a)=0$
-This shows that the nondegenerate bilinear form $b:(x,y)\in\mathbb{D}^{2}\mapsto e(x^{\sigma}\alpha y)$ is an alternator of $\mathcal{S}$.
-Hence $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$.
-Yet $\dim\mathcal{A}_{b,\mathbb{D}}\leq d-1$ because $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive, and we conclude that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$.
+This shows that the nondegenerate bilinear form $b:(x,y)\in\mathbb{D}^{2}\mapsto e(x^{\sigma}\alpha y)$ is an alternator of $S$.
+Hence $S\subseteq\mathcal{A}_{b,\mathbb{D}}$.
+Yet $\dim\mathcal{A}_{b,\mathbb{D}}\leq d-1$ because $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive, and we conclude that $S=\mathcal{A}_{b,\mathbb{D}}$.
 Note here that $b:(x,y)\mapsto e(B(x,y))$ for the $\sigma$-sesquilinear form $B:(x,y)\mapsto x^{\sigma}\alpha y$.
 By point (c) of Proposition 5.4, we see that $b$ is nonisotropic.
-From here, we deduce from point (b) of Proposition 5.4 that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$: it follows that for every nondegenerate bilinear form $b^{\prime}$ on $V$ such that $\mathcal{S}=\mathcal{A}_{b^{\prime},\mathbb{D}}$, we have $b^{\prime}\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ and hence $b^{\prime}\in\mathbb{F}b$.
+From here, we deduce from point (b) of Proposition 5.4 that $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$: it follows that for every nondegenerate bilinear form $b^{\prime}$ on $V$ such that $S=\mathcal{A}_{b^{\prime},\mathbb{D}}$, we have $b^{\prime}\in\operatorname{\operatorname{Alt}}(S)$ and hence $b^{\prime}\in\mathbb{F}b$.
 This completes the proof of Theorem 1.12.
@@
 where $b^{\prime\prime}:(x,y)\mapsto e(B^{\prime\prime}(x,y))$ for the $\sigma$-sesquilinear form $B^{\prime\prime}:(x,y)\mapsto B(\varphi(x),\varphi(y))$.
-Hence $\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is the conjugate of $\mathcal{A}_{b,\mathbb{D}}$ under $\varphi^{-1}$, and more generally $\mathcal{A}_{\alpha b^{\prime\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is conjugated to $\mathcal{A}_{b,\mathbb{D}}$ for all $\alpha\in\mathbb{F}^{\times}$.
-Conversely, assume that $\mathcal{A}_{b,\mathbb{D}}$ is conjugated to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ under an element of $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$.
+Hence $\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is the conjugate of $\mathcal{A}_{b,\mathbb{D}}$ under $\varphi^{-1}$, and more generally $\mathcal{A}_{\alpha b^{\prime\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ is similar to $\mathcal{A}_{b,\mathbb{D}}$ for all $\alpha\in\mathbb{F}^{\times}$.
+Conversely, assume that $\mathcal{A}_{b,\mathbb{D}}$ is similar to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ under an element of $\operatorname{\operatorname{GL}}_{\mathbb{D}}(V)$.
 Then by the above $\mathcal{A}_{b^{\prime},\mathbb{D}}=\mathcal{A}_{b^{\prime\prime},\mathbb{D}}$ for some $b^{\prime\prime}$ of the form $(x,y)\mapsto e(B^{\prime\prime}(x,y))$, where $B^{\prime\prime}$ is equivalent to $B$.
@@
 This completes the proof of Theorem 1.13.
-7 Semiaffine spaces of matrices with rank bounded below
-In this short section, we apply the results on trivial spectrum subspaces to obtain significant ones on $\mathbb{F}$-affine subspaces of matrices with rank bounded below.
-This uses a strategy devised by Meshulam [10].
-7.1 Semiaffine spaces of matrices with rank bounded below
+7 Optimal affine spaces of units
+In this final section, we solve our initial problem of unital affine subspaces of a simple algebra.
+7.1 The greatest possible dimension
+To start with, we have a straightforward application of Theorem 1.8:
 Theorem 7.1.
-Let $n,p,r$ be integers with $\min(n,p)\geq r>0$.
-Let $\mathcal{S}$ be an $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ in which every element has rank at least $r$.
-Assume that $|\mathbb{F}|\geq rd$.
-Then
-$\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq r+d\,\dbinom{r}{2}$,
-where the codimension is to be understood as a codimension over $\mathbb{F}$. Moreover, this inequality is optimal, and equality holds only if $\mathcal{S}$ contains a rank $r$ matrix.
-The proof requires the following weak version of Theorem 1.6:
-Proposition 7.2.
-Let $r\geq 1$ be such that $|\mathbb{F}|\geq rd$, and let $n\geq r$.
-Let $S$ be an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
-Then $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq r+d\dbinom{r}{2}$, with equality attained only if $n=r$.
-For large values of $n$ with respect to $r$, this inequality can be considerably improved with limited effort, but we will not need such refinements.
-Proof of Proposition 7.2.
-We prove the result by induction on $n$.
-The case $n=r$ is a reformulation of Theorem 1.6.
-Assume now that $n>r$.
-Denote by $x$ the first vector of the standard basis of $\mathbb{D}^{n}$.
-Then $\dim_{\mathbb{F}}(Sx)\leq nd-1$ because $S$ is intransitive.
-Denote by $S^{\prime}$ the set of all matrices $N\in\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$ such that $\begin{bmatrix}0&[0]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&N\end{bmatrix}$ belongs to $S$, and note that $S^{\prime}$ is an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$.
-By the rank theorem
-$\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq\operatorname{\operatorname{codim}}_{\mathbb{D}^{n}}(Sx)+\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})}S^{\prime},$
-and by induction $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S^{\prime}\geq r+d\dbinom{r}{2}$.
-Hence $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n}(\mathbb{D})}S\geq 1+r+d\dbinom{r}{2}$, which completes the proof.
-∎
-We can now prove Theorem 7.1.
-Proof of Theorem 7.1.
-The optimality of the inequality is exemplified by taking the space of all matrices of the form
-$\begin{bmatrix}I_{r}+M&[?]_{r\times(p-r)}\\ [?]_{(n-r)\times r}&[?]_{(n-r)\times(p-r)}\end{bmatrix}$
-where $M$ ranges over an $\mathbb{F}$-trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{r}(\mathbb{F})$ with dimension $\alpha(r,d)$ (say, one that is described on page 1.1).
-It is clear that the codimension of this space in $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ is the one of $\mathcal{M}$ in $\operatorname{\operatorname{M}}_{r}(\mathbb{D})$, that is $r+d\,\dbinom{r}{2}$.
-Every matrix in that space has its upper-left $r\times r$ block invertible, and hence its rank is at least $r$.
-Now, denote by $s$ the least possible rank in $\mathcal{S}$ (at this point it may very well be that $s>r$).
-Replacing $\mathcal{S}$ with an equivalent space if necessary, we can then assume that $\mathcal{S}$ contains $J_{s}=\begin{bmatrix}I_{s}&[0]_{s\times(p-s)}\\ [0]_{(n-s)\times s}&[0]_{(n-s)\times(p-s)}\end{bmatrix}$.
-Denote by $\overrightarrow{\mathcal{S}}$ the translation vector space of $\mathcal{S}$, and write every matrix $M$ of it as $M=\begin{bmatrix}A(M)&C(M)\\ B(M)&D(M)\end{bmatrix}$ with $A(M)\in\operatorname{\operatorname{M}}_{s}(\mathbb{F})$ and so on.
-Consider the $\mathbb{F}$-linear subspace $W$ of $\overrightarrow{\mathcal{S}}$ consisting of all the matrices $M\in\overrightarrow{\mathcal{S}}$ such that $B(M)=0$, $C(M)=0$ and $D(M)=0$.
-Hence $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{s}(\mathbb{D})}A(W)$.
-For all $M\in W$, the matrix
-$J_{s}+M=\begin{bmatrix}I_{s}+A(M)&[0]_{s\times(p-s)}\\ [0]_{(n-s)\times s}&[0]_{(n-s)\times(p-s)}\end{bmatrix}$
-has rank at least $s$ because it belongs to $\mathcal{S}$, and hence $I_{s}+A(M)$ is invertible.
-It follows that $A(W)$ is an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{s}(\mathbb{D})$.
-Since $|\mathbb{F}|\geq rd\geq sd$, we deduce from Proposition 7.2 that $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{s}(\mathbb{D})}A(W)\geq r+d\dbinom{r}{2}$, with strict equality if $s>r$, and we conclude that $\operatorname{\operatorname{codim}}_{\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})}\mathcal{S}\geq r+d\dbinom{r}{2}$, with strict inequality if $s>r$.
-This proves all the claimed statements.
-∎
-Remark 7.1.
-If we had used Theorem 1.6 instead of Proposition 7.2, we would have needed the inequality $|\mathbb{F}|\geq d\max(n,p)$ because we have no prior control on the least possible rank in $\mathcal{S}$.
-Clearly, this requirement is not very satisfying, hence our insistence in requiring only that $|\mathbb{F}|\geq rd$.
-7.2 Optimal semiaffine spaces of nonsingular matrices
-At this point, it is natural that we wish to obtain a description of the optimal $\mathbb{F}$-affine subspaces of matrices with rank at least $r$, meaning the ones with the minimal possible codimension in $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$.
-Judging from the length of [16], which covers only the case $\mathbb{D}=\mathbb{F}$, it is safer to leave this problem for future work.
-Nevertheless, we will tackle here the limit case $n=p=r$ as it is can be obtained with few technical difficulties from our results on trivial spectrum linear subspaces.
-Let us start with some terminology and notation.
-Definition 7.1.
-Two subsets $\mathcal{X}$ and $\mathcal{Y}$ of $\operatorname{\operatorname{M}}_{n,p}(\mathbb{D})$ are called equivalent , and we write $\mathcal{X}\sim\mathcal{Y}$, when there exists a pair $(P,Q)\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})\times\operatorname{\operatorname{GL}}_{p}(\mathbb{D})$ such that $\mathcal{Y}=P\mathcal{X}Q$, i.e., when $\mathcal{X}$ and $\mathcal{Y}$ represent the same set of linear mappings from a $p$-dimensional $\mathbb{D}$-vector space to an $n$-dimensional $\mathbb{D}$-vector space, in a different choice of bases of these spaces.
+Let $\mathbb{D}$ be a division ring with finite dimension $d$ over a central subfield $\mathbb{F}$, and let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$.
+Then the greatest possible dimension for a unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is $n(d-1)+d\binom{n}{2}$.
+7.2 The optimal spaces
+For the structure of optimal spaces, we will split the discussion, whether $\mathbb{D}$ has quadratic type over $\mathbb{F}$ or not.
 Theorem 7.2.
@@
 Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$, and that $|\mathbb{F}|\geq nd$.
-Then, up to equivalence, there is exactly one optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
+Then, up to equivalence, there is exactly one optimal unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
 One such space is the $n$-fold joint $(1_{\mathbb{D}}+H)\vee\cdots\vee(1_{\mathbb{D}}+H)$, where $H$ is an arbitrary $\mathbb{F}$-linear hyperplane of $\mathbb{D}$ that does not contain $1_{\mathbb{D}}$.
@@
 Let $n\geq 1$. Assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$, with $|\mathbb{F}|\geq nd$. Let $(\sigma,e)$ be an associated pair.
-Let $\mathcal{M}$ be an optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
+Let $\mathcal{M}$ be an optimal unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
 Then there exists a partition $n=n_{1}+\cdots+n_{p}$ and a list $(P_{1},\dots,P_{p})$ of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$, such that
@@
 Conversely, let $n=n_{1}+\cdots+n_{p}$ be a partition, and $(P_{1},\dots,P_{p})$ be a list of $e$-nonisotropic matrices, with $P_{i}\in\operatorname{\operatorname{GL}}_{n_{i}}(\mathbb{D})$.
-Then $(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D}))\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal $\mathbb{F}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
+Then $(P_{1}+\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D}))\vee\cdots\vee(P_{p}+\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal unital $\mathbb{F}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
 Finally, let $m=m_{1}+\cdots+m_{q}$ be a partition, and $(P^{\prime}_{1},\dots,P^{\prime}_{q})$ be a list of $e$-nonisotropic matrices, with $P^{\prime}_{i}\in\operatorname{\operatorname{GL}}_{m_{i}}(\mathbb{D})$.
@@
 Lemma 7.4.
-Let $\mathcal{S}$ be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, where $V$ is an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$.
-Then $\mathbb{F}x+\mathcal{S}x=V$ for all $x\in V\smallsetminus\{0\}$.
+Let $S$ be an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, where $V$ is an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$.
+Then $\mathbb{F}x+Sx=V$ for all $x\in V\smallsetminus\{0\}$.
 Proof.
-We combine Theorems 1.8 and 1.10 to find that $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$ for some nondegenerate $\mathbb{F}$-bilinear form $b$ of type $(x,y)\mapsto e(B(x,y))$, where $B$ is a nondegenerate $\sigma$-sesquilinear form on $V$.
+We combine Theorems 1.10 and 1.12 to find that $S=\mathcal{A}_{b,\mathbb{D}}$ for some nondegenerate $\mathbb{F}$-bilinear form $b$ of type $(x,y)\mapsto e(B(x,y))$, where $B$ is a nondegenerate $\sigma$-sesquilinear form on $V$.
 Let $x\in V\smallsetminus\{0\}$.
-By point (b) of Proposition 5.4 we find $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$.
-The conclusion is obtained by noting that $x\not\in\mathcal{S}x$ because $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum.
+By point (b) of Proposition 5.4 we find $\dim_{\mathbb{F}}(Sx)=nd-1$.
+The conclusion is obtained by noting that $x\not\in Sx$ because $S$ has trivial spectrum.
 ∎
@@
 $\mathcal{M}\sim(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}}).$
-The second statement is obvious because, for every $k\in\mathopen{[\![}1,p\mathclose{]\!]}$ and every $e$-nonisotropic $P\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ we have $P+\mathcal{M}_{k}\sim I_{k}+P^{-1}\mathcal{M}_{k}$, and $P^{-1}\mathcal{M}_{k}$ has $\mathbb{F}$-trivial spectrum and dimension $\alpha(k,d)$ over $\mathbb{F}$; note also, for the remainder of the proof, that $\operatorname{\operatorname{span}}_{\mathbb{F}}(P+\mathcal{M}_{k})X=\mathbb{D}^{k}$ for all $X\in\mathbb{D}^{k}\smallsetminus\{0\}$, as a consequence of Lemma 7.5.
+The second statement is obvious because, for every $k\in\mathopen{[\![}1,p\mathclose{]\!]}$ and every $e$-nonisotropic $P\in\operatorname{\operatorname{GL}}_{k}(\mathbb{D})$ we have $P+\mathcal{M}_{k}\sim I_{k}+P^{-1}\mathcal{M}_{k}$, and $P^{-1}\mathcal{M}_{k}$ has trivial spectrum and dimension $\alpha(k,d)$ over $\mathbb{F}$; note also, for the remainder of the proof, that $\operatorname{\operatorname{span}}_{\mathbb{F}}(P+\mathcal{M}_{k})X=\mathbb{D}^{k}$ for all $X\in\mathbb{D}^{k}\smallsetminus\{0\}$, as a consequence of Lemma 7.4.
 It remains to prove the uniqueness statement.
 It is better to think in geometrical terms here.
-So, let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, and let $\mathcal{S}$ be an optimal $\mathbb{F}$-affine subspace of automorphisms of $V$.
-Assume that we have bases $(e_{1},\dots,e_{n})$ and $(f_{1},\dots,f_{n})$ of $V$ such that the matrices that represent the elements of $\mathcal{S}$ form the space $(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})$.
+So, let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, and let $S$ be an optimal $\mathbb{F}$-affine subspace of automorphisms of $V$.
+Assume that we have bases $(e_{1},\dots,e_{n})$ and $(f_{1},\dots,f_{n})$ of $V$ such that the matrices that represent the elements of $S$ form the space $(P_{1}+\mathcal{M}_{n_{1}})\vee\cdots\vee(P_{p}+\mathcal{M}_{n_{p}})$.
 We consider the integers $N_{i}:=\sum_{k=1}^{i}n_{k}$ and the spaces $U_{i}:=\operatorname{\operatorname{span}}(e_{1},\dots,e_{N_{i}})$ and $V_{i}:=\operatorname{\operatorname{span}}(f_{1},\dots,f_{N_{i}})$ for $i\in\mathopen{[\![}0,p\mathclose{]\!]}$.
 We shall prove that any other choice of bases will yield the same spaces.
-To see this, we look at the dimension $\dim_{\mathbb{F}}(\mathcal{S}x)$ when $x$ ranges over $V$.
-Using the previous remark, we see that for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ and all $x\in U_{i}\smallsetminus U_{i-1}$, we have $\operatorname{\operatorname{span}}_{\mathbb{F}}(\mathcal{S}x)=V_{i}$.
-Hence the spaces $V_{0},\dots,V_{p}$ are uniquely determined by $\mathcal{S}$ only, as the possible $\mathcal{S}x$ spaces when $x$ ranges over $V$; and then $U_{i}=\{x\in V:\;\mathcal{S}x\subseteq V_{i}\}$ is also uniquely determined by $\mathcal{S}$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$.
+To see this, we look at the dimension $\dim_{\mathbb{F}}(Sx)$ when $x$ ranges over $V$.
+Using the previous remark, we see that for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ and all $x\in U_{i}\smallsetminus U_{i-1}$, we have $\operatorname{\operatorname{span}}_{\mathbb{F}}(Sx)=V_{i}$.
+Hence the spaces $V_{0},\dots,V_{p}$ are uniquely determined by $S$ only, as the possible $Sx$ spaces when $x$ ranges over $V$; and then $U_{i}=\{x\in V:\;Sx\subseteq V_{i}\}$ is also uniquely determined by $S$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$.
 As a consequence, if we have invertible matrices $R,S$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that
@@
 This completes our proof. ∎
+7.3 Making sense of the classification of optimal spaces
 It remains to understand the nature of the mappings $X\in\mathbb{D}^{n}\mapsto e(X^{\star}PX)$ when $n\geq 1$ and $P$ ranges over the $e$-nonisotropic matrices of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$.
-In the special case $d=1$, they are simply the nonisotropic quadratic forms on $\mathbb{D}^{n}$.
-Hence, as seen in [18], the optimal affine subspaces of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are classified up to equivalence by lists of congruence classes of nonisotropic quadratic forms.
-In particular, if $\mathbb{F}=\mathbb{R}$ then there is only one such congruence class of $k$-dimensional forms for each $k\geq 1$, and we obtain that every optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ is equivalent to $(I_{n_{1}}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{R}))\vee\cdots\vee(I_{n_{p}}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{R}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely every space of this form is an optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$.
+In the special case $d=1$, they are simply the nonisotropic quadratic forms on $\mathbb{F}^{n}$.
+Hence, as seen in [12], the optimal unital $\mathbb{F}$-affine subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are classified, up to equivalence, by lists of congruence classes of nonisotropic quadratic forms.
+In particular, if $\mathbb{F}=\mathbb{R}$ then there is only one such congruence class of $k$-dimensional nonisotropic quadratic forms for each $k\geq 1$, and we obtain that every optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ is equivalent to $(I_{n_{1}}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{R}))\vee\cdots\vee(I_{n_{p}}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{R}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely every space of this form is an optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$.
 The situation is not as clear if $d>1$.
@@
 Hence the equivalence class of $P+\mathcal{M}_{n}$ determines $Q$ up to star-congruence and multiplication with an element of $\mathbb{F}^{\times}$.
-As a consequence, optimal affine subspaces of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are classified, up to equivalence, by lists $([Q_{1}],\dots,[Q_{p}])$ of classes of nonisotropic Hermitian matrices over $\mathbb{D}$, where each class is for the equivalence relation where $Q$ and $Q^{\prime}$ are equivalent if and only if $Q^{\prime}=\alpha R^{\star}QR$ for some invertible $R$ and some $\alpha\in\mathbb{F}^{\times}$.
-In particular, if $\mathbb{D}\in\{\mathbb{C},\mathbb{H}\}$ and $\mathbb{F}=\mathbb{R}$, then we recover from the classification of Hermitian forms that every optimal $\mathbb{R}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is equivalent to $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal $\mathbb{R}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for every partition $n=n_{1}+\cdots+n_{p}$.
+As a consequence, optimal unital $\mathbb{F}$-affine subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are classified, up to equivalence, by lists $([Q_{1}],\dots,[Q_{p}])$ of classes of nonisotropic Hermitian matrices over $\mathbb{D}$, under the relation of homo-star-congruence.
+In particular, if $\mathbb{D}\in\{\mathbb{C},\mathbb{H}\}$ and $\mathbb{F}=\mathbb{R}$, then we recover from the classification of Hermitian forms that every optimal unital $\mathbb{R}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is equivalent to $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal unital $\mathbb{R}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for every partition $n=n_{1}+\cdots+n_{p}$.
 For fields of characteristic $2$, Hermitian matrices are skew-Hermitian and hence there is no obvious way to simplify our vision on the classification of the forms $X\mapsto e(X^{\star}PX)$.
 We leave the problem for future work.
-8 Linear subspaces of $\mathbb{F}$-diagonalisable matrices
-In this final section, we give a series of applications of the structure theorem for optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
-8.1 $\mathbb{F}$-diagonalisable subspaces
-Let $V$ be a (finite-dimensional) vector space over $\mathbb{D}$.
-An endomorphism of the $\mathbb{D}$-vector space $V$ is called $\mathbb{F}$-diagonalisable when it is representable over $\mathbb{D}$ by a diagonal matrix whose diagonal entries all belong to $\mathbb{F}$, or equivalently when it is annihilated by a polynomial $p(t)\in\mathbb{F}[t]$ that is split and with simple roots.
-An $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is called $\mathbb{F}$-diagonalisable when all its elements are $\mathbb{F}$-diagonalisable.
-In some sense, being $\mathbb{F}$-diagonalisable is sort of the opposite of having $\mathbb{F}$-trivial spectrum: if an endomorphism $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is $\mathbb{F}$-diagonalisable and has $\mathbb{F}$-trivial spectrum, then it must be zero (because in a diagonalisation, the diagonal entries must equal zero).
-We define $\mathbb{F}$-diagonalisable matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ and $\mathbb{F}$-diagonalisable linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ accordingly.
-Now we introduce three large spaces that are known to have this property.
-The space $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ of all real symmetric matrices; the space $\mathcal{H}_{n}(\mathbb{C})$ of all Hermitian matrices over the complex numbers; and finally the space $\mathcal{H}_{n}(\mathbb{H})$ of all quaternionic Hermitian matrices (see e.g.
-[5] for a proof that the latter is $\mathbb{R}$-diagonalisable).
-More generally, assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ and that $\mathbb{D}$ is not a hyper-radicial extension of $\mathbb{F}$.
-Consider the associated pair $(\sigma,e)$ (so here $e:=\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$) and set
-$\mathcal{H}_{n}(\mathbb{D}):=\{M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D}):\;M^{\star}=M\},$
-where $M^{\star}$ is defined with respect to $\sigma$.
-Then $\mathcal{H}_{n}(\mathbb{D})$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ and its dimension is $n+d\binom{n}{2}$.
-However, it might not be $\mathbb{F}$-diagonalisable (see the case of the symmetric matrices over the complex numbers).
-Nevertheless, we easily have the following result:
-Theorem 8.1.
-Let $\mathcal{V}$ be an $\mathbb{F}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then
-$\dim\mathcal{V}\leq n+d\binom{n}{2}$.
-Proof.
-We know that $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ contains an $\mathbb{F}$-trivial spectrum linear subspace $\mathcal{S}$ with dimension $\alpha(n,d)$.
-By a previous remark we must have $\mathcal{V}\cap\mathcal{S}=\{0\}$ and hence
-$\dim\mathcal{V}\leq\dim_{\mathbb{F}}\operatorname{\operatorname{M}}_{n}(\mathbb{D})-\dim\mathcal{S}=n+d\binom{n}{2}$.
-In most cases, the inequality from Theorem 8.1 is very far from optimality.
-Indeed if $\mathbb{D}=\mathbb{F}=\mathbb{C}$ it is known by the Motzkin-Taussky theorem [11] that the greatest possible dimension for a $\mathbb{C}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ is $n$ (every such subspace is conjugated to the space of all diagonal matrices), and strangely this holds for finite fields as well.
-It would be amazing if the latter result has not been found earlier, but we were unable to find it in the literature on the topic, so we include a quick proof.
-Theorem 8.2 (Motzkin-Taussky theorem for finite fields).
-Assume that $\mathbb{F}$ is finite, let $V$ be a finite-dimensional $\mathbb{F}$-vector space, and $u$ and $v$ be endomorphisms of $V$.
-Assume that every linear combination of $u$ and $v$ with coefficients in $\mathbb{F}$ is diagonalisable.
-Then $u$ and $v$ commute.
-Proof.
-Denote by $q$ the cardinality of $\mathbb{F}$.
-An endomorphism of $V$ is diagonalisable if and only if it is annihilated by the polynomial $P(t):=\prod_{\lambda\in\mathbb{F}}(t-\lambda)=t^{q}-t$.
-Hence $(\lambda u+v)^{q}=\lambda u+v$ for all $\lambda\in\mathbb{F}$.
-Yet there is a list $(a_{1},\dots,a_{p-1})$ of endomorphisms of $V$ such that $(\lambda u+v)^{q}=(\lambda u)^{q}+v^{q}+\sum_{k=1}^{q-1}\lambda^{k}a_{k}$ for all $\lambda\in\mathbb{F}$, where in particular $a_{1}=\sum_{k=0}^{q-1}v^{k}uv^{q-1-k}$.
-Note that $\forall\lambda\in\mathbb{F},\;(\lambda u)^{q}=\lambda u^{q}$.
-Hence $\sum_{k=1}^{q-1}\lambda^{k}a_{k}=0$ for all $\lambda\in\mathbb{F}$.
-Since $|\mathbb{F}|=q$ all the $a_{k}$’s vanish and in particular $a_{1}=0$.
-As a consequence, $\operatorname{\operatorname{Ker}}v$ is invariant under $u$: indeed for all $x\in\operatorname{\operatorname{Ker}}v$ we obtain $0=a_{1}(x)=v^{q-1}(u(x))$, and since $v$ is diagonalisable we classically have $\operatorname{\operatorname{Ker}}v^{q-1}=\operatorname{\operatorname{Ker}}v$, whence $v(u(x))=0$.
-Now, we note that the assumptions are preserved in replacing $v$ with $v-\alpha\operatorname{\operatorname{id}}_{V}$ for an arbitrary $\alpha\in\mathbb{F}$, so the previous step shows that $u$ leaves all the eigenspaces of $v$ invariant, to the effect that $uv$ and $vu$ coincide on all these eigenspaces.
-Since $v$ is diagonalisable, it follows that $uv=vu$.
-∎
-Of course, in case $d=1$ the bound $n$ that is obtained for the field of complex numbers (and more generally for any algebraically closed fields of characteristic 0) and for finite fields is very far from the bound $\dbinom{n+1}{2}$ given in Theorem 8.1.
-Here, we will consider the $\mathbb{F}$-diagonalisable subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ that have dimension $n+d\dbinom{n}{2}$.
-The case $n=1$ is trivial, so we only consider greater values of $n$.
-8.2 Main results
-We recall that the special case $d=1$ is already known (see [3, 13, 23] for various proofs):
-Theorem 8.3.
-Let $n\geq 2$.
-Let $\mathcal{S}$ be an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ such that $\dim\mathcal{S}=n+\binom{n}{2}$.
-Then $\mathbb{F}$ is formally real and $\mathcal{S}$ is similar to $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$.
-To state the generalization, we need an additional notion:
-Definition 8.1.
-Assume that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and denote by $x\mapsto x^{\star}$ its standard involution over $\mathbb{F}$.
-We say that $x\in\mathbb{D}\mapsto xx^{\star}$, which is valued in $\mathbb{F}$, is the standard norm of $\mathbb{D}$ over $\mathbb{F}$.
-Here is the generalization we are aiming at:
-Theorem 8.4.
-Let $n\geq 2$.
-Let $\mathcal{S}$ be an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ such that $\dim\mathcal{S}=n+d\binom{n}{2}$.
-Then $\mathbb{F}$ is formally real and Pythagorean [^6] The field $\mathbb{F}$ is Pythagorean when $\forall(x,y)\in\mathbb{F}^{2},\;\exists z\in\mathbb{F}:x^{2}+y^{2}=z^{2}$.
-(so it has characteristic 0), $\mathbb{D}$ has quadratic type over $\mathbb{F}$, its standard norm is square-valued in $\mathbb{F}$, and $\mathcal{S}$ is similar to $\mathcal{H}_{n}(\mathbb{D})$.
-Hence, if there exists an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ such that $\dim\mathcal{S}=n+d\binom{n}{2}$, then $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and every Hermitian matrix of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable!
-Finally, we note that there is a known characterization of the fields $\mathbb{F}$ over which every symmetric matrix is diagonalisable [28] but there is no real hope of understanding, for a fixed $n\geq 3$, the fields $\mathbb{F}$ over which every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is diagonalisable.
-Nevertheless, if $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and characteristic 0, and if for some $n\geq 2$ every matrix of $\mathcal{H}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable, then so is every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$.
-Conversely, assume that, for some $n\geq 2$, every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is diagonalisable over $\mathbb{F}$, and that $\mathbb{D}$ has quadratic type over $\mathbb{F}$ and its standard norm is square-valued in $\mathbb{F}$.
-Then $\mathbb{F}$ is Pythagorean and formally real, and it is then easy to prove, by using the Gram-Schmidt orthonormalization process, that every matrix of $\mathcal{H}_{n}(\mathbb{D})$ is unitarily similar to a matrix of $\mathcal{H}_{n}(\mathbb{D})$ with entries in $\mathbb{F}$, i.e., to a matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$.
-Hence, $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is $\mathbb{F}$-diagonalisable if and only if $\mathcal{H}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable.
-As a consequence, we obtain:
-Corollary 8.5.
-Let $n\geq 2$. The following conditions are equivalent:
-  1. (i)
-    There exists an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\binom{n}{2}$.
-  2. (ii)
-    Every matrix of $\operatorname{\operatorname{S}}_{n}(\mathbb{F})$ is diagonalisable, $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ and its standard norm is square-valued in $\mathbb{F}$.
-8.3 The orthogonality method
-The strategy of proof of Theorem 8.4 is essentially similar to the one we used in [23]: We will make a direct connection between $\mathbb{F}$-diagonalisable subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\binom{n}{2}$ and $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
-This connection is obtained thanks to a trace-orthogonality argument.
-To obtain it, we consider the traditional trace $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$, defined as in Section 2.1.
-The trace of an element $a\in\mathbb{D}$ is the trace of the endomorphism $x\mapsto ax$ of the $\mathbb{F}$-vector space $\mathbb{D}$.
-It satisfies $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ab)=\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}(ba)$ for all $(a,b)\in\mathbb{D}^{2}$.
-We will simply write $\operatorname{\operatorname{Tr}}$ for $\operatorname{\operatorname{Tr}}_{\mathbb{D}/\mathbb{F}}$ in what follows.
-One major difficulty is that $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}$, and in case $\operatorname{\operatorname{char}}(\mathbb{F})>0$ we might then have $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})=0$, and it might even hold that $\operatorname{\operatorname{Tr}}$ is the zero form.
-In the remainder of this section, we assume that $\operatorname{\operatorname{char}}(\mathbb{F})$ does not divide $d$ (early on in our proof of Theorem 8.4, we will prove that $\operatorname{\operatorname{char}}(\mathbb{F})=0$ under specific assumptions, which will alleviate this difficulty).
-The trace extends to square matrices because
-$\forall(A,B)\in\operatorname{\operatorname{M}}_{n}(\mathbb{D})^{2},\;\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(AB))=\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(BA)),$
-where $\operatorname{\operatorname{tr}}$ stands for the trace of matrices in $\mathbb{D}$ (i.e., the sum of the diagonal entries). Then
-$\langle A,B\rangle:=\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(AB))$
-defines a symmetric $\mathbb{F}$-bilinear form on $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and it satisfies the key identity
-$\langle PAP^{-1},PBP^{-1}\rangle=\langle A,B\rangle.$
-Thanks to $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})\neq 0$, this $\mathbb{F}$-bilinear form is nondegenerate.
-This way, we can define, on any $\mathbb{D}$-vector space $V$ of finite dimension, the $\mathbb{F}$-inner product of two $\mathbb{D}$-endomorphisms $u$ and $v$ of $V$, so that any matrix representation of $u$ and $v$ in a basis yields the same result.
-Now, given an $\mathbb{F}$-linear subspace $\mathcal{V}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ that satisfies the conditions of Theorem 8.4, we shall consider the orthogonal complement $\mathcal{V}^{\bot}$ of $\mathcal{V}$ for the previous inner product: this gives an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $dn^{2}-n-d\binom{n}{2}=\alpha(n,d)$.
-And we will prove that $\mathcal{V}^{\bot}$ has $\mathbb{F}$-trivial spectrum!
-There are two problems with this strategy: the first one is the assumption that $\operatorname{\operatorname{char}}(\mathbb{F})$ does not divide $d$, and the second one is the cardinality assumptions in the theorems that describe the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace and the structure of the optimal spaces.
-Fortunately, we will prove early on that $\operatorname{\operatorname{char}}(\mathbb{F})=0$, and in particular that $\mathbb{F}$ must be infinite (this could of course be immediately derived from Theorem 8.2).
-Hence, no major difficulty will be encountered.
-To obtain these results however, we must use an induction on $n$, where the case $n=2$ is critical.
-Before we go on, we prove a critical result that relates the properties of a space $\mathcal{V}$ to the ones of its orthogonal complement $\mathcal{V}^{\bot}$ when $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})\neq 0$.
-Then we will deal with the case $n=2$, and finally we will be able to prove Theorem 8.4 by using our results on trivial spectrum spaces.
-8.4 Connecting trivial spectrality with a property on rank $1$ idempotents
-Definition 8.2.
-Let $V$ be a (finite-dimensional) $\mathbb{D}$-vector space.
-An $\mathbb{F}$-linear subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ is said to have the full rank $1$ idempotent property whenever for every $1$-dimensional $\mathbb{D}$-linear subspace $E$ of $V$, the space $\mathcal{S}$ contains an idempotent endomorphism with range $E$.
-Lemma 8.6.
-Assume that $\operatorname{\operatorname{char}}(\mathbb{F})$ does not divide $d$.
-Let $V$ be a (finite-dimensional) $\mathbb{D}$-vector space.
-Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that has the full rank $1$ idempotent property.
-Then $\mathcal{S}^{\bot}$ has $\mathbb{F}$-trivial spectrum.
-It can be shown that the converse holds true if $\mathbb{D}=\mathbb{F}$, but we will not need this result here.
-Proof.
-Let $u\in\mathcal{S}^{\bot}$. Assume that there is a nonzero vector $x\in V$ such that $u(x)=x$.
-By assumption $\mathcal{S}$ contains an idempotent endomorphism $p$ with range $x\mathbb{D}$, and we have $\langle p,u\rangle=0$.
-For clarity, we extend $x$ to a basis $(x,e_{2},\dots,e_{n})$ of $V$, and we write the corresponding matrices of $u$ and $p$ as
-$\begin{bmatrix}1_{\mathbb{D}}&[?]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&[?]_{(n-1)\times(n-1)}\end{bmatrix}\quad\text{and}\quad\begin{bmatrix}1_{\mathbb{D}}&[?]_{1\times(n-1)}\\ [0]_{(n-1)\times 1}&[0]_{(n-1)\times(n-1)}\end{bmatrix}$.
-Then $\operatorname{\operatorname{tr}}(up)=\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})=d\,1_{\mathbb{F}}\neq 0$, which contradicts our assumptions.
-Hence $\mathcal{S}^{\bot}$ has $\mathbb{F}$-trivial spectrum.
-∎
-8.5 The case $n=2$
-Lemma 8.7.
-Assume that $\operatorname{\operatorname{M}}_{2}(\mathbb{D})$ contains an $\mathbb{F}$-diagonalisable subspace $\mathcal{V}$ with dimension at least $d+2$.
-Then $\mathbb{F}$ is formally real and Pythagorean.
-Due to the potential non-commutativity of $\mathbb{D}$, the proof is surprisingly difficult because we are not allowed to use the trace and determinant to analyse $\mathbb{F}$-diagonalisable matrices.
-Proof.
-Without loss of generality, we assume that $I_{2}\in\mathcal{V}$ (as we can always replace $\mathcal{V}$ with $\mathcal{V}+\mathbb{F}I_{2}$).
-We shall construct a linear form $t:\mathcal{V}\rightarrow\mathbb{F}$ and a quadratic form $q:\mathcal{V}\rightarrow\mathbb{F}$ such that
-$\forall a\in\mathcal{V},\;a^{2}=t(a)\,a-q(a)\,I_{2}.$
-If $\mathbb{D}$ is commutative it suffices to take for $t$ the trace, and for $q$ the determinant.
-Assume now that $\mathbb{D}$ is noncommutative. Hence $\mathbb{F}$ is infinite by Wedderburn’s theorem. This way, we can use arithmetics on polynomial functions.
-To start with, we take a linear hyperplane $H$ of $\mathcal{V}$ that does not contain $I_{2}$.
-For every $a\in H\smallsetminus\{0\}$, since $\mathcal{V}$ is $\mathbb{F}$-diagonalisable there are scalars $t(a)$ and $q(a)$ in $\mathbb{F}$ such that
-$a^{2}=t(a)\,a-q(a)\,I_{2},$ (6)
-and these scalars are uniquely determined by $a$. We also put $t(0)=0$ and $q(0)=0$. We shall see that $t(a)$ and $q(a)$ are homogeneous polynomial functions of $a$.
-Clearly, as $I_{2}$ and $a$ are $\mathbb{F}$-linearly independent except for $a=0$, $t$ and $q$ are homogeneous rational functions on $H\smallsetminus\{0\}$ with respective degrees $1$ and $2$.
-We claim that they are actually polynomials.
-Split indeed $t=-\frac{r}{s}$ for mutually prime homogeneous polynomials $r$ and $s$, and assume that $s$ is nonconstant.
-Let $\varphi$ be an arbitrary nonzero $\mathbb{F}$-linear form on $\mathcal{V}$ that vanishes at $I_{2}$.
-Then we apply $\varphi$ to identity (6) to obtain $\forall a\in H,\;t(a)\,\varphi(a)=\varphi(a^{2})$.
-Then $a\mapsto s(a)\,\varphi(a^{2})$ divides $r\varphi$, and hence $s$ divides $\varphi$.
-Because $\varphi$ has degree $1$ we deduce that $s=\lambda\varphi$ for some $\lambda\in\mathbb{F}^{\times}$.
-Hence, every $\mathbb{F}$-linear form on $\mathcal{V}$ that vanishes at $I_{2}$ must have its restriction to $H$ linearly dependent on $s$, to the effect that the space of all such linear forms has dimension (at most) $1$.
-Yet by orthogonality theory this space has dimension $\dim\mathcal{V}-1$, which is greater than $1$ because $\dim\mathcal{V}\geq d+2$.
-This contradiction shows that $s$ is constant.
-Hence $t$ is actually a homogeneous polynomial function of degree $1$, and it follows that $q$ is a quadratic form on $H$.
-We can then extend $t$ and $q$ to $\mathcal{V}$ by putting $t(a+\lambda I_{2}):=t(a)+2\lambda$ and $q(a+\lambda I_{2}):=q(a)+\lambda^{2}+\lambda t(a)$ for all $a\in H$ and all $\lambda\in\mathbb{F}$, and now we have the formula
-$\forall a\in\mathcal{V},\;a^{2}=t(a)\,a-q(a)\,I_{2}.$
-Clearly $t$ is a linear form on $\mathcal{V}$, and $q$ is a quadratic form on $\mathcal{V}$.
-Now, we return to the general case (we do not assume anymore that $\mathbb{D}$ is noncommutative).
-Let us take $W:=\operatorname{\operatorname{Ker}}t$, which has dimension at least $d+1\geq 2$.
-Let $a\in W$.
-Then $a^{2}=-q(a)I_{2}$.
-Since $a$ is similar to a diagonal matrix with entries in $\mathbb{F}$, we deduce that q(a)-q(a) is a square in $\mathbb{F}$, and $q(a)=0$ only if $a=0$.
-Hence q-q is a nonisotropic quadratic form on the $\mathbb{F}$-vector space $W$ whose values are all squares in $\mathbb{F}$.
-If either $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$ or $\dim W\geq 3$ then we can take $\mathbb{F}$-linearly independent vectors $a$ and $b$ in $W$ that are orthogonal for the polar form of $q$, and we deduce that $\mathbb{F}$ is Pythagorean and formally real.
-Otherwise $\operatorname{\operatorname{char}}(\mathbb{F})=2$ and $d=1$, but then the only trace zero matrices of $\operatorname{\operatorname{M}}_{2}(\mathbb{D})$ that are diagonalisable are the scalar multiples of $I_{2}$, which would lead to $\dim\mathcal{V}\leq 2$ and would thereby contradict our assumptions.
-∎
-8.6 The proof
-Here is the key step in our proof.
-Lemma 8.8.
-Let $n\geq 2$. Let $\mathcal{V}$ be an $\mathbb{F}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\dbinom{n}{2}$. Then:
-  1. (a)
-    The space $\mathcal{V}$ has the full rank $1$ idempotent property.
-  2. (b)
-    There exists an $\mathbb{F}$-diagonalisable linear subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$ with dimension $(n-1)+d\dbinom{n-1}{2}$.
-Proof.
-Let $E$ be a $1$-dimensional $\mathbb{D}$-linear subspace of $\mathbb{D}^{n}$.
-Conjugating $\mathcal{V}$, we lose no generality in assuming that $E=\mathbb{D}\times\{0_{n-1}\}$.
-For $M\in\mathcal{V}$, write
-$M=\begin{bmatrix}a(M)&R(M)\\ C(M)&K(M)\end{bmatrix}\quad\text{where $a(M)\in\mathbb{D}$, $C(M)\in\mathbb{D}^{n-1}$ and so on.}$
-Set
-$\mathcal{W}:=\{M\in\mathcal{V}:C(M)=0\}.$
-Note that
-$\dim\mathcal{W}\geq\dim\mathcal{V}-(n-1)d.$
-For every $M\in\mathcal{W}$, the scalar $a(M)$ must belong to $\mathbb{F}$ and the matrix $K(M)$ must be $\mathbb{F}$-diagonalisable (use, e.g., the characterization in terms of annihilating polynomials with coefficients in $\mathbb{F}$).
-Hence $K(\mathcal{W})$ is an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n-1}(\mathbb{D})$, and $a(\mathcal{W})\subseteq\mathbb{F}$.
-By the former
-$\dim K(\mathcal{W})\leq(n-1)+d\binom{n-1}{2}$.
-Moreover, for all $M\in\mathcal{W}$, if $K(M)=0$ and $a(M)=0$ then $M$ is nilpotent and hence $R(M)=0$. It follows that the $\mathbb{F}$-linear mapping
-$\varphi:M\in\mathcal{W}\mapsto(a(M),K(M))\in\mathbb{F}\times K(\mathcal{W})$
-is injective, leading to
-$\dim\mathcal{W}\leq 1+\dim K(\mathcal{W})\leq n+d\dbinom{n-1}{2}$.
-Yet $\dim\mathcal{V}-(n-1)d=n+d\dbinom{n-1}{2}$. Hence $\dim\mathcal{W}=n+d\dbinom{n-1}{2}$ and we deduce:
-  • That $\dim K(\mathcal{W})=(n-1)+d\dbinom{n-1}{2}$, which yields point (b);
-  • That $\varphi$ is onto.
-By the latter, we can find an $M\in\mathcal{W}$ such that $a(M)=1$ and $K(M)=0$, and then obviously $M\in\mathcal{V}$ is idempotent with range $E$. This yields point (a). ∎
-As a consequence, by induction we deduce from point (b) and from Lemma 8.7:
-Corollary 8.9.
-Let $n\geq 2$.
-If there exists an $\mathbb{F}$-diagonalisable subspace $\mathcal{V}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\dbinom{n}{2}$, then $\mathbb{F}$ is formally real and consequently $\operatorname{\operatorname{char}}(\mathbb{F})=0$.
-Now, we can finish the proof.
-So, let $n\geq 2$ and $\mathcal{V}$ be an $\mathbb{F}$-diagonalisable subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n+d\dbinom{n}{2}$.
-By the previous corollary, we are on the safe side: the field $\mathbb{F}$ is infinite, and the trace map $\operatorname{\operatorname{Tr}}$ satisfies $\operatorname{\operatorname{Tr}}(1_{\mathbb{D}})\neq 0$.
-Since $\mathbb{F}I_{n}+\mathcal{V}$ remains $\mathbb{F}$-diagonalisable, we deduce from Theorem 8.1 that $\mathcal{V}=\mathbb{F}I_{n}+\mathcal{V}$ and hence $I_{n}\in\mathcal{V}$.
-Now, we consider the standard inner product
-$\langle A,B\rangle:=\operatorname{\operatorname{Tr}}(\operatorname{\operatorname{tr}}(AB))$
-on $\mathcal{V}$ and the orthogonal complement $\mathcal{V}^{\bot}$ of $\mathcal{V}$, which is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ with dimension $n(d-1)+d\dbinom{n}{2}$.
-By Lemma 8.8, $\mathcal{V}$ has the full rank $1$ idempotent property, and hence Lemma 8.6 yields that $\mathcal{V}^{\bot}$ has $\mathbb{F}$-trivial spectrum.
-And finally, $\mathcal{V}^{\bot}$ is optimal as an $\mathbb{F}$-trivial spectrum subspace, because $\mathbb{F}$ is infinite.
-Here is our next step:
-Claim 8.1.
-The space $\mathcal{V}^{\bot}$ is irreducible.
-Proof.
-Assume on the contrary that $\mathcal{V}^{\bot}$ is reducible.
-Because the inner product is invariant under conjugation, we can assume without loss of generality that $\mathbb{D}^{p}\times\{0_{n-p}\}$ is invariant under $\mathcal{V}^{\bot}$ for some $p\in\mathopen{[\![}1,n-1\mathclose{]\!]}$.
-Then clearly the unit matrix $E_{1,n}$ (with zero entries everywhere except on the first row and $n$-th column, where the entry equals $1$) belongs to $(\mathcal{V}^{\bot})^{\bot}=\mathcal{V}$.
-Yet $E_{1,n}$ is nonzero and nilpotent, so it is not $\mathbb{F}$-diagonalisable.
-This is a contradiction.
-Hence $\mathcal{V}^{\bot}$ is irreducible.
-∎
-Now, we can use Theorem 1.8.
-Since $n\geq 2$, the division ring $\mathbb{D}$ must be of quadratic type over $\mathbb{F}$.
-Since $\operatorname{\operatorname{char}}(\mathbb{F})=0$, the hyper-radicial case is discarded, so there are only three cases: either $d=1$, or $d=2$ and $\mathbb{D}$ is a (separable) quadratic extension of $\mathbb{F}$, or $\mathbb{D}$ is a quaternion algebra over $\mathbb{F}$ (in which case $\operatorname{\operatorname{Tr}}$ equals twice the quaternionic trace).
-Denote by $\sigma:x\mapsto x^{\star}$ the standard involution of $\mathbb{D}$ over $\mathbb{F}$, and recall that $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})=\{M\in\operatorname{\operatorname{M}}_{n}(\mathbb{D}):M^{\star}=-M\}$ because $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$.
-By Theorem 1.8 there exists a matrix $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that
-$\mathcal{V}^{\bot}=P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$
-(here we do not care that $P$ should be $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$-nonisotropic).
-As obviously $\mathcal{H}_{n}(\mathbb{D})$ is orthogonal to $\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ for the inner product, whereas $\dim_{\mathbb{F}}\mathcal{H}_{n}(\mathbb{D})+\dim_{\mathbb{F}}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})=dn^{2}$, we find
-$\mathcal{V}=\mathcal{H}_{n}(\mathbb{D})P.$
-The conclusion is near.
-Remembering that $I_{n}\in\mathcal{V}$, we get $P^{-1}\in\mathcal{H}_{n}(\mathbb{D})$, and hence $P\in\mathcal{H}_{n}(\mathbb{D})$.
-Now, note that we can safely replace $P$ with any star-congruent matrix: indeed if we take $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ then
-$Q^{-1}\mathcal{V}Q=Q^{-1}\mathcal{H}_{n}(\mathbb{D})PQ=Q^{-1}\mathcal{H}_{n}(\mathbb{D})(Q^{-1})^{\star}Q^{\star}PQ=\mathcal{H}_{n}(\mathbb{D})(Q^{\star}PQ).$
-In particular, by reduction theory for Hermitian forms (recall that $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$) we can now replace $P$ with a diagonal matrix $\Delta$ with diagonal entries $\delta_{1},\dots,\delta_{n}$ in $\mathbb{F}^{\times}$.
-Hence $\mathcal{H}_{n}(\mathbb{D})\Delta$ is $\mathbb{F}$-diagonalisable.
-Let $i\in\mathopen{[\![}2,n\mathclose{]\!]}$.
-The matrix $M:=E_{1,i}+E_{i,1}$ belongs to $\mathcal{H}_{n}(\mathbb{D})$, so $M\Delta=\delta_{i}E_{1,i}+\delta_{1}E_{i,1}$ must be $\mathbb{F}$-diagonalisable, and since its entries are in $\mathbb{F}$ we can safely use its characteristic polynomial $t^{n-2}(t^{2}-\delta_{1}\delta_{i})$.
-It follows that $\delta_{1}\delta_{i}$ is a nonzero square in $\mathbb{F}$, and hence $\delta_{i}=(\delta_{i})^{2}\,(\delta_{1}\delta_{i})^{-1}\delta_{1}$ is the product of $\delta_{1}$ with a nonzero square in $\mathbb{F}$.
-We conclude that $\Delta$ is star-congruent to $\delta_{1}I_{n}$, and hence $\mathcal{V}$ is conjugated to $\mathcal{H}_{n}(\mathbb{D})(\delta_{1}I_{n})=\mathcal{H}_{n}(\mathbb{D})$.
-In particular $\mathcal{H}_{n}(\mathbb{D})$ is $\mathbb{F}$-diagonalisable.
-Finally, let $x\in\mathbb{D}$.
-Then the matrix $xE_{1,2}+x^{\star}E_{2,1}$ must be diagonalisable over $\mathbb{F}$.
-The square of this matrix is diagonal with diagonal entries $xx^{\star},x^{\star}x,0,\dots,0$, and hence all these values must be squares of elements of $\mathbb{F}$.
-In particular $xx^{\star}=\alpha^{2}$ for some $\alpha\in\mathbb{F}$.
-This completes the proof of Theorem 8.4.
-8.7 Real subspaces of complex-diagonalisable matrices
-Let us give another application of our results on trivial spectrum subspaces.
-This time around, we consider the problem of $\mathbb{R}$-linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ in which all the elements are diagonalisable over $\mathbb{C}$.
-An example of such a subspace if $\mathbb{R}I_{n}\oplus i\mathcal{H}_{n}(\mathbb{C})$, the space of all sums of a real scalar multiple of the identity matrix with a skew-Hermitian matrix.
-As will turn out, this space has the greatest possible dimension for the problem under consideration.
-Theorem 8.10.
-Let $\mathcal{S}$ be an $\mathbb{R}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ in which all the matrices are diagonalisable over $\mathbb{C}$.
-Then $\dim\mathcal{S}\leq 1+n^{2}$.
-Moreover, if $n\geq 3$ and equality holds there exist $\alpha\in\mathbb{C}^{\times}$ and $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{C})$ such that $\mathcal{S}$ is conjugated to $\alpha(\mathbb{R}I_{n}\oplus i\mathcal{H}_{n}(\mathbb{C}))$.
-Interestingly, and in contrast with the earlier results on spaces of diagonalisable matrices, in the special case $n=2$ there are several orbits of such spaces with dimension $5$ under the action of $\mathbb{C}^{\times}\times\operatorname{\operatorname{GL}}_{n}(\mathbb{C})$ by $(\alpha,P).\mathcal{S}:=\alpha P\mathcal{S}P^{-1}$.
-This can be seen heuristically as follows: take a solution $\mathcal{S}$ with dimension $5$.
-Then as $\mathbb{C}I_{2}+\mathcal{S}$ is also a solution, we see that $\mathbb{C}I_{2}\subseteq\mathcal{S}$, and hence $\{M\in\mathcal{S}:\operatorname{\operatorname{tr}}(M)=0\}$ is a $3$-dimensional real linear subspace of $\mathfrak{sl}_{2}(\mathbb{C})$.
-Conversely, given a real linear subspace $\mathcal{M}$ of $\mathfrak{sl}_{2}(\mathbb{C})$, the condition that all the matrices in $\mathcal{M}$ are diagonalisable over $\mathbb{C}$ is equivalent to the condition that all the nonzero matrices in $\mathcal{M}$ are nonsingular.
-Hence the set of all $3$-dimensional real linear subspaces of $\mathfrak{sl}_{2}(\mathbb{C})$ that arise in this way is an open subset in the projective Grassmannian variety of all $3$-dimensional real subspaces of $\mathfrak{sl}_{2}(\mathbb{C})$, to the effect that its topological dimension is 3(63)=93(6-3)=9.
-However, under the said action the orbits are exactly the same ones as under the restricted action of the subgroup $\mathbb{C}^{\times}\times\mathrm{SL}_{2}(\mathbb{C})$, because every matrix of $\operatorname{\operatorname{GL}}_{2}(\mathbb{C})$ splits as $\alpha P$ for some $\alpha\in\mathbb{C}^{\times}$ and some $P\in\mathrm{SL}_{2}(\mathbb{C})$.
-Since the group $\mathbb{C}^{\times}\times\mathrm{SL}_{2}(\mathbb{C})$ has topological dimension $8$, this hints at the presence of several orbits (and even at infinitely many orbits).
-We can actually give a very simple example of a solution that is not conjugated to $\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C})$ under the previous action.
-Proposition 8.11.
-Consider the space $W$ consisting of all the matrices of the form
-$M(x,z)=\begin{bmatrix}ix&-\overline{z}\\ (1+i)z&-ix\end{bmatrix}\quad\text{with $(x,z)\in\mathbb{R}\times\mathbb{C}$.}$
-Then $\mathbb{C}I_{2}\oplus W$ is a $5$-dimensional real linear subspace of $\operatorname{\operatorname{M}}_{2}(\mathbb{C})$ in which all the elements are diagonalisable (over $\mathbb{C}$).
-Moreover, there is no pair $(\alpha,P)\in\mathbb{C}^{\times}\times\operatorname{\operatorname{GL}}_{2}(\mathbb{C})$ such that $\mathbb{C}I_{2}\oplus W=\alpha P(\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C}))P^{-1}$.
-Proof.
-Clearly $W$ is a $3$-dimensional real linear subspace of $\mathfrak{sl}_{2}(\mathbb{C})$.
-For all $(x,z)\in\mathbb{R}\times\mathbb{C}$, we notice that $\det(M(x,z))=x^{2}+|z|^{2}(1+i)$, which vanishes only if $x=z=0$.
-Hence, all the matrices in $W$ are diagonalisable over $\mathbb{C}$.
-Then $\mathcal{S}:=\mathbb{C}I_{2}\oplus W$ is a $5$-dimensional real linear subspace of $\operatorname{\operatorname{M}}_{2}(\mathbb{C})$ in which all the matrices are diagonalisable over $\mathbb{C}$.
-Obviously $W=\mathcal{S}\cap\mathfrak{sl}_{2}(\mathbb{C})$.
-Now, assume that there exists a pair $(\alpha,P)\in\mathbb{C}^{\times}\times\operatorname{\operatorname{GL}}_{2}(\mathbb{C})$ such that $\mathcal{S}=\alpha P(\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C}))P^{-1}$.
-All trace zero matrices in $\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C})$ actually belong to $i\mathcal{H}_{2}(\mathbb{C})$, and hence $W\subseteq\alpha P(i\mathcal{H}_{2}(\mathbb{C}))P^{-1}$.
-Hence the determinant of every element of $W$ is a real multiple of $\alpha^{2}$.
-However, it is clear from the previous computation that no $1$-dimensional real linear subspace of $\mathbb{C}$ contains the determinant of all the elements of $W$.
-Therefore, $\mathbb{C}I_{2}+W$ is not in the orbit of $\mathbb{R}I_{2}\oplus i\mathcal{H}_{2}(\mathbb{C})$.
-∎
-We suspect that classifying the orbits of the optimal solutions for $n=2$ is very difficult, if not intractable.
-Now, we set out to prove Theorem 8.10.
-We start by giving an inductive proof of the first statement, in the form of a theorem on operator spaces.
-The statement is already obvious for $n=1$.
-Let $n\geq 2$, let $V$ be an $n$-dimensional complex vector space, and $\mathcal{S}$ be a real linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{C}}(V)$ in which every element is diagonalisable (as a $\mathbb{C}$-endomorphism).
-Throughout, we will use the basic observation that the only nilpotent element of $\mathcal{S}$ is the zero endomorphism.
-We introduce the trace-orthogonal
-$\mathcal{S}^{\bot}:=\{v\in\operatorname{\operatorname{End}}_{\mathbb{C}}(V):\forall u\in\mathcal{S},\;\operatorname{\operatorname{Re}}(\operatorname{\operatorname{tr}}(uv))=0\},$
-which is a real linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{C}}(V)$ with dimension $2n^{2}-\dim\mathcal{S}$.
-Throughout the proof, we will need the following important observation.
-For a nonzero vector $x$ of $V$, we denote by $\operatorname{\operatorname{EV}}(\mathcal{S},x)$ the space of all $\lambda\in\mathbb{C}$ for which $\lambda x\in\mathcal{S}x$, or in other words the set of all complex numbers that are, for some $u\in\mathcal{S}$, the eigenvalue of $u$ for the eigenvector $x$.
-Obviously $\operatorname{\operatorname{EV}}(\mathcal{S},x)$ is a real linear subspace of $\mathbb{C}$.
-Lemma 8.12.
-Let $x$ be a nonzero vector of $V$.
-Then $\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)$ is orthogonal to $\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)$ for the $\mathbb{R}$-bilinear form $(y,z)\in\mathbb{C}^{2}\mapsto\operatorname{\operatorname{Re}}(yz)$.
-Proof.
-Let indeed $(\lambda,\mu)\in\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)\times\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)$.
-We can choose $u\in\mathcal{S}$ such that $\operatorname{\operatorname{Im}}u\subseteq\mathbb{C}x$ and $u(x)=\lambda x$, and choose $v\in\mathcal{S}^{\bot}$ such that $v(x)=\mu\,x$.
-Then $\operatorname{\operatorname{Im}}(uv)\subseteq\mathbb{C}x$ and $(uv)(x)=\lambda\mu x$.
-It follows that $\operatorname{\operatorname{tr}}(uv)=\lambda\mu$, and we conclude that $\operatorname{\operatorname{Re}}(\lambda\mu)=0$ because $u\in\mathcal{S}$ and $v\in\mathcal{S}^{\bot}$.
-∎
-Now we can move forward.
-Let $x\in V\smallsetminus\{0\}$ be arbitrary.
-We consider the real linear subspace $\mathcal{S}^{\prime}$ of all $u\in\mathcal{S}$ such that $u(x)\in\mathbb{C}x$.
-Every $u\in\mathcal{S}^{\prime}$ induces a complex diagonalisable endomorphism $\overline{u}$ of the quotient space $V/\mathbb{C}x$, and we denote by $\overline{\mathcal{S}}$ the (real) linear subspace of all such endomorphisms.
-By induction $\dim\overline{\mathcal{S}}\leq(n-1)^{2}+1$.
-By the rank theorem applied to $u\mapsto[u(x)]\in V/\mathbb{C}x$, we see that
-$\dim\mathcal{S}\leq 2(n-1)+\dim\mathcal{S}^{\prime}.$
-The kernel of $u\in\mathcal{S}^{\prime}\mapsto\overline{u}\in\overline{\mathcal{S}}$ equals $\mathcal{S}^{\mathbb{C}x}$.
-Every element in this kernel has $x$ as eigenvector.
-Assign to every $u\in\mathcal{S}^{\mathbb{C}x}$ the eigenvalue $\lambda(u)\in\mathbb{C}$ of $u$ associated with the eigenvector $x$, and note that if $\lambda(u)=0$ then $u$ is nilpotent, and hence $u=0$.
-Hence the $\mathbb{R}$-linear mapping $u\in\mathcal{S}^{\mathbb{C}x}\mapsto\lambda(u)$ is injective, and we deduce that
-$\dim(\mathcal{S}^{\mathbb{C}x})=\dim_{\mathbb{R}}(\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x))\leq 2.$
-Finally, the rank theorem yields $\dim\mathcal{S}^{\prime}=\dim(\mathcal{S}^{\mathbb{C}x})+\dim\overline{\mathcal{S}}$. This already yields
-$\dim\mathcal{S}\leq n^{2}+2.$
-Now, let us assume that $\dim\mathcal{S}=n^{2}+2$.
-Then $\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)=\mathbb{C}$, and we deduce from Lemma 8.12 that $\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)=\{0\}$.
-Varying $x$, we deduce that no operator in $\mathcal{S}^{\bot}$ has a nonzero (complex) eigenvalue.
-Let us draw some consequences of this.
-First of all, we deduce that all the operators in $\mathcal{S}^{\bot}$ are nilpotent, and hence we are in the position to use the extension of Gerstenhaber’s theorem to skew fields.
-By theorem 2 of [17], applied with $\mathbb{K}=\mathbb{C}$ and $\mathbb{K}_{0}=\mathbb{R}$, we deduce that $\dim\mathcal{S}^{\bot}\leq 2\dfrac{n(n-1)}{2}=n(n-1)$, and if equality occurs then $\mathcal{S}^{\bot}$ is reducible (because $n>1$).
-However, if $\mathcal{S}^{\bot}$ has a nontrivial invariant $\mathbb{C}$-linear subspace $W$, then by double-orthogonality $\mathcal{S}$ would contain all the $\mathbb{C}$-linear endomorphisms that vanish on $W$ and map into $W$; yet all such endomorphisms are nilpotent, and one of them is nonzero.
-Hence $\mathcal{S}^{\bot}$ is irreducible (as a subset of $\operatorname{\operatorname{End}}_{\mathbb{C}}(V)$).
-It follows that $\dim\mathcal{S}^{\bot}\leq n^{2}-n-1$, to the effect that $\dim\mathcal{S}\geq 2n^{2}-n(n-1)+1=n^{2}+n+1>n^{2}+2$.
-This contradicts an earlier result.
-We conclude that $\dim\mathcal{S}\leq n^{2}+1$.
-From now on, we assume that $\dim\mathcal{S}=n^{2}+1$ and $n\geq 3$.
-Step 1. $\mathcal{S}^{\bot}$ is intransitive.
-To see this, we consider an arbitrary vector $x\in V\smallsetminus\{0\}$.
-By coming back to the above proof, we note that $\dim\operatorname{\operatorname{EV}}(\mathcal{S}^{\mathbb{C}x},x)\geq 1$, for the contrary would yield $\dim\mathcal{S}\leq n^{2}$.
-By Lemma 8.12, we recover $\dim\operatorname{\operatorname{EV}}(\mathcal{S}^{\bot},x)\leq 1$.
-In particular $\mathbb{C}x\not\subset\mathcal{S}^{\bot}x$, and hence $V\neq\mathcal{S}^{\bot}x$.
-Step 2. $\mathcal{S}^{\bot}$ is primitively intransitive.
-Assume the contrary.
-Then we can take a maximal nonzero $\mathbb{C}$-linear subspace $W$ of $V$ such that, for the projection $\pi:V\twoheadrightarrow V/W$, the projected space $\pi\mathcal{S}^{\bot}$ is intransitive.
-By the maximality of $W$, the operator space $\pi\mathcal{S}^{\bot}$ is primitively intransitive.
-Denote by $t$ the dimension of $V/W$ over $\mathbb{C}$, and note that $1\leq t<n$.
-Consider the subspace $\mathcal{S}_{W}$ of $\mathcal{S}$ consisting of its operators that vanish on $W$.
-For every $u\in\mathcal{S}_{W}$, the induced endomorphism $\overline{u}$ of $V/W$ remains $\mathbb{C}$-diagonalisable, and the mapping $u\mapsto\overline{u}$ is injective because its kernel contains only nilpotent operators that belong to $\mathcal{S}$.
-Hence $\dim\mathcal{S}_{W}\leq 1+t^{2}$ by the first part of the proof.
-Besides, it is standard (see e.g.
-lemma 3.4 in [25], for the case of fields) that
-$\dim\mathcal{S}_{W}+\dim(\pi\mathcal{S}^{\bot})=\dim_{\mathbb{R}}\operatorname{\operatorname{Hom}}_{\mathbb{C}}(V/W,V)=2nt.$
-However, since $\pi\mathcal{S}^{\bot}$ is primitively intransitive it is also deeply intransitive (see Proposition 4.5; here $\mathbb{R}$ is infinite) and it follows that $\dim(\pi\mathcal{S}^{\bot})\leq\alpha(t,2)=t^{2}$.
-Therefore $2nt\leq 2t^{2}+1$, which yields $2t(n-t)\leq 1$ and obviously contradicts the fact that $t\in\mathopen{[\![}1,n-1\mathclose{]\!]}$.
-Step 3. Applying the classification theorem for deeply intransitive spaces.
-Since $\mathcal{S}^{\bot}$ is primitively intransitive, it is also deeply intransitive (again, by Proposition 4.5).
-Here $\dim\mathcal{S}^{\bot}=\alpha(n,2)-1$, and we observe that $n-\max(4-2,2)=n-2\geq 1$, to the effect that $\dim\mathcal{S}^{\bot}\geq\alpha(n,2)-n+\max(4-2,2)$.
-Hence we can apply Theorem 1.21, with $e=\operatorname{\operatorname{Re}}$ and $\sigma:z\mapsto\overline{z}$ the complex conjugation.
-This yields a nondegenerate sesquilinear form $B$ on $V$ such that, for $b:(x,y)\in V^{2}\mapsto\operatorname{\operatorname{Re}}(B(x,y))$, one has $\mathcal{S}^{\bot}\subseteq\mathcal{A}_{b,\mathbb{C}}$.
-Let us consider an arbitrary basis $\mathbf{B}$ of $V$, and consider the Gram matrix $P$ of $B$ in $\mathbf{B}$.
-Let $u\in\mathcal{S}^{\bot}$ have representing matrix $M$ in $\mathbf{B}$.
-Then $\forall X\in\mathbb{C}^{n},\;\operatorname{\operatorname{Re}}(X^{\star}PMX)=0$, and we deduce that $PM\in i\mathcal{H}_{n}(\mathbb{C})$.
-Step 4. Back to $\mathcal{S}$.
-Let us now denote by $\mathcal{M}$ the matrix space that represents $\mathcal{S}$ in the basis $\mathbf{B}$.
-Since $\mathcal{S}^{\bot}$ is represented in that basis by a real linear subspace of $iP^{-1}\mathcal{H}_{n}(\mathbb{C})$, we find by double-orthogonality that $\mathcal{M}\supseteq\mathcal{H}_{n}(\mathbb{C})P$.
-Note that $\dim_{\mathbb{R}}\mathcal{M}=n^{2}+1$, whence $\mathcal{H}_{n}(\mathbb{C})P$ is a (real) linear hyperplane of $\mathcal{M}$.
-Taking $Q:=P^{\star}$, we see that $\mathcal{H}_{n}(\mathbb{C})P=Q^{-1}\mathcal{H}_{n}(\mathbb{C})$.
-It remains to analyze the matrix $Q$.
-Noting that $\mathcal{M}+\mathbb{C}I_{n}$ is a real linear subspace of $\mathbb{C}$-diagonalisable matrices, we see from the maximality of the dimension of $\mathcal{S}$ that $\mathbb{C}I_{n}\subseteq\mathcal{M}$.
-It follows in particular that $Q^{-1}\mathcal{H}_{n}(\mathbb{C})$ contains a nonzero scalar matrix.
-Hence $Q=\lambda Q^{\prime}$ for some $\lambda\in\mathbb{C}^{\times}$ and some $Q^{\prime}\in\mathcal{H}_{n}(\mathbb{C})\cap\operatorname{\operatorname{GL}}_{n}(\mathbb{C})$.
-Multiplying $\mathcal{S}$ with $\lambda$, we then lose no generality in assuming that $Q$ is Hermitian.
-As in Section 8.6, we can replace $Q$ at will with a star-congruent matrix.
-Assume that $Q$ is isotropic.
-Then $Q$ is star-congruent to $Q_{1}:=\begin{bmatrix}K&[0]_{2\times(n-2)}\\ [0]_{(n-2)\times 2}&Q_{2}\end{bmatrix}$ for $K:=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$ and some invertible Hermitian matrix $Q_{2}$, and it follows that every matrix in $Q_{1}^{-1}\mathcal{H}_{n}(\mathbb{C})$ is diagonalisable.
-Yet $Q_{1}^{-1}\operatorname{\operatorname{Diag}}(1,0,\dots,0)$ is clearly nilpotent and nonzero, and hence it is not diagonalisable.
-It follows that $Q$ is nonisotropic, and hence star-congruent to $\pm I_{n}$.
-Replacing $\mathcal{M}$ with a similar subspace, we can reduce the situation to the one where $Q=\pm I_{n}$, in which case we have $\mathcal{M}\supseteq\mathcal{H}_{n}(\mathbb{C})$.
-In this reduced situation, $i\mathcal{M}\supseteq\mathbb{C}I_{n}+i\mathcal{H}_{n}(\mathbb{C})$.
-Since the dimensions (over $\mathbb{R}$) are equal we conclude that
-$i\mathcal{M}=\mathbb{C}I_{n}+i\mathcal{H}_{n}(\mathbb{C})=\mathbb{R}I_{n}\oplus i\mathcal{H}_{n}(\mathbb{C}),$
-which completes the proof.
-8.8 Subspaces of semi-simple endomorphisms
-Here we study a variation of the previous problems.
-So far we have investigated the real linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ that consist only of diagonalisable matrices, the real linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ that consist only of diagonalisable matrices with real eigenvalues, and we have just investigated the real linear subspaces $\operatorname{\operatorname{M}}_{n}(\mathbb{C})$ that consist only of diagonalisable matrices.
-Hence it seems natural to study the (real) linear subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ that consist only of matrices that are diagonalisable over the complex numbers.
-The natural generalization of this question over an arbitrary field is to consider, for an $\mathbb{F}$-vector space $V$ with finite dimension $n\geq 1$, the linear subspaces of $\operatorname{\operatorname{End}}(V)$ in which all the elements are semi-simple.
-Recall that an endomorphism $u$ of $V$ is semi-simple if every $u$-invariant subspace $W$ of $V$ has a complementary subspace that is $u$-invariant.
-Classically, this is equivalent to the property that all the irreducible divisors of the minimal polynomial of $u$ have multiplicity $1$, and this property is inherited by induced endomorphisms, meaning that if $u$ is semi-simple and $W$ is a $u$-invariant subspace then the endomorphisms of $W$ and $V/W$ induced by $u$ are also semi-simple.
-Finally, no nonzero nilpotent endomorphism is semi-simple.
-In particular, if $\mathcal{S}$ is a linear subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple, then by taking the intersection of $\mathcal{S}$ with a subspace of nilpotent endomorphisms with dimension $\dbinom{\dim V}{2}$ we see that $\dim\mathcal{S}\leq\dbinom{\dim V+1}{2}$.
-Now, we give a simple example of a large subspace of semi-simple endomorphisms.
-Assume that $V$ has a nonisotropic symmetric bilinear form $b$.
-Consider the space $\mathcal{S}_{b}$ of all $b$-symmetric endomorphisms, i.e., the endomorphisms $u$ of $V$ such that $(x,y)\mapsto b(x,u(y))$ is symmetric (or, equivalently, the $b$-selfadjoint endomorphisms).
-For every such endomorphism $u$ and every $u$-invariant subspace $W$, it is classical that $W^{\bot_{b}}$ is $u$-invariant, while $W\oplus W^{\bot_{b}}=V$ since $b$ is nonisotropic.
-It follows that every element of $\mathcal{S}_{b}$ is semi-simple.
-Moreover $\dim\mathcal{S}_{b}=\dbinom{n+1}{2}$ since $\mathcal{S}_{b}$ is isomorphic to the space of all symmetric bilinear forms on $V$.
-The next theorem shows that these are the only optimal solutions:
-Theorem 8.13.
-Let $\mathbb{F}$ be an arbitrary field, and $V$ be an $n$-dimensional vector space over $\mathbb{F}$, with $n\geq 1$.
-Let $\mathcal{S}$ be a linear subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple.
-Then:
-  1. (a)
-    $\dim\mathcal{S}\leq\dbinom{n+1}{2}$;
-  2. (b)
-    If $\dim\mathcal{S}=\dbinom{n+1}{2}$ then $\mathcal{S}=\mathcal{S}_{b}$ for some nonisotropic symmetric bilinear form $b$ on $V$.
-Note that here no cardinality assumption is made on the field $\mathbb{F}$ under consideration.
-Hence, for fields with characteristic other than $2$, the existence of a subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple, and with dimension $\dbinom{n+1}{2}$, is connected to the so-called u-invariant of $\mathbb{F}$ (the supremum of the dimensions of the nonisotropic quadratic forms over $\mathbb{F}$).
-We have already proved point (a).
-The proof of point (b) is an easy adaptation of the one given previously for diagonalisable subspaces, and here we explain the adaptation.
-First of all, since here $\mathbb{D}=\mathbb{F}$ there is no issue of characteristic when defining the orthogonal complement.
-Thus, let us take $\mathcal{S}$ as a linear subspace of $\operatorname{\operatorname{End}}(V)$ in which every element is semi-simple, and with dimension $\dbinom{n+1}{2}$.
-Then, exactly as previously we obtain that $\mathcal{S}^{\bot}$ has trivial spectrum and is irreducible (this relies upon the observation that the only nilpotent and semi-simple endomorphism is the zero endomorphism, and that semi-simplicity is inherited by induced endomorphisms).
-There is however a difficulty related to the cardinality condition in Theorem 1.7, and it can be circumvented as follows if $n>|\mathbb{F}|$, which requires that $n>3$ if $|\mathbb{F}|>2$, and $n>2$ otherwise.
-Set $k:=3$ if $|\mathbb{F}|>2$, and $k=2$ if $|\mathbb{F}|=2$.
-In that case, we proceed as in Lemma 8.8 to find, by induction, that there is a $k$-dimensional vector space $W$ and a linear subspace $\mathcal{T}$ of $\operatorname{\operatorname{End}}(W)$ with dimension $\dbinom{k+1}{2}$ and in which all the elements are semi-simple.
-Hence by induction $\mathcal{T}=\mathcal{S}_{c}$ for some nonisotropic symmetric bilinear form $c$ on $W$.
-But here $\mathbb{F}$ is finite and:
-  • Either $k=3$, in which case there is no nonisotropic symmetric bilinear form on $W$, because it is known that every quadratic form on $W$ is isotropic;
-  • Or $k=2$, in which case $|\mathbb{F}|=2$ and every symmetric bilinear form on $W$ is isotropic (because the associated quadratic form is actually linear).
-Hence, in any case we find a contradiction.
-It follows that $n\leq|\mathbb{F}|$.
-Hence, we can safely use Theorem 1.7 (actually, we could also have used theorem 3 of [18] for the case $|\mathbb{F}|>2$, but this would still leave out the case of fields with cardinality $2$) and recover that, in some basis $\mathbf{B}$, the space $\mathcal{S}^{\bot}$ is represented by the matrix space $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ for some nonisotropic matrix $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$.
-Then $\mathcal{S}$ is represented by $\operatorname{\operatorname{S}}_{n}(\mathbb{F})P$ in the same basis.
-Once more, we use the maximality of the dimension of $\mathcal{S}$ to observe that $\operatorname{\operatorname{id}}_{V}\in\mathcal{S}$, and we recover that $P^{-1}$ is symmetric, and hence $P$ is symmetric.
-Then $\operatorname{\operatorname{S}}_{n}(\mathbb{F})P=(P^{T})^{-1}(P^{T}\operatorname{\operatorname{S}}_{n}(\mathbb{F})P)=(P^{T})^{-1}\operatorname{\operatorname{S}}_{n}(\mathbb{F})$, and since $P^{T}$ is obviously nonisotropic we conclude that $\mathcal{S}=\mathcal{S}_{b}$ for the nonisotropic symmetric bilinear form $b$ that is represented by $P^{T}$ in $\mathbf{B}$.
-In the special case $\mathbb{F}=\mathbb{R}$, we obtain exactly the same optimal spaces as for the diagonalisability assumption.
-In other words, up to conjugation $\operatorname{\operatorname{S}}_{n}(\mathbb{R})$ is the only linear subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ with dimension $\dbinom{n+1}{2}$ and in which every element is diagonalisable over $\mathbb{C}$.
-Judging from the difficulties encountered in Section 8.7, we suspect that generalizing the previous approach to semilinear subspaces of semi-simple endomorphisms over general division rings could be very difficult.
-8.9 Towards linear subspaces of matrices with no nonzero nilpotent element?
-Let us finish with another variation of the previous problem.
-Instead of considering linear subspaces of semi-simple endomorphisms, why not consider an even looser property, which is the absence of nonzero nilpotent elements in the subspace, still for $\mathbb{F}$-linear endomorphisms of an $\mathbb{F}$-vector space $V$ with finite dimension $n$?
-For such a subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}(V)$, we still have $\dim\mathcal{S}\leq\dbinom{n+1}{2}$ with the same proof as for spaces of diagonalisable endomorphisms (for algebraically closed fields, the basic theory of projective algebraic varieties yields the upper bound $\dim\mathcal{S}\leq n$: apply [27] corollary 1.7 to the coefficients of the characteristic polynomial).
-However, for the analysis of the case of equality, the previous method fails because the property is not inherited by induced endomorphisms anymore (if we have a non-nilpotent endomorphism $u$ that leaves invariant a linear subspace $W$, it is possible that $u_{W}$ be nilpotent and nonzero).
-Hence, it seems difficult to classify the spaces that reach the critical dimension $\dbinom{n+1}{2}$.
-First of all, if we take a nonisotropic bilinear form $b$ on $V$, then the space $\mathcal{S}_{b}$ of all $b$-symmetric endomorphisms contains no nonzero nilpotent endomorphism.
-Indeed, for $u\in\mathcal{S}_{b}$, the kernel and the range of $b$ have no nonzero common vector (for $x\in\operatorname{\operatorname{Ker}}u\cap\operatorname{\operatorname{Im}}u$, we write $x=u(y)$ to get $b(x,x)=b(x,u(y))=b(y,u(x))=0$, and deduce from the nonisotropy of $b$ that $x=0$).
-In contrast with the previous case of spaces of semi-simple endomorphisms, here we do not require that $b$ be symmetric.
-A naive conjecture is that every space with no nonzero nilpotent element and dimension $\dbinom{n+1}{2}$ equals $\mathcal{S}_{b}$ for some nonisotropic bilinear form $b$ on $V$.
-Yet this is disproved in the case of the field of reals by appealing to geometric considerations.
-Indeed, in that case, the set of all $\dbinom{n+1}{2}-$dimensional linear subspaces that contain no nonzero nilpotent matrix is clearly open in the Grassmanian manifold of all $\dbinom{n+1}{2}-$dimensional linear subspaces.
-Yet it is easily proved that the set of all $\mathcal{S}_{b}$ spaces, when $b$ ranges over the nonisotropic bilinear forms on $V$, is a manifold with dimension $n^{2}-1$ (where the “minus 11” comes from the observation that $\mathcal{S}_{b}=\mathcal{S}_{b^{\prime}}$ whenever $b$ and $b^{\prime}$ are linearly dependent).
-Yet the Grassmannian manifold under consideration has dimension $n^{2}\dbinom{n+1}{2}-\dbinom{n+1}{2}^{2}=\dfrac{n^{2}(n^{2}-1)}{4}$, which largely exceeds $n^{2}-1$ when $n>2$.
-In contrast, for $n=2$ it is an easy exercise to validate the conjecture.
-Index